### v1

# Tensor Product Attention Is All You Need

Abstract
Scaling language models to handle longer input sequences typically necessitates large key-value (KV) caches, resulting in substantial memory overhead during inference. In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, significantly shrinking KV cache size at inference time. By factorizing these representations into contextual low-rank components (contextual factorization) and seamlessly integrating with RoPE, TPA achieves improved model quality alongside memory efficiency. Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling. Through extensive empirical evaluation of language modeling tasks, we demonstrate that T6 exceeds the performance of standard Transformer baselines including MHA, MQA, GQA, and MLA across various metrics, including perplexity and a range of renowned evaluation benchmarks. Notably, TPA’s memory efficiency enables the processing of significantly longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models. The code is available at https://github.com/tensorgi/T6.
1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Bubeck et al., 2023). As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completions. However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (Zhang et al., 2023c; Liu et al., 2024c). Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
Figure 1: Tensor Product Attention (TPA) in the Tensor ProducT ATTenTion Transformer (T6). Different from multi-head attention, in each layer, firstly the hidden state goes through different linear layers to get the latent factor matrices $\mathbf{A}$’s and $\mathbf{B}$’s for query, key, and value. We additionally apply RoPE to $\mathbf{B}_{Q}$ and $\mathbf{B}_{K}$ for query and key. Then the multi-head query, key, and value vectors are attained by the tensor product of $\mathbf{A}_{(\cdot)}$ and $\mathbf{B}_{(\cdot)}$. Finally, the output of TPA is produced by scaled dot-product attention followed by linear projection of concatenated results of multiple heads.
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (Child et al., 2019) or token eviction strategies (Zhang et al., 2023c; Xiao et al., 2024; Ribar et al., 2024), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (He & Zhai, 2024), at the expense of increased I/O latency. Attention variants like multi-query attention (MQA) (Shazeer, 2019) and grouped-query attention (GQA) (Ainslie et al., 2023) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (Hu et al., 2022) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates runtime. The recently introduced Multi-head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but needs additional position-encoded parameters per head due to incompatibility with Rotary Position Embedding (RoPE) efficiently (Su et al., 2024b).
In order to overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), as illustrated in Figure 1, a novel architecture that uses higher-order tensors to factorize queries (Q), keys (K), and values (V) during attention computation. By dynamically factorizing activations rather than static weights (e.g., LoRA), TPA constructs low-rank, contextual representations that substantially reduce KV cache memory usage with improved representational capacity. In practice, TPA can reduce the memory overhead by an order of magnitude compared to standard multi-head attention (MHA) with lower pretraining validation loss (perplexity) and improved downstream performance.
A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024b), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023) and Gemma (Team et al., 2024).
Our primary contributions are summarized as follows:
  • We propose Tensor Product Attention (TPA), A mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor-decompositions to achieve 10×10$\times$10 × or more reduction in inference-time KV cache size relative to standard attention mechanism (Vaswani et al., 2017) with improved performance compared to previous methods such as MHA, MQA, GQA, and MLA. In addition, we unify existing attention mechanisms by revealing that MHA, MQA, and GQA all arise naturally as non-contextual variants of TPA.
  • We propose Tensor ProducT ATTenTion Transformer (T6), a new TPA-based model architecture for sequence modeling. On language modeling experiments, T6 consistently improves validation perplexity and downstream evaluation performance with reduced KV cache size.
  • We show TPA integrates seamlessly with RoPE (Su et al., 2024b), facilitating easy adoption in popular foundation model architectures such as LLaMA and Gemma.
(a) Training Loss
(b) Validation Loss
Figure 2: Training loss and validation loss of pretraining large-size (773M) models with different attention mechanisms on the FineWeb-Edu-100B dataset.
2 Background
In this section, we review several classical forms of attention: Scaled Dot-Product Attention, Multi-Head Attention (MHA) (Vaswani et al., 2017), Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped Query Attention (GQA) (Ainslie et al., 2023), as well as Rotary Position Embedding (RoPE, Su et al. (2024b)). We also introduce a recent method called Multi-head Latent Attention (MLA) used in DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b).
Notations. We use bold uppercase letters (e.g., $\mathbf{X}$, $\mathbf{Q}$) for matrices, bold lowercase (e.g., $\mathbf{a}$, $\mathbf{b}$) for vectors, and italic uppercase (e.g., $\bm{W}^{Q}_{i}$) for learnable parameter matrices. We denote by $[n]$ the set $\{1,\ldots,n\}$ for some positive integer $n$. We use $\top$ to denote the transpose of a vector or a matrix. Let $d_{\text{model}}$ be the embedding dimension, $h$ the number of attention heads, $d_{h}$ the dimension per head, $\mathbf{x}_{t}\in\mathbb{R}^{d}$ the input for the $t$-th token at a given attention layer, $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ denotes the input embeddings for $T$ tokens, and $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ denote the queries, keys, and values of $h$ heads for $T$ tokens. With a little abuse of notation, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, $\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ denote the $i$-th head of queries, keys, and values, and $\mathbf{Q}_{t}$, $\mathbf{K}_{t}$, $\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the heads of the query, key, and value for $t$-th token.
Throughout the paper, $\bm{W}^{Q},\bm{W}^{K},\bm{W}^{V}$ denote projection matrices for queries, keys, and values, respectively. In multi-head attention, each head is associated with its own set of $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}$, and each has dimension $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$, where $d_{k}$ is typically set to $d_{h}$, the dimension of each head. [^5] Often, one sets $h\times d_{h}=d_{\text{model}}$, so each head has query/key/value dimension $d_{h}$. Similarly, we have an output projection matrix $\bm{W}^{O}\in\mathbb{R}^{(h\cdot d_{h})\times d_{\text{model}}}$. For methods like MQA and GQA, some of these are shared or partially shared across heads, but their shapes remain consistent.
We define the tensor product of two vectors as follows: for vectors $\mathbf{a}\in\mathbb{R}^{m},\mathbf{b}\in\mathbb{R}^{n}$, the tensor product of $\mathbf{a}$ and $\mathbf{b}$ is:
$\mathbf{a}\otimes\mathbf{b}=\mathbf{C}\in\mathbb{R}^{m\times n},\text{with}~{}C_{ij}=a_{i}b_{j},$
where $a_{i}$ and $b_{j}$ are the $i$-th and $j$-th elements of $\mathbf{a}$ and $\mathbf{b}$ respectively, and $C_{ij}$ is the $(i,j)$-th entry of $\mathbf{C}$. We also define the vectorization of a matrix $\mathbf{C}\in\mathbb{R}^{m\times n}$ by:
$\text{vec}(\mathbf{C})=\mathbf{d}\in\mathbb{R}^{mn},\text{with}~{}d_{i\cdot n+j}=C_{ij},$
where $d_{i\cdot n+j}$ is the $(i\cdot n+j)$-th element of $\mathbf{d}$.
2.1 Scaled Dot-Product Attention
Scaled dot-product attention (Vaswani et al., 2017) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$). It produces a weighted combination of the values ($\mathbf{V}$). Formally, the attention output is:
$\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Softmax}\Bigl{(}\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\Bigr{)}\,\mathbf{V},$
where each of $\mathbf{Q},\mathbf{K},\mathbf{V}$ is an $(n\times d_{k})$ matrix for $n$ tokens and key dimension $d_{k}$. The division by $\sqrt{d_{k}}$ stabilizes training by controlling the scale of the inner products.
2.2 Multi-Head Attention (MHA)
Multi-Head Attention (MHA) extends scaled dot-product attention by dividing the model’s internal representation into several heads. Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information. For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
$\mathbf{Q}_{t,i}=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},\quad\mathbf{K}_{t,i}=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},\quad\mathbf{V}_{t,i}=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
$\textbf{head}_{i}=\operatorname{Attention}\Bigl{(}\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\Bigr{)},$
where $\bm{W}_{i}^{Q},\bm{W}_{i}^{K},\bm{W}_{i}^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ are learnable projection matrices for the $i$-th head, $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$. After computing each head’s attention, the outputs are concatenated and mapped back to the original dimension via another matrix $\bm{W}^{O}\in\mathbb{R}^{hd_{h}\times d_{\text{model}}}$:
$\operatorname{MHA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\,\bm{W}^{O}.$
MHA can capture a rich set of dependencies while each head focuses on different subspaces.
2.3 Multi-Query Attention (MQA)
Multi-Query Attention (MQA) (Shazeer, 2019) significantly reduces memory usage by sharing keys and values across heads, while still preserving unique query projections. For a sequence of embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$,
$\mathbf{Q}_{i}=\mathbf{X}\bm{W}^{Q}_{i},\quad\mathbf{K}_{\text{shared}}=\mathbf{X}\bm{W}^{K}_{\text{shared}},\quad\mathbf{V}_{\text{shared}}=\mathbf{X}\bm{W}^{V}_{\text{shared}}.$
Hence, each head $i$ only has a distinct query $\mathbf{Q}_{i}\in\mathbb{R}^{T\times d_{k}}$, but shares the same key $\mathbf{K}_{\text{shared}}\in\mathbb{R}^{T\times d_{k}}$ and value $\mathbf{V}_{\text{shared}}\in\mathbb{R}^{T\times d_{k}}$. In practice, this means:
$\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},\quad\bm{W}^{K}_{\text{shared}},\bm{W}^{V}_{\text{shared}}\;\in\;\mathbb{R}^{\,d_{\text{model}}\times d_{k}}.$
The resulting MQA operation is:
$\operatorname{MQA}(\mathbf{X})=\operatorname{Concat}\Bigl{(}\textbf{head}_{1},\dots,\textbf{head}_{h}\Bigr{)}\,\bm{W}^{O},$
where
$\textbf{head}_{i}=\operatorname{Attention}\bigl{(}\mathbf{Q}_{i},\mathbf{K}_{\text{shared}},\mathbf{V}_{\text{shared}}\bigr{)}.$
By sharing these key and value projections, MQA cuts down on memory usage (especially for the key-value cache in autoregressive inference) but loses some expressivity since all heads must rely on the same key/value representations.
2.4 Grouped Query Attention (GQA)
Grouped Query Attention (GQA) (Ainslie et al., 2023) generalizes MHA and MQA by grouping heads. Specifically, we partition the $h$ total heads into $G$ groups. Each group has a single set of keys and values, but each individual head within that group still retains its own query projection. Formally, if $g(i)$ maps a head $i\in[h]$ to its group index $g\in[G]$, then:
$\mathbf{K}_{g(i)}=\mathbf{X}\,\bm{W}^{K}_{g(i)},\quad\mathbf{V}_{g(i)}=\mathbf{X}\,\bm{W}^{V}_{g(i)},\quad\mathbf{Q}_{i}=\mathbf{X}\,\bm{W}^{Q}_{i},$
and
$\text{head}_{i}=\operatorname{Attention}\Bigl{(}\mathbf{Q}_{i},\mathbf{K}_{g(i)},\mathbf{V}_{g(i)}\Bigr{)}.$
Again, $\bm{W}^{K}_{g},\bm{W}^{V}_{g}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ for each group $g$, and $\bm{W}^{Q}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$ for each head $i$. The complete output is again a concatenation of all heads:
$\operatorname{GQA}(\mathbf{X})=\operatorname{Concat}\Bigl{(}\text{head}_{1},\ldots,\text{head}_{h}\Bigr{)}\bm{W}^{O}.$
By adjusting $G$ between $1$ and $h$, GQA can interpolate between sharing all key/value projections across heads (i.e., MQA) and having one set of projections per head (i.e., MHA).
2.5 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; Su et al., 2024b) to encode positional information in the query/key vectors. Specifically, let $\mathrm{RoPE}_{t}$ denote the rotation operator $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ corresponding to the $t$-th position. $\mathbf{T}_{t}$ is a block-diagonal matrix, which consists of block-diagonal matrix$\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$, $j\in\{1,\cdots,d_{h}/2\}$, where $\{\theta_{j}\}$ are pre-defined frequency parameters, e.g., $\theta_{j}=1/10000^{2j/d_{h}}$. Then we define
$\operatorname{RoPE}\left(\mathbf{Q}_{t}\right)\triangleq\mathbf{Q}_{t}\mathbf{T}_{t},\quad\text{where }\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}.$
A fundamental property is that
$\mathbf{T}_{t}\,\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (2.1)
which ensures that relative positions $(t-s)$ are preserved, thereby providing a form of translation invariance in the rotary position embedding.
2.6 Multi-head Latent Attention (MLA)
Below, we briefly outline the Multi-head Latent Attention (MLA) approach used by DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b). MLA introduces a low-rank compression of the keys and values to reduce the Key-Value (KV) caching cost at inference.
$\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},\quad(\bm{W}^{DKV}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}}),$
$\operatorname{Concat}\bigl{(}\mathbf{K}_{1}^{C},\mathbf{K}_{2}^{C},\ldots,\mathbf{K}_{h}^{C}\bigr{)}=\mathbf{K}^{C}$ $=\mathbf{C}^{KV}\bm{W}^{UK},\quad(\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times d_{h}h}),$
$\mathbf{K}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{X}\bm{W}^{KR}\bigr{)},\quad(\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}}),$
$\mathbf{K}_{i}$ $=\operatorname{Concat}\bigl{(}\mathbf{K}_{i}^{C},\mathbf{K}^{R}\bigr{)},$
$\operatorname{Concat}\bigl{(}\mathbf{V}_{1}^{C},\mathbf{V}_{2}^{C},\ldots,\mathbf{V}_{h}^{C}\bigr{)}=\mathbf{V}^{C}$ $=\mathbf{C}^{KV}\bm{W}^{UV},\quad(\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times d_{h}h}),$
where $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the compressed KV latent (with $d_{c}\ll d_{h}h$), and $\operatorname{RoPE}(\cdot)$ represents the RoPE transform applied to the separate key embeddings $\mathbf{K}^{R}$ of dimension $d_{h}^{R}$. Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ need to be cached, reducing KV memory usage while largely preserving performance compared to standard MHA (Vaswani et al., 2017).
MLA also compresses the queries, lowering their training-time memory footprint:
$\mathbf{C}^{Q}$ $=\mathbf{X}\bm{W}^{DQ},\quad(\bm{W}^{DQ}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}^{\prime}}),$
$\operatorname{Concat}\bigl{(}\mathbf{Q}_{1}^{C},\mathbf{Q}_{2}^{C},\,\ldots,\mathbf{Q}_{h}^{C}\bigr{)}=\mathbf{Q}^{C}$ $=\mathbf{C}^{Q}\bm{W}^{UQ},\quad(\bm{W}^{UQ}\in\mathbb{R}^{\,d_{c}^{\prime}\times d_{h}h}),$
$\operatorname{Concat}\bigl{(}\mathbf{Q}_{1}^{R},\,\mathbf{Q}_{2}^{R},\,\ldots,\,\mathbf{Q}_{h}^{R}\bigr{)}=\mathbf{Q}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{C}^{Q}\bm{W}^{QR}\bigr{)},\quad(\bm{W}^{QR}\in\mathbb{R}^{d_{c}^{\prime}\times d_{h}^{R}h}),$
$\mathbf{Q}$ $=\operatorname{Concat}\bigl{(}\mathbf{Q}^{C},\mathbf{Q}^{R}\bigr{)}.$
Here, $\mathbf{C}^{Q}\in\mathbb{R}^{T\times d_{c}^{\prime}}$ (with $d_{c}^{\prime}\ll d_{h}h$) is the compressed query latent. As above, each $\bm{W}^{DQ}$, $\bm{W}^{UQ}$, and $\bm{W}^{QR}$ connects these lower-dimensional query latents back to $h$ heads of dimension $d_{h}+d_{h}^{R}$.
Given compressed queries, keys, and values, the final attention output for the $t$-th token is:
$\mathbf{O}_{i}$ $=\operatorname{Softmax}\Bigl{(}\tfrac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\top}}{\sqrt{d_{h}+d_{h}^{R}}}\Bigr{)}\;\mathbf{V}_{i}^{C},$
$\mathbf{U}$ $=\operatorname{Concat}\bigl{(}\mathbf{O}_{1},\mathbf{O}_{2},\ldots,\mathbf{O}_{h}\bigr{)}\bm{W}^{O},$
where $\bm{W}^{O}\in\mathbb{R}^{(d_{h}h)\times\,d_{\text{model}}}$ is the output projection.
In inference time, $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ can be cached to accelerate decoding. In detail, when RoPE is ignored, the inner product $\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ (where $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{d}$) of the $i$-th head between $t$-th and $s$-th tokens can be calculated using the hidden state $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ for $t$-th token and the cached latent state $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ for $s$-th token:
$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (2.2)
where $\bm{W}_{i}^{(\cdot)}$ is the $i$-th head of the original weight, and $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ can be computed previously for faster decoding. However, this process fails when RoPE is considered according to Su (2024). Since RoPE can be considered as multiplication with a block-diagonal matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ (see Section 2.5), with the property (2.1) that $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$, then
[…]
Different from (2.2), acceleration by pre-computing $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t-s}}(\bm{W}^{UK}_{i})^{\top}]$ fails since it varies for different $(t,s)$ position pairs. Therefore, MLA adds the additional $\mathbf{k}_{t}^{R}$ part with a relatively smaller size for RoPE compatibility. In Section 3.2, we will show that TPA addresses the issue of RoPE-incompatibility by applying tensor product.
3 Tensor Product Attention
In this section, we provide a detailed description of our proposed Tensor Product Attention (TPA), which allows contextual low-rank factorization for queries, keys, and values. First, we explain how TPA factorizes queries, keys, and values with explicit tensor shapes. Next, we describe how TPA can be integrated into the multi-head attention framework and how it reduces memory consumption in KV caching at inference time. Finally, we show how RoPE can seamlessly integrate with TPA (including a pre-rotated variant).
3.1 Tensor Factorization of Queries, Keys, and Values
Let $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ for $t=1,\dots,T$ be the hidden-state vector corresponding to the $t$-th token in a sequence of length $T$. A typical multi-head attention block has $h$ heads, each of dimension $d_{h}$, satisfying $d_{\text{model}}=h\times d_{h}$. Standard attention projects the entire sequence into three tensors, $\mathbf{Q},\;\mathbf{K},\;\mathbf{V}\;\in\mathbb{R}^{T\times h\times d_{h}}$, where $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the slices for the $t$-th token.
Contextual Factorization (CF). Instead of forming each head’s query, key, or value via a single linear map, TPA factorizes each $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$ into a sum of (contextual) tensor products whose ranks are $R_{q}$, $R_{k}$, and $R_{v}$, respectively and may differ. Specifically, for each token $t$, with a small abuse of notation, we define:
$\mathbf{Q}_{t}$ $=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}},$ (3.1)
$\mathbf{K}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{K}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}},$ (3.2)
$\mathbf{V}_{t}$ $=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}.$ (3.3)
Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ adds up to form the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Similarly, analogous definitions apply to key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
Latent Factor Maps. Each factor in the tensor product depends on the token’s hidden state $\mathbf{x}_{t}$. For example, for queries, we can write:
$\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}_{r}\,\mathbf{x}_{t}\in\mathbb{R}^{h},\quad\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}_{r}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
and similarly for keys and values.
One often merges the rank index into a single output dimension. For instance, for queries:
$\mathbf{a}^{Q}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot h},\quad\mathbf{b}^{Q}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot d_{h}},$
which are then reshaped into $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{q}\times d_{h}}$. Summing over $R_{q}$ and scaled by $\frac{1}{R_{q}}$ yields
$\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h\times d_{h}}.$
Repeating for all tokens reconstitutes $\mathbf{Q}\in\mathbb{R}^{T\times h\times d_{h}}$. Similar procedures can be applied to obtain $\mathbf{K}$ and $\mathbf{V}$ with ranks $R_{k}$ and $R_{v}$, respectively.
Scaled Dot-Product Attention. Once $\mathbf{Q},\mathbf{K},\mathbf{V}$ are factorized, multi-head attention proceeds as in standard Transformers. For each head $i\in\{1,\dots,h\}$:
$\textbf{head}_{i}=\operatorname{Softmax}\Bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,\mathbf{Q}_{i}\,(\mathbf{K}_{i})^{\top}\Bigr{)}\;\mathbf{V}_{i},$ (3.4)
where $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ are the slices along the head dimension. Concatenating these $h$ heads along the last dimension yields an $\mathbb{R}^{T\times(h\cdot d_{h})}$ tensor, which is projected back to $\mathbb{R}^{T\times d_{\text{model}}}$ by an output weight matrix $\bm{W}^{O}\in\mathbb{R}^{(h\cdot d_{h})\times d_{\text{model}}}$:
$\operatorname{TPA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\bm{W}^{O}.$ (3.5)
Parameter Initialization. We initialize the weight matrices $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, $\bm{W}_{r}^{b^{V}}$ using Xavier initialization (Glorot & Bengio, 2010). Specifically, each entry of the weight matrix is drawn from a uniform distribution with bounds $[-\sqrt{6/(n_{\text{in}}+n_{\text{out}})},\sqrt{6/(n_{\text{in}}+n_{\text{out}})}]$, where $n_{\text{in}}$ and $n_{\text{out}}$ are the input and output dimensions of the respective weight matrices. This initialization strategy helps maintain the variance of activations and gradients across the network.
3.2 RoPE Compatibility and Acceleration
In a typical workflow of adding RoPE to standard multi-head attention, one first computes $\mathbf{Q}_{t},\mathbf{K}_{s}\in\mathbb{R}^{h\times d_{h}}$ of the $t$-th token and $s$-th token and then applies:
$\mathbf{Q}_{t}\mapsto\widetilde{\mathbf{Q}_{t}}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t}),\qquad\mathbf{K}_{s}\mapsto\widetilde{\mathbf{K}_{s}}=\operatorname{RoPE}_{s}(\mathbf{K}_{s}).$
Direct Integration. A useful optimization is to integrate RoPE directly into the TPA factorization. For example, one can pre-rotate the token-dimension factors:
$\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\;\longleftarrow\;\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)},$ (3.6)
yielding a pre-rotated key representation:
$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(s)}(\mathbf{x}_{t})\bigr{)}=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
Thus, each $\mathbf{K}_{t}$ is already rotated before caching, removing the need for explicit rotation at the decoding time and accelerating autoregressive inference. Depending on hardware and performance requirements, one can also adopt different RoPE integration approaches for training and inference.
Theorem 1 (RoPE’s Compatibility with TPA).
Let $\mathbf{Q}_{t}$ be factorized by TPA as
$\mathbf{Q}_{t}=\frac{1}{R_{Q}}\,\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\mathbf{B}_{Q}(\mathbf{x}_{t})\;\in\mathbb{R}^{h\times d_{h}},$
where $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times d_{h}}$. Then we have:
$\operatorname{RoPE}({\mathbf{Q}}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),\qquad\text{where }\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}.$ (3.7)
In addition, assume $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ are factorized by TPA and then rotated by $\operatorname{RoPE}_{t},\operatorname{RoPE}_{s}$. Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$. Then we have
$\operatorname{RoPE}_{t-s}(\mathbf{Q}_{t})\mathbf{K}_{s}^{\top}=\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top},$
Focusing on individual heads $i$, the above matrix equality implies:
$\operatorname{RoPE}_{t-s}\bigl{(}\mathbf{q}_{t,i}\bigr{)}^{\top}\mathbf{k}_{s,i}=\widetilde{\mathbf{q}}_{t,i}^{\top}\,\widetilde{\mathbf{k}}_{s,i}.$
where $\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}}$ is the $i$-th query head of $t$-th token, and $\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}$ is the $j$-th key head of $s$-th token, and
$\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}},\quad\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
Theorem 1 indicates that TPA does not break RoPE’s relative translational property. We prove Theorem 1 in Appendix A. In short, $\operatorname{RoPE}_{t}$ acts as a block-diagonal orthogonal transform (i.e., a matrix $\mathbf{T}_{t}$) on $\mathbf{B}_{Q}(\mathbf{x}_{t})$. Consequently, $\mathbf{A}_{Q}(\mathbf{x}_{t})$ remains unchanged, while each column of $\mathbf{B}_{Q}(\mathbf{x}_{t})$ is rotated appropriately, preserving the TPA structure.
3.3 KV Caching and Memory Reduction
In autoregressive decoding, standard attention caches $\mathbf{K}_{t},\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ for each past token $t$. This accumulates to $\mathbb{R}^{T\times h\times d_{h}}$ for keys and $\mathbb{R}^{T\times h\times d_{h}}$ for values, i.e., $2\,T\,h\,d_{h}$ total.
TPA Factorized KV Caching. Instead of storing the full $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$, TPA stores only their factorized ranks. Specifically, we keep
$\mathbf{A}_{K}(\mathbf{x}_{t}),\,\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\quad\text{and}\quad\mathbf{A}_{V}(\mathbf{x}_{t}),\,\mathbf{B}_{V}(\mathbf{x}_{t}),$
where $\mathbf{A}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times h},\;\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times d_{h}},\;\mathbf{A}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times h},\;\mathbf{B}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times d_{h}}$.
Hence, the memory cost per token is
$\underbrace{R_{K}(h+d_{h})}_{\text{for K}}\;+\;\underbrace{R_{V}(h+d_{h})}_{\text{for V}}=(\,R_{K}+R_{V}\,)\,\bigl{(}h+d_{h}\bigr{)}.$
Compared to the standard caching cost of $2\,h\,d_{h}$, the ratio is:
$\frac{\,(R_{K}+R_{V})\,(h+d_{h})\,}{2\,h\,d_{h}\,}.$
For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll d_{h}$ (e.g., $1$ or $2$) often yields 10×10$\times$10 × or more reduction.
Table 1: Comparison of different attention mechanisms. Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively. Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Section 3.5. For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
Method KV Cache # Parameters # Query Heads # KV Heads
MHA $2hd_{h}$ $4d_{\text{model}}^{2}$ $h$ $h$
MQA $2d_{h}$ $(2+2/h)d_{\text{model}}^{2}$ $h$ $1$
GQA $2gd_{h}$ $(2+2g/h)d_{\text{model}}^{2}$ $h$ $g$
MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $+d_{\text{model}}d_{h}^{R}+d_{c}(d_{\text{model}}+2hd_{h})$ $h$ $h$
TPA $(R_{K}+R_{V})(h+d_{h})$ $d_{\text{model}}(R_{Q}+R_{K}+R_{V})(h+d_{h})+d_{\text{model}}\,hd_{h}$ $h$ $h$
TPA (KVonly) $(R_{K}+R_{V})(h+d_{h})$ $d_{\text{model}}(R_{K}+R_{V})(h+d_{h})+2d_{\text{model}}\,hd_{h}$ $h$ $h$
TPA (Non-contextual A) $(R_{K}+R_{V})d_{h}$ $(R_{Q}+R_{K}+R_{V})(d_{\text{model}}d_{h}+h)+d_{\text{model}}\,hd_{h}$ $h$ $h$
TPA (Non-contextual B) $(R_{K}+R_{V})h$ $(R_{Q}+R_{K}+R_{V})(d_{\text{model}}h+d_{h})+d_{\text{model}}\,hd_{h}$ $h$ $h$
3.4 Unifying MHA, MQA, and GQA as Non-contextual TPA
3.4.1 MHA as Non-contextual TPA
Standard multi-head attention (MHA) can be viewed as a specific instance of TPA in which: 1) the rank is set equal to the number of heads; 2) the head dimension factor is non-contextual (i.e., independent of the $t$-th token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$); 3) the token dimension factor is a linear function of $\mathbf{x}_{t}$.
To match MHA with TPA, let $R_{Q}=R_{K}=R_{V}=h$. Focusing on $\mathbf{Q}_{t}$:
  1. (a)
    Non-contextual head factors. Define
[…]
    so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
  2. (b)
    Contextual token factors. Define
    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$ (3.9)
    where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the per-head query projection defined before, hence $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$ dependent on $\mathbf{x}_{t}$.
Substituting (3.8)–(3.9) into (3.1) gives:
$\mathbf{Q}_{t}=\sum_{i=1}^{h}\Big{[}\mathbf{e}_{i}\otimes\bigl{(}(\bm{W}^{Q}_{i})^{\top}\,\mathbf{x}_{t}\bigr{)}\Big{]}\in\mathbb{R}^{h\times d_{h}}.$ (3.10)
Each term $\mathbf{e}_{i}\otimes\bigl{(}({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\bigr{)}$ in (3.10) contributes only to the $i$-th row, reconstituting the usual MHA form of $\mathbf{Q}_{t}$. Analogous constructions hold for $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$ using $\bm{W}^{K}_{i},\bm{W}^{V}_{i}$. Thus, MHA is a non-contextual, full-rank variant of TPA.
TPA with Non-contextual A. More broadly, TPA can use non-contextual head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}\in\mathbb{R}^{h}$ (i.e., independent of $\mathbf{x}_{t}$), while allowing $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ to remain context-dependent. Then, for keys:
$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}\otimes\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$
and similarly for queries/values. This reduces per-token computations and can be effective when head-dimension relationships are relatively stable across all tokens.
MQA and GQA as Non-Contextual TPA. Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped Query Attention (GQA) (Ainslie et al., 2023) also emerge naturally from TPA by restricting the head-dimension factors to be non-contextual and low-rank:
  • MQA as Rank-1 TPA. In MQA, all heads share a single set of keys/values, corresponding to $R_{K}=R_{V}=1$ along the head dimension. Concretely,
    $\mathbf{K}_{t}=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
    forces every head to use the same $\mathbf{K}_{t},\mathbf{V}_{t}$. Each head retains a distinct query projection, matching the MQA design.
  • GQA as Grouped Rank-1 TPA. GQA partitions $h$ heads into $G$ groups, each sharing keys/values within that group. In TPA form, each group $g$ has a dedicated non-contextual factor pair $\mathbf{a}^{K}_{g},\mathbf{a}^{V}_{g}\in\mathbb{R}^{h}$, which acts as a “mask” for the heads in that group. Varying $G$ from $1$ to $h$ interpolates from MQA to standard MHA.
Hence, by constraining TPA’s head-dimension factors to be constant masks (one for MQA; multiple for GQA), these popular variants are recovered as special cases.
3.5 Other Variants of TPA
TPA with Non-contextual B. Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to $\mathbf{x}_{t}$. For keys:
$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{K}_{r},$
and similarly for keys/values. This arrangement is effective if the token-dimension structure remains mostly uniform across the sequence, while the head-dimension factors capture context.
TPA KV Only. One can preserve a standard query mapping,
$\mathbf{Q}_{t}=\bm{W}^{Q}\,\mathbf{x}_{t}\in\mathbb{R}^{h\times d_{h}},$
and factorize only the keys and values. This leaves the query projection as the original linear transformation while reducing memory usage via factorized KV caching.
TPA KV with Shared $\mathbf{B}$. Another variant is to share the token-dimension factors of keys and values:
$\mathbf{b}^{K}_{r}(\mathbf{x}_{t})=\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$
lowering parameter counts and the KV cache footprint. While it constrains $\mathbf{K}$ and $\mathbf{V}$ to be formed from the same token basis, it can still perform well and provide additional memory savings.
Nonlinear Head Factors. Rather than applying purely linear mappings to the head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}$, one may introduce element-wise nonlinearities such as $\sigma(\cdot)$ or $\operatorname{softmax}(\cdot)$. This effectively yields a Mixture of Heads Attention (MoH Attention), where each component becomes a learned mixture weight modulated by the nonlinearity.
Discussion. These variants illustrate TPA’s versatility in balancing memory cost, computational overhead, and representation power. By choosing which dimensions (heads or tokens) remain contextual and adjusting ranks $(R_{Q},R_{K},R_{V})$, TPA unifies multiple existing attention mechanisms—such as MHA, MQA, and GQA—under one framework, while potentially reducing the KV cache size by an order of magnitude during autoregressive inference.
3.6 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (Shazeer, 2020; Touvron et al., 2023).
TPA QKV Factorization. Let each token’s hidden-state vector be $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, and we follow Section 3.1 to project the entire sequence into three tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\;\mathbb{R}^{T\times h\times d_{h}}$, where $\mathbf{Q}_{t},\;\mathbf{K}_{t},\;\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the slices for the $t$-th token. The factor components $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ are produced by linear transformations on $\mathbf{x}_{t}$. For instance, letting $\bm{W}^{a^{Q}}_{r}\in\mathbb{R}^{h\times d_{\text{model}}}$ and $\bm{W}^{b^{Q}}_{r}\in\mathbb{R}^{d_{h}\times d_{\text{model}}}$, we have:
$\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}_{r}\,\mathbf{x}_{t},\quad\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}_{r}\,\mathbf{x}_{t}.$
In practice, we merge all ranks $r$ into a single dimension of the output, reshape, and sum over rank indices; see Section 3.1 for details. The factorization for K and V follows the same pattern.
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (Su et al., 2024b) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.6) and Theorem 1.
Attention Step and Output Projection. Once we have $\mathbf{Q},\mathbf{K},\mathbf{V}$ factorized per token with RoPE applied on $\mathbf{Q}$ and $\mathbf{K}$, the attention step proceeds for each head $i\in\{1,\dots,h\}$ using (3.4). Finally, concatenating these $h$ heads and then projecting them back using an output weight matrix gives the final attention result, as shown in (3.5).
SwiGLU Feed-Forward Network. Following Shazeer (2020); Touvron et al. (2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN):
$\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3},$
where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters. Note that other activation functions can also be used.
Overall T6 Block Structure. Putting everything together, one T6 block consists of:
$\mathbf{x}$ $\;\leftarrow\;\mathbf{x}+\operatorname{TPA}\bigl{(}\operatorname{RMSNorm}(\mathbf{x})\bigr{)},$
$\mathbf{x}$ $\;\leftarrow\;\mathbf{x}+\operatorname{SwiGLU-FFN}\bigl{(}\operatorname{RMSNorm}(\mathbf{x})\bigr{)}.$
We place norm layers (e.g., RMSNorm) before each sub-layer. Stacking $L$ such blocks yields a T6 model architecture with $L$ layers.
4 Experiments
4.1 Language Modeling Tasks
All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy, 2022), using the FineWeb-Edu 100B dataset (Lozhkov et al., 2024). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024a), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al., 2017) with Multi-Query Attention (MQA; Shazeer, 2019), Grouped Query Attention (GQA; Ainslie et al., 2023), or Multi-head Latent Attention (MLA; Liu et al., 2024a). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), and large (773M). Details on architecture hyperparameters and training hardware appear in Appendix B.1.
Training Setup. We follow the nanoGPT training configuration. In particular, we use the AdamW (Loshchilov, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$. For the small, medium, and large models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, and $2\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $3\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
Training & Validation Curves. Figures 2 and 3 compare training and validation loss curves for the large (773M) and medium (353M) models on FineWeb-Edu-100B. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses. For instance, in Figure 2(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher losses.
Validation Perplexity. Figure 4 shows the validation perplexities of the medium- and large-scale models. Mirroring the loss curves, TPA and TPA-KVonly steadily outperform MHA, MQA, GQA, and MLA over the course of training. By the end of pretraining (around $49$B tokens), TPA-based approaches achieve the lowest perplexities in most configurations.
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), using the lm-evaluation-harness codebase (Gao et al., 2024). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Tables 89 in the appendix present results for small models; Tables 23 for medium models; Tables 45 for large models;
For the medium-size (353M) models (Tables 23), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 48.96%. When given two-shot prompts, TPA again leads with 53.12% average accuracy. A similar trend appears for the large-size (773M) models (Tables 45), where TPA-KVonly attains the highest average (53.52% zero-shot, 55.33% two-shot), closely followed by full TPA.
Our experiments confirm that TPA consistently matches or exceeds the performance of established attention mechanisms (MHA, MQA, GQA, MLA) across medium and large model scales. The fully factorized TPA excels on mid-scale models, while TPA-KVonly can rival or surpass it at larger scales. In both cases, factorizing the attention activations shrinks autoregressive KV cache requirements by up to 5×5$\times$5 ×10×10$\times$10 ×, thus enabling much longer context windows under fixed memory budgets. In summary, tensor product attention provides a flexible, memory-efficient alternative to standard multi-head attention, advancing the scalability of modern language models.
(a) Training Loss
(b) Validation Loss
Figure 3: The training loss and validation loss of medium-size (353M) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
(a) Validation Perplexity of Medium Models
(b) Validation Perplexity of Large Models
Figure 4: The validation perplexity of medium-size (353M) models and large-size (773M) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
Table 2: The evaluation results of medium models with different attention mechanisms pretrained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
MHA 56.52 29.27 58.84 44.06 35.00 68.44 51.07 25.35 76.40 49.44
MQA 55.68 28.24 60.86 44.17 35.20 68.66 52.72 25.14 72.90 49.29
GQA 54.88 29.61 56.36 43.77 35.20 68.82 52.57 25.41 74.80 49.05
MLA 55.30 29.27 58.96 41.92 35.40 67.25 51.78 25.20 75.60 48.96
TPA-KVonly 57.11 30.03 61.25 44.83 34.60 69.04 54.54 23.35 74.60 49.93
TPA 59.30 31.91 60.98 45.57 34.60 69.48 53.91 24.93 77.20 50.88
Table 3: The evaluation results of medium models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
MHA 64.44 32.85 59.05 44.18 33.20 68.72 50.12 26.01 87.40 49.44
MQA 64.27 32.94 57.71 44.36 31.80 68.01 51.70 25.99 86.00 49.29
GQA 61.70 32.17 52.81 43.99 33.80 68.50 53.35 24.44 86.40 50.80
MLA 62.75 30.80 59.17 42.02 34.80 67.08 52.41 26.11 84.80 51.10
TPA-KVonly 65.99 33.70 57.49 44.47 34.20 69.53 53.28 24.23 86.50 49.93
TPA 66.54 34.47 58.96 45.35 33.00 69.21 53.99 24.51 91.30 53.04
Table 4: The evaluation results of large models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
MHA 59.93 33.62 61.93 50.63 36.00 71.06 55.41 22.87 81.20 52.52
MQA 60.73 33.62 57.34 50.09 37.00 69.97 55.49 25.30 79.60 52.13
GQA 61.66 34.30 58.72 49.85 38.40 71.16 53.75 25.23 77.60 52.30
MLA 60.73 31.57 61.74 48.96 35.40 69.59 55.09 26.39 76.70 51.80
TPA-KVonly 63.26 34.13 61.96 50.66 37.20 72.09 55.25 26.06 81.10 53.52
TPA 63.22 35.58 60.03 51.26 36.80 71.44 55.56 24.77 79.60 53.10
Table 5: The evaluation results of large models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
MHA 67.85 36.35 59.82 50.22 35.00 70.67 53.35 23.92 91.10 54.25
MQA 68.86 36.09 53.79 50.50 37.00 70.89 54.70 25.01 88.00 53.87
GQA 69.15 36.09 58.84 50.29 36.20 70.73 54.22 26.08 90.00 54.62
MLA 68.56 35.41 60.12 49.18 38.00 69.21 55.25 25.29 88.20 54.36
TPA-KVonly 71.34 37.71 59.76 51.10 36.00 71.49 54.62 25.83 90.10 55.33
TPA 70.41 37.71 60.06 51.30 34.00 71.06 54.54 25.79 90.30 55.02
5 Related Work
Transformers and Attention. As a sequence-to-sequence architecture Transformer (Vaswani et al., 2017) introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns (Child et al., 2019; Shi et al., 2023; Han et al., 2024; Liang et al., 2024a; Li et al., 2024; Liang et al., 2024b), kernel-based projections (Choromanski et al., 2021), and linearized transformers (Tsai et al., 2019; Katharopoulos et al., 2020; Schlag et al., 2021; Zhang et al., 2023b; Sun et al., 2023; Zhang et al., 2024). To decrease memory usage and circumvent the limitation of memory bandwidth in training, Shazeer (2019) proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle with the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) (Ainslie et al., 2023) divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 (Liu et al., 2024a) applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. In comparison to the approaches above, TPA applied a low-rank tensor product to compute the queries, keys, and values where the cached representations of keys and values are much smaller than those in MHA, achieving better reduction on memory assumption of KV cache in inference time.
Low-Rank Factorizations. Low-rank approximations have been applied to compress model parameters and reduce complexity including LoRA (Hu et al., 2022), which factorizes weight updates during fine-tuning, and its derivatives for other training scenarios such as efficient pretraining (ReLoRA (Lialin et al., 2023), MoRA (Jiang et al., 2024)), long-context training (LongLoRA (Chen et al., 2024), SinkLoRA (Zhang, 2024)), as well as continual training (InfLoRA (Liang & Li, 2024), GS-LoRA (Zhao et al., 2024), I-LoRA (Ren et al., 2024)). These approaches typically produce static low-rank expansions that do not explicitly depend on the input context. And Malladi et al. (2023); Zeng & Lee (2024) provided theoretical proof of the expressiveness of low-rank approximation. For the initialization of factorization matrices, OLoRA (Büyükakyüz, 2024) applied QR-decomposition of pretrained weight to achieve better performance of language models while LoLDU (Shi et al., 2024) used LDU-decomposition to accelerate training of LoRA. Moreover, AdaLoRA (Zhang et al., 2023a) utilized Singular Value Decomposition (SVD) of the pretrained weight and introduced importance score for each parameter as a measurement to achieve dynamic adjustment of rank. TPA, by contrast, constructs Q, K, and V as contextually factorized tensors, enabling dynamic adaptation.
KV Cache Optimization. During the inference time of Transformers, key and value tensors of the previous tokens are repeatedly computed due to their auto-regressive nature. To enhance efficiency, firstly proposed by Ott et al. (2019), these tensors can be cached in memory for future decoding, referred to as the KV cache. However, the KV cache requires additional memory usage and may add to more latencies due to the bandwidth limitation (Adnan et al., 2024). Therefore, previous studies have explored diverse approaches to mitigate these issues, including KV cache eviction to discard less significant tokens (Zhang et al., 2023c; Xiao et al., 2024; Cai et al., 2024; Adnan et al., 2024), dynamic sparse attention among selected keys and values (Ribar et al., 2024; Tang et al., 2024; Singhania et al., 2024), KV cache offloading to CPU (He & Zhai, 2024; Lee et al., 2024; Sun et al., 2024), as well as quantization of KV cache (Xiao et al., 2023; Liu et al., 2024c; Hooper et al., 2024). Besides these methods, it is also effective to reduce the amount of KV cache for each token, by approaches such as reducing the number of KV heads (Ren et al., 2024; Ainslie et al., 2023), cross-layer KV re-usage (Xiao et al., 2019; Mu et al., 2024; Wu et al., 2024), and low-rank KV representation (Saxena et al., 2024). Different from the methods above, TPA reduces the size of the KV cache by using tensor-decomposed keys and values.
6 Conclusion
We introduced Tensor Product Attention (TPA), which factorizes query, key, and value matrices into rank-$R$ tensor products dependent on the token’s hidden state. Storing only the factorized key/value components during autoregressive decoding substantially decreases the kv memory size with improved performance compared with MHA, MQA, GQA, and MLA. The approach is fully compatible with RoPE (and can store pre-rotated keys). Variants of TPA include factorizing only the key/value or sharing basis vectors across tokens. Overall, TPA offers a powerful mechanism for compressing KV storage while improving the model performance, thereby enabling longer sequence contexts under constrained memory.
Appendix A Proofs of Theorems
Proof of Theorem 1.
Proof.
Because RoPE is a linear orthogonal transform, we can write
$\widetilde{\mathbf{Q}}_{t}=\mathbf{Q}_{t}\,\mathbf{T}_{t}=\frac{1}{R_{Q}}\bigl{(}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}\,\mathbf{T}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\,\mathbf{T}_{t}\bigr{)},$
where $\mathbf{T}_{t}$ is the block-diagonal matrix encoding RoPE. This allows us to define
$\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\mathbf{B}_{Q}(\mathbf{x}_{t})\,\mathbf{T}_{t},$
thereby obtaining
$\operatorname{RoPE}(\mathbf{Q}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}).$
Similarly, for the key tensor $\mathbf{K}_{s}$, we have
$\widetilde{\mathbf{K}}_{s}=\mathbf{K}_{s}\,\mathbf{T}_{s}=\frac{1}{R_{K}}\bigl{(}\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\,\mathbf{B}_{K}(\mathbf{x}_{s})\bigr{)}\,\mathbf{T}_{s}=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{s})\,\mathbf{T}_{s}\bigr{)},$
which defines
$\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{s})=\mathbf{B}_{K}(\mathbf{x}_{s})\,\mathbf{T}_{s},$
and thus
$\operatorname{RoPE}(\mathbf{K}_{s})=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{s}).$
Now, consider the product of the rotated queries and keys:
$\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top}$ $=\frac{1}{R_{Q}R_{K}}\left(\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})\right)\left(\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{s})\right)^{\top}$
$=\frac{1}{R_{Q}R_{K}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{s})^{\top}\mathbf{A}_{K}(\mathbf{x}_{s}),$
Since $\mathbf{T}_{t}$ and $\mathbf{T}_{s}$ encode positional rotations, the product $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}$ corresponds to a relative rotation $\mathbf{T}_{t-s}$. Therefore, we can express the above as
$\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top}$ $=\frac{1}{R_{Q}R_{K}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\left(\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{T}_{t}\mathbf{T}_{s}^{\top}\mathbf{B}_{K}(\mathbf{x}_{s})^{\top}\right)\mathbf{A}_{K}(\mathbf{x}_{s})$
$=\frac{1}{R_{Q}R_{K}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\left(\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{T}_{t-s}\mathbf{B}_{K}(\mathbf{x}_{s})^{\top}\right)\mathbf{A}_{K}(\mathbf{x}_{s})$
$=\frac{1}{R_{Q}R_{K}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\left(\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{T}_{t-s}\right)\left(\mathbf{B}_{K}(\mathbf{x}_{s})^{\top}\mathbf{A}_{K}(\mathbf{x}_{s})\right)$
$=\left(\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{T}_{t-s}\right)\left(\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\mathbf{B}_{K}(\mathbf{x}_{s})\right)^{\top},$
This shows that
$\operatorname{RoPE}_{t-s}(\mathbf{Q}_{t})\mathbf{K}_{s}^{\top}=\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top},$
Focusing on individual heads $i$, the above matrix equality implies:
$\operatorname{RoPE}_{t-s}(\mathbf{q}_{t,i})^{\top}\mathbf{k}_{s,i}=\widetilde{\mathbf{q}}_{t,i}^{\top}\widetilde{\mathbf{k}}_{s,i},$
where
$\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}},\quad\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
This equality confirms that the relative positional encoding between queries and keys is preserved under TPA’s factorization and RoPE’s rotation. Thus, TPA maintains compatibility with RoPE. This completes the proof of Theorem 1. ∎
Appendix B More on Experiments
B.1 Experimental Settings
We list the main architecture hyper-parameters and training devices in Table 6. We fix $d_{h}=64$ for all the models. Moreover, we fix the number of KV heads with 2 for GQA models; $d_{h}^{R}=32$ for MLA models; and $R_{k}=R_{v}=2$, $R_{q}=6$ for TPA and TPA-KV only models. Other hyper-parameters are listed in Table 7.
Table 6: The architecture hyper-parameters and training devices of models. Abbreviations: BS. = Batch Size, GAS. = Gradient Accumulation Steps.
Model Size #Param Devices Micro BS. GAS. #Layer $d_{\text{model}}$
Small 124M 4$\times$ A100 GPUs 24 5 12 768
Medium 353M 8$\times$ A100 GPUs 20 3 24 1024
Large 772M 8$\times$ A100 GPUs 15 4 36 1280
Table 7: The architecture hyper-parameters for different models.
Model Size Small Medium Large
$h$ (MHA) 12 16 20
$h$ (MQA) 23 31 39
$h$ (GQA) 22 30 38
$n_{h}$ (MLA) 12 23 34
$h$ (TPA-KVonly) 22 29 37
$h$ (TPA) 34 47 61
$d_{c}$ (MLA) 256 512 512
$d_{c}^{\prime}$ (MLA) 512 1024 1024
B.2 Additional Experimental Results
We display the evaluation results for small-size (124M) models in Tables 8-9.
Table 8: The evaluation results of small models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
MHA 50.63 26.96 59.39 36.18 32.00 64.96 51.85 23.40 70.30 46.19
MQA 49.62 25.34 55.72 35.94 31.40 64.85 51.30 23.37 68.70 45.14
GQA 48.70 25.68 56.15 35.58 31.40 64.91 51.62 23.12 68.20 45.04
MLA 49.66 26.45 61.22 33.94 32.40 62.73 50.43 23.29 71.50 45.74
TPA-KVonly 51.05 26.54 57.25 36.77 32.60 65.02 50.91 23.64 69.70 45.94
TPA 51.26 27.39 57.00 36.68 32.80 64.47 49.72 24.61 72.00 46.21
Table 9: The evaluation results of small models with different attention mechanisms on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
MHA 57.66 28.24 57.28 36.43 29.60 64.09 51.14 26.57 82.00 48.11
MQA 53.79 26.35 44.95 34.18 28.80 62.79 52.01 25.91 78.10 45.21
GQA 55.01 25.94 55.72 35.68 31.80 65.29 51.93 25.27 77.80 47.16
MLA 52.78 26.19 57.25 33.19 29.60 63.98 50.43 24.90 76.00 46.04
TPA-KVonly 54.25 27.90 57.06 36.36 31.80 64.31 53.59 26.18 79.20 47.85
TPA 57.53 28.07 56.33 36.49 31.80 64.36 51.14 25.92 79.70 47.93
B.3 Ablation Studies on Learning Rates
We implement a set of parallel experiments for medium models with learning rate $6\times 10^{-4}$, and the curves for training loss, validation loss and validation perplexity are displayed in Figure 5. We also show the performance of these models on the benchmarks described in Section 4 in Tables 10-11. The results show that TPA and TPA-KVonly models can also outperform other types of attention with different learning rates.
(a) Training Loss
(b) Validation Loss
(c) Validation Perplexity
Figure 5: The training loss, validation loss and validation perplexity of medium-size (353M) models with learning rate $6\times 10^{-4}$ and different attention mechanisms on the FineWeb-Edu 100B dataset.
Table 10: The evaluation results of medium models (learning rate=$6\times 10^{-4}$) with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
MHA 59.51 29.52 59.60 45.68 34.20 68.82 53.43 23.33 76.90 50.11
MQA 57.62 31.91 59.45 45.69 35.40 69.31 53.51 26.47 74.60 50.44
GQA 28.67 31.48 58.29 45.45 35.20 68.50 54.46 24.58 76.50 47.01
MLA 57.49 29.44 59.97 44.09 25.77 68.66 53.04 25.77 76.40 48.96
TPA-KVonly 58.01 30.12 58.01 45.95 35.60 69.10 53.12 25.39 75.10 50.04
TPA 58.38 31.57 59.39 46.83 37.00 70.02 54.06 25.52 79.90 51.41
Table 11: The evaluation results of medium models (learning rate $6\times 10^{-4}$) with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
MHA 64.73 32.42 58.29 45.89 34.20 68.50 53.20 25.86 88.00 52.34
MQA 64.98 33.62 55.02 45.81 34.00 69.59 53.43 24.30 85.20 51.77
GQA 65.24 33.19 56.54 45.41 34.80 69.04 55.72 24.73 87.90 52.51
MLA 63.80 31.06 58.50 44.19 35.40 68.44 51.62 25.22 88.50 51.86
TPA-KVonly 64.69 32.34 59.48 46.23 35.40 70.08 54.06 25.64 86.30 52.69
TPA 67.97 34.56 57.22 46.87 34.60 69.91 52.01 25.07 89.90 53.12

### v2

@@
 Figure 1: Tensor Product Attention (TPA) in the Tensor ProducT ATTenTion Transformer (T6).
-Different from multi-head attention, in each layer, firstly the hidden state goes through different linear layers to get the latent factor matrices $\mathbf{A}$’s and $\mathbf{B}$’s for query, key, and value.
+Different from multi-head attention, in each layer, firstly the hidden state goes through different linear layers to get the latent factor matrices $\mathbf{A}_{(\cdot)}$’s and $\mathbf{B}_{(\cdot)}$’s for query, key, and value.
 We additionally apply RoPE to $\mathbf{B}_{Q}$ and $\mathbf{B}_{K}$ for query and key.
@@
 Meanwhile, low-rank weight factorization methods such as LoRA (Hu et al., 2022) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates runtime.
-The recently introduced Multi-head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but needs additional position-encoded parameters per head due to incompatibility with Rotary Position Embedding (RoPE) efficiently (Su et al., 2024b).
+The recently introduced Multi-head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but needs additional position-encoded parameters per head due to incompatibility with Rotary Position Embedding (RoPE) efficiently (Su et al., 2024).
 In order to overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), as illustrated in Figure 1, a novel architecture that uses higher-order tensors to factorize queries (Q), keys (K), and values (V) during attention computation.
@@
 In practice, TPA can reduce the memory overhead by an order of magnitude compared to standard multi-head attention (MHA) with lower pretraining validation loss (perplexity) and improved downstream performance.
-A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024b), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023) and Gemma (Team et al., 2024).
+A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023) and Gemma (Team et al., 2024).
 Our primary contributions are summarized as follows:
@@
 In addition, we unify existing attention mechanisms by revealing that MHA, MQA, and GQA all arise naturally as non-contextual variants of TPA.
-  • We propose Tensor ProducT ATTenTion Transformer (T6), a new TPA-based model architecture for sequence modeling.
+  • We introduce the Tensor ProducT ATTenTion Transformer (T6), a new TPA-based model architecture for sequence modeling.
 On language modeling experiments, T6 consistently improves validation perplexity and downstream evaluation performance with reduced KV cache size.
-  • We show TPA integrates seamlessly with RoPE (Su et al., 2024b), facilitating easy adoption in popular foundation model architectures such as LLaMA and Gemma.
-(a) Training Loss
-(b) Validation Loss
-Figure 2: Training loss and validation loss of pretraining large-size (773M) models with different attention mechanisms on the FineWeb-Edu-100B dataset.
+  • We show TPA integrates seamlessly with RoPE (Su et al., 2024), facilitating easy adoption in popular foundation model architectures such as LLaMA and Gemma.
 2 Background
-In this section, we review several classical forms of attention: Scaled Dot-Product Attention, Multi-Head Attention (MHA) (Vaswani et al., 2017), Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped Query Attention (GQA) (Ainslie et al., 2023), as well as Rotary Position Embedding (RoPE, Su et al.
-(2024b)).
-We also introduce a recent method called Multi-head Latent Attention (MLA) used in DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b).
+In this section, we review two classical forms of attention: Scaled Dot-Product Attention, and Multi-Head Attention (MHA) (Vaswani et al., 2017).
+More types of attention are detailed in the Appendix E, including Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped Query Attention (GQA) (Ainslie et al., 2023), as well as a recent method called Multi-head Latent Attention (MLA) used in DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b).
+We also introduce Rotary Position Embedding (RoPE, Su et al.
+(2024)), which is commonly used in recent works of large language models.
 Notations.
@@
 For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
-$\mathbf{Q}_{t,i}=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},\quad\mathbf{K}_{t,i}=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},\quad\mathbf{V}_{t,i}=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
+$\mathbf{Q}_{t,i}$ $=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
+$\mathbf{K}_{t,i}$ $=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
+$\mathbf{V}_{t,i}$ $=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
 $\textbf{head}_{i}$ $=\operatorname{Attention}\Bigl{(}\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\Bigr{)},$
@@
 MHA can capture a rich set of dependencies while each head focuses on different subspaces.
-2.3 Multi-Query Attention (MQA)
-Multi-Query Attention (MQA) (Shazeer, 2019) significantly reduces memory usage by sharing keys and values across heads, while still preserving unique query projections.
-For a sequence of embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$,
-$\mathbf{Q}_{i}=\mathbf{X}\bm{W}^{Q}_{i},\quad\mathbf{K}_{\text{shared}}=\mathbf{X}\bm{W}^{K}_{\text{shared}},\quad\mathbf{V}_{\text{shared}}=\mathbf{X}\bm{W}^{V}_{\text{shared}}.$
-Hence, each head $i$ only has a distinct query $\mathbf{Q}_{i}\in\mathbb{R}^{T\times d_{k}}$, but shares the same key $\mathbf{K}_{\text{shared}}\in\mathbb{R}^{T\times d_{k}}$ and value $\mathbf{V}_{\text{shared}}\in\mathbb{R}^{T\times d_{k}}$.
-In practice, this means:
-$\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},\quad\bm{W}^{K}_{\text{shared}},\bm{W}^{V}_{\text{shared}}\;\in\;\mathbb{R}^{\,d_{\text{model}}\times d_{k}}.$
-The resulting MQA operation is:
-$\operatorname{MQA}(\mathbf{X})=\operatorname{Concat}\Bigl{(}\textbf{head}_{1},\dots,\textbf{head}_{h}\Bigr{)}\,\bm{W}^{O},$
-where
-$\textbf{head}_{i}=\operatorname{Attention}\bigl{(}\mathbf{Q}_{i},\mathbf{K}_{\text{shared}},\mathbf{V}_{\text{shared}}\bigr{)}.$
-By sharing these key and value projections, MQA cuts down on memory usage (especially for the key-value cache in autoregressive inference) but loses some expressivity since all heads must rely on the same key/value representations.
-2.4 Grouped Query Attention (GQA)
-Grouped Query Attention (GQA) (Ainslie et al., 2023) generalizes MHA and MQA by grouping heads.
-Specifically, we partition the $h$ total heads into $G$ groups.
-Each group has a single set of keys and values, but each individual head within that group still retains its own query projection.
-Formally, if $g(i)$ maps a head $i\in[h]$ to its group index $g\in[G]$, then:
-$\mathbf{K}_{g(i)}=\mathbf{X}\,\bm{W}^{K}_{g(i)},\quad\mathbf{V}_{g(i)}=\mathbf{X}\,\bm{W}^{V}_{g(i)},\quad\mathbf{Q}_{i}=\mathbf{X}\,\bm{W}^{Q}_{i},$
-and
-$\text{head}_{i}=\operatorname{Attention}\Bigl{(}\mathbf{Q}_{i},\mathbf{K}_{g(i)},\mathbf{V}_{g(i)}\Bigr{)}.$
-Again, $\bm{W}^{K}_{g},\bm{W}^{V}_{g}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ for each group $g$, and $\bm{W}^{Q}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$ for each head $i$.
-The complete output is again a concatenation of all heads:
-$\operatorname{GQA}(\mathbf{X})=\operatorname{Concat}\Bigl{(}\text{head}_{1},\ldots,\text{head}_{h}\Bigr{)}\bm{W}^{O}.$
-By adjusting $G$ between $1$ and $h$, GQA can interpolate between sharing all key/value projections across heads (i.e., MQA) and having one set of projections per head (i.e., MHA).
 2.3 Rotary Position Embedding (RoPE)
-Many recent LLMs use rotary position embedding (RoPE; Su et al., 2024b) to encode positional information in the query/key vectors.
+Many recent LLMs use rotary position embedding (RoPE; Su et al., 2024) to encode positional information in the query/key vectors.
 Specifically, let $\mathrm{RoPE}_{t}$ denote the rotation operator $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ corresponding to the $t$-th position.
@@
 which ensures that relative positions $(t-s)$ are preserved, thereby providing a form of translation invariance in the rotary position embedding.
-2.6 Multi-head Latent Attention (MLA)
-Below, we briefly outline the Multi-head Latent Attention (MLA) approach used by DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b).
-MLA introduces a low-rank compression of the keys and values to reduce the Key-Value (KV) caching cost at inference.
-$\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},\quad(\bm{W}^{DKV}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}}),$
-$\operatorname{Concat}\bigl{(}\mathbf{K}_{1}^{C},\mathbf{K}_{2}^{C},\ldots,\mathbf{K}_{h}^{C}\bigr{)}=\mathbf{K}^{C}$ $=\mathbf{C}^{KV}\bm{W}^{UK},\quad(\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times d_{h}h}),$
-$\mathbf{K}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{X}\bm{W}^{KR}\bigr{)},\quad(\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}}),$
-$\mathbf{K}_{i}$ $=\operatorname{Concat}\bigl{(}\mathbf{K}_{i}^{C},\mathbf{K}^{R}\bigr{)},$
-$\operatorname{Concat}\bigl{(}\mathbf{V}_{1}^{C},\mathbf{V}_{2}^{C},\ldots,\mathbf{V}_{h}^{C}\bigr{)}=\mathbf{V}^{C}$ $=\mathbf{C}^{KV}\bm{W}^{UV},\quad(\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times d_{h}h}),$
-where $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the compressed KV latent (with $d_{c}\ll d_{h}h$), and $\operatorname{RoPE}(\cdot)$ represents the RoPE transform applied to the separate key embeddings $\mathbf{K}^{R}$ of dimension $d_{h}^{R}$.
-Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ need to be cached, reducing KV memory usage while largely preserving performance compared to standard MHA (Vaswani et al., 2017).
-MLA also compresses the queries, lowering their training-time memory footprint:
-$\mathbf{C}^{Q}$ $=\mathbf{X}\bm{W}^{DQ},\quad(\bm{W}^{DQ}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}^{\prime}}),$
-$\operatorname{Concat}\bigl{(}\mathbf{Q}_{1}^{C},\mathbf{Q}_{2}^{C},\,\ldots,\mathbf{Q}_{h}^{C}\bigr{)}=\mathbf{Q}^{C}$ $=\mathbf{C}^{Q}\bm{W}^{UQ},\quad(\bm{W}^{UQ}\in\mathbb{R}^{\,d_{c}^{\prime}\times d_{h}h}),$
-$\operatorname{Concat}\bigl{(}\mathbf{Q}_{1}^{R},\,\mathbf{Q}_{2}^{R},\,\ldots,\,\mathbf{Q}_{h}^{R}\bigr{)}=\mathbf{Q}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{C}^{Q}\bm{W}^{QR}\bigr{)},\quad(\bm{W}^{QR}\in\mathbb{R}^{d_{c}^{\prime}\times d_{h}^{R}h}),$
-$\mathbf{Q}$ $=\operatorname{Concat}\bigl{(}\mathbf{Q}^{C},\mathbf{Q}^{R}\bigr{)}.$
-Here, $\mathbf{C}^{Q}\in\mathbb{R}^{T\times d_{c}^{\prime}}$ (with $d_{c}^{\prime}\ll d_{h}h$) is the compressed query latent.
-As above, each $\bm{W}^{DQ}$, $\bm{W}^{UQ}$, and $\bm{W}^{QR}$ connects these lower-dimensional query latents back to $h$ heads of dimension $d_{h}+d_{h}^{R}$.
-Given compressed queries, keys, and values, the final attention output for the $t$-th token is:
-$\mathbf{O}_{i}$ $=\operatorname{Softmax}\Bigl{(}\tfrac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\top}}{\sqrt{d_{h}+d_{h}^{R}}}\Bigr{)}\;\mathbf{V}_{i}^{C},$
-$\mathbf{U}$ $=\operatorname{Concat}\bigl{(}\mathbf{O}_{1},\mathbf{O}_{2},\ldots,\mathbf{O}_{h}\bigr{)}\bm{W}^{O},$
-where $\bm{W}^{O}\in\mathbb{R}^{(d_{h}h)\times\,d_{\text{model}}}$ is the output projection.
-In inference time, $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ can be cached to accelerate decoding.
-In detail, when RoPE is ignored, the inner product $\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ (where $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{d}$) of the $i$-th head between $t$-th and $s$-th tokens can be calculated using the hidden state $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ for $t$-th token and the cached latent state $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ for $s$-th token:
-$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (2.2)
-where $\bm{W}_{i}^{(\cdot)}$ is the $i$-th head of the original weight, and $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ can be computed previously for faster decoding.
-However, this process fails when RoPE is considered according to Su (2024).
-Since RoPE can be considered as multiplication with a block-diagonal matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ (see Section 2.5), with the property (2.1) that $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$, then
-[…]
-Different from (2.2), acceleration by pre-computing $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t-s}}(\bm{W}^{UK}_{i})^{\top}]$ fails since it varies for different $(t,s)$ position pairs.
-Therefore, MLA adds the additional $\mathbf{k}_{t}^{R}$ part with a relatively smaller size for RoPE compatibility.
-In Section 3.2, we will show that TPA addresses the issue of RoPE-incompatibility by applying tensor product.
 3 Tensor Product Attention
@@
 Specifically, for each token $t$, with a small abuse of notation, we define:
-$\mathbf{Q}_{t}$ $=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}},$ (3.1)
-$\mathbf{K}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{K}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}},$ (3.2)
-$\mathbf{V}_{t}$ $=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}.$ (3.3)
+$\mathbf{Q}_{t}$ $=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),$ (3.1)
+$\mathbf{K}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$ (3.2)
+$\mathbf{V}_{t}$ $=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$ (3.3)
+where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$.
 Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ adds up to form the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$.
@@
 One often merges the rank index into a single output dimension. For instance, for queries:
-$\mathbf{a}^{Q}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot h},\quad\mathbf{b}^{Q}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot d_{h}},$
+$\mathbf{a}^{Q}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot h},~{}\mathbf{b}^{Q}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot d_{h}},$
 which are then reshaped into $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{q}\times d_{h}}$.
@@
 yielding a pre-rotated key representation:
-$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(s)}(\mathbf{x}_{t})\bigr{)}=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
+$\widetilde{\mathbf{K}}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(s)}(\mathbf{x}_{t})\bigr{)}$
+$=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
 Thus, each $\mathbf{K}_{t}$ is already rotated before caching, removing the need for explicit rotation at the decoding time and accelerating autoregressive inference.
@@
 where $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times d_{h}}$. Then we have:
-$\operatorname{RoPE}({\mathbf{Q}}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),\qquad\text{where }\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}.$ (3.7)
+$\operatorname{RoPE}({\mathbf{Q}}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),$ (3.7)
+where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}$.
 In addition, assume $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ are factorized by TPA and then rotated by $\operatorname{RoPE}_{t},\operatorname{RoPE}_{s}$.
@@
 where $\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}}$ is the $i$-th query head of $t$-th token, and $\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}$ is the $j$-th key head of $s$-th token, and
-$\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}},\quad\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
+$\widetilde{\mathbf{q}}_{t,i}$ $=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}}$
+$\quad\widetilde{\mathbf{k}}_{s,i}$ $=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
 Theorem 1 indicates that TPA does not break RoPE’s relative translational property.
-We prove Theorem 1 in Appendix A.
+We prove Theorem 1 in Appendix C.1.
 In short, $\operatorname{RoPE}_{t}$ acts as a block-diagonal orthogonal transform (i.e., a matrix $\mathbf{T}_{t}$) on $\mathbf{B}_{Q}(\mathbf{x}_{t})$.
@@
 $\frac{\,(R_{K}+R_{V})\,(h+d_{h})\,}{2\,h\,d_{h}\,}.$
-For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll d_{h}$ (e.g., $1$ or $2$) often yields 10×10$\times$10 × or more reduction.
+For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll h$ (e.g., rank $1$ or $2$) often yields 10×10$\times$10 × or more reduction.
 Table 1: Comparison of different attention mechanisms.
 Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively.
-Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Section 3.5.
+Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Section F.
 For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@
     Non-contextual head factors. Define
-[…]
-    so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
+    $\mathbf{a}^{Q}_{i}=R_{Q}\mathbf{e}_{i}\in\mathbb{R}^{h},$ (3.8)
+    where $\mathbf{e}_{i}\in\mathbb{R}^{h}$ is the $i$-th standard basis vector, so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
   2. (b)
@@
 MQA and GQA as Non-Contextual TPA.
-Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped Query Attention (GQA) (Ainslie et al., 2023) also emerge naturally from TPA by restricting the head-dimension factors to be non-contextual and low-rank:
+Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped Query Attention (GQA) (Ainslie et al., 2023) [^6] The original definitions of MQA and GQA are presented in Appendix E.1 and E.2, respectively.
+also emerge naturally from TPA by restricting the head-dimension factors to be non-contextual and low-rank:
   • MQA as Rank-1 TPA. In MQA, all heads share a single set of keys/values, corresponding to $R_{K}=R_{V}=1$ along the head dimension. Concretely,
-    $\mathbf{K}_{t}=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
+    $\mathbf{K}_{t}$ $=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),$
+    $\mathbf{V}_{t}$ $=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
     forces every head to use the same $\mathbf{K}_{t},\mathbf{V}_{t}$. Each head retains a distinct query projection, matching the MQA design.
@@
 Hence, by constraining TPA’s head-dimension factors to be constant masks (one for MQA; multiple for GQA), these popular variants are recovered as special cases.
-3.5 Other Variants of TPA
-TPA with Non-contextual B.
-Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to $\mathbf{x}_{t}$.
-For keys:
-$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{K}_{r},$
-and similarly for keys/values. This arrangement is effective if the token-dimension structure remains mostly uniform across the sequence, while the head-dimension factors capture context.
-TPA KV Only. One can preserve a standard query mapping,
-$\mathbf{Q}_{t}=\bm{W}^{Q}\,\mathbf{x}_{t}\in\mathbb{R}^{h\times d_{h}},$
-and factorize only the keys and values. This leaves the query projection as the original linear transformation while reducing memory usage via factorized KV caching.
-TPA KV with Shared $\mathbf{B}$. Another variant is to share the token-dimension factors of keys and values:
-$\mathbf{b}^{K}_{r}(\mathbf{x}_{t})=\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$
-lowering parameter counts and the KV cache footprint.
-While it constrains $\mathbf{K}$ and $\mathbf{V}$ to be formed from the same token basis, it can still perform well and provide additional memory savings.
-Nonlinear Head Factors.
-Rather than applying purely linear mappings to the head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}$, one may introduce element-wise nonlinearities such as $\sigma(\cdot)$ or $\operatorname{softmax}(\cdot)$.
-This effectively yields a Mixture of Heads Attention (MoH Attention), where each component becomes a learned mixture weight modulated by the nonlinearity.
-Discussion.
-These variants illustrate TPA’s versatility in balancing memory cost, computational overhead, and representation power.
-By choosing which dimensions (heads or tokens) remain contextual and adjusting ranks $(R_{Q},R_{K},R_{V})$, TPA unifies multiple existing attention mechanisms—such as MHA, MQA, and GQA—under one framework, while potentially reducing the KV cache size by an order of magnitude during autoregressive inference.
+3.5 Computational Cost.
+For a detailed analysis of the computational cost of TPA, please refer to Appendix A, which shows that the training and inference flops of TPA with optimized implementation (without materializing $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$) are smaller than MHA, GQA, and MLA.
+Specifically, when we set $R_{q}=6$, $R_{k}=R_{v}=2$ (our default setting), TPA is 10×10$\times$10 × or more faster on calculating $\mathbf{Q}\mathbf{K}^{\top}$ than MLA during inference (see Appendix A.8).
 3.6 Model Architectures
 We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention).
-Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version.
+Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version.
 The feed-forward network (FFN) adopts a SwiGLU layer, as in (Shazeer, 2020; Touvron et al., 2023).
@@
 Rotary Positional Embedding (RoPE).
-As discussed in Section 3.2, RoPE (Su et al., 2024b) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
+As discussed in Section 3.2, RoPE (Su et al., 2024) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
 Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.6) and Theorem 1.
@@
 Finally, concatenating these $h$ heads and then projecting them back using an output weight matrix gives the final attention result, as shown in (3.5).
-SwiGLU Feed-Forward Network. Following Shazeer (2020); Touvron et al. (2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN):
-$\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3},$
-where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters.
+SwiGLU Feed-Forward Network.
+Following Shazeer (2020); Touvron et al.
+(2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters.
 Note that other activation functions can also be used.
@@
 The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation.
-We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024a), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al., 2017) with Multi-Query Attention (MQA; Shazeer, 2019), Grouped Query Attention (GQA; Ainslie et al., 2023), or Multi-head Latent Attention (MLA; Liu et al., 2024a).
+We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al., 2017) with Multi-Query Attention (MQA; Shazeer, 2019), Grouped Query Attention (GQA; Ainslie et al., 2023), or Multi-head Latent Attention (MLA; Liu et al., 2024a).
 In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer.
-We train models at four scales: small (124M parameters), medium (353M), and large (773M).
-Details on architecture hyperparameters and training hardware appear in Appendix B.1.
+We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B).
+Details on architecture hyperparameters and training hardware are shown in Appendix G.1.
 Training Setup.
@@
 We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$.
-For the small, medium, and large models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, and $2\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $3\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
+For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
 Training & Validation Curves.
-Figures 2 and 3 compare training and validation loss curves for the large (773M) and medium (353M) models on FineWeb-Edu-100B.
-Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses.
+Figures 2 and 3 compare training and validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B.
+Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see F) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses.
 For instance, in Figure 3(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages.
@@
 Validation Perplexity.
-Figure 4 shows the validation perplexities of the medium- and large-scale models.
+Figure 4 (in the Appendix) shows the validation perplexities of the medium- and large-scale models.
 Mirroring the loss curves, TPA and TPA-KVonly steadily outperform MHA, MQA, GQA, and MLA over the course of training.
@@
 For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy.
-Tables 89 in the appendix present results for small models; Tables 23 for medium models; Tables 45 for large models;
-For the medium-size (353M) models (Tables 23), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 48.96%.
+Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3.
+Zero-shot evaluation of small and XL models are displayed in Tables 6 and 7 in the appendix.
+Moreover, we also present 2-shot evaluation results in Tables 8910 and 11 in the appendix.
+For the medium-size (353M) models (Tables 2 and 9), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%.
 When given two-shot prompts, TPA again leads with 53.12% average accuracy.
-A similar trend appears for the large-size (773M) models (Tables 45), where TPA-KVonly attains the highest average (53.52% zero-shot, 55.33% two-shot), closely followed by full TPA.
+A similar trend appears for the large-size (773M) models (Tables 3), where TPA-KVonly attains the highest average (53.52% zero-shot).
+And for the XL size (1.5B) models (Table 7), TPA-KVonly attains the highest average (55.03% zero-shot).
 Our experiments confirm that TPA consistently matches or exceeds the performance of established attention mechanisms (MHA, MQA, GQA, MLA) across medium and large model scales.
@@
 In summary, tensor product attention provides a flexible, memory-efficient alternative to standard multi-head attention, advancing the scalability of modern language models.
-(a) Training Loss
-(b) Validation Loss
-Figure 3: The training loss and validation loss of medium-size (353M) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
-(a) Validation Perplexity of Medium Models
-(b) Validation Perplexity of Large Models
-Figure 4: The validation perplexity of medium-size (353M) models and large-size (773M) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
-Table 2: The evaluation results of medium models with different attention mechanisms pretrained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
+(a) Medium models (353M)
+(b) Large models (773M)
+(c) XL models (1.5B)
+Figure 2: The training loss of medium-size (353M), large-size (773M) as well as XL-size (1.5B) models, with different attention mechanisms on the FineWeb-Edu 100B dataset.
+(a) Medium models (353M)
+(b) Large models (773M)
+(c) XL models (1.5B)
+Figure 3: The validation loss of medium-size (353M), large-size (773M) as well as XL-size (1.5B) models, with different attention mechanisms on the FineWeb-Edu 100B dataset.
+Table 2: The evaluation results of medium models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
 The best scores in each column are bolded.
@@
 Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
-MHA 56.52 29.27 58.84 44.06 35.00 68.44 51.07 25.35 76.40 49.44
-MQA 55.68 28.24 60.86 44.17 35.20 68.66 52.72 25.14 72.90 49.29
-GQA 54.88 29.61 56.36 43.77 35.20 68.82 52.57 25.41 74.80 49.05
-MLA 55.30 29.27 58.96 41.92 35.40 67.25 51.78 25.20 75.60 48.96
-TPA-KVonly 57.11 30.03 61.25 44.83 34.60 69.04 54.54 23.35 74.60 49.93
-TPA 59.30 31.91 60.98 45.57 34.60 69.48 53.91 24.93 77.20 50.88
-Table 3: The evaluation results of medium models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
-The best scores in each column are bolded.
-Abbreviations: HellaSw.
-= HellaSwag, W.G.
-= WinoGrande.
-Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
-MHA 64.44 32.85 59.05 44.18 33.20 68.72 50.12 26.01 87.40 49.44
-MQA 64.27 32.94 57.71 44.36 31.80 68.01 51.70 25.99 86.00 49.29
-GQA 61.70 32.17 52.81 43.99 33.80 68.50 53.35 24.44 86.40 50.80
-MLA 62.75 30.80 59.17 42.02 34.80 67.08 52.41 26.11 84.80 51.10
-TPA-KVonly 65.99 33.70 57.49 44.47 34.20 69.53 53.28 24.23 86.50 49.93
-TPA 66.54 34.47 58.96 45.35 33.00 69.21 53.99 24.51 91.30 53.04
+MHA 59.51 29.52 59.60 45.68 34.20 68.82 53.43 23.33 76.90 50.11
+MQA 57.62 31.91 59.45 45.69 35.40 69.31 53.51 26.47 74.60 50.44
+GQA 58.67 31.48 58.29 45.45 35.20 68.50 54.46 24.58 76.50 50.35
+MLA 56.65 29.52 57.83 46.05 34.60 69.42 52.80 24.62 79.70 50.13
+TPA-KVonly 58.01 30.12 58.01 45.95 35.60 69.10 53.12 25.39 75.10 50.04
+TPA (non-ctx-A) 58.96 31.48 59.76 45.07 34.80 69.21 53.59 25.42 76.40 50.52
+TPA 58.38 31.57 59.39 46.83 37.00 70.02 54.06 25.52 79.90 51.41
 Table 3: The evaluation results of large models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
@@
 GQA 61.66 34.30 58.72 49.85 38.40 71.16 53.75 25.23 77.60 52.30
-MLA 60.73 31.57 61.74 48.96 35.40 69.59 55.09 26.39 76.70 51.80
+MLA 63.55 32.85 60.95 51.72 38.80 70.51 55.01 24.55 81.90 53.32
 TPA-KVonly 63.26 34.13 61.96 50.66 37.20 72.09 55.25 26.06 81.10 53.52
 TPA 63.22 35.58 60.03 51.26 36.80 71.44 55.56 24.77 79.60 53.10
-Table 5: The evaluation results of large models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
-The best scores in each column are bolded.
-Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
-Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
-MHA 67.85 36.35 59.82 50.22 35.00 70.67 53.35 23.92 91.10 54.25
-MQA 68.86 36.09 53.79 50.50 37.00 70.89 54.70 25.01 88.00 53.87
-GQA 69.15 36.09 58.84 50.29 36.20 70.73 54.22 26.08 90.00 54.62
-MLA 68.56 35.41 60.12 49.18 38.00 69.21 55.25 25.29 88.20 54.36
-TPA-KVonly 71.34 37.71 59.76 51.10 36.00 71.49 54.62 25.83 90.10 55.33
-TPA 70.41 37.71 60.06 51.30 34.00 71.06 54.54 25.79 90.30 55.02
 5 Related Work
@@
 Recently, DeepSeek-V2 (Liu et al., 2024a) applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value.
-In comparison to the approaches above, TPA applied a low-rank tensor product to compute the queries, keys, and values where the cached representations of keys and values are much smaller than those in MHA, achieving better reduction on memory assumption of KV cache in inference time.
-Low-Rank Factorizations.
-Low-rank approximations have been applied to compress model parameters and reduce complexity including LoRA (Hu et al., 2022), which factorizes weight updates during fine-tuning, and its derivatives for other training scenarios such as efficient pretraining (ReLoRA (Lialin et al., 2023), MoRA (Jiang et al., 2024)), long-context training (LongLoRA (Chen et al., 2024), SinkLoRA (Zhang, 2024)), as well as continual training (InfLoRA (Liang & Li, 2024), GS-LoRA (Zhao et al., 2024), I-LoRA (Ren et al., 2024)).
-These approaches typically produce static low-rank expansions that do not explicitly depend on the input context.
-And Malladi et al.
-(2023); Zeng & Lee (2024) provided theoretical proof of the expressiveness of low-rank approximation.
-For the initialization of factorization matrices, OLoRA (Büyükakyüz, 2024) applied QR-decomposition of pretrained weight to achieve better performance of language models while LoLDU (Shi et al., 2024) used LDU-decomposition to accelerate training of LoRA.
-Moreover, AdaLoRA (Zhang et al., 2023a) utilized Singular Value Decomposition (SVD) of the pretrained weight and introduced importance score for each parameter as a measurement to achieve dynamic adjustment of rank.
-TPA, by contrast, constructs Q, K, and V as contextually factorized tensors, enabling dynamic adaptation.
+Concurrently, Hu et al.
+(2024) proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q.
+Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
 KV Cache Optimization.
@@
 Therefore, previous studies have explored diverse approaches to mitigate these issues, including KV cache eviction to discard less significant tokens (Zhang et al., 2023c; Xiao et al., 2024; Cai et al., 2024; Adnan et al., 2024), dynamic sparse attention among selected keys and values (Ribar et al., 2024; Tang et al., 2024; Singhania et al., 2024), KV cache offloading to CPU (He & Zhai, 2024; Lee et al., 2024; Sun et al., 2024), as well as quantization of KV cache (Xiao et al., 2023; Liu et al., 2024c; Hooper et al., 2024).
-Besides these methods, it is also effective to reduce the amount of KV cache for each token, by approaches such as reducing the number of KV heads (Ren et al., 2024; Ainslie et al., 2023), cross-layer KV re-usage (Xiao et al., 2019; Mu et al., 2024; Wu et al., 2024), and low-rank KV representation (Saxena et al., 2024).
-Different from the methods above, TPA reduces the size of the KV cache by using tensor-decomposed keys and values.
+Different from the methods above, TPA reduces the size of the KV cache by using tensor-decomposed KV.
 6 Conclusion
@@
 Overall, TPA offers a powerful mechanism for compressing KV storage while improving the model performance, thereby enabling longer sequence contexts under constrained memory.
-Appendix A Proofs of Theorems
-Proof of Theorem 1.
+Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
+We now explore whether it is possible to compute attention scores $\mathbf{Q}\,\mathbf{K}^{\top}$ of Tensor Product Attention (TPA) directly from their factorized forms, thereby reducing floating-point operations.
+A.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$
+Consider a single head $i$. Each query vector $\mathbf{Q}^{(i)}_{t}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{q}$):
+$\mathbf{Q}^{(i)}_{t}~{}=~{}\sum_{r=1}^{R_{q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
+and each key vector $\mathbf{K}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{k}$):
+$\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{s=1}^{R_{k}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
+Their dot-product for tokens $t,\tau$ is
+$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (A.1)
+A.2 Multi-Head Case
+For multi-head attention with $h$ heads, one repeats the factorization across all heads. The $\mathbf{b}_{q}^{(r)},\mathbf{b}_{k}^{(s)}$ vectors are shared across heads.
+A.3 Complexity Analysis
+We compare the cost of standard multi-head attention versus TPA under two scenarios:
+  1. 1.
+    Naïve: Materialize $\mathbf{Q}$ and $\mathbf{K}$ from factors, then perform the usual batched GEMM.
+  2. 2.
+    Specialized: Attempt to compute $\mathbf{Q}\,\mathbf{K}^{\top}$ directly from the rank-$(R_{q},R_{k})$ factors without explicitly forming $\mathbf{Q},\mathbf{K}$.
+Standard Multi-Head Attention. For batch size $B$ and sequence length $T$:
+  • Projection cost: $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}^{2}\bigr{)}$ or $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}\,d_{h}\bigr{)}$.
+  • Dot-product: $\mathbf{Q}\,(\mathbf{K})^{\top}\in\mathbb{R}^{(B\,h)\times T\times T}$ costs $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$.
+For large $T$, the $\mathcal{O}(B\,T^{2}\,d_{\text{model}})$ term dominates.
+TPA: Naïve Implementation.
+  • Constructing factors: $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}\,\times R_{q}(h+d_{h})+R_{k}(h+d_{h})+R_{v}(h+d_{h})\bigr{)}$.
+  • Materializing $\mathbf{Q},\mathbf{K}$: $\mathcal{O}\bigl{(}B\,T\,(R_{q}\,h\,d_{h}+R_{k}\,h\,d_{h})\bigr{)}$.
+  • Dot-product $\mathbf{Q}\,(\mathbf{K})^{\top}$: $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$.
+Typically $R_{q},R_{k},R_{v}\ll h$, so the overhead of constructing factors is small relative to $\mathcal{O}(T^{2}\,d_{\text{model}})$. Meanwhile, we still gain KV caching benefits.
+TPA: Specialized Implementation.
+If we bypass explicitly forming $\mathbf{Q},\mathbf{K}$, each dot product $\mathbf{Q}_{t}\cdot\mathbf{K}_{\tau}$ is a double sum over rank indices.
+Below we detail its complexity.
+A.4 Complexity Analysis for the Specialized Implementation
+Single-Head Complexity. A single attention head of dimension $d_{h}$. For each query:
+$\mathbf{Q}^{(i)}_{t}~{}=~{}\sum_{r=1}^{R_{q}}a_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
+and for each key:
+$\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{s=1}^{R_{k}}a_{k,i}^{(s)}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
+Their dot product:
+$\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}\Bigl{[}a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})\Bigr{]}~{}\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$
+For each pair $(r,s)$, we pay:
+  1. 1.
+    $\mathcal{O}(1)$ for multiplying two scalars,
+  2. 2.
+    $\mathcal{O}(d_{h})$ for the dot product $\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t})\cdot\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})$.
+Since $(r,s)$ runs over $R_{q}\times R_{k}$, each token-pair $(t,\tau)$ costs roughly
+$\mathcal{O}\Bigl{(}R_{q}\,R_{k}\,\bigl{(}1+d_{h}\bigr{)}\Bigr{)}~{}\approx~{}\mathcal{O}(R_{q}\,R_{k}\,d_{h}).$
+For $T$ queries and $T$ keys, that is $\mathcal{O}(T^{2}\,R_{q}\,R_{k}\,d_{h})$ for a single head.
+Multi-Head and Batches (Reusing $\mathbf{b}$-Dot Products).
+When extending to $h$ heads, each head $i$ has its own scalar factors $\mathbf{a}_{q,i}^{(r)}(\mathbf{x}_{t})$ and $\mathbf{a}_{k,i}^{(s)}(\mathbf{x}_{\tau})$, but the $\mathbf{b}$-vectors $\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t})$ and $\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})$ can still be shared across all heads (assuming the same rank-$R$ factors for every head).
+Hence, one can split the total cost into two stages:
+  1. 1.
+    $\mathbf{b}$-Dot-Product Stage:
+    For each token pair $(t,\tau)$ and each rank pair $(r,s)$, compute the dot product
+    $\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),~{}\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}~{}\in~{}\mathbb{R}.$
+    Since each dot product is $\mathcal{O}(d_{h})$ and there are $R_{q}R_{k}$ rank pairs as well as $T^{2}$ token pairs, this stage costs:
+    $\mathcal{O}\bigl{(}T^{2}\,R_{q}R_{k}\,d_{h}\bigr{)}.$
+    Crucially, these $\mathbf{b}$-dot products need only be computed once and can be cached for reuse by all heads.
+  2. 2.
+    Per-Head Scalar Multiplications:
+    After the $\mathbf{b}$-dot products are precomputed (and cached), each head $i$ only needs to multiply each stored dot product by the corresponding scalars $\mathbf{a}_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{a}_{k,i}^{(s)}(\mathbf{x}_{\tau})$.
+Since this scalar multiplication is $\mathcal{O}(1)$ per pair, and there are $T^{2}$ token pairs and $R_{q}R_{k}$ rank pairs for each of the $h$ heads, this step costs:
+    $\mathcal{O}\bigl{(}h\,T^{2}\,R_{q}R_{k}\bigr{)}.$
+Putting these together, for batch size $B$, the total cost is
+$\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}R_{k}\,d_{h}\bigr{)}~{}+~{}\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}R_{k}\bigr{)}~{}=~{}\mathcal{O}\Bigl{(}B\,T^{2}\,R_{q}R_{k}\bigl{(}d_{h}+h\bigr{)}\Bigr{)}.$
+By contrast, the standard multi-head attention dot-product step is $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$. Hence, for the specialized TPA approach to reduce flops,
+$R_{q}R_{k}\,(d_{h}+h)~{}\leq~{}h\,d_{h}.$
+Thus a practical guideline is to ensure $\,R_{q}R_{k}<h\,\tfrac{d_{h}}{d_{h}+h}\,$. When that holds, bypassing explicit materialization of $\mathbf{Q}$ and $\mathbf{K}$ can be beneficial.
+A.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
+We have explored a two-step procedure for computing $\mathbf{Q}\,\mathbf{K}^{\top}$ directly from factorized queries and keys without materializing $\mathbf{Q}$ or $\mathbf{K}$.
+Here, we extend this idea to also avoid explicitly forming $\mathbf{V}$.
+That is, all three activations $\mathbf{Q},\mathbf{K},\mathbf{V}$ remain factorized throughout the attention pipeline.
+We present a single-head formulation below, and then discuss multi-head and batch extensions.
+Extending the Two-Step Approach to Avoid $\mathbf{V}$ Materialization.
+After we obtain $\mathbf{Q}\mathbf{K}^{\top}$, we apply $\alpha_{t,\tau}=\operatorname{softmax}\bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,(QK^{\top})_{t,\tau}\bigr{)}$.
+The final attention output at token $t$ (single head) is
+$\operatorname{head}(t)~{}=~{}\sum_{\tau=1}^{T}\alpha_{t,\tau}~{}\mathbf{V}_{\tau}.$
+Using the factorization $\mathbf{V}_{\tau}=\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$, we write:
+$\operatorname{head}(t)~{}=~{}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
+Rearrange sums:
+$\operatorname{head}(t)~{}=~{}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
+We still do not explicitly form $\mathbf{V}_{\tau}$. Instead:
+Stage 1: Calculating $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for all tokens.
+We simply observe that each output $\operatorname{head}(t)$ can be computed by summing vectors $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\in\mathbb{R}^{d_{h}}$ weighted by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$.
+The complexity for constructing $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ $\forall u,\tau$ is $\mathcal{O}(T\,R_{v}\,d_{h})$.
+Stage 2: Weighted Summation by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$. For each token $t$, the final attention head output is
+$\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})~{}=~{}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
+We still never explicitly materialize $\mathbf{V}$.
+Instead, for each pair $(t,u)$, we must accumulate the sum of $T$ vectors $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\in\mathbb{R}^{d_{h}}$, each scaled by the scalar $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$.
+Because each vector is $d_{h}$-dimensional, each $(t,u)$ summation costs $\mathcal{O}(T\,d_{h})$.
+Summed over $t=1\ldots T$ and $u=1\ldots R_{v},$ the total work is $\mathcal{O}(T^{2}\,R_{v}\,d_{h})$ for the entire sequence.
+In practice, one precomputes all $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for $\tau=1\ldots N$, so each accumulation can be implemented as a simple “scalar‐times‐vector add” in a tight loop.
+This cost is usually smaller than the $\mathbf{Q}\mathbf{K}^{\top}$ factorized cost if $R_{v}\ll h$.
+A.6 Overall Complexity for Single-Head
+Combining the four bullet‐point stages from above (ignoring smaller overheads like the softmax) yields:
+  1. (i)
+    QK $\mathbf{b}$-Dot Product Stage: $\mathcal{O}(T^{2}\,R_{q}\,R_{k}\,d_{h})$.
+  2. (ii)
+    QK Scalar-Multiply Stage: $\mathcal{O}(T^{2}\,R_{q}\,R_{k})$.
+  3. (iii)
+    Computing $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for all tokens: $\mathcal{O}(T\,R_{v}\,d_{h})$.
+  4. (iv)
+    Weighted Summation by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$: $\mathcal{O}(T^{2}\,R_{v}\,d_{h})$.
+Hence, for a single head, the total cost is:
+$\mathcal{O}\Bigl{(}T^{2}\,R_{q}\,R_{k}\,d_{h}~{}+~{}T^{2}\,R_{q}\,R_{k}~{}+~{}T\,R_{v}\,d_{h}~{}+~{}T^{2}\,R_{v}\,d_{h}\Bigr{)}.$
+In many cases (especially for large $T$), the $\mathcal{O}(T^{2})$ terms dominate, so one often focuses on
+$\mathcal{O}\Bigl{(}T^{2}\,R_{q}\,R_{k}\,d_{h}~{}+~{}T^{2}\,R_{q}\,R_{k}~{}+~{}T^{2}\,R_{v}\,d_{h}\Bigr{)}.$
+A.7 Multi-Head and Batch Extensions (Reuse of $\mathbf{b}$-Dot Products)
+When extending to $h$ heads and batch size $B$, all sequence‐length‐dependent terms are multiplied by $\sim B\,h$. However, crucial $\mathbf{b}$-dot products can be shared across heads:
+QK $\mathbf{b}$-Dot Products. Since each head has distinct scalar factors $a_{q,i},\,a_{k,i}$ but the same $\mathbf{b}_{q}^{(r)},\,\mathbf{b}_{k}^{(s)}$ across heads, each pairwise dot product
+$\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\rangle$
+is computed just once per batch. That cost remains
+$\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}\,R_{k}\,d_{h}\bigr{)},$
+not multiplied by $h$.
+After caching these dot products, each of the $h$ heads pays $\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}\,R_{k}\bigr{)}$ total for the head‐specific scalar multiplications (the “$\alpha_{t,\tau}$”–like factors).
+V $\mathbf{b}$-Evaluations.
+Likewise, the $\mathbf{b}_{v}^{(u)}$ factors are shared across heads (i.e.
+one set of $\mathbf{b}_{v}$-vectors for all heads).
+Hence, computing all $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for $\tau=1\ldots T$ (across the batch) is a one‐time cost:
+$\mathcal{O}\bigl{(}B\,T\,R_{v}\,d_{h}\bigr{)}.$
+Then each head $i$ has its own scalar factors $a_{v,i}^{(u)}(\mathbf{x}_{\tau})$, so the final accumulation $\sum_{\tau=1}^{T}\alpha_{t,\tau}\,a_{v,i}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ costs $\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{v}\,d_{h}\bigr{)}$ in total (for all $t,u$).
+Putting it all together, the total flops for multi‐head attention with batch size $B$ are:
+$\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}\,R_{k}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{QK $\mathbf{b}$-dot products}\\ \text{(shared across heads)}\end{subarray}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}\,R_{k}\bigr{)}}_{\text{per-head QK scalar mult.}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T\,R_{v}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{Compute $\mathbf{b}_{v}$ for all tokens}\\ \text{(shared across heads)}\end{subarray}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{v}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{final accumulations}\\ \text{(per head)}\end{subarray}}.$
+Discussion.
+By contrast, standard multi-head attention typically requires $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$ flops for the $\mathbf{Q}\mathbf{K}^{\top}$ dot product (plus a similar $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$ for multiplying by $\mathbf{V}$).
+The factorization can yield savings provided $R_{q}R_{k}\ll h$ (for QK) and $R_{v}\ll h$ (for V), though actual speedups depend on how well these multi‐stage kernels are implemented and on hardware efficiency.
+By retaining $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}$ in factorized form, one can forgo the usual steps:
+$\mathbf{x}_{t}~{}\mapsto~{}\mathbf{Q}_{t},~{}\mathbf{K}_{\tau}~{}\mapsto~{}(\mathbf{Q}\,\mathbf{K}^{\top})~{}\mapsto~{}\operatorname{softmax}(\mathbf{Q}\,\mathbf{K}^{\top})\,\mathbf{V}~{}\mapsto~{}\text{final output}.$
+Instead, the large $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors (of size $T\times d_{h}$) are never materialized.
+The cost is replaced by rank-based $\mathbf{b}$-dot‐product computations plus per‐head scalar multiplications.
+The main challenge is to keep the factor ranks $(R_{q},R_{k},R_{v})$ sufficiently small relative to $d_{h}$ and to implement the necessary multi‐stage kernels efficiently.
+When $R_{q},R_{k},R_{v}\ll h$, fully factorized QKV attention can yield substantial gains in both computation and memory footprint.
+A.8 Decoding Speed during Inference Time of MHA, MQA, GQA, MLA, and TPA
+Suppose we are in an autoregressive setting, decoding the current token $\mathbf{x}_{T}$ given cached keys and values (KV) from all previous tokens $\mathbf{x}_{1},\dots,\mathbf{x}_{T-1}$.
+For each attention head $i\in\{1,\dots,h\}$, we store $\mathbf{K}_{i}\in\mathbb{R}^{T\times d_{h}},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$.
+Below, we compare the flops needed by MHA, MQA, GQA, MLA, and TPA to compute the next-token logits during inference.
+MHA, MQA, and GQA. Despite sharing or grouping keys/values in MQA and GQA, the decoding cost for MHA, MQA, and GQA remains of the same order. Specifically, for each head $i$, we compute:
+$\mathbf{Q}_{i}(\mathbf{x}_{T})~{}\in~{}\mathbb{R}^{d_{h}},\quad\mathbf{K}_{i}~{}\in~{}\mathbb{R}^{T\times d_{h}},\quad\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}~{}\in~{}\mathbb{R}^{1\times T},\quad\text{and}\quad\operatorname{Softmax}\bigl{(}\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}\bigr{)}\,\mathbf{V}_{i}~{}\in~{}\mathbb{R}^{d_{h}}.$
+Hence, the flops scale linearly in $h$, $d_{h}$, and $T$. For example, forming $\mathbf{Q}_{i}(\mathbf{x}_{T})\mathbf{K}_{i}^{\top}$ for each head $i$ costs roughly $\mathcal{O}(h\,d_{h}\,T)$.
+MLA.
+During inference, MLA can be seen as MQA but uses a larger head dimension to accommodate both RoPE and compressed representations (e.g., $d_{h}^{\prime}=d_{\text{rope}}+d_{c}$).
+In typical configurations, $d_{\text{rope}}+d_{c}$ can be significantly larger (e.g., $d_{h}^{\prime}=576$ rather than $d_{h}=64$ or $128$), thus inflating the dot-product cost by roughly 4.5×4.5$\times$4.5 × to 9×9$\times$9 × compared to MHA/MQA/GQA.
+TPA.
+Recall that TPA factorizes $\mathbf{Q}$ and $\bm{K}$ into rank-$(R_{q},R_{k})$ terms (see Section A), potentially avoiding large $\mathbf{Q},\bm{K}$ materializations.
+At inference, TPA’s dot-product cost can be broken into two parts:
+$\underbrace{R_{q}\,R_{k}\,d_{h}\,T}_{\text{QK $\mathbf{b}$-dot products (shared across all heads)}}~{}+~{}2\,\underbrace{R_{q}\,R_{k}\,h\,T}_{\text{per-head scalar multiplications}},$
+where $T$ is the current sequence length. For concrete values $d_{h}=128$, $h=64$, $R_{q}=8$, and $R_{k}=2$ (or $R_{q}=16$, $R_{k}=1$), we obtain:
+[…]
+[…]
+[…]
+Thus, in this setup, TPA can significantly reduce the flops needed for computing the $\mathbf{Q}(\mathbf{x}_{T})\bm{K}^{\top}$ operation at each decoding step.
+The actual end-to-end wall-clock speedup also depends on kernel fusion, caching strategies, and hardware implementation details, but the factorized formulation offers a pathway to more efficient decoding than standard attention.
+Appendix B Higher-Order Tensor Product Attention
+All prior discussions have focused on a second-order factorization in which each rank-$R_{Q}$ (and similarly $R_{K}$, $R_{V}$) component is the outer product of two vectors: one in $\mathbb{R}^{h}$ (the “head” dimension) and one in $\mathbb{R}^{d_{h}}$.
+We now generalize this by introducing an additional latent factor, yielding a third-order (or higher) factorization reminiscent of canonical polyadic (CP) decomposition.
+Concretely, for a single token $t$, we write
+$\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\,\otimes\,\operatorname{vec}\bigl{(}\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\,\otimes\,\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\bigr{)},$
+where the newly introduced factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{\mathbf{c}}}$ can be viewed as a learnable gate or modulation term.
+Analogous expansions apply to $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$.
+In practice, these triple (or higher-order) products still collapse into a matrix in $\mathbb{R}^{h\times d_{h}}$.
+One straightforward way to achieve this collapse is to split the feature dimension $d_{h}$ such that $d_{b}\times d_{c}=d_{h}$,
+$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}},\quad\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}},\quad\operatorname{vec}\bigl{(}\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\bigr{)}\in\mathbb{R}^{d_{h}}.$
+This additional factor can enhance expressiveness without necessarily increasing the base rank.
+Conceptually, it can act as a learnable nonlinearity or gating mechanism.
+One could also tie or share $\mathbf{c}^{Q}_{r}$ across queries, keys, and values, to reduce parameter overhead.
+A similar setup holds for keys (with rank $R_{K}$) and values (with rank $R_{V}$).
+Although this extra dimension adds to the parameter count, it can reduce the required rank to achieve a certain level of representational power.
+From a memory perspective, higher-order TPA still leverages factorized KV caching: only the factors $\mathbf{a}(\mathbf{x}_{t}),\mathbf{b}(\mathbf{x}_{t})$, and $\mathbf{c}(\mathbf{x}_{t})$ for each past token are cached.
+As usual, a trade-off arises between model capacity and the overhead of memory and computing.
+Nonetheless, moving from a rank-$\bigl{(}R_{Q},R_{K},R_{V}\bigr{)}$ matrix factorization to a higher-order tensor decomposition can provide additional flexibility and increased capacity.
+B.1 RoPE Compatibility in Higher-Order TPA
+Rotary positional embeddings (RoPE) remain compatible even under higher-order factorizations.
+In second-order TPA, RoPE can be treated as an invertible blockwise linear map acting on the last dimension of $\mathbf{Q}_{t}$ or $\mathbf{K}_{t}$.
+The same argument carries over when a third factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})$ is present.
+Suppose RoPE acts on the $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})$ portion (of dimension size $d_{b}$), we have the following theorem.
+Theorem 2 (RoPE Compatibility in Higher-Order TPA).
+Consider the higher-order (3-order) Tensor Product Attention (TPA) query factorization
+$\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\in\mathbb{R}^{h\times d_{h}},$
+where $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$, $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$, with $d_{c}=\tfrac{d_{h}}{d_{b}}$.
+Define the RoPE-transformed query as $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}=\mathbf{Q}_{t}\mathbf{T}_{t}$, where
+$\mathbf{T}_{t}=\mathbf{R}_{t}\otimes\mathbf{I}_{d_{c}}=\begin{pmatrix}\mathbf{R}_{t}&\cdots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{R}_{t}&\cdots&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&\mathbf{R}_{t}\\ \end{pmatrix}\in\mathbb{R}^{d_{h}\times d_{h}},$
+and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ ($d_{b}\in\mathbb{Z}_{+}$ is even) is a block-diagonal matrix composed of $2\times 2$ rotation matrices:
+$\mathbf{R}_{t}=\begin{pmatrix}\cos(t\theta_{1})&-\sin(t\theta_{1})&&&\\ \sin(t\theta_{1})&\cos(t\theta_{1})&&&\\ &&\cos(t\theta_{2})&-\sin(t\theta_{2})&\\ &&\sin(t\theta_{2})&\cos(t\theta_{2})&\\ &&&&\ddots\\ &&&&&\cos(t\theta_{d_{b}/2})&-\sin(t\theta_{d_{b}/2})\\ &&&&&\sin(t\theta_{d_{b}/2})&\cos(t\theta_{d_{b}/2})\\ \end{pmatrix},$
+for $t\in\{1,\dots,T\}$ and $j\in\{1,\dots,d_{b}/2\}$.
+This construction ensures that RoPE rotates only the coordinates corresponding to $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged.
+Under these conditions, the RoPE-transformed query $\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}$ admits a higher-order TPA factorization of the same rank $R_{Q}$.
+Specifically, we have
+$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$ (B.1)
+where $\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
+Please see Appendix C.2 for the proof. For fourth-order or higher, this result still holds.
+Appendix C Proofs of Theorems
+C.1 Proof of Theorem 1
 Proof.
@@
 ∎
-Appendix B More on Experiments
-B.1 Experimental Settings
+C.2 Proof of Theorem 2
+Proof.
+We begin by observing that each term $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}$ is an element of $\mathbb{R}^{h}\otimes\mathbb{R}^{d_{h}}$.
+Here, $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$, $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$, with $d_{c}=\tfrac{d_{h}}{d_{b}}$.
+Consequently, the tensor product $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ forms a $d_{b}\times d_{c}$ matrix, and its vectorization lies in $\mathbb{R}^{d_{b}\cdot d_{c}}=\mathbb{R}^{d_{h}}$.
+Applying the RoPE transformation to a single summand yields
+$\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\mapsto\mathbf{T}_{t}\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
+Since $\mathbf{T}_{t}$ is defined as the Kronecker product $\mathbf{R}_{t}\otimes\mathbf{I}_{d_{c}}$, where $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ and $\mathbf{I}_{d_{c}}$ is the identity matrix of size $d_{c}\times d_{c}$, it follows that
+$\mathbf{T}_{t}\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}=\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
+This is because the Kronecker product with an identity matrix effectively applies the rotation $\mathbf{R}_{t}$ to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged.
+Therefore, the RoPE transformation of a single summand becomes
+$\operatorname{RoPE}_{t}\Bigl{(}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\Bigr{)}=\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
+Importantly, this transformation does not mix the components $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ and $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$; it solely rotates $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ via $\mathbf{R}_{t}$.
+Summing over all ranks $r=1,\dots,R_{Q}$, we obtain
+$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$
+which retains the same higher-order TPA structure with rank $R_{Q}$.
+Thus, the RoPE transformation is fully compatible with higher-order TPA, preserving the factorization rank and maintaining the structure by only rotating the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ components while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged.
+∎
+Appendix D More Related Works
+Low-Rank Factorizations.
+Low-rank approximations have been applied to compress model parameters and reduce complexity including LoRA (Hu et al., 2022), which factorizes weight updates during fine-tuning, and its derivatives for other training scenarios such as efficient pretraining (ReLoRA (Lialin et al., 2023), MoRA (Jiang et al., 2024)), long-context training (LongLoRA (Chen et al., 2024), SinkLoRA (Zhang, 2024)), as well as continual training (InfLoRA (Liang & Li, 2024), GS-LoRA (Zhao et al., 2024), I-LoRA (Ren et al., 2024)).
+These approaches typically produce static low-rank expansions that do not explicitly depend on the input context.
+And Malladi et al.
+(2023); Zeng & Lee (2024) provided theoretical proof of the expressiveness of low-rank approximation.
+For the initialization of factorization matrices, OLoRA (Büyükakyüz, 2024) applied QR-decomposition of pretrained weight to achieve better performance of language models while LoLDU (Shi et al., 2024) used LDU-decomposition to accelerate training of LoRA.
+Moreover, AdaLoRA (Zhang et al., 2023a) utilized Singular Value Decomposition (SVD) of the pretrained weight and introduced importance score for each parameter as a measurement to achieve dynamic adjustment of rank.
+TPA, by contrast, constructs Q, K, and V as contextually factorized tensors, enabling dynamic adaptation.
+Appendix E More on Attention Mechanisms
+E.1 Multi-Query Attention (MQA)
+Multi-Query Attention (MQA) (Shazeer, 2019) significantly reduces memory usage by sharing keys and values across heads, while still preserving unique query projections.
+For a sequence of embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$,
+$\mathbf{Q}_{i}=\mathbf{X}\bm{W}^{Q}_{i},\quad\mathbf{K}_{\text{shared}}=\mathbf{X}\bm{W}^{K}_{\text{shared}},\quad\mathbf{V}_{\text{shared}}=\mathbf{X}\bm{W}^{V}_{\text{shared}}.$
+Hence, each head $i$ only has a distinct query $\mathbf{Q}_{i}\in\mathbb{R}^{T\times d_{h}}$, but shares the same key $\mathbf{K}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$ and value $\mathbf{V}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$.
+In practice, this means:
+$\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}},\quad\bm{W}^{K}_{\text{shared}},\bm{W}^{V}_{\text{shared}}\;\in\;\mathbb{R}^{\,d_{\text{model}}\times d_{h}}.$
+The resulting MQA operation is:
+$\operatorname{MQA}(\mathbf{X})=\operatorname{Concat}\Bigl{(}\textbf{head}_{1},\dots,\textbf{head}_{h}\Bigr{)}\,\bm{W}^{O},$
+where
+$\textbf{head}_{i}=\operatorname{Attention}\bigl{(}\mathbf{Q}_{i},\mathbf{K}_{\text{shared}},\mathbf{V}_{\text{shared}}\bigr{)}.$
+By sharing these key and value projections, MQA cuts down on memory usage (especially for the key-value cache in autoregressive inference) but loses some expressivity since all heads must rely on the same key/value representations.
+E.2 Grouped Query Attention (GQA)
+Grouped Query Attention (GQA) (Ainslie et al., 2023) generalizes MHA and MQA by grouping heads.
+Specifically, we partition the $h$ total heads into $G$ groups.
+Each group has a single set of keys and values, but each individual head within that group still retains its own query projection.
+Formally, if $g(i)$ maps a head $i\in[h]$ to its group index $g\in[G]$, then:
+$\mathbf{K}_{g(i)}=\mathbf{X}\,\bm{W}^{K}_{g(i)},\quad\mathbf{V}_{g(i)}=\mathbf{X}\,\bm{W}^{V}_{g(i)},\quad\mathbf{Q}_{i}=\mathbf{X}\,\bm{W}^{Q}_{i},$
+and
+$\text{head}_{i}=\operatorname{Attention}\Bigl{(}\mathbf{Q}_{i},\mathbf{K}_{g(i)},\mathbf{V}_{g(i)}\Bigr{)}.$
+Again, $\bm{W}^{K}_{g},\bm{W}^{V}_{g}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ for each group $g$, and $\bm{W}^{Q}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{h}}$ for each head $i$.
+The complete output is again a concatenation of all heads:
+$\operatorname{GQA}(\mathbf{X})=\operatorname{Concat}\Bigl{(}\text{head}_{1},\ldots,\text{head}_{h}\Bigr{)}\bm{W}^{O}.$
+By adjusting $G$ between $1$ and $h$, GQA can interpolate between sharing all key/value projections across heads (i.e., MQA) and having one set of projections per head (i.e., MHA).
+E.3 Multi-head Latent Attention (MLA)
+Below, we briefly outline the Multi-head Latent Attention (MLA) approach used by DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b).
+MLA introduces a low-rank compression of the keys and values to reduce the Key-Value (KV) caching cost at inference.
+$\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},$
+$\operatorname{Concat}\bigl{(}\mathbf{K}_{1}^{C},\mathbf{K}_{2}^{C},\ldots,\mathbf{K}_{h}^{C}\bigr{)}$ $=\mathbf{K}^{C}=\mathbf{C}^{KV}\bm{W}^{UK},$
+$\mathbf{K}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{X}\bm{W}^{KR}\bigr{)},$
+$\mathbf{K}_{i}$ $=\operatorname{Concat}\bigl{(}\mathbf{K}_{i}^{C},\mathbf{K}^{R}\bigr{)},$
+$\operatorname{Concat}\bigl{(}\mathbf{V}_{1}^{C},\mathbf{V}_{2}^{C},\ldots,\mathbf{V}_{h}^{C}\bigr{)}$ $=\mathbf{V}^{C}=\mathbf{C}^{KV}\bm{W}^{UV},$
+where $\bm{W}^{DKV}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}},\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times d_{h}h},\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}},\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times d_{h}h}$, and $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the compressed KV latent (with $d_{c}\ll d_{h}h$), and $\operatorname{RoPE}(\cdot)$ represents the RoPE transform applied to the separate key embeddings $\mathbf{K}^{R}$ of dimension $d_{h}^{R}$.
+Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ need to be cached, reducing KV memory usage while largely preserving performance compared to standard MHA (Vaswani et al., 2017).
+MLA also compresses the queries, lowering their training-time memory footprint:
+$\mathbf{C}^{Q}$ $=\mathbf{X}\bm{W}^{DQ},$
+$\operatorname{Concat}\bigl{(}\mathbf{Q}_{1}^{C},\mathbf{Q}_{2}^{C},\,\ldots,\mathbf{Q}_{h}^{C}\bigr{)}$ $=\mathbf{Q}^{C}=\mathbf{C}^{Q}\bm{W}^{UQ},$
+$\operatorname{Concat}\bigl{(}\mathbf{Q}_{1}^{R},\,\mathbf{Q}_{2}^{R},\,\ldots,\,\mathbf{Q}_{h}^{R}\bigr{)}$ $=\mathbf{Q}^{R}=\operatorname{RoPE}\bigl{(}\mathbf{C}^{Q}\bm{W}^{QR}\bigr{)},$
+$\mathbf{Q}$ $=\operatorname{Concat}\bigl{(}\mathbf{Q}^{C},\mathbf{Q}^{R}\bigr{)}.$
+where $\bm{W}^{DQ}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}^{\prime}},\bm{W}^{UQ}\in\mathbb{R}^{\,d_{c}^{\prime}\times d_{h}h},\bm{W}^{QR}\in\mathbb{R}^{d_{c}^{\prime}\times d_{h}^{R}h}$.
+Here, $\mathbf{C}^{Q}\in\mathbb{R}^{T\times d_{c}^{\prime}}$ (with $d_{c}^{\prime}\ll d_{h}h$) is the compressed query latent.
+As above, each $\bm{W}^{DQ}$, $\bm{W}^{UQ}$, and $\bm{W}^{QR}$ connects these lower-dimensional query latents back to $h$ heads of dimension $d_{h}+d_{h}^{R}$.
+Given compressed queries, keys, and values, the final attention output for the $t$-th token is:
+$\mathbf{O}_{i}$ $=\operatorname{Softmax}\Bigl{(}\tfrac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\top}}{\sqrt{d_{h}+d_{h}^{R}}}\Bigr{)}\;\mathbf{V}_{i}^{C},$
+$\mathbf{U}$ $=\operatorname{Concat}\bigl{(}\mathbf{O}_{1},\mathbf{O}_{2},\ldots,\mathbf{O}_{h}\bigr{)}\bm{W}^{O},$
+where $\bm{W}^{O}\in\mathbb{R}^{(d_{h}h)\times\,d_{\text{model}}}$ is the output projection.
+In inference time, $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ can be cached to accelerate decoding.
+In detail, when RoPE is ignored, the inner product $\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ (where $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{d}$) of the $i$-th head between $t$-th and $s$-th tokens can be calculated using the hidden state $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ for $t$-th token and the cached latent state $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ for $s$-th token:
+$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (E.1)
+$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (E.2)
+where $\bm{W}_{i}^{(\cdot)}$ is the $i$-th head of the original weight, and $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ can be computed previously for faster decoding.
+However, this process fails when RoPE is considered according to Su (2024).
+Since RoPE can be considered as multiplication with a block-diagonal matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ (see Section 2.3), with the property (2.1) that $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$, then
+[…]
+Different from (E.2), acceleration by pre-computing $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t-s}}(\bm{W}^{UK}_{i})^{\top}]$ fails since it varies for different $(t,s)$ position pairs.
+Therefore, MLA adds the additional $\mathbf{k}_{t}^{R}$ part with a relatively smaller size for RoPE compatibility.
+In Section 3.2, we will show that TPA addresses the issue of RoPE-incompatibility by applying tensor product.
+$\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},$
+$\operatorname{Concat}\bigl{(}\mathbf{K}_{1}^{C},\mathbf{K}_{2}^{C},\ldots,\mathbf{K}_{h}^{C}\bigr{)}$ $=\mathbf{K}^{C}=\mathbf{C}^{KV}\bm{W}^{UK},$
+$\mathbf{K}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{X}\bm{W}^{KR}\bigr{)},$
+$\mathbf{K}_{i}$ $=\operatorname{Concat}\bigl{(}\mathbf{K}_{i}^{C},\mathbf{K}^{R}\bigr{)},$
+$\operatorname{Concat}\bigl{(}\mathbf{V}_{1}^{C},\mathbf{V}_{2}^{C},\ldots,\mathbf{V}_{h}^{C}\bigr{)}$ $=\mathbf{V}^{C}=\mathbf{C}^{KV}\bm{W}^{UV},$
+E.4 Multi-matrix Factorization Attention (MFA)
+Hu et al. (2024) proposed Multi-matrix Factorization Attention (MFA), which can be seen as Multi-Query Attention (MQA) with dimension of each head equals $d_{C}$, and low-rank factorized Q:
+$\mathbf{Q}_{i}=\mathbf{X}\bm{W}^{DQ}\bm{W}^{UQ}_{i},\quad\mathbf{K}_{\text{shared}}=\mathbf{X}\bm{W}^{K}_{\text{shared}},\quad\mathbf{V}_{\text{shared}}=\mathbf{X}\bm{W}^{V}_{\text{shared}},$
+where
+$\bm{W}^{DQ}\in\mathbb{R}^{d_{\text{model}}\times d_{c}},\quad\bm{W}^{UQ}_{i}\in\mathbb{R}^{d_{c}\times d_{c}},\quad\bm{W}^{K}_{\text{shared}},\bm{W}^{V}_{\text{shared}}\;\in\;\mathbb{R}^{\,d_{\text{model}}\times d_{c}}.$
+Appendix F Other Variants of TPA
+TPA with Non-contextual B.
+Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to $\mathbf{x}_{t}$.
+For keys:
+$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{K}_{r},$
+and similarly for values. This arrangement is effective if the token-dimension structure remains mostly uniform across the sequence, while the head-dimension factors capture context.
+TPA KV Only. One can preserve a standard query mapping,
+$\mathbf{Q}_{t}=\bm{W}^{Q}\,\mathbf{x}_{t}\in\mathbb{R}^{h\times d_{h}},$
+and factorize only the keys and values. This leaves the query projection as the original linear transformation while reducing memory usage via factorized KV caching.
+TPA KV with Shared $\mathbf{B}$. Another variant is to share the token-dimension factors of keys and values:
+$\mathbf{b}^{K}_{r}(\mathbf{x}_{t})=\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$
+lowering parameter counts and the KV cache footprint.
+While it constrains $\mathbf{K}$ and $\mathbf{V}$ to be formed from the same token basis, it can still perform well and provide additional memory savings.
+Nonlinear Head Factors.
+Rather than applying purely linear mappings to the head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}$, one may introduce element-wise nonlinearities such as $\sigma(\cdot)$ or $\operatorname{softmax}(\cdot)$.
+This effectively yields a Mixture of Heads Attention (MoH Attention), where each component becomes a learned mixture weight modulated by the nonlinearity.
+Discussion.
+These variants illustrate TPA’s versatility in balancing memory cost, computational overhead, and representation power.
+By choosing which dimensions (heads or tokens) remain contextual and adjusting ranks $(R_{Q},R_{K},R_{V})$, TPA unifies multiple existing attention mechanisms—such as MHA, MQA, and GQA—under one framework, while potentially reducing the KV cache size by an order of magnitude during autoregressive inference.
+Appendix G More on Experiments
+G.1 Experimental Settings
 We list the main architecture hyper-parameters and training devices in Table 4.
@@
 Large 772M 8$\times$ A100 GPUs 15 4 36 1280
+XL 1.55B 8$\times$ A100 GPUs 6 10 48 1600
 Table 5: The architecture hyper-parameters for different models.
-Model Size Small Medium Large
-$h$ (MHA) 12 16 20
-$h$ (MQA) 23 31 39
-$h$ (GQA) 22 30 38
-$n_{h}$ (MLA) 12 23 34
-$h$ (TPA-KVonly) 22 29 37
-$h$ (TPA) 34 47 61
-$d_{c}$ (MLA) 256 512 512
-$d_{c}^{\prime}$ (MLA) 512 1024 1024
-B.2 Additional Experimental Results
-We display the evaluation results for small-size (124M) models in Tables 8-9.
+Model Size Small Medium Large XL
+$h$ (MHA) 12 16 20 25
+$h$ (MQA) 23 31 39 49
+$h$ (GQA) 22 30 38 48
+$h$ (MLA) 12 23 34 49
+$h$ (TPA-KVonly) 22 29 37 47
+$h$ (TPA) 34 47 61 78
+$d_{c}$ (MLA) 256 512 512 512
+$d_{c}^{\prime}$ (MLA) 512 1024 1024 1024
+G.2 Additional Experimental Results
+G.2.1 Perplexity Curves
+We display the perplexity curves for medium, large and XL size of models in Figure 4.
+(a) Validation Perplexity
+(b) Validation Perplexity
+(c) Validation Perplexity
+Figure 4: The validation perplexity of medium-size (353M) models, large-size (773M), and XL-size (1.5B) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
+G.2.2 Ablation Study on Different Ranks
+Figure 5 shows the training loss, validation loss, and validation perplexity curves of XL-size (1.5B) T6 models with different ranks trained on the FineWeb-Edu 100B dataset, and the evaluation results are displayed in Table 7.
+It can be observed that increase in rank can improve the performances of large language models.
+(a) Training Loss
+(b) Validation Loss
+(c) Validation Perplexity
+Figure 5: The training loss, validation loss and validation perplexity curves of XL-size (1.5B) T6 models with different ranks on the FineWeb-Edu 100B dataset.
+G.2.3 0-shot Evaluation with lm-evaluation-harness
+For the evaluation, We show the 0-shot performances with lm-evaluation-harness for small-size (124M) and XL-size (1.5B) models in Tables 6 and 7.
 Table 6: The evaluation results of small models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
@@
 GQA 48.70 25.68 56.15 35.58 31.40 64.91 51.62 23.12 68.20 45.04
-MLA 49.66 26.45 61.22 33.94 32.40 62.73 50.43 23.29 71.50 45.74
+MLA 50.21 26.71 58.01 36.25 32.80 64.69 50.59 24.67 71.90 46.20
 TPA-KVonly 51.05 26.54 57.25 36.77 32.60 65.02 50.91 23.64 69.70 45.94
+TPA (non-ctx-A) 50.17 25.60 57.95 36.13 31.40 64.80 49.57 24.88 64.80 45.03
 TPA 51.26 27.39 57.00 36.68 32.80 64.47 49.72 24.61 72.00 46.21
+Table 7: The evaluation results of XL models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
+The best scores in each column are bolded.
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+If not specified, TPA and TPA-KVonly set $R_{K}=R_{V}=2$.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+MHA 64.81 35.41 61.90 54.32 37.20 72.74 55.80 25.44 82.80 54.49
+MQA 64.10 36.01 62.26 54.38 39.00 72.58 56.43 23.70 81.90 54.48
+GQA 63.68 35.92 60.46 54.17 38.40 73.56 56.27 24.77 81.70 54.33
+MLA 64.14 35.92 60.12 53.60 39.20 72.25 55.17 24.71 81.60 54.08
+TPA-KVonly 65.61 36.77 63.02 54.17 37.00 73.34 54.62 25.02 81.60 54.57
+TPA-KVonly ($R_{K,V}=4$) 64.52 37.03 63.27 54.89 39.80 72.91 56.51 24.74 81.60 55.03
+TPA-KVonly ($R_{K,V}=6$) 65.78 35.92 61.71 54.86 38.60 72.69 57.93 25.59 82.20 55.03
+TPA 66.71 36.52 61.38 54.03 40.40 72.52 56.83 24.49 82.20 55.01
+G.2.4 2-shot Evaluation with lm-evaluation-harness
+We also show 2-shot performances in Tables 8910 and 11.
 Table 8: The evaluation results of small models with different attention mechanisms on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
@@
 GQA 55.01 25.94 55.72 35.68 31.80 65.29 51.93 25.27 77.80 47.16
-MLA 52.78 26.19 57.25 33.19 29.60 63.98 50.43 24.90 76.00 46.04
+MLA 54.76 27.13 58.07 36.13 31.40 65.07 51.30 25.90 78.90 47.63
 TPA-KVonly 54.25 27.90 57.06 36.36 31.80 64.31 53.59 26.18 79.20 47.85
+TPA (non-ctx-A) 55.09 27.65 53.82 36.24 30.20 64.53 50.75 26.01 78.60 46.99
 TPA 57.53 28.07 56.33 36.49 31.80 64.36 51.14 25.92 79.70 47.93
-B.3 Ablation Studies on Learning Rates
-We implement a set of parallel experiments for medium models with learning rate $6\times 10^{-4}$, and the curves for training loss, validation loss and validation perplexity are displayed in Figure 5.
-We also show the performance of these models on the benchmarks described in Section 4 in Tables 10-11.
-The results show that TPA and TPA-KVonly models can also outperform other types of attention with different learning rates.
-(a) Training Loss
-(b) Validation Loss
-(c) Validation Perplexity
-Figure 5: The training loss, validation loss and validation perplexity of medium-size (353M) models with learning rate $6\times 10^{-4}$ and different attention mechanisms on the FineWeb-Edu 100B dataset.
-Table 10: The evaluation results of medium models (learning rate=$6\times 10^{-4}$) with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
-The best scores in each column are bolded.
-Abbreviations: HellaSw.
-= HellaSwag, W.G.
-= WinoGrande.
-Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
-MHA 59.51 29.52 59.60 45.68 34.20 68.82 53.43 23.33 76.90 50.11
-MQA 57.62 31.91 59.45 45.69 35.40 69.31 53.51 26.47 74.60 50.44
-GQA 28.67 31.48 58.29 45.45 35.20 68.50 54.46 24.58 76.50 47.01
-MLA 57.49 29.44 59.97 44.09 25.77 68.66 53.04 25.77 76.40 48.96
-TPA-KVonly 58.01 30.12 58.01 45.95 35.60 69.10 53.12 25.39 75.10 50.04
-TPA 58.38 31.57 59.39 46.83 37.00 70.02 54.06 25.52 79.90 51.41
-Table 11: The evaluation results of medium models (learning rate $6\times 10^{-4}$) with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+Table 9: The evaluation results of medium models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
 The best scores in each column are bolded.
@@
 GQA 65.24 33.19 56.54 45.41 34.80 69.04 55.72 24.73 87.90 52.51
-MLA 63.80 31.06 58.50 44.19 35.40 68.44 51.62 25.22 88.50 51.86
+MLA 64.98 33.62 53.52 45.94 33.00 68.55 51.85 25.46 89.10 51.78
 TPA-KVonly 64.69 32.34 59.48 46.23 35.40 70.08 54.06 25.64 86.30 52.69
+TPA (non-ctx-A) 65.45 33.79 56.88 45.23 33.60 68.61 54.22 25.00 85.00 51.98
 TPA 67.97 34.56 57.22 46.87 34.60 69.91 52.01 25.07 89.90 53.12
+Table 10: The evaluation results of large models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+The best scores in each column are bolded.
+Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
+Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
+MHA 67.85 36.35 59.82 50.22 35.00 70.67 53.35 23.92 91.10 54.25
+MQA 68.86 36.09 53.79 50.50 37.00 70.89 54.70 25.01 88.00 53.87
+GQA 69.15 36.09 58.84 50.29 36.20 70.73 54.22 26.08 90.00 54.62
+MLA 70.54 38.74 61.50 51.86 36.00 70.89 54.22 25.47 92.40 55.74
+TPA-KVonly 71.34 37.71 59.76 51.10 36.00 71.49 54.62 25.83 90.10 55.33
+TPA 70.41 37.71 60.06 51.30 34.00 71.06 54.54 25.79 90.30 55.02
+Table 11: The evaluation results of XL models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+The best scores in each column are bolded.
+Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
+If not specified, We set $R_{K}=R_{V}=2$ for TPA and TPA-KVonly.
+Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
+MHA 70.83 39.93 59.85 54.05 36.20 72.52 55.17 25.42 91.70 56.18
+MQA 71.34 39.76 58.93 54.27 39.40 72.96 57.38 24.74 91.90 56.74
+GQA 71.17 39.08 60.18 54.05 37.40 73.07 56.35 24.87 92.20 56.49
+MLA 70.79 37.54 50.83 53.33 40.00 72.09 56.51 24.93 91.80 55.31
+TPA-KVonly 72.85 39.68 60.92 53.81 37.00 73.34 56.83 26.19 91.30 56.88
+TPA-KVonly ($R_{K,V}=4$) 72.98 40.27 60.15 54.88 36.80 73.29 56.43 25.50 92.10 56.93
+TPA-KVonly ($R_{K,V}=6$) 73.95 39.76 58.99 54.73 36.80 72.91 59.04 24.93 92.90 57.11
+TPA 71.76 39.16 61.25 53.74 37.80 72.80 55.49 23.86 90.70 56.28
+G.3 Ablation Studies on Learning Rates
+We implement a set of parallel experiments for medium models with learning rate $3\times 10^{-4}$, and the curves for training loss, validation loss, and validation perplexity are displayed in Figure 6.
+We also show the performance of these models on the benchmarks described in Section 4 in Tables 12-13.
+The results show that TPA and TPA-KVonly models can also outperform other types of attention with different learning rates.
+(a) Training Loss
+(b) Validation Loss
+(c) Validation Perplexity
+Figure 6: The training loss, validation loss, and validation perplexity of medium-size (353M) models (learning rate $3\times 10^{-4}$) and different attention mechanisms on the FineWeb-Edu 100B dataset.
+Table 12: The evaluation results of medium models (learning rate $3\times 10^{-4}$) with different attention mechanisms pretrained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
+The best scores in each column are bolded.
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+MHA 56.52 29.27 58.84 44.06 35.00 68.44 51.07 25.35 76.40 49.44
+MQA 55.68 28.24 60.86 44.17 35.20 68.66 52.72 25.14 72.90 49.29
+GQA 54.88 29.61 56.36 43.77 35.20 68.82 52.57 25.41 74.80 49.05
+MLA 59.64 29.78 60.73 45.17 34.20 68.66 52.80 25.34 75.70 50.22
+TPA-KVonly 57.11 30.03 61.25 44.83 34.60 69.04 54.54 23.35 74.60 49.93
+TPA 59.30 31.91 60.98 45.57 34.60 69.48 53.91 24.93 77.20 50.88
+Table 13: The evaluation results of medium models (learning rate $3\times 10^{-4}$) with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+The best scores in each column are bolded.
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+MHA 64.44 32.85 59.05 44.18 33.20 68.72 50.12 26.01 87.40 49.44
+MQA 64.27 32.94 57.71 44.36 31.80 68.01 51.70 25.99 86.00 49.29
+GQA 61.70 32.17 52.81 43.99 33.80 68.50 53.35 24.44 86.40 50.80
+MLA 65.95 31.48 50.98 44.99 32.20 68.93 51.93 25.89 88.80 51.24
+TPA-KVonly 65.99 33.70 57.49 44.47 34.20 69.53 53.28 24.23 86.50 49.93
+TPA 66.54 34.47 58.96 45.35 33.00 69.21 53.99 24.51 91.30 53.04

### v3

@@
 Hence, by constraining TPA’s head-dimension factors to be constant masks (one for MQA; multiple for GQA), these popular variants are recovered as special cases.
-3.5 Computational Cost.
+3.5 Computational Cost
 For a detailed analysis of the computational cost of TPA, please refer to Appendix A, which shows that the training and inference flops of TPA with optimized implementation (without materializing $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$) are smaller than MHA, GQA, and MLA.
@@
 We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$.
-For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
+For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $6\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $3\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
 Training & Validation Curves.
 Figures 2 and 3 compare training and validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B.
-Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see F) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses.
+Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix F) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses.
 For instance, in Figure 3(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages.
@@
 Consider a single head $i$. Each query vector $\mathbf{Q}^{(i)}_{t}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{q}$):
-$\mathbf{Q}^{(i)}_{t}~{}=~{}\sum_{r=1}^{R_{q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
+$\mathbf{Q}^{(i)}_{t}~{}=~{}\frac{1}{R_{q}}\sum_{r=1}^{R_{q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
 and each key vector $\mathbf{K}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{k}$):
-$\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{s=1}^{R_{k}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
+$\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{k}}\sum_{s=1}^{R_{k}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
 Their dot-product for tokens $t,\tau$ is
-$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (A.1)
+$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}\frac{1}{R_{q}R_{k}}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (A.1)
 A.2 Multi-Head Case
@@
 Single-Head Complexity. A single attention head of dimension $d_{h}$. For each query:
-$\mathbf{Q}^{(i)}_{t}~{}=~{}\sum_{r=1}^{R_{q}}a_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
+$\mathbf{Q}^{(i)}_{t}~{}=~{}\frac{1}{R_{q}}\sum_{r=1}^{R_{q}}a_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
 and for each key:
-$\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{s=1}^{R_{k}}a_{k,i}^{(s)}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
+$\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{k}}\sum_{s=1}^{R_{k}}a_{k,i}^{(s)}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
 Their dot product:
-$\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}~{}=~{}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}\Bigl{[}a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})\Bigr{]}~{}\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$
+$\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{q}R_{k}}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}\Bigl{[}a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})\Bigr{]}~{}\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$
 For each pair $(r,s)$, we pay:
@@
 $\operatorname{head}(t)~{}=~{}\sum_{\tau=1}^{T}\alpha_{t,\tau}~{}\mathbf{V}_{\tau}.$
-Using the factorization $\mathbf{V}_{\tau}=\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$, we write:
-$\operatorname{head}(t)~{}=~{}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
+Using the factorization $\mathbf{V}_{\tau}=\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$, we write:
+$\operatorname{head}(t)~{}=~{}\frac{1}{R_{v}}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
 Rearrange sums:
-$\operatorname{head}(t)~{}=~{}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
+$\operatorname{head}(t)~{}=~{}\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
 We still do not explicitly form $\mathbf{V}_{\tau}$. Instead:
@@
 Stage 2: Weighted Summation by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$. For each token $t$, the final attention head output is
-$\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})~{}=~{}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
+$\frac{1}{R_{v}}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})~{}=~{}\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
 We still never explicitly materialize $\mathbf{V}$.
@@
 MHA, MQA, and GQA. Despite sharing or grouping keys/values in MQA and GQA, the decoding cost for MHA, MQA, and GQA remains of the same order. Specifically, for each head $i$, we compute:
-$\mathbf{Q}_{i}(\mathbf{x}_{T})~{}\in~{}\mathbb{R}^{d_{h}},\quad\mathbf{K}_{i}~{}\in~{}\mathbb{R}^{T\times d_{h}},\quad\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}~{}\in~{}\mathbb{R}^{1\times T},\quad\text{and}\quad\operatorname{Softmax}\bigl{(}\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}\bigr{)}\,\mathbf{V}_{i}~{}\in~{}\mathbb{R}^{d_{h}}.$
+$\mathbf{Q}_{i}(\mathbf{x}_{T})~{}\in~{}\mathbb{R}^{d_{h}},\quad\mathbf{K}_{i}~{}\in~{}\mathbb{R}^{T\times d_{h}},\quad\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}~{}\in~{}\mathbb{R}^{1\times T},$
+$\quad\text{and}\quad\operatorname{Softmax}\bigl{(}\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}\bigr{)}\,\mathbf{V}_{i}~{}\in~{}\mathbb{R}^{d_{h}}.$
 Hence, the flops scale linearly in $h$, $d_{h}$, and $T$. For example, forming $\mathbf{Q}_{i}(\mathbf{x}_{T})\mathbf{K}_{i}^{\top}$ for each head $i$ costs roughly $\mathcal{O}(h\,d_{h}\,T)$.
@@
 G.2.1 Perplexity Curves
-We display the perplexity curves for medium, large and XL size of models in Figure 4.
+We display the perplexity curves for medium, large, and XL size models in Figure 4.
 (a) Validation Perplexity
@@
 Figure 5 shows the training loss, validation loss, and validation perplexity curves of XL-size (1.5B) T6 models with different ranks trained on the FineWeb-Edu 100B dataset, and the evaluation results are displayed in Table 7.
-It can be observed that increase in rank can improve the performances of large language models.
+It can be observed that an increase in rank can improve the performance of large language models.
 (a) Training Loss

### v4

@@
 Scaling language models to handle longer input sequences typically necessitates large key-value (KV) caches, resulting in substantial memory overhead during inference.
-In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, significantly shrinking KV cache size at inference time.
-By factorizing these representations into contextual low-rank components (contextual factorization) and seamlessly integrating with RoPE, TPA achieves improved model quality alongside memory efficiency.
+In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, substantially shrinking the KV cache size at inference time.
+By factorizing these representations into contextual low-rank components and seamlessly integrating with Rotary Position Embedding (RoPE), TPA achieves improved model quality alongside memory efficiency.
 Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling.
-Through extensive empirical evaluation of language modeling tasks, we demonstrate that T6 exceeds the performance of standard Transformer baselines including MHA, MQA, GQA, and MLA across various metrics, including perplexity and a range of renowned evaluation benchmarks.
-Notably, TPA’s memory efficiency enables the processing of significantly longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models.
+Through extensive empirical evaluation on language modeling tasks, we demonstrate that T6 surpasses or matches the performance of standard Transformer baselines, including Multi-Head Attention (MHA), Multi-Query Attention (MQA), Grouped-Query Attention (GQA), and Multi-Head Latent Attention (MLA) across various metrics, including perplexity and a range of established evaluation benchmarks.
+Notably, TPA’s memory efficiency and computational efficiency at the decoding stage enable processing longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models.
 The code is available at https://github.com/tensorgi/T6.
 1 Introduction
-Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Bubeck et al., 2023).
-As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completions.
-However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (Zhang et al., 2023c; Liu et al., 2024c).
+Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (brown2020language, ; chowdhery2022palm, ; touvron2023llama, ; bubeck2023sparks, ).
+As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completion.
+However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (zhang2023h2o, ; liu2024kivi, ).
 Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
-Figure 1: Tensor Product Attention (TPA) in the Tensor ProducT ATTenTion Transformer (T6).
-Different from multi-head attention, in each layer, firstly the hidden state goes through different linear layers to get the latent factor matrices $\mathbf{A}_{(\cdot)}$’s and $\mathbf{B}_{(\cdot)}$’s for query, key, and value.
-We additionally apply RoPE to $\mathbf{B}_{Q}$ and $\mathbf{B}_{K}$ for query and key.
-Then the multi-head query, key, and value vectors are attained by the tensor product of $\mathbf{A}_{(\cdot)}$ and $\mathbf{B}_{(\cdot)}$.
-Finally, the output of TPA is produced by scaled dot-product attention followed by linear projection of concatenated results of multiple heads.
+Figure 1: Tensor Product Attention (TPA) within the Tensor ProducT ATTenTion Transformer (T6).
+In each TPA layer, the input hidden state $\mathbf{x}_{t}$ is processed by linear layers to produce latent factor matrices for query (e.g., $\mathbf{A}_{Q}(\mathbf{x}_{t}),\mathbf{B}_{Q}(\mathbf{x}_{t})$), key (e.g., $\mathbf{A}_{K}(\mathbf{x}_{t}),\mathbf{B}_{K}(\mathbf{x}_{t})$), and value (e.g., $\mathbf{A}_{V}(\mathbf{x}_{t}),\mathbf{B}_{V}(\mathbf{x}_{t})$).
+Rotary Position Embedding (RoPE) is applied to the $\mathbf{B}_{Q}(\mathbf{x}_{t})$ and $\mathbf{B}_{K}(\mathbf{x}_{t})$ factors.
+The query, key, and value tensors for each attention head are then formed by the tensor product of these factor matrices (e.g., $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$).
+Finally, the TPA output is computed using scaled dot-product attention, followed by a linear projection of the concatenated results from all heads.
 A variety of solutions have been explored to address this memory bottleneck.
-Some approaches compress or selectively prune cached states through sparse attention patterns (Child et al., 2019) or token eviction strategies (Zhang et al., 2023c; Xiao et al., 2024; Ribar et al., 2024), though such methods risk discarding tokens that may later prove important.
-Other work proposes off-chip storage of key-value states (He & Zhai, 2024), at the expense of increased I/O latency.
-Attention variants like multi-query attention (MQA) (Shazeer, 2019) and grouped-query attention (GQA) (Ainslie et al., 2023) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications.
-Meanwhile, low-rank weight factorization methods such as LoRA (Hu et al., 2022) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates runtime.
-The recently introduced Multi-head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but needs additional position-encoded parameters per head due to incompatibility with Rotary Position Embedding (RoPE) efficiently (Su et al., 2024).
-In order to overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), as illustrated in Figure 1, a novel architecture that uses higher-order tensors to factorize queries (Q), keys (K), and values (V) during attention computation.
-By dynamically factorizing activations rather than static weights (e.g., LoRA), TPA constructs low-rank, contextual representations that substantially reduce KV cache memory usage with improved representational capacity.
-In practice, TPA can reduce the memory overhead by an order of magnitude compared to standard multi-head attention (MHA) with lower pretraining validation loss (perplexity) and improved downstream performance.
-A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023) and Gemma (Team et al., 2024).
-Our primary contributions are summarized as follows:
-  • We propose Tensor Product Attention (TPA), A mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor-decompositions to achieve 10×10$\times$10 × or more reduction in inference-time KV cache size relative to standard attention mechanism (Vaswani et al., 2017) with improved performance compared to previous methods such as MHA, MQA, GQA, and MLA.
-In addition, we unify existing attention mechanisms by revealing that MHA, MQA, and GQA all arise naturally as non-contextual variants of TPA.
+Some approaches compress or selectively prune cached states through sparse attention patterns (child2019generating, ) or token eviction strategies (zhang2023h2o, ; xiao2023efficient, ; ribar2023sparq, ), though such methods risk discarding tokens that may later prove important.
+Other work proposes off-chip storage of key-value states (he2024fastdecode, ), at the expense of increased I/O latency.
+Attention variants like Multi-Query Attention (MQA) (shazeer2019fast, ) and Grouped-Query Attention (GQA) (ainslie2023gqa, ) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications.
+Meanwhile, low-rank weight factorization methods such as LoRA (hu2021lora, ) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime.
+The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (liu2024deepseek, ) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (su2024roformer, ) integration, necessitating additional position-encoded parameters per head.
+To overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), illustrated in Figure 1.
+TPA is a novel attention mechanism that employs tensor factorizations for queries (Q), keys (K), and values (V).
+By dynamically factorizing activations rather than static weights (as in LoRA), TPA constructs low-rank, contextual representations.
+This approach substantially reduces KV cache memory usage while offering improved representational capacity.
+In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance.
+A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer, ), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama, ) and Gemma (team2024gemma, ).
+Our main contributions are summarized as follows:
+  1. 1.
+    We propose Tensor Product Attention (TPA), a mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor decompositions.
+This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (vaswani2017attention, ), MHA, MQA, GQA, and MLA, while also improving performance.
+In addition, we analyze existing attention mechanisms and reveal that MHA, MQA, and GQA can be expressed as non-contextual variants of TPA.
+  2. 2.
     We introduce the Tensor ProducT ATTenTion Transformer (T6), a new TPA-based model architecture for sequence modeling.
-On language modeling experiments, T6 consistently improves validation perplexity and downstream evaluation performance with reduced KV cache size.
-  • We show TPA integrates seamlessly with RoPE (Su et al., 2024), facilitating easy adoption in popular foundation model architectures such as LLaMA and Gemma.
+In language modeling experiments, T6 consistently improves or matches validation perplexity and downstream evaluation performance, all while maintaining a reduced KV cache size.
+  3. 3.
+    We demonstrate that TPA integrates seamlessly with RoPE (su2024roformer, ), facilitating its easy adoption in popular foundation model architectures like LLaMA and Gemma.
+  4. 4.
+    We develop FlashTPA Decoding, an efficient autoregressive inference algorithm for TPA.
+Our empirical results show that FlashTPA Decoding can be faster than optimized MHA, MQA, GQA, and MLA decoding methods, particularly for long sequences.
 2 Background
-In this section, we review two classical forms of attention: Scaled Dot-Product Attention, and Multi-Head Attention (MHA) (Vaswani et al., 2017).
-More types of attention are detailed in the Appendix E, including Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped Query Attention (GQA) (Ainslie et al., 2023), as well as a recent method called Multi-head Latent Attention (MLA) used in DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b).
-We also introduce Rotary Position Embedding (RoPE, Su et al.
-(2024)), which is commonly used in recent works of large language models.
+In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention, ) and introduce key notations.
+Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast, ), Grouped Query Attention (GQA) (ainslie2023gqa, ), Multi-head Latent Attention (MLA) (liu2024deepseek, ; liu2024deepseekv3, ), and Rotary Position Embedding (RoPE) (su2024roformer, ) are further discussed in Appendix H.
 Notations.
@@
 In multi-head attention, each head is associated with its own set of $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}$, and each has dimension $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$, where $d_{k}$ is typically set to $d_{h}$, the dimension of each head.
-[^5] Often, one sets $h\times d_{h}=d_{\text{model}}$, so each head has query/key/value dimension $d_{h}$.
+[^5] Often, $h\times d_{h}=d_{\text{model}}$, so each head has query/key/value dimension $d_{h}$.
 Similarly, we have an output projection matrix $\bm{W}^{O}\in\mathbb{R}^{(h\cdot d_{h})\times d_{\text{model}}}$.
-For methods like MQA and GQA, some of these are shared or partially shared across heads, but their shapes remain consistent.
-We define the tensor product of two vectors as follows: for vectors $\mathbf{a}\in\mathbb{R}^{m},\mathbf{b}\in\mathbb{R}^{n}$, the tensor product of $\mathbf{a}$ and $\mathbf{b}$ is:
-$\mathbf{a}\otimes\mathbf{b}=\mathbf{C}\in\mathbb{R}^{m\times n},\text{with}~{}C_{ij}=a_{i}b_{j},$
-where $a_{i}$ and $b_{j}$ are the $i$-th and $j$-th elements of $\mathbf{a}$ and $\mathbf{b}$ respectively, and $C_{ij}$ is the $(i,j)$-th entry of $\mathbf{C}$.
-We also define the vectorization of a matrix $\mathbf{C}\in\mathbb{R}^{m\times n}$ by:
-$\text{vec}(\mathbf{C})=\mathbf{d}\in\mathbb{R}^{mn},\text{with}~{}d_{i\cdot n+j}=C_{ij},$
-where $d_{i\cdot n+j}$ is the $(i\cdot n+j)$-th element of $\mathbf{d}$.
+For methods like MQA and GQA, some of these projection matrices are shared or partially shared across heads, but their shapes remain consistent.
+We define the tensor product of two vectors as follows: for vectors $\mathbf{a}\in\mathbb{R}^{m},\mathbf{b}\in\mathbb{R}^{n}$, the tensor product of $\mathbf{a}$ and $\mathbf{b}$ is: $\mathbf{a}\otimes\mathbf{b}=\mathbf{C}\in\mathbb{R}^{m\times n},\text{with}~{}C_{ij}=a_{i}b_{j}$, where $a_{i}$ is the $i$-th element of $\mathbf{a}$, $b_{j}$ is the $j$-th element of $\mathbf{b}$, and $C_{ij}$ is the $(i,j)$-th entry of $\mathbf{C}$.
+The vectorization of a matrix $\mathbf{C}\in\mathbb{R}^{m\times n}$, denoted $\text{vec}(\mathbf{C})\in\mathbb{R}^{mn}$, stacks the columns of $\mathbf{C}$ into a single column vector.
+For example, if $\mathbf{C}=[\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{n}]$ where $\mathbf{c}_{j}$ are columns, then $\text{vec}(\mathbf{C})=[\mathbf{c}_{1}^{\top},\mathbf{c}_{2}^{\top},\ldots,\mathbf{c}_{n}^{\top}]^{\top}$.
 2.1 Scaled Dot-Product Attention
-Scaled dot-product attention (Vaswani et al., 2017) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$).
+Scaled dot-product attention (vaswani2017attention, ) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$).
 It produces a weighted combination of the values ($\mathbf{V}$).
@@
 where each of $\mathbf{Q},\mathbf{K},\mathbf{V}$ is an $(n\times d_{k})$ matrix for $n$ tokens and key dimension $d_{k}$.
-The division by $\sqrt{d_{k}}$ stabilizes training by controlling the scale of the inner products.
 2.2 Multi-Head Attention (MHA)
-Multi-Head Attention (MHA) extends scaled dot-product attention by dividing the model’s internal representation into several heads.
-Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information.
+Multi-Head Attention (MHA) (vaswani2017attention, ) extends scaled dot-product attention by dividing the model’s internal representation into several heads.
+Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information from different representational subspaces.
 For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
-$\mathbf{Q}_{t,i}$ $=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
-$\mathbf{K}_{t,i}$ $=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
-$\mathbf{V}_{t,i}$ $=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
+$\mathbf{Q}_{t,i}=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},~{}\mathbf{K}_{t,i}$ $=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},~{}\mathbf{V}_{t,i}=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
 $\textbf{head}_{i}$ $=\operatorname{Attention}\Bigl{(}\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\Bigr{)},$
-where $\bm{W}_{i}^{Q},\bm{W}_{i}^{K},\bm{W}_{i}^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ are learnable projection matrices for the $i$-th head, $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$.
-After computing each head’s attention, the outputs are concatenated and mapped back to the original dimension via another matrix $\bm{W}^{O}\in\mathbb{R}^{hd_{h}\times d_{\text{model}}}$:
-$\operatorname{MHA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\,\bm{W}^{O}.$
-MHA can capture a rich set of dependencies while each head focuses on different subspaces.
-2.3 Rotary Position Embedding (RoPE)
-Many recent LLMs use rotary position embedding (RoPE; Su et al., 2024) to encode positional information in the query/key vectors.
-Specifically, let $\mathrm{RoPE}_{t}$ denote the rotation operator $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ corresponding to the $t$-th position.
-$\mathbf{T}_{t}$ is a block-diagonal matrix, which consists of block-diagonal matrix$\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$, $j\in\{1,\cdots,d_{h}/2\}$, where $\{\theta_{j}\}$ are pre-defined frequency parameters, e.g., $\theta_{j}=1/10000^{2j/d_{h}}$.
-Then we define
-$\operatorname{RoPE}\left(\mathbf{Q}_{t}\right)\triangleq\mathbf{Q}_{t}\mathbf{T}_{t},\quad\text{where }\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}.$
-A fundamental property is that
-$\mathbf{T}_{t}\,\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (2.1)
-which ensures that relative positions $(t-s)$ are preserved, thereby providing a form of translation invariance in the rotary position embedding.
+where $\bm{W}_{i}^{Q},\bm{W}_{i}^{K},\bm{W}_{i}^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ are learnable projection matrices for the $i$-th head, and $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ are the query, key, and value matrices for the $i$-th head over $T$ tokens.
+After computing each head’s attention output, the results are concatenated and mapped back to the model’s original dimension via another learnable linear projection matrix $\bm{W}^{O}\in\mathbb{R}^{hd_{h}\times d_{\text{model}}}$:
+$\operatorname{MHA}(\mathbf{X})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\,\bm{W}^{O}.$
+MHA enables the model to capture a rich set of dependencies by allowing each head to focus on different aspects of the input sequence.
+We also discuss how MHA relates to TPA in Appendix A.
+Further discussion on Transformers and attention mechanisms, KV cache optimization techniques, and low-rank factorization methods is available in Appendix G.
 3 Tensor Product Attention
-In this section, we provide a detailed description of our proposed Tensor Product Attention (TPA), which allows contextual low-rank factorization for queries, keys, and values.
-First, we explain how TPA factorizes queries, keys, and values with explicit tensor shapes.
-Next, we describe how TPA can be integrated into the multi-head attention framework and how it reduces memory consumption in KV caching at inference time.
-Finally, we show how RoPE can seamlessly integrate with TPA (including a pre-rotated variant).
+In this section, we provide a detailed description of our proposed Tensor Product Attention (TPA), which enables contextual low-rank factorization for queries, keys, and values.
+First, we explain how TPA factorizes these components, specifying tensor shapes.
+Next, we describe TPA’s integration into the multi-head attention framework and its benefits for reducing KV cache memory consumption during inference.
+Finally, we demonstrate RoPE’s seamless integration with TPA, including a pre-rotated variant for efficiency.
 3.1 Tensor Factorization of Queries, Keys, and Values
@@
 Standard attention projects the entire sequence into three tensors, $\mathbf{Q},\;\mathbf{K},\;\mathbf{V}\;\in\mathbb{R}^{T\times h\times d_{h}}$, where $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the slices for the $t$-th token.
-Contextual Factorization (CF).
+Contextual Factorization.
 Instead of forming each head’s query, key, or value via a single linear map, TPA factorizes each $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$ into a sum of (contextual) tensor products whose ranks are $R_{q}$, $R_{k}$, and $R_{v}$, respectively and may differ.
 Specifically, for each token $t$, with a small abuse of notation, we define:
-$\mathbf{Q}_{t}$ $=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),$ (3.1)
-$\mathbf{K}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$ (3.2)
-$\mathbf{V}_{t}$ $=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$ (3.3)
+[…]
+$\hskip 77.49976pt\mathbf{V}_{t}=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$ (3.1)
 where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$.
-Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ adds up to form the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$.
-Similarly, analogous definitions apply to key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
+Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ (an outer product) contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$.
+Analogous definitions apply to the key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
 Latent Factor Maps. Each factor in the tensor product depends on the token’s hidden state $\mathbf{x}_{t}$. For example, for queries, we can write:
 $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}_{r}\,\mathbf{x}_{t}\in\mathbb{R}^{h},\quad\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}_{r}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
-and similarly for keys and values.
+where $\bm{W}^{a^{Q}}_{r}\in\mathbb{R}^{h\times d_{\text{model}}}$ and $\bm{W}^{b^{Q}}_{r}\in\mathbb{R}^{d_{h}\times d_{\text{model}}}$ are learnable weight matrices.
+Similar linear maps produce the factors for keys and values.
 One often merges the rank index into a single output dimension. For instance, for queries:
 $\mathbf{a}^{Q}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot h},~{}\mathbf{b}^{Q}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot d_{h}},$
-which are then reshaped into $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{q}\times d_{h}}$.
-Summing over $R_{q}$ and scaled by $\frac{1}{R_{q}}$ yields
+which are then reshaped into $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times d_{h}}$ (where each row of $\mathbf{A}_{Q}(\mathbf{x}_{t})$ corresponds to an $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})^{\top}$ and each row of $\mathbf{B}_{Q}(\mathbf{x}_{t})$ to a $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})^{\top}$).
+The query tensor for token $t$ can then be expressed as:
 $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h\times d_{h}}.$
+This operation is equivalent to $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})(\mathbf{b}_{r}^{Q}(\mathbf{x}_{t}))^{\top}$, where $\mathbf{a}_{r}^{Q}$ is the $r$-th column of $\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}$ and $(\mathbf{b}_{r}^{Q})^{\top}$ is the $r$-th row of $\mathbf{B}_{Q}(\mathbf{x}_{t})$.
 Repeating for all tokens reconstitutes $\mathbf{Q}\in\mathbb{R}^{T\times h\times d_{h}}$.
-Similar procedures can be applied to obtain $\mathbf{K}$ and $\mathbf{V}$ with ranks $R_{k}$ and $R_{v}$, respectively.
+Similar procedures are applied to obtain $\mathbf{K}$ and $\mathbf{V}$ with ranks $R_{K}$ and $R_{V}$, respectively.
 Scaled Dot-Product Attention. Once $\mathbf{Q},\mathbf{K},\mathbf{V}$ are factorized, multi-head attention proceeds as in standard Transformers. For each head $i\in\{1,\dots,h\}$:
@@
 $\operatorname{TPA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\bm{W}^{O}.$ (3.3)
-Parameter Initialization.
-We initialize the weight matrices $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, $\bm{W}_{r}^{b^{V}}$ using Xavier initialization (Glorot & Bengio, 2010).
-Specifically, each entry of the weight matrix is drawn from a uniform distribution with bounds $[-\sqrt{6/(n_{\text{in}}+n_{\text{out}})},\sqrt{6/(n_{\text{in}}+n_{\text{out}})}]$, where $n_{\text{in}}$ and $n_{\text{out}}$ are the input and output dimensions of the respective weight matrices.
-This initialization strategy helps maintain the variance of activations and gradients across the network.
+Parameter Initialization. We use Xavier initialization (glorot2010understanding, ) for the factor weight matrices; details are in Appendix I.
 3.2 RoPE Compatibility and Acceleration
 In a typical workflow of adding RoPE to standard multi-head attention, one first computes $\mathbf{Q}_{t},\mathbf{K}_{s}\in\mathbb{R}^{h\times d_{h}}$ of the $t$-th token and $s$-th token and then applies:
-$\mathbf{Q}_{t}\mapsto\widetilde{\mathbf{Q}_{t}}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t}),\qquad\mathbf{K}_{s}\mapsto\widetilde{\mathbf{K}_{s}}=\operatorname{RoPE}_{s}(\mathbf{K}_{s}).$
+$\mathbf{Q}_{t}\mapsto\widetilde{\mathbf{Q}_{t}}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t}),\qquad\mathbf{K}_{s}\mapsto\widetilde{\mathbf{K}_{s}}=\operatorname{RoPE}_{s}(\mathbf{K}_{s}).$ (3.4)
 Direct Integration. A useful optimization is to integrate RoPE directly into the TPA factorization. For example, one can pre-rotate the token-dimension factors:
@@
 yielding a pre-rotated key representation:
-$\widetilde{\mathbf{K}}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(s)}(\mathbf{x}_{t})\bigr{)}$
-$=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
-Thus, each $\mathbf{K}_{t}$ is already rotated before caching, removing the need for explicit rotation at the decoding time and accelerating autoregressive inference.
-Depending on hardware and performance requirements, one can also adopt different RoPE integration approaches for training and inference.
-Theorem 1 (RoPE’s Compatibility with TPA).
+$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(r)}(\mathbf{x}_{t})\bigr{)}=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
+Here, $\operatorname{RoPE}_{t}$ is applied to each row of $\mathbf{B}_{K}(\mathbf{x}_{t})$ (i.e., to each $\mathbf{b}^{K}_{(r)}(\mathbf{x}_{t})$ vector).
+Thus, each $\mathbf{K}_{t}$ is effectively rotated before caching.
+This removes the need for explicit rotation at decoding time, accelerating autoregressive inference.
+Depending on hardware and performance requirements, different RoPE integration strategies can be adopted for training and inference.
+Theorem 3.1 (RoPE’s Compatibility with TPA).
 Let $\mathbf{Q}_{t}$ be factorized by TPA as
@@
 $\operatorname{RoPE}({\mathbf{Q}}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),$ (3.6)
-where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}$.
-In addition, assume $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ are factorized by TPA and then rotated by $\operatorname{RoPE}_{t},\operatorname{RoPE}_{s}$.
-Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$.
-Then we have
+where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}$ (RoPE applied row-wise to $\mathbf{B}_{Q}(\mathbf{x}_{t})$).
+Furthermore, let $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ be factorized by TPA.
+Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$ be their RoPE-transformed versions.
+The relative positional encoding property of RoPE is preserved:
 $\operatorname{RoPE}_{t-s}(\mathbf{Q}_{t})\mathbf{K}_{s}^{\top}=\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top},$
-Focusing on individual heads $i$, the above matrix equality implies:
-$\operatorname{RoPE}_{t-s}\bigl{(}\mathbf{q}_{t,i}\bigr{)}^{\top}\mathbf{k}_{s,i}=\widetilde{\mathbf{q}}_{t,i}^{\top}\,\widetilde{\mathbf{k}}_{s,i}.$
-where $\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}}$ is the $i$-th query head of $t$-th token, and $\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}$ is the $j$-th key head of $s$-th token, and
-$\widetilde{\mathbf{q}}_{t,i}$ $=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}}$
-$\quad\widetilde{\mathbf{k}}_{s,i}$ $=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
-Theorem 1 indicates that TPA does not break RoPE’s relative translational property.
-We prove Theorem 1 in Appendix C.1.
-In short, $\operatorname{RoPE}_{t}$ acts as a block-diagonal orthogonal transform (i.e., a matrix $\mathbf{T}_{t}$) on $\mathbf{B}_{Q}(\mathbf{x}_{t})$.
-Consequently, $\mathbf{A}_{Q}(\mathbf{x}_{t})$ remains unchanged, while each column of $\mathbf{B}_{Q}(\mathbf{x}_{t})$ is rotated appropriately, preserving the TPA structure.
+where $\operatorname{RoPE}_{t-s}$ denotes applying RoPE with relative position $t-s$.
+Focusing on individual heads $i$, if $\mathbf{q}_{t,i}$ and $\mathbf{k}_{s,i}$ are the $i$-th head vectors from $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ respectively (as column vectors of dimension $d_{h}$), and $\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i},t)$ and $\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i},s)$, then the dot product for attention scores satisfies:
+$\widetilde{\mathbf{q}}_{t,i}^{\top}\,\widetilde{\mathbf{k}}_{s,i}=\mathbf{q}_{t,i}^{\top}\operatorname{RoPE}(\cdot,t-s)\mathbf{k}_{s,i}.$
+Theorem 3.1 indicates that TPA does not break RoPE’s relative translational property.
+We prove Theorem 3.1 in Appendix F.1.
+In essence, $\operatorname{RoPE}_{t}$ applies a linear transformation (rotation matrix $\mathbf{R}_{t}$) to the $d_{h}$-dimensional space.
+Since $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$, applying RoPE yields $\mathbf{Q}_{t}\mathbf{R}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}(\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{R}_{t})$.
+Thus, RoPE effectively transforms $\mathbf{B}_{Q}(\mathbf{x}_{t})$ to $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{R}_{t}$, where $\mathbf{R}_{t}$ acts on each row of $\mathbf{B}_{Q}(\mathbf{x}_{t})$.
+The $\mathbf{A}_{Q}(\mathbf{x}_{t})$ factor remains unchanged, preserving the TPA structure.
 3.3 KV Caching and Memory Reduction
@@
 This accumulates to $\mathbb{R}^{T\times h\times d_{h}}$ for keys and $\mathbb{R}^{T\times h\times d_{h}}$ for values, i.e., $2\,T\,h\,d_{h}$ total.
-TPA Factorized KV Caching. Instead of storing the full $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$, TPA stores only their factorized ranks. Specifically, we keep
+TPA Factorized KV Caching. Instead of storing the full $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$, TPA stores only their factor components. Specifically, for each past token $t$, we cache:
 $\mathbf{A}_{K}(\mathbf{x}_{t}),\,\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\quad\text{and}\quad\mathbf{A}_{V}(\mathbf{x}_{t}),\,\mathbf{B}_{V}(\mathbf{x}_{t}),$
-where $\mathbf{A}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times h},\;\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times d_{h}},\;\mathbf{A}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times h},\;\mathbf{B}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times d_{h}}$.
-Hence, the memory cost per token is
-$\underbrace{R_{K}(h+d_{h})}_{\text{for K}}\;+\;\underbrace{R_{V}(h+d_{h})}_{\text{for V}}=(\,R_{K}+R_{V}\,)\,\bigl{(}h+d_{h}\bigr{)}.$
-Compared to the standard caching cost of $2\,h\,d_{h}$, the ratio is:
-$\frac{\,(R_{K}+R_{V})\,(h+d_{h})\,}{2\,h\,d_{h}\,}.$
-For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll h$ (e.g., rank $1$ or $2$) often yields 10×10$\times$10 × or more reduction.
+where $\mathbf{A}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times h},\;\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\in\mathbb{R}^{R_{K}\times d_{h}}(\text{pre-rotated}),\;\mathbf{A}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times h},\;\mathbf{B}_{V}(\mathbf{x}_{t})\in\mathbb{R}^{R_{V}\times d_{h}}$.
+Hence, the memory cost per token is $\underbrace{R_{K}(h+d_{h})}_{\text{for K}}\;+\;\underbrace{R_{V}(h+d_{h})}_{\text{for V}}=(\,R_{K}+R_{V}\,)\,\bigl{(}h+d_{h}\bigr{)}$.
+Compared to the standard caching cost of $2\,h\,d_{h}$, the ratio is $\frac{\,(R_{K}+R_{V})\,(h+d_{h})\,}{2\,h\,d_{h}\,}$.
+For large $h$ and $d_{h}$ (typically $d_{h}=64$ or $128$), setting $R_{K},R_{V}\ll h$ (e.g., rank $1$ or $2$) often yields substantial reduction of KV cache size.
+Table 1 provides a comparative overview of different attention mechanisms, including TPA and its variants, focusing on KV cache size per token and the number of parameters in an attention layer.
 Table 1: Comparison of different attention mechanisms.
 Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively.
-Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Section F.
+Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Appendix I.
 For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@
 TPA (Non-contextual B) $(R_{K}+R_{V})h$ $(R_{Q}+R_{K}+R_{V})(d_{\text{model}}h+d_{h})+d_{\text{model}}\,hd_{h}$ $h$ $h$
-3.4 Unifying MHA, MQA, and GQA as Non-contextual TPA
-3.4.1 MHA as Non-contextual TPA
-Standard multi-head attention (MHA) can be viewed as a specific instance of TPA in which: 1) the rank is set equal to the number of heads; 2) the head dimension factor is non-contextual (i.e., independent of the $t$-th token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$); 3) the token dimension factor is a linear function of $\mathbf{x}_{t}$.
-To match MHA with TPA, let $R_{Q}=R_{K}=R_{V}=h$. Focusing on $\mathbf{Q}_{t}$:
-  1. (a)
-    Non-contextual head factors. Define
-    $\mathbf{a}^{Q}_{i}=R_{Q}\mathbf{e}_{i}\in\mathbb{R}^{h},$ (3.8)
-    where $\mathbf{e}_{i}\in\mathbb{R}^{h}$ is the $i$-th standard basis vector, so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
-  2. (b)
-    Contextual token factors. Define
-    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$ (3.9)
-    where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the per-head query projection defined before, hence $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$ dependent on $\mathbf{x}_{t}$.
-Substituting (3.8)–(3.9) into (3.1) gives:
-$\mathbf{Q}_{t}=\sum_{i=1}^{h}\Big{[}\mathbf{e}_{i}\otimes\bigl{(}(\bm{W}^{Q}_{i})^{\top}\,\mathbf{x}_{t}\bigr{)}\Big{]}\in\mathbb{R}^{h\times d_{h}}.$ (3.10)
-Each term $\mathbf{e}_{i}\otimes\bigl{(}({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\bigr{)}$ in (3.10) contributes only to the $i$-th row, reconstituting the usual MHA form of $\mathbf{Q}_{t}$.
-Analogous constructions hold for $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$ using $\bm{W}^{K}_{i},\bm{W}^{V}_{i}$.
-Thus, MHA is a non-contextual, full-rank variant of TPA.
-TPA with Non-contextual A.
-More broadly, TPA can use non-contextual head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}\in\mathbb{R}^{h}$ (i.e., independent of $\mathbf{x}_{t}$), while allowing $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ to remain context-dependent.
-Then, for keys:
-$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}\otimes\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$
-and similarly for queries/values. This reduces per-token computations and can be effective when head-dimension relationships are relatively stable across all tokens.
-MQA and GQA as Non-Contextual TPA.
-Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped Query Attention (GQA) (Ainslie et al., 2023) [^6] The original definitions of MQA and GQA are presented in Appendix E.1 and E.2, respectively.
-also emerge naturally from TPA by restricting the head-dimension factors to be non-contextual and low-rank:
-  • MQA as Rank-1 TPA. In MQA, all heads share a single set of keys/values, corresponding to $R_{K}=R_{V}=1$ along the head dimension. Concretely,
-    $\mathbf{K}_{t}$ $=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),$
-    $\mathbf{V}_{t}$ $=(1,\ldots,1)^{\top}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
-    forces every head to use the same $\mathbf{K}_{t},\mathbf{V}_{t}$. Each head retains a distinct query projection, matching the MQA design.
-  • GQA as Grouped Rank-1 TPA.
-GQA partitions $h$ heads into $G$ groups, each sharing keys/values within that group.
-In TPA form, each group $g$ has a dedicated non-contextual factor pair $\mathbf{a}^{K}_{g},\mathbf{a}^{V}_{g}\in\mathbb{R}^{h}$, which acts as a “mask” for the heads in that group.
-Varying $G$ from $1$ to $h$ interpolates from MQA to standard MHA.
-Hence, by constraining TPA’s head-dimension factors to be constant masks (one for MQA; multiple for GQA), these popular variants are recovered as special cases.
-3.5 Computational Cost
-For a detailed analysis of the computational cost of TPA, please refer to Appendix A, which shows that the training and inference flops of TPA with optimized implementation (without materializing $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$) are smaller than MHA, GQA, and MLA.
-Specifically, when we set $R_{q}=6$, $R_{k}=R_{v}=2$ (our default setting), TPA is 10×10$\times$10 × or more faster on calculating $\mathbf{Q}\mathbf{K}^{\top}$ than MLA during inference (see Appendix A.8).
+3.4 Expressing MHA, MQA, and GQA as Non-contextual TPA
+We demonstrate in Appendix A that several existing attention mechanisms, including MHA, MQA, and GQA, can be elegantly expressed as non-contextual variants of TPA by imposing specific constraints on the TPA factors.
 3.5 Model Architectures
 We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention).
-Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version.
-The feed-forward network (FFN) adopts a SwiGLU layer, as in (Shazeer, 2020; Touvron et al., 2023).
-TPA QKV Factorization.
-Let each token’s hidden-state vector be $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, and we follow Section 3.1 to project the entire sequence into three tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\;\mathbb{R}^{T\times h\times d_{h}}$, where $\mathbf{Q}_{t},\;\mathbf{K}_{t},\;\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the slices for the $t$-th token.
-The factor components $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ are produced by linear transformations on $\mathbf{x}_{t}$.
-For instance, letting $\bm{W}^{a^{Q}}_{r}\in\mathbb{R}^{h\times d_{\text{model}}}$ and $\bm{W}^{b^{Q}}_{r}\in\mathbb{R}^{d_{h}\times d_{\text{model}}}$, we have:
-$\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}_{r}\,\mathbf{x}_{t},\quad\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}_{r}\,\mathbf{x}_{t}.$
-In practice, we merge all ranks $r$ into a single dimension of the output, reshape, and sum over rank indices; see Section 3.1 for details. The factorization for K and V follows the same pattern.
+Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (touvron2023llama, ) while changing the self-attention block to our TPA-based version.
+The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu, ; touvron2023llama, ).
+The specific mechanisms for TPA-based query, key, and value factorization, as well as the attention computation within T6, directly follow the comprehensive descriptions provided in Section 3 (particularly Section 3.1 for factorization, and Equations 3.2 and 3.3 for attention computation).
+A brief recap of these steps within the T6 model context is provided in Appendix B.
 Rotary Positional Embedding (RoPE).
-As discussed in Section 3.2, RoPE (Su et al., 2024) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
-Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.6) and Theorem 1.
-Attention Step and Output Projection.
-Once we have $\mathbf{Q},\mathbf{K},\mathbf{V}$ factorized per token with RoPE applied on $\mathbf{Q}$ and $\mathbf{K}$, the attention step proceeds for each head $i\in\{1,\dots,h\}$ using (3.4).
-Finally, concatenating these $h$ heads and then projecting them back using an output weight matrix gives the final attention result, as shown in (3.5).
+As discussed in Section 3.2, RoPE (su2024roformer, ) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
+Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.5) and Theorem 3.1.
 SwiGLU Feed-Forward Network.
-Following Shazeer (2020); Touvron et al.
-(2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters.
+Following shazeer2020glu ; touvron2023llama , our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl{[}\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr{]}\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters.
 Note that other activation functions can also be used.
@@
 We place norm layers (e.g., RMSNorm) before each sub-layer. Stacking $L$ such blocks yields a T6 model architecture with $L$ layers.
+4 FlashTPA Decoding Algorithm
+$\mathbf{B}_{Q}$
+$(R_{Q},D)$
+$\bm{b}^{K}_{\text{cache}}$
+$(M,D)$
+$\sum_{D}$
+$S^{(1)}$
+$(M,R_{Q})$
+$\mathbf{A}_{Q}$
+$(H,R_{Q})$
+$\sum_{R_{Q}}$
+$S^{(2)}$
+$(M,H)$
+$\mathbf{a}^{K}_{\text{cache}}$
+$(M,H)$
+$\odot$
+$\mathcal{L}$
+$(H,M)$
+Softmax
+$\bm{\alpha}$
+$(H,M)$
+$\mathbf{a}^{V}_{\text{cache}}$
+$(M,H)$
+$\odot$
+$\mathbf{O}^{(A)}$
+$(M,H)$
+$\bm{b}^{V}_{\text{cache}}$
+$(M,E)$
+$\sum_{M}$
+$\mathbf{O}$
+$(H,E)$
+Figure 2: Data flow diagram for FlashTPA Decoding.
+Rectangles represent tensors (blue for inputs, yellow for intermediates, red for final output), circles with $\sum$ or $\odot$ denote Einstein summation contractions or element-wise products respectively, and the green rounded rectangle is the softmax operation.
+Shapes are shown for a single query ($N=1$) interacting with $M$ cached items.
+$H$ is the number of heads, $R_{Q}$ is the query rank, and $D,E$ are respective feature dimensions for the $\mathbf{B}_{Q}/\bm{b}^{K}_{\text{cache}}$ and $\bm{b}^{V}_{\text{cache}}$ factors.
+Scaling factors in softmax are omitted for visual clarity.
+For efficient autoregressive inference with Tensor Product Attention (TPA), we introduce FlashTPA Decoding.
+This algorithm is optimized for generating one token at a time by leveraging the factorized representation of queries, keys, and values.
+The core idea, illustrated in Figure 2, is to perform attention computations using a sequence of Einstein summations (“einsum”) that operate directly on these factorized components.
+This avoids materializing the full query, key, and value tensors, which is particularly beneficial as the Key-Value (KV) cache grows with sequence length.
+The detailed definitions of the input factorized components and the step-by-step pseudocode for FlashTPA Decoding are provided in Algorithm 1 (see Appendix D for details).
+An optimized Triton kernel implementation: Algorithm 3 is also outlined in Appendix D.
+This sequence of factorized operations allows FlashTPA Decoding to compute the attention output efficiently.
+It minimizes memory allocations for large intermediate tensors and reduces the overall computational load compared to materializing full query, key, and value tensors, especially for long sequences.
+Consequently, TPA is not only memory-efficient due to its smaller KV cache footprint but can also be computationally efficient during inference.
+The experimental results for FlashTPA decoding time are presented in Section 5.2.
+Algorithm 1 FlashTPA Decoding Algorithm
+1:$\mathbf{A}_{Q}$: Query A tensor $(B,1,H,R_{Q})$
+2:$\mathbf{B}_{Q}$: Query B tensor $(B,1,R_{Q},D)$
+3:$\mathbf{a}^{K}_{\text{cache}}$: Cached Key A tensor $(B,M,H)$
+4:$\bm{b}^{K}_{\text{cache}}$: Cached Key B tensor $(B,M,D)$
+5:$\mathbf{a}^{V}_{\text{cache}}$: Cached Value A tensor $(B,M,H)$
+6:$\bm{b}^{V}_{\text{cache}}$: Cached Value B tensor $(B,M,E)$
+7:$s_{\text{total}},s_{Q},s_{K},s_{V}$: Scaling factors (with defaults $s_{\text{total}}\leftarrow 1/\sqrt{D}$, $s_{Q}\leftarrow 1/R_{Q}$)
+8:$\mathbf{O}$: Output tensor $(B,1,H,E)$
+9$\triangleright$ Step 1: Interaction between $\mathbf{B}_{Q}$ and $\bm{b}^{K}_{\text{cache}}$ components
+10:$S^{(1)}\leftarrow\text{einsum}(\text{``bnrd, bmd}\rightarrow\text{bnmr''},\mathbf{B}_{Q},\bm{b}^{K}_{\text{cache}})$triangleright Shape: $(B,1,M,R_{Q})$.
+$S^{(1)}_{b,n,m,r}=\sum_{d}(\mathbf{B}_{Q})_{b,n,r,d}\cdot(\bm{b}^{K}_{\text{cache}})_{b,m,d}$
+11$\triangleright$ Step 2: Incorporate $\mathbf{A}_{Q}$ component
+12:$S^{(2)}\leftarrow\text{einsum}(\text{``bnhr, bnmr}\rightarrow\text{bnmh''},\mathbf{A}_{Q},S^{(1)})$triangleright Shape: $(B,1,M,H)$.
+$S^{(2)}_{b,n,m,h}=\sum_{r}(\mathbf{A}_{Q})_{b,n,h,r}\cdot S^{(1)}_{b,n,m,r}$
+13$\triangleright$ Step 3: Incorporate $\mathbf{a}^{K}_{\text{cache}}$ component to get full logits
+14:einsum(“bnmh, $\mathcal{L}\leftarrow\text{einsum}(\text{``bnmh, bmh}\rightarrow\text{bhnm''},S^{(2)},\mathbf{a}^{K}_{\text{cache}})$triangleright Shape: $(B,H,1,M)$.
+$\mathcal{L}_{b,h,n,m}=S^{(2)}_{b,n,m,h}\cdot(\mathbf{a}^{K}_{\text{cache}})_{b,m,h}$
+15$\triangleright$ Step 4: Apply scaling and Softmax
+16:$\bm{\alpha}\leftarrow\operatorname{Softmax}_{M}(\mathcal{L}\cdot s_{Q}\cdot s_{K}\cdot s_{\text{total}})$triangleright Shape: $(B,H,1,M)$. Softmax over cache dimension $M$
+17$\triangleright$ Step 5: Compute weighted sum with $\mathbf{a}^{V}_{\text{cache}}$ component
+18:$\mathbf{O}^{(A)}\leftarrow\text{einsum}(\text{``bhnm, bmh}\rightarrow\text{bnmh''},\bm{\alpha},\mathbf{a}^{V}_{\text{cache}})$triangleright Shape: $(B,1,M,H)$.
+$\mathbf{O}^{(A)}_{b,n,m,h}=\bm{\alpha}_{b,h,n,m}\cdot(\mathbf{a}^{V}_{\text{cache}})_{b,m,h}$
+19$\triangleright$ Step 6: Incorporate $\bm{b}^{V}_{\text{cache}}$ component and final scaling
+[…]
+21:return $\mathbf{O}$
 5 Experiments
 5.1 Language Modeling Tasks
-All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy, 2022), using the FineWeb-Edu 100B dataset (Lozhkov et al., 2024).
+All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy2022, ), using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu, ).
 The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation.
-We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al., 2017) with Multi-Query Attention (MQA; Shazeer, 2019), Grouped Query Attention (GQA; Ainslie et al., 2023), or Multi-head Latent Attention (MLA; Liu et al., 2024a).
+We compare T6 against the baseline Llama architecture (touvron2023llama, ) with SwiGLU activation (shazeer2020glu, ) and RoPE embeddings (su2024roformer, ), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention ) with Multi-Query Attention (MQA; shazeer2019fast ), Grouped Query Attention (GQA; ainslie2023gqa ), or Multi-head Latent Attention (MLA; liu2024deepseek ).
 In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer.
 We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B).
-Details on architecture hyperparameters and training hardware are shown in Appendix G.1.
-Training Setup.
-We follow the nanoGPT training configuration.
-In particular, we use the AdamW (Loshchilov, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$.
-We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler (Loshchilov & Hutter, 2016) with $2{,}000$ warmup steps and a (total) global batch size of $480$.
-For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $6\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $3\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
+Details on architecture hyperparameters and training hardware are shown in Appendix J.1.
 Training & Validation Curves.
-Figures 2 and 3 compare training and validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B.
-Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix F) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final losses.
+Figure 4 compares validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B.
+Training loss curves are provided in Appendix Figure 3.
+Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix I) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses.
 For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages.
-Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher losses.
+Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek, ) (blue curves) generally trains more slowly and yields higher validation losses.
 Validation Perplexity.
@@
 Downstream Evaluation.
-We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), using the lm-evaluation-harness codebase (Gao et al., 2024).
+We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC, ), BoolQ (BoolQ, ), HellaSwag (hellaswag, ), OBQA (OpenBookQA, ), PIQA (PIQA, ), WinoGrande (WinoGrande, ), and MMLU (MMLU, ), using the lm-evaluation-harness codebase (eval-harness, ).
 For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy.
 Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3.
-Zero-shot evaluation of small and XL models are displayed in Tables 6 and 7 in the appendix.
-Moreover, we also present 2-shot evaluation results in Tables 8910 and 11 in the appendix.
-For the medium-size (353M) models (Tables 2 and 9), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%.
+Zero-shot evaluation of small and XL models are displayed in Tables 12 and 13 in the appendix.
+Moreover, we also present 2-shot evaluation results in Tables 141516 and 17 in the appendix.
+For the medium-size (353M) models (Table 2 for 0-shot and Table 15 in appendix for 2-shot), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%.
 When given two-shot prompts, TPA again leads with 53.12% average accuracy.
-A similar trend appears for the large-size (773M) models (Tables 3), where TPA-KVonly attains the highest average (53.52% zero-shot).
-And for the XL size (1.5B) models (Table 7), TPA-KVonly attains the highest average (55.03% zero-shot).
+A similar trend appears for the large-size (773M) models (Table 3), where TPA-KVonly attains the highest average (53.52% zero-shot).
+For the XL size models (1.5B) (Table 13 in the appendix), TPA-KV only achieves the highest average (55.03% zero-shot).
 Our experiments confirm that TPA consistently matches or exceeds the performance of established attention mechanisms (MHA, MQA, GQA, MLA) across medium and large model scales.
-The fully factorized TPA excels on mid-scale models, while TPA-KVonly can rival or surpass it at larger scales.
-In both cases, factorizing the attention activations shrinks autoregressive KV cache requirements by up to 5×5$\times$5 ×10×10$\times$10 ×, thus enabling much longer context windows under fixed memory budgets.
-In summary, tensor product attention provides a flexible, memory-efficient alternative to standard multi-head attention, advancing the scalability of modern language models.
 (a) Medium models (353M)
@@
 TPA-KVonly 58.01 30.12 58.01 45.95 35.60 69.10 53.12 25.39 75.10 50.04
-TPA (non-ctx-A) 58.96 31.48 59.76 45.07 34.80 69.21 53.59 25.42 76.40 50.52
 TPA 58.38 31.57 59.39 46.83 37.00 70.02 54.06 25.52 79.90 51.41
@@
 TPA 63.22 35.58 60.03 51.26 36.80 71.44 55.56 24.77 79.60 53.10
-5 Related Work
-Transformers and Attention.
-As a sequence-to-sequence architecture Transformer (Vaswani et al., 2017) introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies.
-Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns (Child et al., 2019; Shi et al., 2023; Han et al., 2024; Liang et al., 2024a; Li et al., 2024; Liang et al., 2024b), kernel-based projections (Choromanski et al., 2021), and linearized transformers (Tsai et al., 2019; Katharopoulos et al., 2020; Schlag et al., 2021; Zhang et al., 2023b; Sun et al., 2023; Zhang et al., 2024).
-To decrease memory usage and circumvent the limitation of memory bandwidth in training, Shazeer (2019) proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head.
-To tackle with the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) (Ainslie et al., 2023) divides queries into several groups, and each group of queries shares a single key head and value head.
-Recently, DeepSeek-V2 (Liu et al., 2024a) applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value.
-Concurrently, Hu et al.
-(2024) proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q.
-Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
-KV Cache Optimization.
-During the inference time of Transformers, key and value tensors of the previous tokens are repeatedly computed due to their auto-regressive nature.
-To enhance efficiency, firstly proposed by Ott et al.
-(2019), these tensors can be cached in memory for future decoding, referred to as the KV cache.
-However, the KV cache requires additional memory usage and may add to more latencies due to the bandwidth limitation (Adnan et al., 2024).
-Therefore, previous studies have explored diverse approaches to mitigate these issues, including KV cache eviction to discard less significant tokens (Zhang et al., 2023c; Xiao et al., 2024; Cai et al., 2024; Adnan et al., 2024), dynamic sparse attention among selected keys and values (Ribar et al., 2024; Tang et al., 2024; Singhania et al., 2024), KV cache offloading to CPU (He & Zhai, 2024; Lee et al., 2024; Sun et al., 2024), as well as quantization of KV cache (Xiao et al., 2023; Liu et al., 2024c; Hooper et al., 2024).
-Different from the methods above, TPA reduces the size of the KV cache by using tensor-decomposed KV.
+5.2 Experimental Results on FlashTPA Decoding
+(a) Batch Size=1
+(b) Batch Size=2
+(c) Batch Size=4
+(d) Batch Size=8
+(e) Batch Size=16
+Figure 5: Decoding time comparison of different attention mechanisms with an embedding dimension of 2048 and $d_{h}=64$.
+The y-axis represents $\log_{2}(\text{time})$ in seconds, and the x-axis represents $\log_{2}(\text{sequence length})$.
+Each subfigure corresponds to a different batch size.
+This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms.
+We benchmark FlashTPA against FlashMHA (shah2024flashattention, ), FlashGQA, FlashMQA, and FlashMLA (flashmla2025, ).
+It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton, ).
+While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential.
+Development of a CUDA-based FlashTPA kernel is ongoing and is expected to yield further performance improvements.
+The evaluations were performed with batch sizes selected from $\{1,2,4,8,16\}$, model embedding dimensions ($d_{\text{model}}$) chosen from $\{1024,2048,3072\}$, and sequence lengths ranging from $2^{12}$ (4,096) to $2^{19}$ (524,288).
+For all experiments, the dimension per head ($d_{h}$) was fixed at 64.
+The ranks for TPA’s factorized components ($R_{Q},R_{K},R_{V}$) were set to $(16,1,1)$, and for GQA configurations, the number of key-value head groups was 4.
+The decoding time per token, measured as $\log_{2}(\text{time})$ in seconds, is plotted against
+$\log_{2}(\text{sequence length})$.
+Lower values on the y-axis indicate faster decoding times.
+Results are presented in Figure 5 for an embedding dimension of 2048 (corresponding to 32 attention heads).
+Additional results for embedding dimensions of 1024 (16 heads, Figure 8) and 3072 (48 heads, Figure 7) are provided in Appendix D.
+Figure 5 depicts these speed comparisons for an embedding dimension of 2048.
+The results indicate that FlashTPA (blue line) is highly competitive and often outperforms other attention mechanisms, especially as the sequence length increases.
+A detailed breakdown of these comparisons across different batch sizes for the $d_{\text{model}}=2048$ case (Figure 5) is provided in Appendix D.
+These findings underscore the computational efficiency of the FlashTPA decoding algorithm, particularly its favorable scaling with increasing sequence lengths.
 6 Conclusion
@@
 Overall, TPA offers a powerful mechanism for compressing KV storage while improving the model performance, thereby enabling longer sequence contexts under constrained memory.
-Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
-We now explore whether it is possible to compute attention scores $\mathbf{Q}\,\mathbf{K}^{\top}$ of Tensor Product Attention (TPA) directly from their factorized forms, thereby reducing floating-point operations.
-A.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$
-Consider a single head $i$. Each query vector $\mathbf{Q}^{(i)}_{t}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{q}$):
-$\mathbf{Q}^{(i)}_{t}~{}=~{}\frac{1}{R_{q}}\sum_{r=1}^{R_{q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
-and each key vector $\mathbf{K}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{k}$):
-$\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{k}}\sum_{s=1}^{R_{k}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
+Appendix A Expressing MHA, MQA, and GQA as Non-contextual TPA
+This appendix section details how standard Multi-Head Attention (MHA), Multi-Query Attention (MQA), and Grouped-Query Attention (GQA) can be expressed as special, non-contextual variants of Tensor Product Attention (TPA).
+A.1 MHA as Non-contextual TPA
+Standard multi-head attention (MHA) can be viewed as a specific instance of TPA in which: 1) the rank is set equal to the number of heads; 2) the head dimension factor is non-contextual (i.e., independent of the $t$-th token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$); 3) the token dimension factor is a linear function of $\mathbf{x}_{t}$.
+More precisely, if we set $R_{Q}=R_{K}=R_{V}=h$ (the number of heads) and define the factors appropriately, MHA emerges.
+To match MHA with TPA, let $R_{Q}=R_{K}=R_{V}=h$.
+Focusing on $\mathbf{Q}_{t}$:
+  1. (a)
+    Non-contextual head factors. Define
+    $\mathbf{a}^{Q}_{i}=h\mathbf{e}_{i}\in\mathbb{R}^{h},$ (A.1)
+    where $\mathbf{e}_{i}\in\mathbb{R}^{h}$ is the $i$-th standard basis vector, so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
+  2. (b)
+    Contextual token factors. Define
+    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$ (A.2)
+    where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the per-head query projection matrix, hence $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$ is dependent on $\mathbf{x}_{t}$.
+[…]
+$\mathbf{Q}_{t}=\sum_{i=1}^{h}\Big{[}\mathbf{e}_{i}\otimes\bigl{(}(\bm{W}^{Q}_{i})^{\top}\,\mathbf{x}_{t}\bigr{)}\Big{]}\in\mathbb{R}^{h\times d_{h}}.$ (A.3)
+[…]
+TPA with Non-contextual A factors.
+More broadly, TPA can use non-contextual head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}\in\mathbb{R}^{h}$ (i.e., independent of $\mathbf{x}_{t}$), while allowing the token-dimension factors $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ to remain context-dependent.
+Then, for keys (using the original $1/R_{K}$ scaling):
+$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}\otimes\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$
+and similarly for queries and values.
+This version, referred to as TPA (Non-contextual A), reduces per-token computations for the A factors and can be effective when head-dimension relationships are relatively stable across tokens.
+The KV cache for this variant would store $\mathbf{b}^{K}_{r}(\mathbf{x}_{t})$ and $\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ for each token, while $\mathbf{a}^{K}_{r}$ and $\mathbf{a}^{V}_{r}$ are fixed.
+A.2 MQA and GQA as Non-contextual TPA
+Multi-Query Attention (MQA) [45] and Grouped Query Attention (GQA) [2] (detailed in Appendix H.1 and H.2) also emerge naturally from TPA by restricting the head-dimension factors ($\mathbf{a}^{K},\mathbf{a}^{V}$) to be non-contextual and low-rank:
+  • MQA as Rank-1 TPA (for K and V).
+In MQA, all heads share a single set of keys and values.
+This corresponds to TPA with $R_{K}=R_{V}=1$, where the head-dimension factors for keys and values are non-contextual vectors of all ones (or any constant non-zero vector, as it’s a shared projection).
+For example:
+    $\mathbf{K}_{t}=\frac{1}{1}\mathbf{1}_{h}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=\frac{1}{1}\mathbf{1}_{h}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
+    where $\mathbf{1}_{h}\in\mathbb{R}^{h}$ is a vector of ones.
+This forces every head to use the same key and value vectors $\mathbf{b}^{K}(\mathbf{x}_{t})$ and $\mathbf{b}^{V}(\mathbf{x}_{t})$ (which are effectively $\mathbf{X}\bm{W}^{K}_{\text{shared}}$ and $\mathbf{X}\bm{W}^{V}_{\text{shared}}$ from MQA’s definition).
+Each head $i$ would still use its distinct query $\mathbf{Q}_{t,i}$, which in TPA could be modeled with $R_{Q}=h$ and $\mathbf{a}^{Q}_{i}=\mathbf{e}_{i}$ as in the MHA case if queries are not factorized, or a different TPA structure for queries.
+  • GQA as Grouped Rank-1 TPA (for K and V within groups).
+GQA partitions $h$ heads into $G$ groups, with each group sharing keys and values.
+In TPA form, this means for keys and values, the rank $R_{K}$ and $R_{V}$ would be $G$.
+For each group $j\in[G]$, there would be a non-contextual head-dimension factor $\mathbf{a}^{K}_{j}\in\mathbb{R}^{h}$ (and similarly $\mathbf{a}^{V}_{j}$) that acts as a "mask" or selector for the heads in that group (e.g., $\mathbf{a}^{K}_{j}$ has ones for heads in group $j$ and zeros elsewhere).
+Then:
+    $\mathbf{K}_{t}=\frac{1}{G}\sum_{j=1}^{G}\mathbf{a}^{K}_{j}\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=\frac{1}{G}\sum_{j=1}^{G}\mathbf{a}^{V}_{j}\otimes\mathbf{b}^{V}_{j}(\mathbf{x}_{t}),$
+    where $\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$ is the shared key vector for group $j$.
+Varying $G$ from $1$ (MQA) to $h$ (MHA, if $\mathbf{b}^{K}_{j}$ are distinct per original head) shows GQA as an interpolation.
+Hence, by constraining TPA’s head-dimension factors $\mathbf{a}^{K},\mathbf{a}^{V}$ to be specific constant vectors (or masks) and choosing appropriate ranks, these popular attention variants can be recovered as special cases of TPA with non-contextual A factors for keys and values.
+This underscores the flexibility and unifying nature of the TPA framework.
+Appendix B T6 Model Architecture Details
+This section provides a recap of how Tensor Product Attention (TPA) components are integrated within the Tensor ProducT ATTenTion Transformer (T6) architecture, building upon the detailed exposition in Section 3.
+TPA QKV Factorization Recap.
+As defined in Section 3.1, for each token’s hidden-state vector $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, TPA projects the input into three tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\;\mathbb{R}^{T\times h\times d_{h}}$.
+The slices for the $t$-th token, $\mathbf{Q}_{t},\;\mathbf{K}_{t},\;\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$, are formed by sums of tensor products.
+The factor components $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ are produced by linear transformations on $\mathbf{x}_{t}$.
+For instance, for queries, with learnable weight matrices $\bm{W}^{a^{Q}}_{r}\in\mathbb{R}^{h\times d_{\text{model}}}$ and $\bm{W}^{b^{Q}}_{r}\in\mathbb{R}^{d_{h}\times d_{\text{model}}}$, these factors are:
+$\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}_{r}\,\mathbf{x}_{t},\quad\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}_{r}\,\mathbf{x}_{t}.$
+In practice, these are often computed by merging ranks into a single linear layer followed by reshaping, as detailed in Section 3.1.
+The factorization for keys ($\mathbf{K}_{t}$) and values ($\mathbf{V}_{t}$) follows the same pattern using their respective ranks $R_{K},R_{V}$ and projection matrices.
+Attention Step and Output Projection Recap.
+Once the factorized $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors are obtained for all tokens (with RoPE applied to the appropriate factors of $\mathbf{Q}$ and $\mathbf{K}$ as per Section 3.2), the attention output for each head $i\in\{1,\dots,h\}$ is computed using the scaled dot-product attention formula (Equation 3.2 from the main paper):
+$\textbf{head}_{i}=\operatorname{Softmax}\Bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,\mathbf{Q}_{i}\,(\mathbf{K}_{i})^{\top}\Bigr{)}\;\mathbf{V}_{i}.$
+Finally, the outputs from all $h$ heads are concatenated and projected back to the model dimension using an output weight matrix $\bm{W}^{O}$, as shown in Equation 3.3 from the main paper:
+$\operatorname{TPA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\bm{W}^{O}.$
+This constitutes the output of the TPA sub-layer in a T6 block.
+Appendix C Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
+In TPA, the query, key, and value tensors ($\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$) are constructed as sums of $R_{Q},R_{K},R_{V}$ rank-one tensors, respectively, as detailed in Equation 3.1.
+These sums are typically normalized by scaling factors $s_{Q}=1/R_{Q}$, $s_{K}=1/R_{K}$, and $s_{V}=1/R_{V}$.
+This section explores the computational advantages of performing attention calculations directly on these factorized representations, including their scaling factors, without explicitly materializing the full $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors.
+This approach can potentially reduce floating-point operations (FLOPs) and memory bandwidth.
+C.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$
+Consider a single head $i$. Each query vector $\mathbf{Q}^{(i)}_{t}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{Q}$ and scale $s_{Q}=1/R_{Q}$):
+$\mathbf{Q}^{(i)}_{t}~{}=~{}s_{Q}\sum_{r=1}^{R_{Q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
+and each key vector $\mathbf{K}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{K}$ and scale $s_{K}=1/R_{K}$):
+$\mathbf{K}^{(i)}_{\tau}~{}=~{}s_{K}\sum_{s=1}^{R_{K}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
 Their dot-product for tokens $t,\tau$ is
-$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}\frac{1}{R_{q}R_{k}}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (A.1)
-A.2 Multi-Head Case
-For multi-head attention with $h$ heads, one repeats the factorization across all heads. The $\mathbf{b}_{q}^{(r)},\mathbf{b}_{k}^{(s)}$ vectors are shared across heads.
-A.3 Complexity Analysis
+$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}s_{Q}s_{K}\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (C.1)
+Similarly, the value vector $\mathbf{V}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{V}$ and scale $s_{V}=1/R_{V}$):
+$\mathbf{V}^{(i)}_{\tau}~{}=~{}s_{V}\sum_{u=1}^{R_{V}}a^{(u)}_{v,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
+C.2 Multi-Head Case
+For multi-head attention with $h$ heads, one repeats the factorization across all heads.
+The $\mathbf{b}_{q}^{(r)},\mathbf{b}_{k}^{(s)},\mathbf{b}_{v}^{(u)}$ vectors are shared across heads.
+The scaling factors $s_{Q},s_{K},s_{V}$ are applied per head as part of the definition.
+C.3 Complexity Analysis
 We compare the cost of standard multi-head attention versus TPA under two scenarios:
   1. 1.
-    Naïve: Materialize $\mathbf{Q}$ and $\mathbf{K}$ from factors, then perform the usual batched GEMM.
+    Naïve: Materialize $\mathbf{Q},\mathbf{K},\mathbf{V}$ from factors (including their scales), then perform the usual batched GEMM.
   2. 2.
-    Specialized: Attempt to compute $\mathbf{Q}\,\mathbf{K}^{\top}$ directly from the rank-$(R_{q},R_{k})$ factors without explicitly forming $\mathbf{Q},\mathbf{K}$.
+    Specialized: Attempt to compute $\mathbf{Q}\,\mathbf{K}^{\top}$ and the final attention output directly from the rank-$(R_{Q},R_{K},R_{V})$ factors and their scales, without explicitly forming full $\mathbf{Q},\mathbf{K},\mathbf{V}$.
 Standard Multi-Head Attention. For batch size $B$ and sequence length $T$:
@@
 TPA: Naïve Implementation.
-  • Constructing factors: $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}\,\times R_{q}(h+d_{h})+R_{k}(h+d_{h})+R_{v}(h+d_{h})\bigr{)}$.
-  • Materializing $\mathbf{Q},\mathbf{K}$: $\mathcal{O}\bigl{(}B\,T\,(R_{q}\,h\,d_{h}+R_{k}\,h\,d_{h})\bigr{)}$.
-  • Dot-product $\mathbf{Q}\,(\mathbf{K})^{\top}$: $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$.
-Typically $R_{q},R_{k},R_{v}\ll h$, so the overhead of constructing factors is small relative to $\mathcal{O}(T^{2}\,d_{\text{model}})$. Meanwhile, we still gain KV caching benefits.
+  • Constructing factors:   $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}\,\times(R_{Q}(h+d_{h})+R_{K}(h+d_{h})+R_{V}(h+d_{h}))\bigr{)}$.
+  • Materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$ (including scaling):   $\mathcal{O}\bigl{(}B\,T\,(R_{Q}\,h\,d_{h}+R_{K}\,h\,d_{h}+R_{V}\,h\,d_{h})\bigr{)}$.
+The scaling is a per-element multiplication.
+  • Dot-product $\mathbf{Q}\,(\mathbf{K})^{\top}$ and multiply by $\mathbf{V}$:   $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$ for scores, and $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$ for value aggregation.
+Typically $R_{Q},R_{K},R_{V}\ll h,d_{h}$, so the overhead of constructing factors and materializing is influenced by these ranks.
+The key benefit of TPA is in reducing KV cache size and potentially computation if specialized kernels are used.
 TPA: Specialized Implementation.
-If we bypass explicitly forming $\mathbf{Q},\mathbf{K}$, each dot product $\mathbf{Q}_{t}\cdot\mathbf{K}_{\tau}$ is a double sum over rank indices.
-Below we detail its complexity.
-A.4 Complexity Analysis for the Specialized Implementation
-Single-Head Complexity. A single attention head of dimension $d_{h}$. For each query:
-$\mathbf{Q}^{(i)}_{t}~{}=~{}\frac{1}{R_{q}}\sum_{r=1}^{R_{q}}a_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
-and for each key:
-$\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{k}}\sum_{s=1}^{R_{k}}a_{k,i}^{(s)}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
-Their dot product:
-$\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{q}R_{k}}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}\Bigl{[}a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})\Bigr{]}~{}\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$
+If we bypass explicitly forming $\mathbf{Q},\mathbf{K},\mathbf{V}$, the computation involves the sums over ranks and the scaling factors.
+Below, we detail its complexity.
+Algorithm 2 Specialized TPA Computation (Without Materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$)
+1:Factorized Query components: $\mathbf{A}_{Q}(T,H,R_{Q})$, $\mathbf{B}_{Q}(T,R_{Q},d_{h})$
+2:Factorized Key components: $\mathbf{A}_{K}(T,H,R_{K})$, $\mathbf{B}_{K}(T,R_{K},d_{h})$
+3:Factorized Value components: $\mathbf{A}_{V}(T,H,R_{V})$, $\mathbf{B}_{V}(T,R_{V},d_{v})$ (Note: $d_{v}$ is value head dim, often $d_{v}=d_{h}$)
+4:Sequence length $T$, Number of heads $H$, Ranks $R_{Q},R_{K},R_{V}$, Head feature dimension $d_{h}$
+5:Scaling factors: $s_{Q}=1/R_{Q},s_{K}=1/R_{K},s_{V}=1/R_{V}$ (or other pre-defined scales)
+6:Output tensor $\mathbf{O}(T,H,d_{v})$
+7:// Phase 1: Attention Score Computation
+8$\triangleright$ Step 1.1: Compute shared $\mathbf{B}$-factor dot products $P$
+9:$P\leftarrow\text{einsum}(\text{"tqrd, tkrd}\rightarrow\text{tqtkrqrk"},\mathbf{B}_{Q},\mathbf{B}_{K})$
+10$\triangleright$ Indices: $t_{q}$=query token, $t_{k}$=key token, $r_{q}$=query rank, $r_{k}$=key rank, $d$=feature dim
+11$\triangleright$ Step 1.2: Combine $P$ with $\mathbf{A}_{Q},\mathbf{A}_{K}$ and apply $s_{Q},s_{K}$ to form scaled logits $\mathcal{L}^{\prime}$
+12:$\mathcal{L}_{\text{unscaled}}\leftarrow\text{einsum}(\text{"tqhrq, tkhrk, tqtkrqrk}\rightarrow\text{tqtkh"},\mathbf{A}_{Q},\mathbf{A}_{K},P)$
+13:$\mathcal{L}^{\prime}\leftarrow s_{Q}\cdot s_{K}\cdot\mathcal{L}_{\text{unscaled}}$
+14$\triangleright$ Indices: $h$=head
+15$\triangleright$ Step 1.3: Apply attention scaling ($1/\sqrt{d_{h}}$) and Softmax to get probabilities $\bm{\alpha}$
+16:$\mathcal{L}^{\prime\prime}_{\text{scaled}}\leftarrow\mathcal{L}^{\prime}/\sqrt{d_{h}}$
+17:$\bm{\alpha}\leftarrow\operatorname{Softmax}_{T_{k}}(\mathcal{L}^{\prime\prime}_{\text{scaled}})$triangleright Softmax over key tokens $T_{k}$ for each query $T_{q}$ and head $H$
+18$\triangleright$ Output $\bm{\alpha}(t_{q},t_{k},h)$
+19:
+20:// Phase 2: Value Aggregation
+21$\triangleright$ Step 2.1: Conceptually weight $\mathbf{A}_{V}$ by $\bm{\alpha}$ to form $W_{\mathbf{A}_{V}}$
+22:$W_{\mathbf{A}_{V}}\leftarrow\text{einsum}(\text{"tqtkh, tkhrv}\rightarrow\text{tqtkhrv"},\bm{\alpha},\mathbf{A}_{V})$
+23$\triangleright$ Indices: $r_{v}$=value rank
+24$\triangleright$ Step 2.2: Aggregate $W_{\mathbf{A}_{V}}$ with $\mathbf{B}_{V}$ and apply $s_{V}$ to get final output $\mathbf{O}$
+25:$\mathbf{O}_{\text{unscaled}}\leftarrow\text{einsum}(\text{"tqtkhrv, tkrvd}\rightarrow\text{tqhd"},W_{\mathbf{A}_{V}},\mathbf{B}_{V})$triangleright $d$ is value feature dim $d_{v}$
+26:$\mathbf{O}\leftarrow s_{V}\cdot\mathbf{O}_{\text{unscaled}}$
+27:return $\mathbf{O}$
+C.4 Complexity Analysis for the Specialized Implementation
+Complexity for a Single Query Token (Autoregressive Decoding).
+The dot product $\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}$ from Eq.
+(C.1) includes the $s_{Q}s_{K}$ scaling.
 For each pair $(r,s)$, we pay:
   1. 1.
-    $\mathcal{O}(1)$ for multiplying two scalars,
+    $\mathcal{O}(1)$ for multiplying $a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})$,
   2. 2.
-    $\mathcal{O}(d_{h})$ for the dot product $\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t})\cdot\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})$.
-Since $(r,s)$ runs over $R_{q}\times R_{k}$, each token-pair $(t,\tau)$ costs roughly
-$\mathcal{O}\Bigl{(}R_{q}\,R_{k}\,\bigl{(}1+d_{h}\bigr{)}\Bigr{)}~{}\approx~{}\mathcal{O}(R_{q}\,R_{k}\,d_{h}).$
-For $T$ queries and $T$ keys, that is $\mathcal{O}(T^{2}\,R_{q}\,R_{k}\,d_{h})$ for a single head.
-Multi-Head and Batches (Reusing $\mathbf{b}$-Dot Products).
-When extending to $h$ heads, each head $i$ has its own scalar factors $\mathbf{a}_{q,i}^{(r)}(\mathbf{x}_{t})$ and $\mathbf{a}_{k,i}^{(s)}(\mathbf{x}_{\tau})$, but the $\mathbf{b}$-vectors $\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t})$ and $\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})$ can still be shared across all heads (assuming the same rank-$R$ factors for every head).
-Hence, one can split the total cost into two stages:
+    $\mathcal{O}(d_{h})$ for the dot product $\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\rangle$.
+The $s_{Q}s_{K}$ scaling is a single multiplication after the sums.
+Since $(r,s)$ runs over $R_{Q}\times R_{K}$, each token-pair $(t,\tau)$ costs roughly $\mathcal{O}\Bigl{(}R_{Q}\,R_{K}\,\bigl{(}1+d_{h}\bigr{)}\Bigr{)}\approx\mathcal{O}(R_{Q}\,R_{K}\,d_{h})$.
+For a single query token $t$ and $M$ cached key tokens $\tau=1\dots M$, this is $\mathcal{O}(M\,R_{Q}\,R_{K}\,d_{h})$ for a single head to compute all unscaled dot products.
+The scaling by $s_{Q}s_{K}$ is $\mathcal{O}(M)$ for the score vector.
+Multi-Head and Batches (Full Sequence Processing - reusing $\mathbf{b}$-Dot Products as in Algorithm 2). The $\mathbf{b}$-vectors can be shared across heads. Let $T$ be the sequence length.
   1. 1.
-    $\mathbf{b}$-Dot-Product Stage:
-    For each token pair $(t,\tau)$ and each rank pair $(r,s)$, compute the dot product
-    $\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),~{}\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}~{}\in~{}\mathbb{R}.$
-    Since each dot product is $\mathcal{O}(d_{h})$ and there are $R_{q}R_{k}$ rank pairs as well as $T^{2}$ token pairs, this stage costs:
-    $\mathcal{O}\bigl{(}T^{2}\,R_{q}R_{k}\,d_{h}\bigr{)}.$
-    Crucially, these $\mathbf{b}$-dot products need only be computed once and can be cached for reuse by all heads.
+    $\mathbf{b}$-Dot-Product Stage ($P$):
+    Compute $P(t_{q},\tau_{k},r_{q},r_{k})=\langle\mathbf{b}_{q}^{(r_{q})}(\mathbf{x}_{t_{q}}),\mathbf{b}_{k}^{(r_{k})}(\mathbf{x}_{\tau_{k}})\rangle$.
+Cost: $\mathcal{O}\bigl{(}T^{2}\,R_{Q}R_{K}\,d_{h}\bigr{)}$.
+This is shared across heads.
   2. 2.
-    Per-Head Scalar Multiplications:
-    After the $\mathbf{b}$-dot products are precomputed (and cached), each head $i$ only needs to multiply each stored dot product by the corresponding scalars $\mathbf{a}_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{a}_{k,i}^{(s)}(\mathbf{x}_{\tau})$.
-Since this scalar multiplication is $\mathcal{O}(1)$ per pair, and there are $T^{2}$ token pairs and $R_{q}R_{k}$ rank pairs for each of the $h$ heads, this step costs:
-    $\mathcal{O}\bigl{(}h\,T^{2}\,R_{q}R_{k}\bigr{)}.$
-Putting these together, for batch size $B$, the total cost is
-$\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}R_{k}\,d_{h}\bigr{)}~{}+~{}\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}R_{k}\bigr{)}~{}=~{}\mathcal{O}\Bigl{(}B\,T^{2}\,R_{q}R_{k}\bigl{(}d_{h}+h\bigr{)}\Bigr{)}.$
-By contrast, the standard multi-head attention dot-product step is $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$. Hence, for the specialized TPA approach to reduce flops,
-$R_{q}R_{k}\,(d_{h}+h)~{}\leq~{}h\,d_{h}.$
-Thus a practical guideline is to ensure $\,R_{q}R_{k}<h\,\tfrac{d_{h}}{d_{h}+h}\,$. When that holds, bypassing explicit materialization of $\mathbf{Q}$ and $\mathbf{K}$ can be beneficial.
-A.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
-We have explored a two-step procedure for computing $\mathbf{Q}\,\mathbf{K}^{\top}$ directly from factorized queries and keys without materializing $\mathbf{Q}$ or $\mathbf{K}$.
-Here, we extend this idea to also avoid explicitly forming $\mathbf{V}$.
-That is, all three activations $\mathbf{Q},\mathbf{K},\mathbf{V}$ remain factorized throughout the attention pipeline.
-We present a single-head formulation below, and then discuss multi-head and batch extensions.
-Extending the Two-Step Approach to Avoid $\mathbf{V}$ Materialization.
-After we obtain $\mathbf{Q}\mathbf{K}^{\top}$, we apply $\alpha_{t,\tau}=\operatorname{softmax}\bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,(QK^{\top})_{t,\tau}\bigr{)}$.
-The final attention output at token $t$ (single head) is
-$\operatorname{head}(t)~{}=~{}\sum_{\tau=1}^{T}\alpha_{t,\tau}~{}\mathbf{V}_{\tau}.$
-Using the factorization $\mathbf{V}_{\tau}=\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$, we write:
-$\operatorname{head}(t)~{}=~{}\frac{1}{R_{v}}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
-Rearrange sums:
-$\operatorname{head}(t)~{}=~{}\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
-We still do not explicitly form $\mathbf{V}_{\tau}$. Instead:
-Stage 1: Calculating $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for all tokens.
-We simply observe that each output $\operatorname{head}(t)$ can be computed by summing vectors $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\in\mathbb{R}^{d_{h}}$ weighted by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$.
-The complexity for constructing $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ $\forall u,\tau$ is $\mathcal{O}(T\,R_{v}\,d_{h})$.
-Stage 2: Weighted Summation by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$. For each token $t$, the final attention head output is
-$\frac{1}{R_{v}}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})~{}=~{}\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
-We still never explicitly materialize $\mathbf{V}$.
-Instead, for each pair $(t,u)$, we must accumulate the sum of $T$ vectors $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\in\mathbb{R}^{d_{h}}$, each scaled by the scalar $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$.
-Because each vector is $d_{h}$-dimensional, each $(t,u)$ summation costs $\mathcal{O}(T\,d_{h})$.
-Summed over $t=1\ldots T$ and $u=1\ldots R_{v},$ the total work is $\mathcal{O}(T^{2}\,R_{v}\,d_{h})$ for the entire sequence.
-In practice, one precomputes all $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for $\tau=1\ldots N$, so each accumulation can be implemented as a simple “scalar‐times‐vector add” in a tight loop.
-This cost is usually smaller than the $\mathbf{Q}\mathbf{K}^{\top}$ factorized cost if $R_{v}\ll h$.
-A.6 Overall Complexity for Single-Head
-Combining the four bullet‐point stages from above (ignoring smaller overheads like the softmax) yields:
+    Per-Head Score Computation:
+    For each head $h$, compute $\mathcal{L}_{t_{q},\tau_{k},h}=s_{Q}s_{K}\sum_{r_{q},r_{k}}a_{q,h}^{(r_{q})}(\mathbf{x}_{t_{q}})\,a_{k,h}^{(r_{k})}(\mathbf{x}_{\tau_{k}})\,P(t_{q},\tau_{k},r_{q},r_{k})$.
+    Cost: $\mathcal{O}\bigl{(}H\,T^{2}\,R_{Q}R_{K}\bigr{)}$.
+Putting these together for batch size $B$, the total cost for scores is
+$\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}R_{K}\,d_{h}\bigr{)}+\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}R_{K}\bigr{)}=\mathcal{O}\Bigl{(}B\,T^{2}\,R_{Q}R_{K}\bigl{(}d_{h}+H\bigr{)}\Bigr{)}$.
+This is for computing the scaled logits before softmax.
+The standard attention dot-product step is $\mathcal{O}\bigl{(}B\,T^{2}\,H\,d_{h}\bigr{)}$.
+For the specialized TPA to reduce flops in score computation: $R_{Q}R_{K}\,(d_{h}+H)<H\,d_{h}$.
+This implies $R_{Q}R_{K}<H/(1+H/d_{h})$.
+If $H\approx d_{h}$, then $R_{Q}R_{K}\approx H/2$.
+C.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
+We extend the idea to also avoid explicitly forming $\mathbf{V}$.
+After obtaining scaled logits $\mathcal{L}^{\prime}_{t_{q},\tau_{k}}=s_{Q}s_{K}(\mathbf{Q}_{t_{q}}\mathbf{K}_{\tau_{k}}^{\top})$ (per head), we apply $\alpha_{t_{q},\tau_{k}}=\operatorname{softmax}_{\tau_{k}}\bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,\mathcal{L}^{\prime}_{t_{q},\tau_{k}}\bigr{)}$.
+The final attention output at token $t_{q}$ (single head) using $\mathbf{V}_{\tau_{k}}=s_{V}\sum_{u=1}^{R_{V}}a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})$ is:
+$\operatorname{head}(t_{q})~{}=~{}\sum_{\tau_{k}=1}^{T}\alpha_{t_{q},\tau_{k}}~{}\mathbf{V}_{\tau_{k}}~{}=~{}s_{V}\sum_{\tau_{k}=1}^{T}\alpha_{t_{q},\tau_{k}}\sum_{u=1}^{R_{V}}a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}}).$
+Rearranging sums:
+$\operatorname{head}(t_{q})~{}=~{}s_{V}\sum_{u=1}^{R_{V}}\left[\sum_{\tau_{k}=1}^{T}\bigl{(}\alpha_{t_{q},\tau_{k}}\,a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\right].$
+The computation of $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})$ from $\mathbf{x}_{\tau_{k}}$ is $\mathcal{O}(T\,R_{V}\,d_{\text{model}}d_{v})$ if contextual, or loaded if fixed (where $d_{v}$ is the dimension of $\mathbf{b}_{v}$).
+Assuming $\mathbf{b}_{v}^{(u)}$ are available, the weighted summation for each $(t_{q},u)$ pair (inner sum over $\tau_{k}$) costs $\mathcal{O}(T\,d_{v})$.
+Summed over $t_{q}=1\ldots T$ and $u=1\ldots R_{V}$, this stage is $\mathcal{O}(T^{2}\,R_{V}\,d_{v})$.
+The final scaling by $s_{V}$ is $\mathcal{O}(Td_{v})$.
+C.6 Overall Complexity for Single-Head (Specialized)
+The dominant costs for a single head, assuming factors $\mathbf{A},\mathbf{B}$ are already computed from $\mathbf{X}$:
   1. (i)
-    QK $\mathbf{b}$-Dot Product Stage: $\mathcal{O}(T^{2}\,R_{q}\,R_{k}\,d_{h})$.
+    QK $\mathbf{B}$-Dot Product Stage ($P$): $\mathcal{O}(T^{2}\,R_{Q}\,R_{K}\,d_{h})$ (shared calculation if multi-head).
   2. (ii)
-    QK Scalar-Multiply Stage: $\mathcal{O}(T^{2}\,R_{q}\,R_{k})$.
+    QK $\mathbf{A}$-Factor Combination and Scaling Stage ($\mathcal{L}^{\prime}$): $\mathcal{O}(T^{2}\,R_{Q}\,R_{K})$. (Scaling by $s_{Q}s_{K}$ is part of this per-element operation on scores).
   3. (iii)
-    Computing $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for all tokens: $\mathcal{O}(T\,R_{v}\,d_{h})$.
-  4. (iv)
-    Weighted Summation by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$: $\mathcal{O}(T^{2}\,R_{v}\,d_{h})$.
-Hence, for a single head, the total cost is:
-$\mathcal{O}\Bigl{(}T^{2}\,R_{q}\,R_{k}\,d_{h}~{}+~{}T^{2}\,R_{q}\,R_{k}~{}+~{}T\,R_{v}\,d_{h}~{}+~{}T^{2}\,R_{v}\,d_{h}\Bigr{)}.$
-In many cases (especially for large $T$), the $\mathcal{O}(T^{2})$ terms dominate, so one often focuses on
-$\mathcal{O}\Bigl{(}T^{2}\,R_{q}\,R_{k}\,d_{h}~{}+~{}T^{2}\,R_{q}\,R_{k}~{}+~{}T^{2}\,R_{v}\,d_{h}\Bigr{)}.$
-A.7 Multi-Head and Batch Extensions (Reuse of $\mathbf{b}$-Dot Products)
-When extending to $h$ heads and batch size $B$, all sequence‐length‐dependent terms are multiplied by $\sim B\,h$. However, crucial $\mathbf{b}$-dot products can be shared across heads:
-QK $\mathbf{b}$-Dot Products. Since each head has distinct scalar factors $a_{q,i},\,a_{k,i}$ but the same $\mathbf{b}_{q}^{(r)},\,\mathbf{b}_{k}^{(s)}$ across heads, each pairwise dot product
-$\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\rangle$
-is computed just once per batch. That cost remains
-$\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}\,R_{k}\,d_{h}\bigr{)},$
-not multiplied by $h$.
-After caching these dot products, each of the $h$ heads pays $\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}\,R_{k}\bigr{)}$ total for the head‐specific scalar multiplications (the “$\alpha_{t,\tau}$”–like factors).
-V $\mathbf{b}$-Evaluations.
-Likewise, the $\mathbf{b}_{v}^{(u)}$ factors are shared across heads (i.e.
-one set of $\mathbf{b}_{v}$-vectors for all heads).
-Hence, computing all $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for $\tau=1\ldots T$ (across the batch) is a one‐time cost:
-$\mathcal{O}\bigl{(}B\,T\,R_{v}\,d_{h}\bigr{)}.$
-Then each head $i$ has its own scalar factors $a_{v,i}^{(u)}(\mathbf{x}_{\tau})$, so the final accumulation $\sum_{\tau=1}^{T}\alpha_{t,\tau}\,a_{v,i}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ costs $\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{v}\,d_{h}\bigr{)}$ in total (for all $t,u$).
-Putting it all together, the total flops for multi‐head attention with batch size $B$ are:
-$\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}\,R_{k}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{QK $\mathbf{b}$-dot products}\\ \text{(shared across heads)}\end{subarray}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}\,R_{k}\bigr{)}}_{\text{per-head QK scalar mult.}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T\,R_{v}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{Compute $\mathbf{b}_{v}$ for all tokens}\\ \text{(shared across heads)}\end{subarray}}~{}+~{}\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{v}\,d_{h}\bigr{)}}_{\begin{subarray}{c}\text{final accumulations}\\ \text{(per head)}\end{subarray}}.$
+    Value Aggregation Stage (including $s_{V}$ scaling): $\mathcal{O}(T^{2}\,R_{V}\,d_{v})$.
+Total for single head, assuming $P$ is computed once: $\mathcal{O}\Bigl{(}T^{2}\,R_{Q}\,R_{K}\,d_{h}+T^{2}\,R_{Q}\,R_{K}+T^{2}\,R_{V}\,d_{v}\Bigr{)}$.
+C.7 Multi-Head and Batch Extensions (Reuse of $\mathbf{B}$-Dot Products)
+QK $\mathbf{B}$-Dot Products ($P$): Shared across heads. Cost: $\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}\,R_{K}\,d_{h}\bigr{)}$.
+Per-Head QK $\mathbf{A}$-Factor Combination and Scaling ($\mathcal{L}^{\prime}$): Each head combines its $\mathbf{A}_{Q},\mathbf{A}_{K}$ with $P$ and applies $s_{Q}s_{K}$.
+Cost: $\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}\,R_{K}\bigr{)}$.
+Value Aggregation (including $s_{V}$): Assuming $\mathbf{B}_{V}$ factors are shared and $\mathbf{A}_{V}$ are per-head.
+Cost: $\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{V}\,d_{h}\bigr{)}$.
+Assuming $d_{v}=d_{h}$.
+Total FLOPs for specialized multi-head TPA (full sequence processing) with batch size $B$:
+$\mathcal{F}_{\text{TPA-specialized}}=\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}\,R_{K}\,d_{h}\bigr{)}}_{\text{QK $\mathbf{B}$-dot products }P}+\underbrace{\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}\,R_{K}\bigr{)}}_{\text{Per-head QK }\mathbf{A}\text{-comb.
++ }s_{Q}s_{K}\text{ scale}}+\underbrace{\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{V}\,d_{h}\bigr{)}}_{\text{Value Aggregation+ }s_{V}\text{ scale}}.$
 Discussion.
-By contrast, standard multi-head attention typically requires $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$ flops for the $\mathbf{Q}\mathbf{K}^{\top}$ dot product (plus a similar $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$ for multiplying by $\mathbf{V}$).
-The factorization can yield savings provided $R_{q}R_{k}\ll h$ (for QK) and $R_{v}\ll h$ (for V), though actual speedups depend on how well these multi‐stage kernels are implemented and on hardware efficiency.
-By retaining $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}$ in factorized form, one can forgo the usual steps:
-$\mathbf{x}_{t}~{}\mapsto~{}\mathbf{Q}_{t},~{}\mathbf{K}_{\tau}~{}\mapsto~{}(\mathbf{Q}\,\mathbf{K}^{\top})~{}\mapsto~{}\operatorname{softmax}(\mathbf{Q}\,\mathbf{K}^{\top})\,\mathbf{V}~{}\mapsto~{}\text{final output}.$
-Instead, the large $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors (of size $T\times d_{h}$) are never materialized.
-The cost is replaced by rank-based $\mathbf{b}$-dot‐product computations plus per‐head scalar multiplications.
-The main challenge is to keep the factor ranks $(R_{q},R_{k},R_{v})$ sufficiently small relative to $d_{h}$ and to implement the necessary multi‐stage kernels efficiently.
-When $R_{q},R_{k},R_{v}\ll h$, fully factorized QKV attention can yield substantial gains in both computation and memory footprint.
-A.8 Decoding Speed during Inference Time of MHA, MQA, GQA, MLA, and TPA
-Suppose we are in an autoregressive setting, decoding the current token $\mathbf{x}_{T}$ given cached keys and values (KV) from all previous tokens $\mathbf{x}_{1},\dots,\mathbf{x}_{T-1}$.
-For each attention head $i\in\{1,\dots,h\}$, we store $\mathbf{K}_{i}\in\mathbb{R}^{T\times d_{h}},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$.
-Below, we compare the flops needed by MHA, MQA, GQA, MLA, and TPA to compute the next-token logits during inference.
-MHA, MQA, and GQA. Despite sharing or grouping keys/values in MQA and GQA, the decoding cost for MHA, MQA, and GQA remains of the same order. Specifically, for each head $i$, we compute:
-$\mathbf{Q}_{i}(\mathbf{x}_{T})~{}\in~{}\mathbb{R}^{d_{h}},\quad\mathbf{K}_{i}~{}\in~{}\mathbb{R}^{T\times d_{h}},\quad\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}~{}\in~{}\mathbb{R}^{1\times T},$
-$\quad\text{and}\quad\operatorname{Softmax}\bigl{(}\mathbf{Q}_{i}(\mathbf{x}_{T})\,\mathbf{K}_{i}^{\top}\bigr{)}\,\mathbf{V}_{i}~{}\in~{}\mathbb{R}^{d_{h}}.$
-Hence, the flops scale linearly in $h$, $d_{h}$, and $T$. For example, forming $\mathbf{Q}_{i}(\mathbf{x}_{T})\mathbf{K}_{i}^{\top}$ for each head $i$ costs roughly $\mathcal{O}(h\,d_{h}\,T)$.
+By contrast, standard multi-head attention typically requires $\mathcal{F}_{\text{MHA}}=\mathcal{O}\bigl{(}2B\,T^{2}\,H\,d_{h}\bigr{)}$ FLOPs (for $\mathbf{Q}\mathbf{K}^{\top}$ and attention-weighted $\mathbf{V}$).
+The specialized TPA can be more efficient if $R_{Q}R_{K}d_{h}+HR_{Q}R_{K}+HR_{V}d_{h}<2Hd_{h}$.
+Dividing by $Hd_{h}$, this condition becomes $(R_{Q}R_{K}/H)+(R_{Q}R_{K}/d_{h})+R_{V}<2$.
+For example, if $R_{Q}=R_{K}=R_{V}=1$, the condition is $21/H+1/d_{h}+1<$ 2, or $11/H+1/d_{h}<$ 1, which is generally true for typical $H,d_{h}\geq 2$.
+Thus, with small ranks, TPA can offer computational savings.
+Actual speedups also depend critically on memory access patterns and kernel implementations, as demonstrated by FlashAttention-style approaches.
+By retaining $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}$ in factorized form, one can bypass the explicit materialization of these large tensors:
+$\mathbf{x}_{t}\rightarrow\mathbf{Q}_{t},~{}\mathbf{K}_{\tau},\mathbf{V}_{\tau}\quad(\text{materialized})\rightarrow(\mathbf{Q}\,\mathbf{K}^{\top})\rightarrow\operatorname{softmax}(\mathbf{Q}\,\mathbf{K}^{\top})\,\mathbf{V}\rightarrow\text{final output}.$
+Instead, the large $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors (each of size $B\times T\times H\times d_{h}$) are not explicitly formed.
+The computation is restructured into rank-based $\mathbf{B}$-dot-product computations, per-head $\mathbf{A}$-factor combinations, and appropriate scaling.
+The challenge lies in choosing ranks $(R_{Q},R_{K},R_{V})$ small enough to ensure computational benefits while maintaining model quality, and in implementing the multi-stage kernels efficiently.
+When ranks are sufficiently small, this specialized approach can lead to gains in both computation and memory footprint.
+[…]
+Figure 6: Data flow diagram for specialized TPA computation, avoiding materialization of full $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors.
+Phase 1 (Top): Attention Score Computation.
+Factorized $\mathbf{B}_{Q},\mathbf{B}_{K}$ produce shared dot-products $P$ (summing over feature dimension $d$).
+$P$ is combined with factorized $\mathbf{A}_{Q},\mathbf{A}_{K}$ and scaled by $s_{Q}s_{K}$ (summing over ranks $r_{q},r_{k}$) to form scaled logits $\mathcal{L}^{\prime}$.
+These are then scaled by $1/\sqrt{d_{h}}$ and passed through Softmax to yield attention probabilities $\bm{\alpha}$.
+Phase 2 (Bottom): Value Aggregation.
+$\bm{\alpha}$ weights the factorized $\mathbf{A}_{V}$ (conceptually forming $W_{\mathbf{A}_{V}}$ via an einsum-like combination), which then aggregates with $\mathbf{B}_{V}$ and is scaled by $s_{V}$ (summing over key tokens $t_{k}$ and value ranks $r_{v}$) to produce the final output $\mathbf{O}$.
+(Diagram simplifies batch dimensions $B$; $T$ denotes sequence length; $H$ is heads; $R_{X}$ are ranks $R_{Q},R_{K},R_{V}$ corresponding to indices $r_{q},r_{k},r_{v}$; $d_{h}$ is the feature dimension $d$ for $\mathbf{B}$ factors.
+Operations $\sum_{X}$ denote contractions over index $X$, and $\odot$ denotes an einsum-like combination.
+Scaling factors $s_{Q},s_{K},s_{V}$ are incorporated as described.).
+C.8 Inference Time Decoding of Different Attention Mechanisms
+In autoregressive decoding, we generate the output for the current token $\mathbf{x}_{M}$ (where $M$ is the current sequence length, previously denoted $T$ in this section for a generic single token, but $M$ is used in Alg.
+1) given cached keys and values from $M-1$ previous tokens.
+We analyze the FLOPs for computing the attention output for this single query token.
+For all mechanisms, we consider the FLOPs for computing scaled logits and aggregating values, excluding the initial projection of the current token’s hidden state $\mathbf{x}_{M}$ into Q, K, V representations.
+MHA, MQA, and GQA. For Multi-Head Attention (MHA), with $H$ query heads and $H$ distinct Key/Value heads:
+  • Logits: Each of the $H$ query vectors (dim $d_{h}$) interacts with its corresponding K-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
+  • Value Aggregation: Each of the $H$ attention patterns interacts with its V-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
+  • Total MHA: $2MHd_{h}$.
+Multi-Query Attention (MQA) uses $H$ query heads but shares a single Key/Value head ($H_{kv}=1$).
+  • Logits: Each of the $H$ query vectors (dim $d_{h}$) interacts with the single shared K-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
+  • Value Aggregation: Each of the $H$ attention patterns interacts with the single shared V-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
+  • Total MQA: $2MHd_{h}$.
+Grouped-Query Attention (GQA) uses $H$ query heads and $N_{g}$ Key/Value head groups ($H_{kv}=N_{g}$).
+The $H$ query heads are partitioned into $N_{g}$ groups, each containing $H/N_{g}$ query heads that share one K/V head pair within that group.
+  • Logits: For each of the $N_{g}$ groups, $H/N_{g}$ query vectors interact with the group’s K-cache.
+Cost per group: $(H/N_{g})\cdot(M\cdot d_{h})$.
+Total for $N_{g}$ groups: $N_{g}\cdot(H/N_{g})\cdot Md_{h}=MHd_{h}$.
+  • Value Aggregation: Similarly, $MHd_{h}$.
+  • Total GQA: $2MHd_{h}$.
+MQA and GQA significantly reduce the KV cache size and memory bandwidth requirements compared to MHA.
+While the arithmetic FLOP count for the core attention computation (dot products and weighted sums) is $2MHd_{h}$ for all three if they have the same number of query heads $H$ and head dimension $d_{h}$, the practical speedups for MQA/GQA often arise from better memory access patterns due to smaller K/V caches.
 MLA.
-During inference, MLA can be seen as MQA but uses a larger head dimension to accommodate both RoPE and compressed representations (e.g., $d_{h}^{\prime}=d_{\text{rope}}+d_{c}$).
-In typical configurations, $d_{\text{rope}}+d_{c}$ can be significantly larger (e.g., $d_{h}^{\prime}=576$ rather than $d_{h}=64$ or $128$), thus inflating the dot-product cost by roughly 4.5×4.5$\times$4.5 × to 9×9$\times$9 × compared to MHA/MQA/GQA.
+Multi-Head Latent Attention (MLA), as described in Appendix H.3, uses $H$ heads.
+For each head, the key $\mathbf{K}_{i}$ has dimension $d_{k}^{\prime}=d_{c}+d_{h}^{R}$ (compressed part + RoPE part), and the value $\mathbf{V}_{i}^{C}$ has dimension $d_{v}^{\prime}=d_{c}$ (only compressed part).
+  • Logits: $H$ query heads (dim $d_{k}^{\prime}$) interact with K-caches (dim $d_{k}^{\prime}$). Cost: $MHd_{k}^{\prime}$.
+  • Value Aggregation: $H$ attention patterns interact with V-caches (dim $d_{v}^{\prime}$). Cost: $MHd_{v}^{\prime}$.
+  • Total MLA: $MH(d_{k}^{\prime}+d_{v}^{\prime})=MH(d_{c}+d_{h}^{R}+d_{c})=MH(2d_{c}+d_{h}^{R})$.
 TPA.
-Recall that TPA factorizes $\mathbf{Q}$ and $\bm{K}$ into rank-$(R_{q},R_{k})$ terms (see Section A), potentially avoiding large $\mathbf{Q},\bm{K}$ materializations.
-At inference, TPA’s dot-product cost can be broken into two parts:
-$\underbrace{R_{q}\,R_{k}\,d_{h}\,T}_{\text{QK $\mathbf{b}$-dot products (shared across all heads)}}~{}+~{}2\,\underbrace{R_{q}\,R_{k}\,h\,T}_{\text{per-head scalar multiplications}},$
-where $T$ is the current sequence length. For concrete values $d_{h}=128$, $h=64$, $R_{q}=8$, and $R_{k}=2$ (or $R_{q}=16$, $R_{k}=1$), we obtain:
+We use the FlashTPA Decoding algorithm (Algorithm 1) for FLOPs analysis, with $N=1$ query token, $M$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$).
+For ranks $(R_{Q},R_{K},R_{V})$:
+  • Logits computation (Steps 1-3 of Alg. 1): $\mathcal{O}(M(R_{Q}R_{K}D+HR_{Q}R_{K}+HR_{K}))$.
+  • Value aggregation (Steps 5-6 of Alg. 1): $\mathcal{O}(M(HR_{V}+HR_{V}E))$.
+  • Total for TPA decoding: $\mathcal{O}(M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
+Example Comparison.
+Let’s use the ranks from FlashTPA experiments (Section 5.2): $(R_{Q},R_{K},R_{V})=(16,1,1)$, and typical dimensions $D=E=d_{h}=64$, $H=32$ (e.g., for a 2048 $d_{\text{model}}$).
+MHA Logits: $\quad M\cdot H\cdot d_{h}=M\cdot 32\cdot 64=2048M$
+MHA Value Agg.: $\quad M\cdot H\cdot d_{h}=M\cdot 32\cdot 64=2048M$
+MHA Total: $\quad 4096M$
+TPA $\text{TPA Logits}(R_{K}=1,R_{Q}=16):$ $\quad M(R_{Q}d_{h}+HR_{Q}+H)=M(16\cdot 64+32\cdot 16+32)$
+$=M(1024+512+32)=1568M$
 […]
-[…]
-[…]
-Thus, in this setup, TPA can significantly reduce the flops needed for computing the $\mathbf{Q}(\mathbf{x}_{T})\bm{K}^{\top}$ operation at each decoding step.
-The actual end-to-end wall-clock speedup also depends on kernel fusion, caching strategies, and hardware implementation details, but the factorized formulation offers a pathway to more efficient decoding than standard attention.
-Appendix B Higher-Order Tensor Product Attention
-All prior discussions have focused on a second-order factorization in which each rank-$R_{Q}$ (and similarly $R_{K}$, $R_{V}$) component is the outer product of two vectors: one in $\mathbb{R}^{h}$ (the “head” dimension) and one in $\mathbb{R}^{d_{h}}$.
-We now generalize this by introducing an additional latent factor, yielding a third-order (or higher) factorization reminiscent of canonical polyadic (CP) decomposition.
-Concretely, for a single token $t$, we write
+TPA Total: $\quad 1568M+2080M=3648M$
+For MQA and GQA, assuming $H=32$ query heads and $d_{h}=64$:
+MQA/GQA Total: $\quad 2MHd_{h}=2\cdot M\cdot 32\cdot 64=4096M$
+For MLA, with $H=32,d_{c}=256,d_{h}^{R}=32$:
+MLA Logits: $\quad MH(d_{c}+d_{h}^{R})=M\cdot 32\cdot(256+32)=M\cdot 32\cdot 288=9216M$
+MLA Value Agg.: $\quad MHd_{c}=M\cdot 32\cdot 256=8192M$
+MLA Total: $\quad M\cdot 32\cdot(2\cdot 256+32)=M\cdot 32\cdot 544=17408M$
+In this configuration, TPA requires fewer FLOPs ($3648M$) than MHA ($4096M$).
+The FLOPs for scaled logits in TPA ($1568M$) are less than MHA’s ($2048M$).
+The value aggregation FLOPs for TPA ($2080M$) are comparable to MHA’s ($2048M$).
+This reduction in FLOPs, particularly when $R_{K}$ and $R_{V}$ are small (e.g., 1), combined with optimized memory access patterns in algorithms like FlashTPA Decoding, contributes to TPA’s competitive decoding speed, especially for long sequences.
+The actual end-to-end wall-clock speedup also depends on kernel fusion, hardware specifics, and the efficiency of einsum implementations, but the factorized formulation offers a clear path to reduced computational load.
+Appendix D More on FlashTPA Decoding Algorithm
+This section provides further details on the FlashTPA Decoding algorithm (see Algorithm 1), including the setup of factorized components.
+Factorized Component Setup for FlashTPA Decoding.
+Let $B$ be the batch size, $N$ the number of query tokens (which is 1 for decoding, $N=1$), $M$ the current length of the KV cache, $H$ the number of attention heads, and $R_{Q}$ the rank of the query factorization.
+The feature dimensions for the factorized components are $D$ (for query $\mathbf{B}_{Q}$ and key $\bm{b}^{K}$) and $E$ (for value $\bm{b}^{V}$).
+Typically, $D$ and $E$ correspond to the head dimension $d_{h}$.
+The query for the current token $\mathbf{x}_{t}$ (where $t$ is the current time step, $N=1$) is factorized as $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$, where $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{H\times R_{Q}}$ (derived from the input tensor ‘Aq’ of shape $(B,1,H,R_{Q})$ for the current token) and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times D}$ (similarly from ‘Bq’ of shape $(B,1,R_{Q},D)$).
+The cached keys $\mathbf{K}_{s}$ and values $\mathbf{V}_{s}$ for past tokens $s\in[1,M]$ are stored in their factorized form.
+For clarity in the data flow diagram (Figure 2 in the main paper) and the pseudo-code in Algorithm 1, we present the case where cached keys and values use ranks $R_{K}=1$ and $R_{V}=1$.
+The general TPA formulation allows for $R_{K}>1$ and $R_{V}>1$ for cached items (by storing, for example, $\mathbf{A}_{K}(\mathbf{x}_{s})\in\mathbb{R}^{R_{K}\times H}$ and $\mathbf{B}_{K}(\mathbf{x}_{s})\in\mathbb{R}^{R_{K}\times D}$).
+The einsum operations detailed in Algorithm 1 can be extended to accommodate higher ranks for keys and values by including summations over these additional rank dimensions.
+For $R_{K}=1$ and $R_{V}=1$, the cached $s$-th key and value are: $\mathbf{K}_{s}=\mathbf{a}^{K}(\mathbf{x}_{s})\otimes\bm{b}^{K}(\mathbf{x}_{s})\quad\in\mathbb{R}^{H\times D}$, $\mathbf{V}_{s}=\mathbf{a}^{V}(\mathbf{x}_{s})\otimes\bm{b}^{V}(\mathbf{x}_{s})\quad\in\mathbb{R}^{H\times E}$.
+Here, $\mathbf{a}^{K}(\mathbf{x}_{s})\in\mathbb{R}^{H}$ (from the cached tensor $\mathbf{a}^{K}_{\text{cache}}$ of shape $(B,M,H)$), $\bm{b}^{K}(\mathbf{x}_{s})\in\mathbb{R}^{D}$ (from $\bm{b}^{K}_{\text{cache}}$ of shape $(B,M,D)$), $\mathbf{a}^{V}(\mathbf{x}_{s})\in\mathbb{R}^{H}$ (from $\mathbf{a}^{V}_{\text{cache}}$ of shape $(B,M,H)$), and $\bm{b}^{V}(\mathbf{x}_{s})\in\mathbb{R}^{E}$ (from $\bm{b}^{V}_{\text{cache}}$ of shape $(B,M,E)$).
+D.1 Detailed Computation Steps of FlashTPA Decoding Algorithm
+The algorithm proceeds through a series of einsum operations, which correspond to specific tensor contractions:
+  1. 1.
+    Compute $\mathbf{B}_{Q}$-$\bm{b}^{K}$ inner products ($S^{(1)}$): For each query (batch $b$, query index $n=0$), each query rank component $r$, and each cached key position $m$, we compute the dot product of the $D$-dimensional feature vector of the $r$-th query factor $(\mathbf{B}_{Q})_{b,n,r,:}$ with the $D$-dimensional feature vector of the $m$-th cached key $(\bm{b}^{K}_{\text{cache}})_{b,m,:}$.
+    $S^{(1)}_{b,n,m,r}=\sum_{d=1}^{D}(\mathbf{B}_{Q})_{b,n,r,d}\cdot(\bm{b}^{K}_{\text{cache}})_{b,m,d}.$
+    This step captures the similarity between the feature parts of the query factors and the cached key factors.
+  2. 2.
+    Combine with query factor $\mathbf{A}_{Q}$ ($S^{(2)}$): The result $S^{(1)}$ is then combined with the head-specific query factors $\mathbf{A}_{Q}$.
+For each head $h$, the contributions from all query ranks $r$ are summed up.
+    $S^{(2)}_{b,n,m,h}=\sum_{r=1}^{R_{Q}}(\mathbf{A}_{Q})_{b,n,h,r}\cdot S^{(1)}_{b,n,m,r}.$
+    This yields an intermediate score for each query, cached key, and head.
+  3. 3.
+    Incorporate cached key factor $\mathbf{a}^{K}_{\text{cache}}$ for full logits ($\mathcal{L}$): The head-specific factor of the cached keys, $\mathbf{a}^{K}_{\text{cache}}$, is multiplied with $S^{(2)}$ to produce the final unscaled attention logits.
+    $\mathcal{L}_{b,h,n,m}=S^{(2)}_{b,n,m,h}\cdot(\mathbf{a}^{K}_{\text{cache}})_{b,m,h}.$
+    This logit $\mathcal{L}_{b,h,n,m}$ represents $s_{Q}^{-1}\cdot(\mathbf{Q}_{t}\mathbf{K}_{m}^{\top})_{h}$, i.e., the dot product between the query $\mathbf{Q}_{t}$ and the $m$-th cached key $\mathbf{K}_{m}$ for head $h$, scaled by $R_{Q}$ (since $s_{Q}=1/R_{Q}$).
+The ‘einsum‘ operation also rearranges dimensions to $(B,H,N,M)$ for the subsequent softmax operation.
+  4. 4.
+    Apply Scaling and Softmax ($\bm{\alpha}$): The logits $\mathcal{L}$ are scaled by the product of $s_{Q}$ (e.g., $1/R_{Q}$), $s_{K}$ (e.g., $1/R_{K}$, which is 1 if $R_{K}=1$), and $s_{\text{total}}$ (e.g., $1/\sqrt{D}$, where $D$ is the dimension used for dot products, typically $d_{h}$).
+Softmax is then applied across the $M$ cached key positions for each head independently.
+    $\alpha_{b,h,n,m}=\operatorname{Softmax}_{m}\left(\mathcal{L}_{b,h,n,m}\cdot s_{Q}\cdot s_{K}\cdot s_{total}\right).$
+    These are the attention probabilities.
+  5. 5.
+    Compute weighted sum with cached value factor $\mathbf{a}^{V}_{\text{cache}}$ ($\mathbf{O}^{(A)}$): The attention probabilities $\bm{\alpha}$ are used to weight the head-specific factors of the cached values, $\mathbf{a}^{V}_{\text{cache}}$.
+    $\mathbf{O}^{(A)}_{b,n,m,h}=\alpha_{b,h,n,m}\cdot(\mathbf{a}^{V}_{\text{cache}})_{b,m,h}.$
+    This step prepares the value components for the final aggregation. The ‘einsum‘ operation also reorders dimensions to $(B,N,M,H)$.
+  6. 6.
+    Incorporate cached value factor $\bm{b}^{V}_{\text{cache}}$ and apply final scaling ($\mathbf{O}$): Finally, the intermediate weighted value factors $\mathbf{O}^{(A)}$ are combined with the feature-specific factors of the cached values, $\bm{b}^{V}_{\text{cache}}$, by summing over the $M$ cached positions.
+The result is then scaled by $s_{V}$ (e.g., $1/R_{V}$, which is 1 if $R_{V}=1$).
+    $\mathbf{O}_{b,n,h,e}=\left(\sum_{m=1}^{M}\mathbf{O}^{(A)}_{b,n,m,h}\cdot(\bm{b}^{V}_{\text{cache}})_{b,m,e}\right)\cdot s_{V}.$
+    This produces the final output tensor of shape $(B,1,H,E)$.
+D.2 Triton FlashTPA Decoding Kernel
+We implement the experiments using Triton language [56], and the detailed implementation pseudo code is displayed in Algorithm 3.
+The Triton FlashTPA decoding kernel supports parallelism across the number of heads, rank, and head dimensions.
+And in 3, we will show the experiment results with this kernel.
+Algorithm 3 Triton FlashTPA Decoding Kernel
+1:Input Tensors: $\mathbf{A}_{Q}(B,N,H,R_{Q})$, $\mathbf{a}^{K}(B,M,H)$, $\mathbf{a}^{V}(B,M,H)$, $\mathbf{B}_{Q}(B,N,R_{Q},D)$, $\bm{b}^{K}(B,M,D)$, $\bm{b}^{V}(B,M,E)$
+2:Scaling factors: $s_{\text{total}},s_{Q},s_{K},s_{V}$; Dimensions: $B,N(=1),M,H,R_{Q},D,E$
+3:Kernel Block dims: $B_{H},B_{R},B_{D},B_{E}$; Sequence Blocking: $M_{\text{block}},M_{\text{chunk}}$
+4:Program IDs: $p_{id_{B}},p_{id_{H}},p_{id_{M}}$
+5:Partial Output $\mathbf{O}_{\text{partial}}(B,\text{Num}_{M},N,H,E)$, Log-Sum-Exp $\mathbf{LSE}_{\text{partial}}(B,\text{Num}_{M},H)$
+6:
+7:$b\leftarrow p_{id_{B}}$; $h_{\text{start}}\leftarrow p_{id_{H}}\cdot B_{H}$
+8:$m_{\text{block}\_\text{start}}\leftarrow p_{id_{M}}\cdot M_{\text{block}}$; $m_{\text{block}\_\text{end}}\leftarrow\min((p_{id_{M}}+1)\cdot M_{\text{block}},M)$
+9$\triangleright$ $B_{H},B_{R},B_{D},B_{E}$ are tile sizes for dimensions H, R, D, E respectively.
+10:
+11$\triangleright$ Initialize accumulators for the head block
+12:$\mathbf{o}_{\text{accum}}\leftarrow\mathbf{0}^{(E\times B_{H})}$; $\mathbf{m}_{\text{max}}\leftarrow-\infty^{(B_{H})}$; $\mathbf{s}_{\text{exp\_sum}}\leftarrow\mathbf{0}^{(B_{H})}$; $c_{\text{scale}}\leftarrow s_{\text{total}}\cdot s_{Q}\cdot s_{K}$
+13:
+14$\triangleright$ Load query factors (fixed for this program as N=1)
+15:Load $\mathbf{A}_{Q,\text{local}}^{(R_{Q}\times B_{H})}$ from $\mathbf{A}_{Q}[b,0,h_{\text{start}}\dots,:]$
+16:Load $\mathbf{B}_{Q,\text{local}}^{(D\times R_{Q})}$ from $\mathbf{B}_{Q}[b,0,:,:]$triangleright Dimensions may be transposed after loading for matmul
+17:
+18$\triangleright$ Iterate over $M_{\text{chunk}}$-sized chunks within the K/V block
+19:for $m_{\text{chunk\_start}}$ from $m_{\text{block}\_\text{start}}$ to $m_{\text{block}\_\text{end}}-1$ step $M_{\text{chunk}}$ do
+20:    $m_{\text{chunk\_\text{end}}}\leftarrow\min(m_{\text{chunk\_start}}+M_{\text{chunk}},m_{\text{block}\_\text{end}})$
+21:    $M_{\text{curr}\_\text{chunk}}\leftarrow m_{\text{chunk\_\text{end}}}-m_{\text{chunk\_start}}$
+22:    $\triangleright$ Load K/V factors for the current chunk
+23:    Load $\mathbf{a}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\mathbf{a}^{V}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\bm{b}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},D)$; $\bm{b}^{V}_{\text{chunk}}(E,M_{\text{curr}\_\text{chunk}})$triangleright Layouts optimized for memory access and matmuls
+24:    $\bm{b}^{V}_{\text{chunk}}\leftarrow\bm{b}^{V}_{\text{chunk}}\cdot s_{V}$
+25:    $\triangleright$ Core TPA Score Calculation for the chunk
+26:    $S1_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{K}_{\text{chunk}},\mathbf{B}_{Q,\text{local}})$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},R_{Q})$
+27:    $S2_{\text{chunk}}\leftarrow\text{MatMul}(S1_{\text{chunk}},\mathbf{A}_{Q,\text{local}})$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
+28:    $S3_{\text{chunk}}\leftarrow S2_{\text{chunk}}\odot\mathbf{a}^{K}_{\text{chunk}}\cdot c_{\text{scale}}$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
+29:    $\triangleright$ Online Softmax Update for the chunk
+30:    $\mathbf{m}_{\text{max\_local}}\leftarrow\max_{axis=0}(S3_{\text{chunk}})$triangleright Shape: $(B_{H})$
+31:    $\mathbf{m}_{\text{max\_new}}\leftarrow\max(\mathbf{m}_{\text{max}},\mathbf{m}_{\text{max\_local}})$
+32:    $\mathbf{p}_{\text{num}}\leftarrow\exp(S3_{\text{chunk}}-\mathbf{m}_{\text{max\_new}}[\text{None},:])$
+33:    $\mathbf{s}_{\text{exp\_sum\_local}}\leftarrow\sum_{axis=0}(\mathbf{p}_{\text{num}})$
+34:    $\mathbf{p}_{\text{weighted\_av}}\leftarrow(\mathbf{p}_{\text{num}}/\mathbf{s}_{\text{exp\_sum\_local}}[\text{None},:])\odot\mathbf{a}^{V}_{\text{chunk}}$
+35:    $\mathbf{o}_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{V}_{\text{chunk}},\mathbf{p}_{\text{weighted\_av}})$triangleright Shape: $(E,B_{H})$
+36:    $\triangleright$ Update global (M-block level) accumulators
+37:    $\mathbf{s}_{\text{exp\_sum\_prev\_rescaled}}\leftarrow\mathbf{s}_{\text{exp\_sum}}\cdot\exp(\mathbf{m}_{\text{max}}-\mathbf{m}_{\text{max\_new}})$
+38:    $\mathbf{s}_{\text{exp\_sum}}\leftarrow\mathbf{s}_{\text{exp\_sum\_prev\_rescaled}}+\mathbf{s}_{\text{exp\_sum\_local}}$
+39:    $\text{ratio}\leftarrow\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum}}$triangleright This is $\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum\_new}}$
+40:    $\mathbf{o}_{\text{accum}}\leftarrow(1-\text{ratio})\cdot\mathbf{o}_{\text{accum}}+\text{ratio}\cdot\mathbf{o}_{\text{chunk}}$
+41:    $\mathbf{m}_{\text{max}}\leftarrow\mathbf{m}_{\text{max\_new}}$
+42:end for
+43:
+44$\triangleright$ Store partial results for this program’s (batch, head_block, M_block)
+45:Store $\mathbf{o}_{\text{accum}}$ into $\mathbf{O}_{\text{partial}}[b,p_{id_{M}},0,h_{\text{start}}\dots,:]$
+46:$\mathbf{LSE}_{\text{val}}\leftarrow\log(\mathbf{s}_{\text{exp\_sum}})+\mathbf{m}_{\text{max}}$
+47:Store $\mathbf{LSE}_{\text{val}}$ into $\mathbf{LSE}_{\text{partial}}[b,p_{id_{M}},h_{\text{start}}\dots]$
+D.3 Additional Experimental Results
+The following figures present additional speed comparisons for different embedding dimensions, with $d_{h}=64$ maintained.
+The y-axis represents $\log_{2}(\text{time})$ in seconds (lower is faster), and the x-axis represents $\log_{2}(\text{sequence length})$.
+Detailed Analysis of Figure 5 (Main Text; Embedding Dimension 2048): Figure 5 in the main paper depicts speed comparisons for an embedding dimension of 2048.
+The results indicate that FlashTPA (blue line) is highly competitive.
+Across all tested batch sizes (1 to 16) for $d_{\text{model}}=2048$:
+  • MHA (orange line) is consistently the slowest mechanism, with its decoding time increasing most rapidly with sequence length.
+  • MQA (green line) and GQA (red line) offer significant speedups over MHA and perform very similarly to each other, often overlapping in the plots.
+  • MLA (purple line) demonstrates strong performance, generally being faster than MQA/GQA, particularly at longer sequence lengths.
+  • FlashTPA shows excellent scalability.
+While at very short sequence lengths (e.g., $2^{12}$ to $2^{13}$), its performance is comparable to MQA/GQA and MLA, its decoding time increases at a notably slower rate with sequence length.
+Consequently, FlashTPA becomes significantly faster than MQA/GQA for sequences longer than approximately $2^{14}$.
+  • Compared to MLA, FlashTPA is consistently among the top two performers.
+In many instances, particularly at sequence lengths greater than $2^{14}$ or $2^{15}$, FlashTPA matches or slightly surpasses MLA in speed.
+The logarithmic scale for time suggests that these differences can be substantial in practice for very long contexts.
+For example, at a sequence length of $2^{19}$ across various batch sizes, FlashTPA often shows a visible advantage over MLA.
+Figure 7 (Embedding Dimension 3072): With a larger embedding dimension of 3072, the relative performance trends observed in Figure 5 largely persist.
+  • FlashTPA (blue line) remains one of the most efficient decoding methods.
+MHA (orange line) is consistently the slowest, while MQA (green line) and GQA (red line) offer considerable improvements over MHA.
+  • MLA (purple line) and FlashTPA are the top two performers.
+FlashTPA consistently matches or exceeds the speed of MLA, particularly at longer sequence lengths (e.g., beyond $2^{15}$ or $2^{16}$ depending on the batch size).
+Its advantage often becomes more pronounced at the longest sequences tested ($2^{19}$).
+For instance, in batch size 1, TPA is clearly faster than MLA for sequence lengths $2^{16}$ and above.
+A similar trend is seen across other batch sizes, where TPA maintains a competitive edge or becomes superior at longer contexts.
+This suggests that FlashTPA’s efficiency is well-maintained even as the model’s embedding dimension increases.
+Figure 8 (Embedding Dimension 1024): For a smaller embedding dimension of 1024, similar trends are observed:
+  • FlashTPA (blue line) is highly competitive. MHA (orange line) remains the least performant. MQA (green line) and GQA (red line) are faster than MHA.
+  • At very short sequence lengths (around $2^{12}$ to $2^{13}$), MQA/GQA can be slightly faster than or comparable to TPA and MLA, especially for smaller batch sizes (e.g., Batch Size 1).
+  • However, as sequence length increases, both MLA (purple line) and FlashTPA demonstrate superior scalability.
+FlashTPA generally matches or outperforms MLA, particularly for sequences longer than $2^{15}$.
+For example, with a batch size of 16, TPA shows a clear speed advantage over MLA for sequence lengths $2^{16}$ and greater.
+These results across different embedding dimensions highlight the robustness of FlashTPA’s decoding speed advantages, especially for long sequences where it consistently ranks as one of the fastest, if not the fastest, attention mechanisms among those tested.
+(a) Batch Size=1
+(b) Batch Size=2
+(c) Batch Size=4
+(d) Batch Size=8
+(e) Batch Size=16
+Figure 7: Decoding time comparison of different attention mechanisms with an embedding dimension of 3072 and $d_{h}=64$.
+(a) Batch Size=1
+(b) Batch Size=2
+(c) Batch Size=4
+(d) Batch Size=8
+(e) Batch Size=16
+Figure 8: Decoding time comparison of different attention mechanisms with an embedding dimension of 1024 and $d_{h}=64$.
+Appendix E Higher-Order Tensor Product Attention
+All prior discussions have focused on TPA where the query, key, and value matrices (e.g., $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$) are formed as a sum of $R_{Q}$ components.
+Each component is an outer product of two context-dependent vectors, one spanning the head dimension ($\mathbb{R}^{h}$) and the other spanning the feature-per-head dimension ($\mathbb{R}^{d_{h}}$), as detailed in Section 3.1 (e.g., $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$ implies $\mathbf{Q}_{t}=\sum_{r}\mathbf{a}_{r}\bm{b}_{r}^{\top}$ where $\mathbf{a}_{r}$ are columns of $\mathbf{A}_{Q}^{\top}$ and $\bm{b}_{r}^{\top}$ are rows of $\mathbf{B}_{Q}$).
+We now generalize this by introducing additional latent factors in the construction of the feature-per-head vectors, leading to what we term higher-order TPA.
+This approach allows for more complex interactions in forming these feature vectors.
+For instance, in a third-order factorization, the query tensor $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$ for a single token $t$ is constructed as:
 $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\,\otimes\,\operatorname{vec}\bigl{(}\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\,\otimes\,\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\bigr{)},$
-where the newly introduced factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{\mathbf{c}}}$ can be viewed as a learnable gate or modulation term.
+where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$.
+The term $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$ and the newly introduced factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$ first form a matrix $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}\times d_{c}}$ via an outer product (as defined in Section 2).
+This matrix is then vectorized by $\operatorname{vec}(\cdot)$ into a column vector of dimension $d_{h}=d_{b}d_{c}$.
+The final query $\mathbf{Q}_{t}$ is formed by the sum of outer products between $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})$ and these resulting $d_{h}$-dimensional vectors.
 Analogous expansions apply to $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$.
-In practice, these triple (or higher-order) products still collapse into a matrix in $\mathbb{R}^{h\times d_{h}}$.
-One straightforward way to achieve this collapse is to split the feature dimension $d_{h}$ such that $d_{b}\times d_{c}=d_{h}$,
-$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}},\quad\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}},\quad\operatorname{vec}\bigl{(}\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\bigr{)}\in\mathbb{R}^{d_{h}}.$
-This additional factor can enhance expressiveness without necessarily increasing the base rank.
-Conceptually, it can act as a learnable nonlinearity or gating mechanism.
-One could also tie or share $\mathbf{c}^{Q}_{r}$ across queries, keys, and values, to reduce parameter overhead.
-A similar setup holds for keys (with rank $R_{K}$) and values (with rank $R_{V}$).
-Although this extra dimension adds to the parameter count, it can reduce the required rank to achieve a certain level of representational power.
-From a memory perspective, higher-order TPA still leverages factorized KV caching: only the factors $\mathbf{a}(\mathbf{x}_{t}),\mathbf{b}(\mathbf{x}_{t})$, and $\mathbf{c}(\mathbf{x}_{t})$ for each past token are cached.
-As usual, a trade-off arises between model capacity and the overhead of memory and computing.
-Nonetheless, moving from a rank-$\bigl{(}R_{Q},R_{K},R_{V}\bigr{)}$ matrix factorization to a higher-order tensor decomposition can provide additional flexibility and increased capacity.
-B.1 RoPE Compatibility in Higher-Order TPA
-Rotary positional embeddings (RoPE) remain compatible even under higher-order factorizations.
-In second-order TPA, RoPE can be treated as an invertible blockwise linear map acting on the last dimension of $\mathbf{Q}_{t}$ or $\mathbf{K}_{t}$.
-The same argument carries over when a third factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})$ is present.
-Suppose RoPE acts on the $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})$ portion (of dimension size $d_{b}$), we have the following theorem.
-Theorem 2 (RoPE Compatibility in Higher-Order TPA).
+The additional factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})$ can be viewed as a learnable, context-dependent modulation or gating term for the features generated by $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})$.
+$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}},\quad\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}},\quad d_{h}=d_{b}d_{c}.$
+This higher-order construction can enhance expressiveness.
+While introducing $\mathbf{c}^{Q}_{r}$ increases the parameter count for the factors, it might allow for the use of smaller base ranks ($R_{Q},R_{K},R_{V}$) to achieve comparable representational power, thus offering a different design choice.
+One could also explore tying or sharing $\mathbf{c}^{Q}_{r}$ across queries, keys, and values to manage parameter overhead.
+From a memory perspective during inference, higher-order TPA maintains the benefit of factorized KV caching.
+Only the constituent factors $\mathbf{a}_{K}(\mathbf{x}_{t}),\mathbf{b}_{K}(\mathbf{x}_{t}),\mathbf{c}_{K}(\mathbf{x}_{t})$ (and similarly for values) for each past token need to be stored.
+A trade-off arises between model capacity and the overhead of memory and computation.
+Higher-order tensor decompositions can provide additional flexibility and potentially increased capacity.
+E.1 RoPE Compatibility in Higher-Order TPA
+Rotary positional embeddings (RoPE) remain compatible with higher-order factorizations.
+In second-order TPA, RoPE applies rotations to the $d_{h}$-dimensional feature vectors.
+This compatibility extends to higher-order TPA.
+Consider the case where RoPE is intended to primarily rotate feature pairs derived from the $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})$ components, while the structural influence of $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})$ components on the $d_{h}$-dimensional vector is preserved.
+More formally, RoPE acts on the $d_{h}$-dimensional vector $\operatorname{vec}(\mathbf{b}_{r}^{Q}\otimes\mathbf{c}_{r}^{Q})$ such that the transformation is equivalent to rotating $\mathbf{b}_{r}^{Q}$ to $\widetilde{\mathbf{b}}_{r}^{Q}=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}$ (where $\mathbf{R}_{t}$ is the RoPE rotation matrix for $d_{b}$ dimensions) and then forming $\operatorname{vec}(\widetilde{\mathbf{b}}_{r}^{Q}\otimes\mathbf{c}_{r}^{Q})$.
+This is achieved by a specific RoPE transformation matrix $\mathbf{T}_{t}$ acting on the full $d_{h}$-dimensional vector, as stated in the following theorem.
+Theorem E.1 (RoPE Compatibility in Higher-Order TPA).
 Consider the higher-order (3-order) Tensor Product Attention (TPA) query factorization
 $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\in\mathbb{R}^{h\times d_{h}},$
-where $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$, $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$, with $d_{c}=\tfrac{d_{h}}{d_{b}}$.
+where $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$, $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$, with $d_{h}=d_{b}d_{c}$.
 Define the RoPE-transformed query as $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}=\mathbf{Q}_{t}\mathbf{T}_{t}$, where
-$\mathbf{T}_{t}=\mathbf{R}_{t}\otimes\mathbf{I}_{d_{c}}=\begin{pmatrix}\mathbf{R}_{t}&\cdots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{R}_{t}&\cdots&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&\mathbf{R}_{t}\\ \end{pmatrix}\in\mathbb{R}^{d_{h}\times d_{h}},$
-and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ ($d_{b}\in\mathbb{Z}_{+}$ is even) is a block-diagonal matrix composed of $2\times 2$ rotation matrices:
+$\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}=\begin{pmatrix}(\mathbf{R}_{t})^{\top}&\cdots&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&(\mathbf{R}_{t})^{\top}&\cdots&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&(\mathbf{R}_{t})^{\top}\\ \end{pmatrix}\in\mathbb{R}^{d_{h}\times d_{h}},$
+$\mathbf{I}_{d_{c}}$ is the identity matrix of size $d_{c}\times d_{c}$, and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ ($d_{b}\in\mathbb{Z}_{+}$ is even) is the standard RoPE block-diagonal matrix composed of $2\times 2$ rotation matrices:
 $\mathbf{R}_{t}=\begin{pmatrix}\cos(t\theta_{1})&-\sin(t\theta_{1})&&&\\ \sin(t\theta_{1})&\cos(t\theta_{1})&&&\\ &&\cos(t\theta_{2})&-\sin(t\theta_{2})&\\ &&\sin(t\theta_{2})&\cos(t\theta_{2})&\\ &&&&\ddots\\ &&&&&\cos(t\theta_{d_{b}/2})&-\sin(t\theta_{d_{b}/2})\\ &&&&&\sin(t\theta_{d_{b}/2})&\cos(t\theta_{d_{b}/2})\\ \end{pmatrix},$
 for $t\in\{1,\dots,T\}$ and $j\in\{1,\dots,d_{b}/2\}$.
-This construction ensures that RoPE rotates only the coordinates corresponding to $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged.
-Under these conditions, the RoPE-transformed query $\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}$ admits a higher-order TPA factorization of the same rank $R_{Q}$.
-Specifically, we have
-$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$ (B.1)
+The transformation $\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}$ operates on the $d_{h}$-dimensional vectorized features by post-multiplication.
+This structure of $\mathbf{T}_{t}$ ensures that the rotation effectively applied to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component (which is a column vector) corresponds to a pre-multiplication by $\mathbf{R}_{t}$, as detailed in the proof (Appendix F.2).
+This preserves the structure induced by $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ while rotating $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
+Under these conditions, the RoPE-transformed query $\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}$ admits a higher-order TPA factorization of the same rank $R_{Q}$:
+$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$ (E.1)
 where $\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
-Please see Appendix C.2 for the proof. For fourth-order or higher, this result still holds.
-Appendix C Proofs of Theorems
-C.1 Proof of Theorem 1
+Please see Appendix F.2 for the proof. For fourth-order or higher, this result still holds.
+To assess its empirical performance, we implemented third-order TPA.
+Table 4 lists the evaluation results for a small model.
+These results provide an initial indication of its viability.
+A comprehensive comparison with second-order TPA variants of similar parameter counts or ranks would be necessary to fully evaluate the trade-offs.
+Table 4: The evaluation results of small models with third-order TPA pre-trained using FineWeb-Edu 100B dataset with lm-evaluation-harness. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
+Few-shot ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+0-shot 49.24 24.91 57.06 34.01 31.80 63.33 50.59 23.23 66.90 44.56
+2-shot 53.37 25.34 48.78 34.00 29.20 62.79 52.33 26.41 75.30 45.28
+Table 5: The evaluation results of medium models with third-order TPA pre-trained using FineWeb-Edu 100B dataset with lm-evaluation-harness. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
+Few-shot ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+0-shot 56.19 29.44 58.36 43.74 33.80 67.95 52.80 24.07 77.20 49.28
+2-shot 62.84 30.89 55.69 43.16 33.00 68.01 51.78 26.57 86.90 50.98
+Appendix F Proofs of Theorems
+F.1 Proof of Theorem 3.1
 Proof.
@@
 Thus, TPA maintains compatibility with RoPE.
-This completes the proof of Theorem 1.
+This completes the proof of Theorem 3.1.
 ∎
-C.2 Proof of Theorem 2
+F.2 Proof of Theorem E.1
+Theorem E.1 addresses the compatibility of RoPE with higher-order (specifically, 3rd-order) Tensor Product Attention. The theorem considers the query factorization:
+$\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\in\mathbb{R}^{h\times d_{h}},$
+where $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{h}$ (column vector), $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$ (column vector), $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$ (column vector), and $d_{h}=d_{b}d_{c}$.
+The term $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ is interpreted as the matrix $\mathbf{M}_{r}=\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})(\mathbf{c}_{r}^{Q}(\mathbf{x}_{t}))^{\top}\in\mathbb{R}^{d_{b}\times d_{c}}$.
+The notation $\mathbf{a}\otimes\mathbf{v}$ for $\mathbf{a}\in\mathbb{R}^{h}$ and $\mathbf{v}\in\mathbb{R}^{d_{h}}$ (column vectors) implies the outer product $\mathbf{a}\mathbf{v}^{\top}$.
+Thus, $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})(\operatorname{vec}(\mathbf{M}_{r}))^{\top}$.
+The RoPE-transformed query is defined as $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)}=\mathbf{Q}_{t}\mathbf{T}_{t}$.
+Crucially, for the theorem’s conclusion to hold as intended (i.e., that the $\mathbf{b}_{r}^{Q}$ component is transformed by pre-multiplication with the standard RoPE matrix $\mathbf{R}_{t}$), the global transformation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ (that post-multiplies $\mathbf{Q}_{t}$) is given by:
+$\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top},$
+where $\mathbf{I}_{d_{c}}$ is the $d_{c}\times d_{c}$ identity matrix, and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ is the standard RoPE block-diagonal matrix that pre-multiplies $d_{b}$-dimensional column vectors (as defined explicitly in the theorem statement in Section E).
+The theorem claims that, under these conditions, $\widetilde{\mathbf{Q}}_{t}$ admits a higher-order TPA factorization:
+$\widetilde{\mathbf{Q}}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)},$
+where $\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
 Proof.
-We begin by observing that each term $\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}$ is an element of $\mathbb{R}^{h}\otimes\mathbb{R}^{d_{h}}$.
-Here, $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$, $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$, with $d_{c}=\tfrac{d_{h}}{d_{b}}$.
-Consequently, the tensor product $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ forms a $d_{b}\times d_{c}$ matrix, and its vectorization lies in $\mathbb{R}^{d_{b}\cdot d_{c}}=\mathbb{R}^{d_{h}}$.
-Applying the RoPE transformation to a single summand yields
-$\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\mapsto\mathbf{T}_{t}\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
-Since $\mathbf{T}_{t}$ is defined as the Kronecker product $\mathbf{R}_{t}\otimes\mathbf{I}_{d_{c}}$, where $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ and $\mathbf{I}_{d_{c}}$ is the identity matrix of size $d_{c}\times d_{c}$, it follows that
-$\mathbf{T}_{t}\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}=\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
-This is because the Kronecker product with an identity matrix effectively applies the rotation $\mathbf{R}_{t}$ to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged.
-Therefore, the RoPE transformation of a single summand becomes
-$\operatorname{RoPE}_{t}\Bigl{(}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}\Bigr{)}=\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}.$
-Importantly, this transformation does not mix the components $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ and $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$; it solely rotates $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ via $\mathbf{R}_{t}$.
-Summing over all ranks $r=1,\dots,R_{Q}$, we obtain
-$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl{(}\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$
-which retains the same higher-order TPA structure with rank $R_{Q}$.
-Thus, the RoPE transformation is fully compatible with higher-order TPA, preserving the factorization rank and maintaining the structure by only rotating the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ components while leaving $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ unchanged.
+Let $\mathbf{a}_{r}^{Q}\equiv\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})$, $\mathbf{b}_{r}^{Q}\equiv\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$, and $\mathbf{c}_{r}^{Q}\equiv\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ for brevity.
+Let $\mathbf{M}_{r}=\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top}\in\mathbb{R}^{d_{b}\times d_{c}}$.
+Let $\mathbf{v}_{r}=\operatorname{vec}(\mathbf{M}_{r})\in\mathbb{R}^{d_{h}}$ be the column vector obtained by stacking the columns of $\mathbf{M}_{r}$.
+The query tensor is $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{v}_{r})^{\top}$.
+The RoPE transformation is $\widetilde{\mathbf{Q}}_{t}=\mathbf{Q}_{t}\mathbf{T}_{t}$. Substituting the factorization and the revised definition of $\mathbf{T}_{t}$:
+$\widetilde{\mathbf{Q}}_{t}$ $=\left(\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{v}_{r})^{\top}\right)(\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})$
+$=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}\left((\mathbf{v}_{r})^{\top}(\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})\right).$
+Let’s analyze the transformed vector part for the $r$-th component: $(\mathbf{v}_{r})^{\top}(\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})$.
+This row vector is the transpose of $((\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})^{\top}\mathbf{v}_{r})$.
+Let’s compute the pre-multiplying matrix:
+$((\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top})^{\top}=(\mathbf{I}_{d_{c}})^{\top}\otimes((\mathbf{R}_{t})^{\top})^{\top}=\mathbf{I}_{d_{c}}\otimes\mathbf{R}_{t}.$
+So, the column vector transformation is $(\mathbf{I}_{d_{c}}\otimes\mathbf{R}_{t})\mathbf{v}_{r}$.
+Substitute $\mathbf{v}_{r}=\operatorname{vec}(\mathbf{M}_{r})=\operatorname{vec}(\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top})$:
+$(\mathbf{I}_{d_{c}}\otimes\mathbf{R}_{t})\operatorname{vec}(\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top}).$
+We use the Kronecker product identity: $(\mathbf{B_{0}}^{\top}\otimes\mathbf{A_{0}})\operatorname{vec}(\mathbf{X_{0}})=\operatorname{vec}(\mathbf{A_{0}}\mathbf{X_{0}}\mathbf{B_{0}})$.
+To match our expression $(\mathbf{I}_{d_{c}}\otimes\mathbf{R}_{t})\operatorname{vec}(\mathbf{M}_{r})$, we identify: $\mathbf{A_{0}}=\mathbf{R}_{t}$, $\mathbf{B_{0}}^{\top}=\mathbf{I}_{d_{c}}\implies\mathbf{B_{0}}=\mathbf{I}_{d_{c}}$, $\mathbf{X_{0}}=\mathbf{M}_{r}=\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top}$.
+Applying the identity, we get:
+$\operatorname{vec}\left(\mathbf{R}_{t}(\mathbf{b}_{r}^{Q}(\mathbf{c}_{r}^{Q})^{\top})\mathbf{I}_{d_{c}}\right)=\operatorname{vec}\left((\mathbf{R}_{t}\mathbf{b}_{r}^{Q})(\mathbf{c}_{r}^{Q})^{\top}\right).$
+Let $\widetilde{\mathbf{b}}^{Q}_{r}=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}$.
+This is precisely the transformation for the $\mathbf{b}_{r}^{Q}$ component as claimed in the theorem.
+So the transformed column vector is $\operatorname{vec}(\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{c}_{r}^{Q})^{\top})$.
+The corresponding row vector in the sum for $\widetilde{\mathbf{Q}}_{t}$ is therefore $(\operatorname{vec}(\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{c}_{r}^{Q})^{\top}))^{\top}$.
+Substituting this back into the expression for $\widetilde{\mathbf{Q}}_{t}$:
+$\widetilde{\mathbf{Q}}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\operatorname{vec}(\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{c}_{r}^{Q})^{\top}))^{\top}.$
+This is equivalent to the theorem’s claimed factorization, using the definition $\mathbf{a}\otimes\mathbf{col\_vec}=\mathbf{a}(\mathbf{col\_vec})^{\top}$:
+$\widetilde{\mathbf{Q}}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}\otimes\operatorname{vec}\bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}\otimes\mathbf{c}_{r}^{Q}\bigr{)},$
+where $\widetilde{\mathbf{b}}^{Q}_{r}=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}$.
+This completes the proof, showing that RoPE can be consistently applied to higher-order TPA representations if the global RoPE transformation matrix $\mathbf{T}_{t}$ (that post-multiplies $\mathbf{Q}_{t}$) is appropriately defined as $\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}$, ensuring that the standard RoPE matrix $\mathbf{R}_{t}$ effectively pre-multiplies the $\mathbf{b}_{r}^{Q}$ component.
 ∎
-Appendix D More Related Works
+Appendix G More Related Work
+Transformers and Attention.
+As a sequence-to-sequence architecture, Transformer [59] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies.
+Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [9, 48, 15, 29, 26, 30], kernel-based projections [10], and linearized transformers [58, 24, 43, 68, 53, 66].
+To decrease memory usage and circumvent the limitation of memory bandwidth in training, [45] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head.
+To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [2] divides queries into several groups, and each group of queries shares a single key head and value head.
+Recently, DeepSeek-V2 [31] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value.
+Concurrently, [20] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q.
+Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
+KV Cache Optimization.
+During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [39].
+This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [1].
+Consequently, various studies have explored methods to mitigate these issues.
+These include KV cache eviction strategies that discard less significant tokens [69, 61, 7, 1], dynamic sparse attention mechanisms focusing on selected keys and values [41, 54, 49], offloading the KV cache to CPU memory [16, 25, 52], and quantizing the KV cache [60, 33, 18].
+In contrast to these approaches, TPA focuses on reducing the intrinsic size of the KV cache by employing tensor-decomposed key and value representations.
 Low-Rank Factorizations.
-Low-rank approximations have been applied to compress model parameters and reduce complexity including LoRA (Hu et al., 2022), which factorizes weight updates during fine-tuning, and its derivatives for other training scenarios such as efficient pretraining (ReLoRA (Lialin et al., 2023), MoRA (Jiang et al., 2024)), long-context training (LongLoRA (Chen et al., 2024), SinkLoRA (Zhang, 2024)), as well as continual training (InfLoRA (Liang & Li, 2024), GS-LoRA (Zhao et al., 2024), I-LoRA (Ren et al., 2024)).
-These approaches typically produce static low-rank expansions that do not explicitly depend on the input context.
-And Malladi et al.
-(2023); Zeng & Lee (2024) provided theoretical proof of the expressiveness of low-rank approximation.
-For the initialization of factorization matrices, OLoRA (Büyükakyüz, 2024) applied QR-decomposition of pretrained weight to achieve better performance of language models while LoLDU (Shi et al., 2024) used LDU-decomposition to accelerate training of LoRA.
-Moreover, AdaLoRA (Zhang et al., 2023a) utilized Singular Value Decomposition (SVD) of the pretrained weight and introduced importance score for each parameter as a measurement to achieve dynamic adjustment of rank.
-TPA, by contrast, constructs Q, K, and V as contextually factorized tensors, enabling dynamic adaptation.
-Appendix E More on Attention Mechanisms
-E.1 Multi-Query Attention (MQA)
-Multi-Query Attention (MQA) (Shazeer, 2019) significantly reduces memory usage by sharing keys and values across heads, while still preserving unique query projections.
-For a sequence of embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$,
+Low-rank approximations are widely used to compress model parameters and reduce computational complexity.
+Notable examples include LoRA [19], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [27], MoRA [21]), long-context training (LongLoRA [8], SinkLoRA [65]), and continual training (InfLoRA [28], GS-LoRA [70], I-LoRA [40]).
+These methods generally produce static low-rank expansions that are independent of the input context.
+Theoretical justifications for the expressiveness of low-rank approximations have been provided by [37, 64].
+Initialization strategies for these factorization matrices have also been explored: OLoRA [6] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [47] employs LDU-decomposition to accelerate LoRA training.
+Furthermore, AdaLoRA [67] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks.
+TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.
+Appendix H More on Attention Mechanisms
+H.1 Multi-Query Attention (MQA)
+Multi-Query Attention (MQA) [45] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection.
+Given a sequence of input embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$, the query, shared key, and shared value tensors are computed as:
 $\mathbf{Q}_{i}=\mathbf{X}\bm{W}^{Q}_{i},\quad\mathbf{K}_{\text{shared}}=\mathbf{X}\bm{W}^{K}_{\text{shared}},\quad\mathbf{V}_{\text{shared}}=\mathbf{X}\bm{W}^{V}_{\text{shared}}.$
-Hence, each head $i$ only has a distinct query $\mathbf{Q}_{i}\in\mathbb{R}^{T\times d_{h}}$, but shares the same key $\mathbf{K}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$ and value $\mathbf{V}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$.
-In practice, this means:
+Thus, each head $i$ uses a distinct query projection $\mathbf{Q}_{i}\in\mathbb{R}^{T\times d_{h}}$ but shares the common key $\mathbf{K}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$ and value $\mathbf{V}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$ tensors.
+The weight matrices are:
 $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}},\quad\bm{W}^{K}_{\text{shared}},\bm{W}^{V}_{\text{shared}}\;\in\;\mathbb{R}^{\,d_{\text{model}}\times d_{h}}.$
@@
 $\textbf{head}_{i}=\operatorname{Attention}\bigl{(}\mathbf{Q}_{i},\mathbf{K}_{\text{shared}},\mathbf{V}_{\text{shared}}\bigr{)}.$
-By sharing these key and value projections, MQA cuts down on memory usage (especially for the key-value cache in autoregressive inference) but loses some expressivity since all heads must rely on the same key/value representations.
-E.2 Grouped Query Attention (GQA)
-Grouped Query Attention (GQA) (Ainslie et al., 2023) generalizes MHA and MQA by grouping heads.
-Specifically, we partition the $h$ total heads into $G$ groups.
-Each group has a single set of keys and values, but each individual head within that group still retains its own query projection.
-Formally, if $g(i)$ maps a head $i\in[h]$ to its group index $g\in[G]$, then:
+By sharing key and value projections, MQA substantially reduces memory demands, especially for the KV cache during autoregressive inference.
+However, this comes at the cost of reduced model expressivity, as all heads must utilize the same key and value representations.
+H.2 Grouped Query Attention (GQA)
+Grouped Query Attention (GQA) [2] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups.
+Within each group, heads share a common key and value projection, while each head maintains its own unique query projection.
+Formally, let $g(i)$ denote the group index for head $i\in\{1,\dots,h\}$, where $g(i)\in\{1,\dots,G\}$.
+The projections are:
 $\mathbf{K}_{g(i)}=\mathbf{X}\,\bm{W}^{K}_{g(i)},\quad\mathbf{V}_{g(i)}=\mathbf{X}\,\bm{W}^{V}_{g(i)},\quad\mathbf{Q}_{i}=\mathbf{X}\,\bm{W}^{Q}_{i},$
@@
 $\text{head}_{i}=\operatorname{Attention}\Bigl{(}\mathbf{Q}_{i},\mathbf{K}_{g(i)},\mathbf{V}_{g(i)}\Bigr{)}.$
-Again, $\bm{W}^{K}_{g},\bm{W}^{V}_{g}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ for each group $g$, and $\bm{W}^{Q}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{h}}$ for each head $i$.
+Here, $\bm{W}^{K}_{g}$ and $\bm{W}^{V}_{g}$ are the shared weight matrices for group $g$, each in $\mathbb{R}^{d_{\text{model}}\times d_{h}}$, and $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the query weight matrix for head $i$.
 The complete output is again a concatenation of all heads:
 $\operatorname{GQA}(\mathbf{X})=\operatorname{Concat}\Bigl{(}\text{head}_{1},\ldots,\text{head}_{h}\Bigr{)}\bm{W}^{O}.$
-By adjusting $G$ between $1$ and $h$, GQA can interpolate between sharing all key/value projections across heads (i.e., MQA) and having one set of projections per head (i.e., MHA).
-E.3 Multi-head Latent Attention (MLA)
-Below, we briefly outline the Multi-head Latent Attention (MLA) approach used by DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b).
-MLA introduces a low-rank compression of the keys and values to reduce the Key-Value (KV) caching cost at inference.
+By varying $G$ from $1$ (equivalent to MQA) to $h$ (equivalent to MHA), GQA offers a trade-off between memory efficiency and model capacity.
+H.3 Multi-head Latent Attention (MLA)
+Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [31] and DeepSeek-V3 [32], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
 $\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},$
@@
 $\operatorname{Concat}\bigl{(}\mathbf{V}_{1}^{C},\mathbf{V}_{2}^{C},\ldots,\mathbf{V}_{h}^{C}\bigr{)}$ $=\mathbf{V}^{C}=\mathbf{C}^{KV}\bm{W}^{UV},$
-where $\bm{W}^{DKV}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}},\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times d_{h}h},\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}},\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times d_{h}h}$, and $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the compressed KV latent (with $d_{c}\ll d_{h}h$), and $\operatorname{RoPE}(\cdot)$ represents the RoPE transform applied to the separate key embeddings $\mathbf{K}^{R}$ of dimension $d_{h}^{R}$.
-Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ need to be cached, reducing KV memory usage while largely preserving performance compared to standard MHA (Vaswani et al., 2017).
+Here, $\bm{W}^{DKV}\in\mathbb{R}^{d_{\text{model}}\times d_{c}}$ projects to a compressed dimension $d_{c}$, $\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times(d_{h}h)}$ up-projects the compressed keys, $\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}}$ projects to a residual key component for RoPE, and $\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times(d_{h}h)}$ up-projects the compressed values.
+$\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the shared compressed KV latent (where $d_{c}\ll d_{h}h$).
+The RoPE transformation is applied to a separate key embedding $\mathbf{K}^{R}\in\mathbb{R}^{T\times d_{h}^{R}}$.
+Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached, reducing KV memory usage while largely preserving performance compared to standard MHA [59].
 MLA also compresses the queries, lowering their training-time memory footprint:
@@
 $\mathbf{Q}$ $=\operatorname{Concat}\bigl{(}\mathbf{Q}^{C},\mathbf{Q}^{R}\bigr{)}.$
-where $\bm{W}^{DQ}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}^{\prime}},\bm{W}^{UQ}\in\mathbb{R}^{\,d_{c}^{\prime}\times d_{h}h},\bm{W}^{QR}\in\mathbb{R}^{d_{c}^{\prime}\times d_{h}^{R}h}$.
-Here, $\mathbf{C}^{Q}\in\mathbb{R}^{T\times d_{c}^{\prime}}$ (with $d_{c}^{\prime}\ll d_{h}h$) is the compressed query latent.
-As above, each $\bm{W}^{DQ}$, $\bm{W}^{UQ}$, and $\bm{W}^{QR}$ connects these lower-dimensional query latents back to $h$ heads of dimension $d_{h}+d_{h}^{R}$.
+The weight matrices are $\bm{W}^{DQ}\in\mathbb{R}^{d_{\text{model}}\times d_{c}^{\prime}}$, $\bm{W}^{UQ}\in\mathbb{R}^{d_{c}^{\prime}\times(d_{h}h)}$, and $\bm{W}^{QR}\in\mathbb{R}^{d_{c}^{\prime}\times(d_{h}^{R}h)}$.
+Here, $\mathbf{C}^{Q}\in\mathbb{R}^{T\times d_{c}^{\prime}}$ (where $d_{c}^{\prime}\ll d_{h}h$) is the compressed query latent.
+The final query $\mathbf{Q}_{i}$ for each head, formed by concatenating $\mathbf{Q}_{i}^{C}$ and $\mathbf{Q}_{i}^{R}$, has a dimension of $d_{h}+d_{h}^{R}$.
 Given compressed queries, keys, and values, the final attention output for the $t$-th token is:
@@
 $\mathbf{U}$ $=\operatorname{Concat}\bigl{(}\mathbf{O}_{1},\mathbf{O}_{2},\ldots,\mathbf{O}_{h}\bigr{)}\bm{W}^{O},$
-where $\bm{W}^{O}\in\mathbb{R}^{(d_{h}h)\times\,d_{\text{model}}}$ is the output projection.
-In inference time, $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ can be cached to accelerate decoding.
-In detail, when RoPE is ignored, the inner product $\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ (where $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{d}$) of the $i$-th head between $t$-th and $s$-th tokens can be calculated using the hidden state $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ for $t$-th token and the cached latent state $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ for $s$-th token:
-$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (E.1)
-$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (E.2)
-where $\bm{W}_{i}^{(\cdot)}$ is the $i$-th head of the original weight, and $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ can be computed previously for faster decoding.
-However, this process fails when RoPE is considered according to Su (2024).
-Since RoPE can be considered as multiplication with a block-diagonal matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ (see Section 2.3), with the property (2.1) that $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$, then
+where $\mathbf{V}_{i}$ is typically $\mathbf{V}_{i}^{C}$ as no residual value component is explicitly defined, and $\bm{W}^{O}\in\mathbb{R}^{(d_{h}h)\times d_{\text{model}}}$ is the output projection.
+During inference, $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached to accelerate decoding.
+In detail, if RoPE were ignored for the compressed components, the inner product $\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ (where $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}$) of the $i$-th head between $t$-th token query and $s$-th token key could be calculated using the current hidden state $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ and the cached latent state $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ for the $s$-th token:
+$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (H.1)
+$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (H.2)
+where $\bm{W}_{i}^{(\cdot)}$ denotes the $i$-th head’s portion of the respective weight matrix.
+The term $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding.
+However, as noted by [50], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations.
+RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section H.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation H.4).
+If RoPE were applied to the up-projected $Q^{C}$ and $K^{C}$:
 […]
-Different from (E.2), acceleration by pre-computing $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t-s}}(\bm{W}^{UK}_{i})^{\top}]$ fails since it varies for different $(t,s)$ position pairs.
-Therefore, MLA adds the additional $\mathbf{k}_{t}^{R}$ part with a relatively smaller size for RoPE compatibility.
-In Section 3.2, we will show that TPA addresses the issue of RoPE-incompatibility by applying tensor product.
-$\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},$
-$\operatorname{Concat}\bigl{(}\mathbf{K}_{1}^{C},\mathbf{K}_{2}^{C},\ldots,\mathbf{K}_{h}^{C}\bigr{)}$ $=\mathbf{K}^{C}=\mathbf{C}^{KV}\bm{W}^{UK},$
-$\mathbf{K}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{X}\bm{W}^{KR}\bigr{)},$
-$\mathbf{K}_{i}$ $=\operatorname{Concat}\bigl{(}\mathbf{K}_{i}^{C},\mathbf{K}^{R}\bigr{)},$
-$\operatorname{Concat}\bigl{(}\mathbf{V}_{1}^{C},\mathbf{V}_{2}^{C},\ldots,\mathbf{V}_{h}^{C}\bigr{)}$ $=\mathbf{V}^{C}=\mathbf{C}^{KV}\bm{W}^{UV},$
-E.4 Multi-matrix Factorization Attention (MFA)
-Hu et al. (2024) proposed Multi-matrix Factorization Attention (MFA), which can be seen as Multi-Query Attention (MQA) with dimension of each head equals $d_{C}$, and low-rank factorized Q:
+Unlike Equation (H.2), acceleration by pre-computing the term $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}\mathbf{T}_{t-s}(\bm{W}^{UK}_{i})^{\top}]$ is not possible because it depends on the relative position $(t-s)$ and thus varies for different $(t,s)$ pairs.
+To maintain RoPE compatibility while benefiting from compression, MLA introduces an additional, smaller key component $\mathbf{K}^{R}$ (and similarly $\mathbf{Q}^{R}$) to which RoPE is applied, while the main compressed components $\mathbf{K}^{C}$ and $\mathbf{V}^{C}$ (derived from $\mathbf{C}^{KV}$) remain RoPE-free.
+As we will demonstrate in Section 3.2 of the main paper, TPA offers a different approach to integrate RoPE efficiently with factorized attention through its tensor product formulation.
+H.4 Multi-matrix Factorization Attention (MFA)
+[20] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
 $\mathbf{Q}_{i}=\mathbf{X}\bm{W}^{DQ}\bm{W}^{UQ}_{i},\quad\mathbf{K}_{\text{shared}}=\mathbf{X}\bm{W}^{K}_{\text{shared}},\quad\mathbf{V}_{\text{shared}}=\mathbf{X}\bm{W}^{V}_{\text{shared}},$
@@
 $\bm{W}^{DQ}\in\mathbb{R}^{d_{\text{model}}\times d_{c}},\quad\bm{W}^{UQ}_{i}\in\mathbb{R}^{d_{c}\times d_{c}},\quad\bm{W}^{K}_{\text{shared}},\bm{W}^{V}_{\text{shared}}\;\in\;\mathbb{R}^{\,d_{\text{model}}\times d_{c}}.$
-Appendix F Other Variants of TPA
+H.5 Rotary Position Embedding (RoPE)
+Many recent LLMs use rotary position embedding (RoPE; 51) to encode positional information in the query/key vectors.
+Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head).
+$\mathbf{T}_{t}$ is a block-diagonal matrix composed of $d/2$ rotation blocks of the form $\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$ for $j\in\{1,\dots,d/2\}$.
+The frequencies $\{\theta_{j}\}$ are typically defined as $\theta_{j}=\text{base}^{-2j/d}$, with a common base like $10000$.
+If $\mathbf{q}_{t}\in\mathbb{R}^{d}$ is a query (or key) row vector for a specific head at position $t$, RoPE is applied as:
+$\operatorname{RoPE}(\mathbf{q}_{t})\triangleq\mathbf{q}_{t}\mathbf{T}_{t}.$
+A key property of RoPE is that the inner product between RoPE-transformed vectors depends only on their relative position.
+For a query $\mathbf{q}_{t}$ and key $\mathbf{k}_{s}$: $(\mathbf{q}_{t}\mathbf{T}_{t})(\mathbf{k}_{s}\mathbf{T}_{s})^{\top}=\mathbf{q}_{t}\mathbf{T}_{t}\mathbf{T}_{s}^{\top}\mathbf{k}_{s}^{\top}=\mathbf{q}_{t}\mathbf{T}_{t-s}\mathbf{k}_{s}^{\top}$.
+This relies on the property:
+$\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (H.4)
+which embeds relative positional information $(t-s)$ into the attention scores.
+Appendix I More on TPA
+Parameter Initialization for TPA Factors.
+We initialize the weight matrices for TPA factors, such as $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, and $\bm{W}_{r}^{b^{V}}$ (or their combined forms $\bm{W}^{a^{Q}}$, $\bm{W}^{b^{Q}}$, etc.), using Xavier initialization [14].
+Specifically, each entry of a weight matrix is drawn from a uniform distribution $\mathcal{U}(-bound,bound)$, where $bound=\sqrt{6/(n_{\text{in}}+n_{\text{out}})}$.
+Here, $n_{\text{in}}$ and $n_{\text{out}}$ are the input and output dimensions of the respective weight matrix.
+This initialization strategy is chosen to help maintain the variance of activations and gradients as they propagate through the network layers, contributing to stable training.
 TPA with Non-contextual B.
-Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to $\mathbf{x}_{t}$.
-For keys:
-$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{K}_{r},$
-and similarly for values. This arrangement is effective if the token-dimension structure remains mostly uniform across the sequence, while the head-dimension factors capture context.
-TPA KV Only. One can preserve a standard query mapping,
-$\mathbf{Q}_{t}=\bm{W}^{Q}\,\mathbf{x}_{t}\in\mathbb{R}^{h\times d_{h}},$
-and factorize only the keys and values. This leaves the query projection as the original linear transformation while reducing memory usage via factorized KV caching.
-TPA KV with Shared $\mathbf{B}$. Another variant is to share the token-dimension factors of keys and values:
-$\mathbf{b}^{K}_{r}(\mathbf{x}_{t})=\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$
-lowering parameter counts and the KV cache footprint.
-While it constrains $\mathbf{K}$ and $\mathbf{V}$ to be formed from the same token basis, it can still perform well and provide additional memory savings.
+In Section A.1, we have introduced TPA with non-contextual A, where head-dimension factors $\mathbf{a}_{r}^{Q},\mathbf{a}_{r}^{K},\mathbf{a}_{r}^{V}\in\mathbb{R}^{h}$ are fixed.
+Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to the input token $\mathbf{x}_{t}$.
+The key tensor for token $t$, $\mathbf{K}_{t}\in\mathbb{R}^{h\times d_{h}}$, would then be constructed as:
+$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{K}_{r}.$
+A similar formulation applies to values.
+This configuration might be effective if the fundamental token-level features (captured by $\mathbf{b}_{r}$) are relatively stable, while their combination across heads (captured by $\mathbf{a}_{r}(\mathbf{x}_{t})$) needs to adapt to the context.
+Performance comparisons for TPA with non-contextual A factors versus non-contextual B factors on small and medium-sized models are presented in Tables 6, 7, 8, and 9.
+Table 6: Evaluation results of small models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+TPA (non-ctx-A) 50.17 25.60 57.95 36.13 31.40 64.80 49.57 24.88 64.80 45.03
+TPA (non-ctx-B) 47.39 26.37 54.8 32.71 30.2 63.38 50.2 23.13 64.8 43.66
+Table 7: Evaluation results of small models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+TPA (non-ctx-A) 55.09 27.65 53.82 36.24 30.20 64.53 50.75 26.01 78.60 46.99
+TPA (non-ctx-B) 50.8   26.96 57.65 32.4 29.4 63.22 49.57 23.96 66.4 44.48
+Table 8: Evaluation results of medium models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+TPA (non-ctx-A) 58.96 31.48 59.76 45.07 34.80 69.21 53.59 25.42 76.40 50.52
+TPA (non-ctx-B) 55.43 29.69 58.32 40.77 34.40 66.92 51.38 25.66 71.10 48.19
+Table 9: Evaluation results of medium models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+TPA (non-ctx-A) 65.45 33.79 56.88 45.23 33.60 68.61 54.22 25.00 85.00 51.98
+TPA (non-ctx-B) 61.20 30.20 55.93 40.45 34.40 68.23 51.78 26.11 78.10 49.60
+TPA KV Only. A simpler variant involves using a standard linear projection for queries,
+$\mathbf{Q}_{t}=\mathbf{W}^{Q}\mathbf{x}_{t}\in\mathbb{R}^{h\times d_{h}},$
+and factorize only the key and value tensors ($\mathbf{K}_{t},\mathbf{V}_{t}$).
+This approach, termed TPA-KVonly, maintains the standard query projection mechanism but still achieves significant KV cache reduction through factorized key and value representations.
+TPA KV with Shared $\mathbf{B}$. Further parameter reduction can be achieved by sharing the token-dimension factors $\mathbf{b}_{r}$ between keys and values:
+[…]
+This sharing reduces both parameter count and the KV cache footprint.
+Although it constrains $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$ to be constructed from the same token-level basis vectors, this variant can still offer strong performance with additional memory savings.
 Nonlinear Head Factors.
-Rather than applying purely linear mappings to the head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}$, one may introduce element-wise nonlinearities such as $\sigma(\cdot)$ or $\operatorname{softmax}(\cdot)$.
-This effectively yields a Mixture of Heads Attention (MoH Attention), where each component becomes a learned mixture weight modulated by the nonlinearity.
+Instead of using purely linear transformations to derive the contextual head-dimension factors $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$, one can introduce element-wise nonlinearities (e.g., sigmoid $\sigma(\cdot)$ or softmax).
+Applying softmax, for instance, to the coefficients that generate $\mathbf{a}_{r}(\mathbf{x}_{t})$ could be interpreted as a form of Mixture-of-Heads, where the network learns to dynamically weight different head configurations based on the input context.
 Discussion.
-These variants illustrate TPA’s versatility in balancing memory cost, computational overhead, and representation power.
-By choosing which dimensions (heads or tokens) remain contextual and adjusting ranks $(R_{Q},R_{K},R_{V})$, TPA unifies multiple existing attention mechanisms—such as MHA, MQA, and GQA—under one framework, while potentially reducing the KV cache size by an order of magnitude during autoregressive inference.
-Appendix G More on Experiments
-G.1 Experimental Settings
+These variants highlight the flexibility of the TPA framework, allowing for different trade-offs between memory efficiency, computational cost, and model expressiveness.
+By carefully choosing which factor components (head-dimension or token-dimension) are contextual versus non-contextual, and by adjusting the ranks $(R_{Q},R_{K},R_{V})$, TPA can not only unify existing mechanisms like MHA, MQA, and GQA but also significantly reduce KV cache size—potentially by an order of magnitude—during autoregressive inference.
+Appendix J More on Experiments
+J.1 Experimental Settings
 We list the main architecture hyper-parameters and training devices in Table 10.
-We fix $d_{h}=64$ for all the models.
-Moreover, we fix the number of KV heads with 2 for GQA models; $d_{h}^{R}=32$ for MLA models; and $R_{k}=R_{v}=2$, $R_{q}=6$ for TPA and TPA-KV only models.
-Other hyper-parameters are listed in Table 5.
+For all models, the head dimension $d_{h}$ is fixed at 64.
+Specific architectural choices include: 2 KV heads for GQA models; a residual key dimension $d_{h}^{R}=32$ for MLA models; and ranks $R_{K}=R_{V}=2$ and $R_{Q}=6$ for TPA and TPA-KVonly models, unless otherwise specified.
+Other relevant hyper-parameters are listed in Table 11.
+Training Setup Details.
+We follow the nanoGPT training configuration [23].
+In particular, we use the AdamW [34] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$.
+We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [35] with $2{,}000$ warmup steps and a (total) global batch size of $480$.
+For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
 Table 10: The architecture hyper-parameters and training devices of models. Abbreviations: BS. = Batch Size, GAS. = Gradient Accumulation Steps.
-Model Size #Param Devices Micro BS. GAS. #Layer $d_{\text{model}}$
+Model Size Parameters Devices Micro BS. GAS. #Layers $d_{\text{model}}$
 Small 124M 4$\times$ A100 GPUs 24 5 12 768
@@
 $d_{c}^{\prime}$ (MLA) 512 1024 1024 1024
-G.2 Additional Experimental Results
-G.2.1 Perplexity Curves
+J.2 Additional Experimental Results
+J.2.1 Perplexity Curves
 We display the perplexity curves for medium, large, and XL size models in Figure 9.
@@
 Figure 9: The validation perplexity of medium-size (353M) models, large-size (773M), and XL-size (1.5B) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
-G.2.2 Ablation Study on Different Ranks
-Figure 5 shows the training loss, validation loss, and validation perplexity curves of XL-size (1.5B) T6 models with different ranks trained on the FineWeb-Edu 100B dataset, and the evaluation results are displayed in Table 7.
-It can be observed that an increase in rank can improve the performance of large language models.
+J.2.2 Ablation Study on Different Ranks
+Figure 10 illustrates the training loss, validation loss, and validation perplexity for XL-sized (1.5B parameters) TPA models with varying key/value ranks ($R_{K}=R_{V}=R$, as indicated in the figure legend), trained on the FineWeb-Edu 100B dataset.
+Corresponding 0-shot evaluation results are presented in Table 13 (rows for TPA-KVonly with different $R_{K,V}$).
+These results indicate that increasing the ranks for key and value factorizations generally improves the performance of the TPA models.
 (a) Training Loss
@@
 (c) Validation Perplexity
-Figure 5: The training loss, validation loss and validation perplexity curves of XL-size (1.5B) T6 models with different ranks on the FineWeb-Edu 100B dataset.
-G.2.3 0-shot Evaluation with lm-evaluation-harness
-For the evaluation, We show the 0-shot performances with lm-evaluation-harness for small-size (124M) and XL-size (1.5B) models in Tables 6 and 7.
-Table 6: The evaluation results of small models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
+Figure 10: The training loss, validation loss and validation perplexity curves of XL-size (1.5B) TPA models with different key/value ranks ($R_{K}=R_{V}=R$) on the FineWeb-Edu 100B dataset.
+J.2.3 0-shot Evaluation with lm-evaluation-harness
+We present 0-shot evaluation results using the lm-evaluation-harness for small (124M parameters) and XL (1.5B parameters) models in Tables 12 and 13, respectively.
+Table 12: Evaluation results of small models (124M) with different attention mechanisms, pre-trained on FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
 The best scores in each column are bolded.
@@
 TPA-KVonly 51.05 26.54 57.25 36.77 32.60 65.02 50.91 23.64 69.70 45.94
-TPA (non-ctx-A) 50.17 25.60 57.95 36.13 31.40 64.80 49.57 24.88 64.80 45.03
 TPA 51.26 27.39 57.00 36.68 32.80 64.47 49.72 24.61 72.00 46.21
-Table 7: The evaluation results of XL models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
+Table 13: Evaluation results of XL models (1.5B) with different attention mechanisms, pre-trained on the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
 The best scores in each column are bolded.
@@
 = WinoGrande.
-If not specified, TPA and TPA-KVonly set $R_{K}=R_{V}=2$.
+If not specified, TPA and TPA-KVonly models use $R_{K}=R_{V}=2$.
 Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
@@
 TPA 66.71 36.52 61.38 54.03 40.40 72.52 56.83 24.49 82.20 55.01
-G.2.4 2-shot Evaluation with lm-evaluation-harness
-We also show 2-shot performances in Tables 8910 and 11.
-Table 8: The evaluation results of small models with different attention mechanisms on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+J.2.4 2-shot Evaluation with lm-evaluation-harness
+Similarly, 2-shot evaluation results are provided in Tables 14 (Small), 15 (Medium), 16 (Large), and 17 (XL).
+Table 14: Evaluation results of small models (124M) with different attention mechanisms, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
 The best scores in each column are bolded.
@@
 TPA-KVonly 54.25 27.90 57.06 36.36 31.80 64.31 53.59 26.18 79.20 47.85
-TPA (non-ctx-A) 55.09 27.65 53.82 36.24 30.20 64.53 50.75 26.01 78.60 46.99
 TPA 57.53 28.07 56.33 36.49 31.80 64.36 51.14 25.92 79.70 47.93
-Table 9: The evaluation results of medium models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+Table 15: Evaluation results of medium models (353M) with different attention mechanisms, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness, default LR $6\times 10^{-4}$).
 The best scores in each column are bolded.
@@
 TPA-KVonly 64.69 32.34 59.48 46.23 35.40 70.08 54.06 25.64 86.30 52.69
-TPA (non-ctx-A) 65.45 33.79 56.88 45.23 33.60 68.61 54.22 25.00 85.00 51.98
 TPA 67.97 34.56 57.22 46.87 34.60 69.91 52.01 25.07 89.90 53.12
-Table 10: The evaluation results of large models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+Table 16: Evaluation results of large models (772M) with different attention mechanisms, pre-trained on the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
 The best scores in each column are bolded.
-Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
-Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
 MHA 67.85 36.35 59.82 50.22 35.00 70.67 53.35 23.92 91.10 54.25
@@
 TPA 70.41 37.71 60.06 51.30 34.00 71.06 54.54 25.79 90.30 55.02
-Table 11: The evaluation results of XL models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
+Table 17: Evaluation results of XL models (1.5B) with different attention mechanisms, pre-trained on the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
 The best scores in each column are bolded.
-Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
-If not specified, We set $R_{K}=R_{V}=2$ for TPA and TPA-KVonly.
-Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
+Abbreviations: HellaSw.
+= HellaSwag, W.G.
+= WinoGrande.
+If not specified, $R_{K}=R_{V}=2$ for TPA and TPA-KVonly models.
+Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
 MHA 70.83 39.93 59.85 54.05 36.20 72.52 55.17 25.42 91.70 56.18
@@
 TPA 71.76 39.16 61.25 53.74 37.80 72.80 55.49 23.86 90.70 56.28
-G.3 Ablation Studies on Learning Rates
-We implement a set of parallel experiments for medium models with learning rate $3\times 10^{-4}$, and the curves for training loss, validation loss, and validation perplexity are displayed in Figure 6.
-We also show the performance of these models on the benchmarks described in Section 4 in Tables 12-13.
-The results show that TPA and TPA-KVonly models can also outperform other types of attention with different learning rates.
+J.3 Ablation Studies on Learning Rates
+To assess sensitivity to learning rates, we conducted parallel experiments on medium-sized models using a learning rate of $3\times 10^{-4}$ (compared to the default $6\times 10^{-4}$ used for other medium model results).
+The training loss, validation loss, and validation perplexity curves are shown in Figure 11.
+Performance on standard benchmarks for these models trained with the $3\times 10^{-4}$ learning rate are reported in Tables 18 (0-shot) and 19 (2-shot).
+The results demonstrate that TPA and TPA-KVonly maintain their performance advantages over other attention mechanisms even with this alternative learning rate.
 (a) Training Loss
@@
 (c) Validation Perplexity
-Figure 6: The training loss, validation loss, and validation perplexity of medium-size (353M) models (learning rate $3\times 10^{-4}$) and different attention mechanisms on the FineWeb-Edu 100B dataset.
+Figure 11: The training loss, validation loss, and validation perplexity of medium-size (353M) models (learning rate $3\times 10^{-4}$) with different attention mechanisms on the FineWeb-Edu 100B dataset.
 Table 18: The evaluation results of medium models (learning rate $3\times 10^{-4}$) with different attention mechanisms pretrained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
@@
 MHA 64.44 32.85 59.05 44.18 33.20 68.72 50.12 26.01 87.40 49.44
-MQA 64.27 32.94 57.71 44.36 31.80 68.01 51.70 25.99 86.00 49.29
+MQA 64.27 32.94 57.71 44.36 31.80 68.01 51.70 25.99 86.00 51.42
 GQA 61.70 32.17 52.81 43.99 33.80 68.50 53.35 24.44 86.40 50.80
 MLA 65.95 31.48 50.98 44.99 32.20 68.93 51.93 25.89 88.80 51.24
-TPA-KVonly 65.99 33.70 57.49 44.47 34.20 69.53 53.28 24.23 86.50 49.93
+TPA-KVonly 65.99 33.70 57.49 44.47 34.20 69.53 53.28 24.23 86.50 52.15
 TPA 66.54 34.47 58.96 45.35 33.00 69.21 53.99 24.51 91.30 53.04

### v5

@@
 Notably, TPA’s memory efficiency and computational efficiency at the decoding stage enable processing longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models.
-The code is available at https://github.com/tensorgi/T6.
+Project Page: https://github.com/tensorgi/TPA.
 1 Introduction
@@
 In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance.
-A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer, ), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama, ) and Gemma (team2024gemma, ).
+A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer, ), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama, ), Qwen (bai2023qwen, ), and Gemma (team2024gemma, ).
 Our main contributions are summarized as follows:
@@
 2 Background
-In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention, ) and introduce key notations.
-Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast, ), Grouped Query Attention (GQA) (ainslie2023gqa, ), Multi-head Latent Attention (MLA) (liu2024deepseek, ; liu2024deepseekv3, ), and Rotary Position Embedding (RoPE) (su2024roformer, ) are further discussed in Appendix H.
+In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention, ), and introduce key notations.
+Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast, ), Grouped Query Attention (GQA) (ainslie2023gqa, ), Multi-head Latent Attention (MLA) (liu2024deepseek, ; liu2024deepseekv3, ), and Rotary Position Embedding (RoPE) (su2024roformer, ) are further discussed in the Appendix F.
 Notations.
@@
 MHA enables the model to capture a rich set of dependencies by allowing each head to focus on different aspects of the input sequence.
-We also discuss how MHA relates to TPA in Appendix A.
-Further discussion on Transformers and attention mechanisms, KV cache optimization techniques, and low-rank factorization methods is available in Appendix G.
+We also discuss how MHA, MQA, and GQA relate to TPA in the Section 4.
 3 Tensor Product Attention
@@
 Specifically, for each token $t$, with a small abuse of notation, we define:
-[…]
+$\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\hskip 21.52771pt\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$
 $\hskip 77.49976pt\mathbf{V}_{t}=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$ (3.1)
-where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$.
+where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$.
 Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ (an outer product) contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$.
@@
 $\operatorname{TPA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl(\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr)\bm{W}^{O}.$ (3.3)
-Parameter Initialization. We use Xavier initialization (glorot2010understanding, ) for the factor weight matrices; details are in Appendix I.
+Parameter Initialization. We use Xavier initialization (glorot2010understanding, ) for the factor weight matrices; details are in the Appendix G.
 3.2 RoPE Compatibility and Acceleration
@@
 $\operatorname{RoPE}_{t-s}(\mathbf{Q}_{t})\mathbf{K}_{s}^{\top}=\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top},$
-where $\operatorname{RoPE}_{t-s}$ denotes applying RoPE with relative position $t-s$.
+where $\operatorname{RoPE}_{t-s}$ denotes applying RoPE with relative position tst-s.
 Focusing on individual heads $i$, if $\mathbf{q}_{t,i}$ and $\mathbf{k}_{s,i}$ are the $i$-th head vectors from $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ respectively (as column vectors of dimension $d_{h}$), and $\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i},t)$ and $\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i},s)$, then the dot product for attention scores satisfies:
 $\widetilde{\mathbf{q}}_{t,i}^{\top}\,\widetilde{\mathbf{k}}_{s,i}=\mathbf{q}_{t,i}^{\top}\operatorname{RoPE}(\cdot,t-s)\mathbf{k}_{s,i}.$
-Theorem 3.1 indicates that TPA does not break RoPE’s relative translational property.
-We prove Theorem 3.1 in Appendix F.1.
-In essence, $\operatorname{RoPE}_{t}$ applies a linear transformation (rotation matrix $\mathbf{R}_{t}$) to the $d_{h}$-dimensional space.
-Since $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$, applying RoPE yields $\mathbf{Q}_{t}\mathbf{R}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}(\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{R}_{t})$.
-Thus, RoPE effectively transforms $\mathbf{B}_{Q}(\mathbf{x}_{t})$ to $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{R}_{t}$, where $\mathbf{R}_{t}$ acts on each row of $\mathbf{B}_{Q}(\mathbf{x}_{t})$.
-The $\mathbf{A}_{Q}(\mathbf{x}_{t})$ factor remains unchanged, preserving the TPA structure.
+Theorem 3.1 indicates that TPA does not break RoPE’s relative translational property. We prove it in the Appendix D.1.
 3.3 KV Caching and Memory Reduction
@@
 Here, $R_{Q}$, $R_{K}$, and $R_{V}$ denote the ranks for queries, keys, and values in TPA, respectively.
-Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in Appendix I.
+Variants of TPA, such as TPA (KVonly), TPA (Non-contextual A), and TPA (Non-contextual B), are detailed in the Appendix G.
 For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
@@
 MHA $2hd_{h}$ $4d_{\text{model}}^{2}$ $h$ $h$
-MQA $2d_{h}$ $(2+2/h)d_{\text{model}}^{2}$ $h$ $1$
+MQA $2d_{h}$ $(2+2/h)d_{\text{model}}^{2}$ $h$ 11
 GQA $2gd_{h}$ $(2+2g/h)d_{\text{model}}^{2}$ $h$ $g$
-MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $+d_{\text{model}}d_{h}^{R}+d_{c}(d_{\text{model}}+2hd_{h})$ $h$ $h$
+MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $d_{\text{model}}d_{h}^{R}+d_{c}(d_{\text{model}}+2hd_{h})$ $h$ $h$
 TPA $(R_{K}+R_{V})(h+d_{h})$ $d_{\text{model}}(R_{Q}+R_{K}+R_{V})(h+d_{h})+d_{\text{model}}\,hd_{h}$ $h$ $h$
@@
 TPA (Non-contextual B) $(R_{K}+R_{V})h$ $(R_{Q}+R_{K}+R_{V})(d_{\text{model}}h+d_{h})+d_{\text{model}}\,hd_{h}$ $h$ $h$
-3.4 Expressing MHA, MQA, and GQA as Non-contextual TPA
-We demonstrate in Appendix A that several existing attention mechanisms, including MHA, MQA, and GQA, can be elegantly expressed as non-contextual variants of TPA by imposing specific constraints on the TPA factors.
+4 Expressing MHA, MQA, GQA as Non-contextual TPA
+We demonstrate that standard Multi-Head Attention (MHA), Multi-Query Attention (MQA), and Grouped-Query Attention (GQA) can be expressed as special, non-contextual variants of Tensor Product Attention (TPA).
+This is achieved by imposing specific constraints on the TPA factors, particularly by making the head-dimension factors ($\mathbf{a}$) independent of the input token ($\mathbf{x}_{t}$).
+4.1 MHA as Non-contextual TPA
+Standard Multi-Head Attention (MHA) can be precisely formulated as a TPA where the rank is equal to the number of heads ($R_{Q}=R_{K}=R_{V}=h$), and the head-dimension factors are fixed, non-contextual basis vectors.
+To recover MHA, we set the rank $R_{Q}=h$ and define the factors for each head $i\in[h]$ as follows:
+  • Contextual token factor: This is the standard linear projection for the $i$-th head’s query:
+    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=(\bm{W}^{Q}_{i})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}}$
+  • Non-contextual head factor: This factor is a scaled standard basis vector, independent of $\mathbf{x}_{t}$:
+    $\mathbf{a}^{Q}_{i}=h\cdot\mathbf{e}_{i}\in\mathbb{R}^{h}$
+    where $\mathbf{e}_{i}$ is the $i$-th standard basis vector (a vector of zeros with a one at the $i$-th position).
+Substituting these into the TPA equation, the $1/R_{Q}=1/h$ scaling factor cancels with the scaling of the $\mathbf{a}^{Q}_{i}$ factor:
+$\mathbf{Q}_{t}$ $=\frac{1}{h}\sum_{i=1}^{h}\left(h\cdot\mathbf{e}_{i}\right)\otimes\left((\bm{W}^{Q}_{i})^{\top}\mathbf{x}_{t}\right)=\sum_{i=1}^{h}\mathbf{e}_{i}\otimes\left((\bm{W}^{Q}_{i})^{\top}\mathbf{x}_{t}\right)$
+The resulting tensor product, $\mathbf{e}_{i}\otimes\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$, produces an $h\times d_{h}$ matrix where only the $i$-th row is non-zero and contains the vector $(\mathbf{b}^{Q}_{i}(\mathbf{x}_{t}))^{\top}$.
+Summing these matrices for $i=1,\dots,h$ assembles the complete query tensor $\mathbf{Q}_{t}$, where the $i$-th row is precisely the query vector for the $i$-th head in standard MHA.
+An analogous construction applies to the key ($\mathbf{K}_{t}$) and value ($\mathbf{V}_{t}$) tensors.
+Thus, MHA is equivalent to a non-contextual TPA where the head-dimension factors are fixed and orthogonal, effectively assigning a dedicated rank component to each attention head.
+4.2 MQA and GQA as Non-contextual TPA
+Similarly, Multi-Query Attention (MQA) and Grouped-Query Attention (GQA) can be seen as non-contextual TPAs where the key and value tensors are formed with a rank lower than the number of heads.
+  • MQA as Rank-1 TPA (for K and V).
+In MQA, all $h$ query heads share a single key and value.
+This corresponds to a TPA with ranks $R_{K}=1$ and $R_{V}=1$.
+The key tensor $\mathbf{K}_{t}$ is formed using a single, non-contextual head-dimension factor $\mathbf{a}^{K}=\mathbf{1}_{h}$ (a vector of all ones) and a single contextual token-dimension factor $\mathbf{b}^{K}(\mathbf{x}_{t})=(\bm{W}^{K})^{\top}\mathbf{x}_{t}$:
+    $\mathbf{K}_{t}=\frac{1}{1}\left(\mathbf{1}_{h}\otimes\mathbf{b}^{K}(\mathbf{x}_{t})\right)$
+    This creates an $h\times d_{h}$ matrix where every row is the same shared key vector $(\mathbf{b}^{K}(\mathbf{x}_{t}))^{\top}$.
+The same logic applies to the value tensor $\mathbf{V}_{t}$.
+The queries remain full-rank ($R_{Q}=h$) as in MHA.
+  • GQA as Rank-G TPA (for K and V).
+GQA is an intermediate approach where $h$ heads are divided into $G$ groups, with heads in the same group sharing a key and value.
+This is equivalent to a TPA with ranks $R_{K}=G$ and $R_{V}=G$.
+The key tensor is formed by summing $G$ components:
+    $\mathbf{K}_{t}=\frac{1}{G}\sum_{j=1}^{G}\mathbf{a}^{K}_{j}\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$
+    Here, $\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$ is the shared key vector for group $j$.
+The non-contextual factor $\mathbf{a}^{K}_{j}$ is a scaled mask vector, defined as $\mathbf{a}^{K}_{j}=G\cdot\text{mask}_{j}$, where the $\text{mask}_{j}$ vector has ones for heads belonging to group $j$ and zeros elsewhere.
+This scaling cancels the $1/G$ pre-factor:
+    $\mathbf{K}_{t}=\frac{1}{G}\sum_{j=1}^{G}(G\cdot\text{mask}_{j})\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t})=\sum_{j=1}^{G}\text{mask}_{j}\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$
+    For example, with $h=8$ heads and $G=2$ groups ($2$ KV heads), the factor for the first group of 4 heads would be $\mathbf{a}^{K}_{1}=2\cdot[1,1,1,1,0,0,0,0]^{\top}$.
+This construction correctly assembles the final key tensor by broadcasting each group’s shared key to its designated heads without any unintended extra scaling.
+This perspective highlights that MHA, MQA, and GQA are specific instances of a more general TPA framework, where expressiveness and parameter sharing are controlled by the rank and the nature (contextual vs.
+non-contextual) of the tensor factors.
 4.3 Model Architectures
 We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention).
-Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilize the overall architecture of LLaMA (touvron2023llama, ) while changing the self-attention block to our TPA-based version.
+Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (touvron2023llama, ) while changing the self-attention block to our TPA-based version.
 The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu, ; touvron2023llama, ).
-The specific mechanisms for TPA-based query, key, and value factorization, as well as the attention computation within T6, directly follow the comprehensive descriptions provided in Section 3 (particularly Section 3.1 for factorization, and Equations 3.2 and 3.3 for attention computation).
-A brief recap of these steps within the T6 model context is provided in Appendix B.
 Rotary Positional Embedding (RoPE).
 As discussed in Section 3.2, RoPE (su2024roformer, ) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
-Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see (3.5) and Theorem 3.1.
+Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see Equation (3.5) and Theorem 3.1.
 SwiGLU Feed-Forward Network.
@@
 5 FlashTPA Decoding Algorithm
+For efficient autoregressive inference with Tensor Product Attention (TPA), we introduce FlashTPA Decoding.
+This algorithm is optimized for generating one token at a time by leveraging the factorized representation of queries, keys, and values.
+The core idea, illustrated in Figure 2, is to perform attention computations using a sequence of Einstein summations (“einsum”) that operate directly on these factorized components.
+This avoids materializing the full query, key, and value tensors, which is particularly beneficial as the Key-Value (KV) cache grows with sequence length.
+The detailed definitions of the input factorized components and the step-by-step pseudo-code for FlashTPA Decoding are provided in Algorithm 2.
+An optimized Triton kernel implementation is outlined inAlgorithm 3 (see Appendix B.1).
 $\mathbf{B}_{Q}$
@@
 Scaling factors in softmax are omitted for visual clarity.
-For efficient autoregressive inference with Tensor Product Attention (TPA), we introduce FlashTPA Decoding.
-This algorithm is optimized for generating one token at a time by leveraging the factorized representation of queries, keys, and values.
-The core idea, illustrated in Figure 2, is to perform attention computations using a sequence of Einstein summations (“einsum”) that operate directly on these factorized components.
-This avoids materializing the full query, key, and value tensors, which is particularly beneficial as the Key-Value (KV) cache grows with sequence length.
-The detailed definitions of the input factorized components and the step-by-step pseudocode for FlashTPA Decoding are provided in Algorithm 1 (see Appendix D for details).
-An optimized Triton kernel implementation: Algorithm 3 is also outlined in Appendix D.
 This sequence of factorized operations allows FlashTPA Decoding to compute the attention output efficiently.
-It minimizes memory allocations for large intermediate tensors and reduces the overall computational load compared to materializing full query, key, and value tensors, especially for long sequences.
 Consequently, TPA is not only memory-efficient due to its smaller KV cache footprint but can also be computationally efficient during inference.
 The experimental results for FlashTPA decoding time are presented in Section 6.2.
-Algorithm 1 FlashTPA Decoding Algorithm
-1:$\mathbf{A}_{Q}$: Query A tensor $(B,1,H,R_{Q})$
-2:$\mathbf{B}_{Q}$: Query B tensor $(B,1,R_{Q},D)$
-3:$\mathbf{a}^{K}_{\text{cache}}$: Cached Key A tensor $(B,M,H)$
-4:$\bm{b}^{K}_{\text{cache}}$: Cached Key B tensor $(B,M,D)$
-5:$\mathbf{a}^{V}_{\text{cache}}$: Cached Value A tensor $(B,M,H)$
-6:$\bm{b}^{V}_{\text{cache}}$: Cached Value B tensor $(B,M,E)$
-7:$s_{\text{total}},s_{Q},s_{K},s_{V}$: Scaling factors (with defaults $s_{\text{total}}\leftarrow 1/\sqrt{D}$, $s_{Q}\leftarrow 1/R_{Q}$)
-8:$\mathbf{O}$: Output tensor $(B,1,H,E)$
-9$\triangleright$ Step 1: Interaction between $\mathbf{B}_{Q}$ and $\bm{b}^{K}_{\text{cache}}$ components
-10:$S^{(1)}\leftarrow\text{einsum}(\text{``bnrd, bmd}\rightarrow\text{bnmr''},\mathbf{B}_{Q},\bm{b}^{K}_{\text{cache}})$triangleright Shape: $(B,1,M,R_{Q})$.
-$S^{(1)}_{b,n,m,r}=\sum_{d}(\mathbf{B}_{Q})_{b,n,r,d}\cdot(\bm{b}^{K}_{\text{cache}})_{b,m,d}$
-11$\triangleright$ Step 2: Incorporate $\mathbf{A}_{Q}$ component
-12:$S^{(2)}\leftarrow\text{einsum}(\text{``bnhr, bnmr}\rightarrow\text{bnmh''},\mathbf{A}_{Q},S^{(1)})$triangleright Shape: $(B,1,M,H)$.
-$S^{(2)}_{b,n,m,h}=\sum_{r}(\mathbf{A}_{Q})_{b,n,h,r}\cdot S^{(1)}_{b,n,m,r}$
-13$\triangleright$ Step 3: Incorporate $\mathbf{a}^{K}_{\text{cache}}$ component to get full logits
-14:einsum(“bnmh, $\mathcal{L}\leftarrow\text{einsum}(\text{``bnmh, bmh}\rightarrow\text{bhnm''},S^{(2)},\mathbf{a}^{K}_{\text{cache}})$triangleright Shape: $(B,H,1,M)$.
-$\mathcal{L}_{b,h,n,m}=S^{(2)}_{b,n,m,h}\cdot(\mathbf{a}^{K}_{\text{cache}})_{b,m,h}$
-15$\triangleright$ Step 4: Apply scaling and Softmax
-16:$\bm{\alpha}\leftarrow\operatorname{Softmax}_{M}(\mathcal{L}\cdot s_{Q}\cdot s_{K}\cdot s_{\text{total}})$triangleright Shape: $(B,H,1,M)$. Softmax over cache dimension $M$
-17$\triangleright$ Step 5: Compute weighted sum with $\mathbf{a}^{V}_{\text{cache}}$ component
-18:$\mathbf{O}^{(A)}\leftarrow\text{einsum}(\text{``bhnm, bmh}\rightarrow\text{bnmh''},\bm{\alpha},\mathbf{a}^{V}_{\text{cache}})$triangleright Shape: $(B,1,M,H)$.
-$\mathbf{O}^{(A)}_{b,n,m,h}=\bm{\alpha}_{b,h,n,m}\cdot(\mathbf{a}^{V}_{\text{cache}})_{b,m,h}$
-19$\triangleright$ Step 6: Incorporate $\bm{b}^{V}_{\text{cache}}$ component and final scaling
-[…]
-21:return $\mathbf{O}$
 6 Experiments
 6.1 Language Modeling Tasks
-All experiments reported in this paper are implemented on the nanoGPT code base (Karpathy2022, ), using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu, ).
+All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy2022, ), and we pretrain our models using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu, ).
 The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation.
@@
 We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B).
-Details on architecture hyperparameters and training hardware are shown in Appendix J.1.
+We pretrain all models with 50B tokens using 1 epoch.
+Details on architecture hyperparameters and training hardware are shown in Appendix H.1.
 Training & Validation Curves.
@@
 Training loss curves are provided in Appendix Figure 3.
-Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix I) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses.
+Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix G) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses.
 For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages.
@@
 Mirroring the loss curves, TPA and TPA-KVonly steadily outperform MHA, MQA, GQA, and MLA over the course of training.
-By the end of pretraining (around $49$B tokens), TPA-based approaches achieve the lowest perplexities in most configurations.
+By the end of pretraining (around 4949B tokens), TPA-based approaches achieve the lowest perplexities in most configurations.
 Downstream Evaluation.
@@
 Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3.
-Zero-shot evaluation of small and XL models are displayed in Tables 12 and 13 in the appendix.
-Moreover, we also present 2-shot evaluation results in Tables 141516 and 17 in the appendix.
+Zero-shot evaluation of small and XL models are displayed in Tables 11 and 12 in the appendix.
+Moreover, we also present 2-shot evaluation results in Tables 131415 and 16 in the appendix.
 For the medium-size (353M) models (Table 2 for 0-shot and Table 14 in appendix for 2-shot), TPA generally ties or outperforms all competing methods, achieving, for example, an average of 51.41% in zero-shot mode versus MHA’s 50.11%, MQA’s 50.44%, and MLA’s 50.13%.
@@
 6.2 Experimental Results on FlashTPA Decoding
+This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms.
+We benchmark FlashTPA against FlashMHA (shah2024flashattention, ), FlashGQA, FlashMQA, and FlashMLA (flashmla2025, ).
+It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton, ).
+While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential.
+Development of a CUDA-based FlashTPA kernel is ongoing and is expected to yield further performance improvements.
 (a) Batch Size=1
@@
 Each subfigure corresponds to a different batch size.
-This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms.
-We benchmark FlashTPA against FlashMHA (shah2024flashattention, ), FlashGQA, FlashMQA, and FlashMLA (flashmla2025, ).
-It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton, ).
-While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential.
-Development of a CUDA-based FlashTPA kernel is ongoing and is expected to yield further performance improvements.
 The evaluations were performed with batch sizes selected from $\{1,2,4,8,16\}$, model embedding dimensions ($d_{\text{model}}$) chosen from $\{1024,2048,3072\}$, and sequence lengths ranging from $2^{12}$ (4,096) to $2^{19}$ (524,288).
@@
 The ranks for TPA’s factorized components ($R_{Q},R_{K},R_{V}$) were set to $(16,1,1)$, and for GQA configurations, the number of key-value head groups was 4.
-The decoding time per token, measured as $\log_{2}(\text{time})$ in seconds, is plotted against
-$\log_{2}(\text{sequence length})$.
+The decoding time per token, measured as $\log_{2}(\text{time})$ in seconds, is plotted against $\log_{2}(\text{sequence length})$.
 Lower values on the y-axis indicate faster decoding times.
 Results are presented in Figure 5 for an embedding dimension of 2048 (corresponding to 32 attention heads).
-Additional results for embedding dimensions of 1024 (16 heads, Figure 8) and 3072 (48 heads, Figure 7) are provided in Appendix D.
+Additional results for embedding dimensions of 1024 (16 heads, Figure 8) and 3072 (48 heads, Figure 7) are provided in Appendix B.
 Figure 5 depicts these speed comparisons for an embedding dimension of 2048.
 The results indicate that FlashTPA (blue line) is highly competitive and often outperforms other attention mechanisms, especially as the sequence length increases.
-A detailed breakdown of these comparisons across different batch sizes for the $d_{\text{model}}=2048$ case (Figure 5) is provided in Appendix D.
-These findings underscore the computational efficiency of the FlashTPA decoding algorithm, particularly its favorable scaling with increasing sequence lengths.
 7 Conclusion
@@
 Overall, TPA offers a powerful mechanism for compressing KV storage while improving the model performance, thereby enabling longer sequence contexts under constrained memory.
-Appendix A Expressing MHA, MQA, and GQA as Non-contextual TPA
-This appendix section details how standard Multi-Head Attention (MHA), Multi-Query Attention (MQA), and Grouped-Query Attention (GQA) can be expressed as special, non-contextual variants of Tensor Product Attention (TPA).
-A.1 MHA as Non-contextual TPA
-Standard multi-head attention (MHA) can be viewed as a specific instance of TPA in which: 1) the rank is set equal to the number of heads; 2) the head dimension factor is non-contextual (i.e., independent of the $t$-th token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$); 3) the token dimension factor is a linear function of $\mathbf{x}_{t}$.
-More precisely, if we set $R_{Q}=R_{K}=R_{V}=h$ (the number of heads) and define the factors appropriately, MHA emerges.
-To match MHA with TPA, let $R_{Q}=R_{K}=R_{V}=h$.
-Focusing on $\mathbf{Q}_{t}$:
-  1. (a)
-    Non-contextual head factors. Define
-    $\mathbf{a}^{Q}_{i}=h\mathbf{e}_{i}\in\mathbb{R}^{h},$ (A.1)
-    where $\mathbf{e}_{i}\in\mathbb{R}^{h}$ is the $i$-th standard basis vector, so that $\mathbf{e}_{i}\otimes\cdot$ corresponds to the $i$-th head of $\mathbf{Q}_{t}$.
-  2. (b)
-    Contextual token factors. Define
-    $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})=({\bm{W}^{Q}_{i}})^{\top}\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$ (A.2)
-    where $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the per-head query projection matrix, hence $\mathbf{b}^{Q}_{i}(\mathbf{x}_{t})$ is dependent on $\mathbf{x}_{t}$.
+Acknowledgements
+We thank the anonymous reviewers and area chair for their helpful comments. We acknowledge the compute credits provided by Fetch.ai.
+Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
+Our objective in this section is to compute attention without explicitly forming $\mathbf{Q},\mathbf{K},\mathbf{V}$, by contracting their factorized representations in a cache- and throughput-friendly order.
+Recall from Section˜3.1 that each per-token slice $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ is a sum of rank-$1$ outer products.
+Unless otherwise stated we use the per-factor normalizations $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$.
+We make the batch/time/head/rank/value dimensions explicit and introduce the shorthands $D:=d_{h}$ and $E:=d_{v}$ (typically E=DE{=}D):
+$\mathbf{A}_{Q}\in\mathbb{R}^{B\times T_{q}\times h\times R_{Q}},\quad\mathbf{B}_{Q}\in\mathbb{R}^{B\times T_{q}\times R_{Q}\times D},\qquad\mathbf{A}_{K}\in\mathbb{R}^{B\times T_{k}\times h\times R_{K}},\quad\mathbf{B}_{K}\in\mathbb{R}^{B\times T_{k}\times R_{K}\times D},$
+$\mathbf{A}_{V}\in\mathbb{R}^{B\times T_{k}\times h\times R_{V}},\qquad\mathbf{B}_{V}\in\mathbb{R}^{B\times T_{k}\times R_{V}\times E}.$
+Indices $b,q,k,h,r,s,u,d,e$ denote batch, query position, key position, head, query-rank, key-rank, value-rank, feature ($D$), and value feature ($E$).
+We write $T:=T_{q}\!=\!T_{k}$ for full-sequence attention; in decoding, $T_{q}\!=\!1$ and we denote the cache length by $M\!=\!T_{k}$.
+High-level idea.
+We first compute head-shared feature-space dot products between $\mathbf{B}_{Q}$ and $\mathbf{B}_{K}$, then mix them with head-specific $\mathbf{A}_{Q},\mathbf{A}_{K}$ to obtain logits, apply the masked softmax, and finally aggregate values via $\mathbf{A}_{V},\mathbf{B}_{V}$.
+This ordering avoids materializing any $T_{q}\times h\times D$ queries/keys/values.
 […]
-$\mathbf{Q}_{t}=\sum_{i=1}^{h}\Big{[}\mathbf{e}_{i}\otimes\bigl{(}(\bm{W}^{Q}_{i})^{\top}\,\mathbf{x}_{t}\bigr{)}\Big{]}\in\mathbb{R}^{h\times d_{h}}.$ (A.3)
+Figure 6: Specialized TPA computation without materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$.
+Phase 1 (top): compute head-shared feature-space dot products $P[b,q,k,r,s]{=}\langle\mathbf{B}_{Q}[b,q,r,:],\,\mathbf{B}_{K}[b,k,s,:]\rangle$ and mix them with head-specific factors $\mathbf{A}_{Q},\mathbf{A}_{K}$ to obtain logits $\mathcal{L}[b,h,q,k]$.
+Phase 2 (bottom): apply the causal/padding mask and softmax to get $\alpha[b,h,q,k]$, then aggregate values via $\mathbf{A}_{V},\mathbf{B}_{V}$.
+Scalings $s_{Q},s_{K},s_{V}$ and $1/\sqrt{D}$ are folded into the corresponding phases.
+Dropout is omitted for clarity.
+Batch $B$, heads $H$, ranks $R_{Q},R_{K},R_{V}$, and feature dims $D,E$ are indicated in the nodes.
+A.1 Direct computation in factor space
+Single head. For a fixed head $h\!\in\![H]$ and token indices $(q,k)$, using $s_{Q}{=}1/R_{Q}$ and $s_{K}{=}1/R_{K}$ we have
+$\bigl[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}\bigr]_{q,k}\;=\;\frac{1}{R_{Q}R_{K}}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\,a^{(r)}_{q,h}(\mathbf{x}_{q})\,a^{(s)}_{k,h}(\mathbf{x}_{k})\;\big\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{q}),\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{k})\big\rangle,$ (A.1)
+and for values (with $s_{V}{=}1/R_{V}$), $\mathbf{V}^{(h)}_{k}=\frac{1}{R_{V}}\sum_{u=1}^{R_{V}}a^{(u)}_{k,h}(\mathbf{x}_{k})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{k})$.
+The per-head attention output at query position $q$ is then $\sum\nolimits_{k}\operatorname{softmax}\!\big(\frac{1}{\sqrt{D}}[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}]_{q,:}\big)_{k}\,\mathbf{V}^{(h)}_{k}$.
+Multi-head with head-shared feature dot-products.
+Define head-shared feature-space dot products $P[b,q,k,r,s]\!=\!\langle\mathbf{B}_{Q}[b,q,r,:],\mathbf{B}_{K}[b,k,s,:]\rangle$.
+With $\mathcal{S}\!=\!\frac{s_{Q}s_{K}}{\sqrt{D}}$, we compute
+$\mathcal{L}[b,h,q,k]$ $=\mathcal{S}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\mathbf{A}_{Q}[b,q,h,r]\;\mathbf{A}_{K}[b,k,h,s]\;P[b,q,k,r,s],$ (A.2)
+$\alpha[b,h,q,k]$ $=\operatorname{Softmax}_{k}\!\left(\mathcal{L}[b,h,q,k]+\log\mathsf{mask}[q,k]\right),$
+$\mathbf{O}[b,h,q,e]$ $=s_{V}\sum_{k=1}^{T_{k}}\sum_{u=1}^{R_{V}}\alpha[b,h,q,k]\;\mathbf{A}_{V}[b,k,h,u]\;\mathbf{B}_{V}[b,k,u,e].$ (A.3)
+Here $\mathsf{mask}[q,k]\!\in\!\{0,-\infty\}$ enforces causality and padding.
+Eqs.
+(A.2)–(A.3) make explicit that (i) feature-space dot products $P$ are head-shared, and (ii) all $1/R_{(\cdot)}$ scalings can be folded into a single multiplicative factor per phase without affecting numerical stability (softmax shift invariance).
+A.2 Complexity: materialized vs. specialized computation
+We compare two execution strategies.
+(i) Naïve/materialized: form $\mathbf{Q},\mathbf{K},\mathbf{V}$ explicitly and call standard kernels.
+(ii) Specialized: compute via Eq.
+(A.2)–(A.3) using head-shared feature-dot products and per-head rank contractions.
+Standard MHA (baseline).
+Ignoring projections, full-sequence attention uses $\Theta(BHT^{2}D)$ FLOPs for scores and $\Theta(BHT^{2}D)$ for value aggregation, i.e., $\mathcal{F}_{\text{MHA}}\;=\;2\,\Theta(BHT^{2}D)$.
+TPA (materialized).
+Forming $\mathbf{Q},\mathbf{K},\mathbf{V}$ from factors costs $\Theta(BTHD(R_{Q}{+}R_{K}{+}R_{V}))$ after the linear projections; subsequent attention uses the same $2\,\Theta(BHT^{2}D)$ as MHA.
+TPA (specialized).
+Using Eqs. (A.2)–(A.3) and writing $T_{q}{=}T_{k}{=}T$, the dominant FLOPs are
+$\underbrace{\Theta(B\,T^{2}\,R_{Q}R_{K}\,D)}_{\text{feature dots }P}\;+\;\underbrace{\Theta(B\,H\,T^{2}\,R_{Q}R_{K})}_{\text{per-head rank combine}}\;+\;\underbrace{\Theta(B\,H\,T^{2}\,R_{V}\,E)}_{\text{value aggregation}}.$
+Compared to $\mathcal{F}_{\text{MHA}}$, the specialized path reduces FLOPs whenever
+$R_{Q}R_{K}\,D\;+\;H\,R_{Q}R_{K}\;+\;H\,R_{V}\,E\;<\;2\,H\,D.$ (A.4)
+Dividing by HDHD yields $(R_{Q}R_{K}/H)+(R_{Q}R_{K}/D)+R_{V}(E/D)<2$.
+For E=DE{=}D and small ranks (e.g., $R_{Q}{=}R_{K}{=}R_{V}{=}1$), the inequality holds for typical $H,D\!\geq\!2$ and the benefit grows with larger $H$ or $D$.
+Memory traffic and peak working set.
+For full-sequence attention the naive path streams $\mathbf{Q},\mathbf{K},\mathbf{V}$ of size $\Theta(BTHD)$ each.
+The specialized path streams factors only and needs the head-shared $P$ tiles of size $\Theta(B\,T_{q}^{\text{tile}}\,T_{k}^{\text{tile}}\,R_{Q}R_{K})$ plus per-head tiles for the rank combine/value aggregation.
+In decoding with cache length $M$, the factorized KV cache uses $(R_{K}{+}R_{V})(h{+}D)$ numbers per token (cf.
+Section 3.3), vs.
+$2hD$ for MHA; this reduction directly lowers memory bandwidth pressure.
+Algorithm 1 Specialized TPA (no explicit $\mathbf{Q},\mathbf{K},\mathbf{V}$; causal)
 […]
-TPA with Non-contextual A factors.
-More broadly, TPA can use non-contextual head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}\in\mathbb{R}^{h}$ (i.e., independent of $\mathbf{x}_{t}$), while allowing the token-dimension factors $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ to remain context-dependent.
-Then, for keys (using the original $1/R_{K}$ scaling):
-$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}\otimes\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$
-and similarly for queries and values.
-This version, referred to as TPA (Non-contextual A), reduces per-token computations for the A factors and can be effective when head-dimension relationships are relatively stable across tokens.
-The KV cache for this variant would store $\mathbf{b}^{K}_{r}(\mathbf{x}_{t})$ and $\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ for each token, while $\mathbf{a}^{K}_{r}$ and $\mathbf{a}^{V}_{r}$ are fixed.
-A.2 MQA and GQA as Non-contextual TPA
-Multi-Query Attention (MQA) [45] and Grouped Query Attention (GQA) [2] (detailed in Appendix H.1 and H.2) also emerge naturally from TPA by restricting the head-dimension factors ($\mathbf{a}^{K},\mathbf{a}^{V}$) to be non-contextual and low-rank:
-  • MQA as Rank-1 TPA (for K and V).
-In MQA, all heads share a single set of keys and values.
-This corresponds to TPA with $R_{K}=R_{V}=1$, where the head-dimension factors for keys and values are non-contextual vectors of all ones (or any constant non-zero vector, as it’s a shared projection).
-For example:
-    $\mathbf{K}_{t}=\frac{1}{1}\mathbf{1}_{h}\;\otimes\;\mathbf{b}^{K}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=\frac{1}{1}\mathbf{1}_{h}\;\otimes\;\mathbf{b}^{V}(\mathbf{x}_{t}),$
-    where $\mathbf{1}_{h}\in\mathbb{R}^{h}$ is a vector of ones.
-This forces every head to use the same key and value vectors $\mathbf{b}^{K}(\mathbf{x}_{t})$ and $\mathbf{b}^{V}(\mathbf{x}_{t})$ (which are effectively $\mathbf{X}\bm{W}^{K}_{\text{shared}}$ and $\mathbf{X}\bm{W}^{V}_{\text{shared}}$ from MQA’s definition).
-Each head $i$ would still use its distinct query $\mathbf{Q}_{t,i}$, which in TPA could be modeled with $R_{Q}=h$ and $\mathbf{a}^{Q}_{i}=\mathbf{e}_{i}$ as in the MHA case if queries are not factorized, or a different TPA structure for queries.
-  • GQA as Grouped Rank-1 TPA (for K and V within groups).
-GQA partitions $h$ heads into $G$ groups, with each group sharing keys and values.
-In TPA form, this means for keys and values, the rank $R_{K}$ and $R_{V}$ would be $G$.
-For each group $j\in[G]$, there would be a non-contextual head-dimension factor $\mathbf{a}^{K}_{j}\in\mathbb{R}^{h}$ (and similarly $\mathbf{a}^{V}_{j}$) that acts as a "mask" or selector for the heads in that group (e.g., $\mathbf{a}^{K}_{j}$ has ones for heads in group $j$ and zeros elsewhere).
-Then:
-    $\mathbf{K}_{t}=\frac{1}{G}\sum_{j=1}^{G}\mathbf{a}^{K}_{j}\otimes\mathbf{b}^{K}_{j}(\mathbf{x}_{t}),\quad\mathbf{V}_{t}=\frac{1}{G}\sum_{j=1}^{G}\mathbf{a}^{V}_{j}\otimes\mathbf{b}^{V}_{j}(\mathbf{x}_{t}),$
-    where $\mathbf{b}^{K}_{j}(\mathbf{x}_{t})$ is the shared key vector for group $j$.
-Varying $G$ from $1$ (MQA) to $h$ (MHA, if $\mathbf{b}^{K}_{j}$ are distinct per original head) shows GQA as an interpolation.
-Hence, by constraining TPA’s head-dimension factors $\mathbf{a}^{K},\mathbf{a}^{V}$ to be specific constant vectors (or masks) and choosing appropriate ranks, these popular attention variants can be recovered as special cases of TPA with non-contextual A factors for keys and values.
-This underscores the flexibility and unifying nature of the TPA framework.
-Appendix B T6 Model Architecture Details
-This section provides a recap of how Tensor Product Attention (TPA) components are integrated within the Tensor ProducT ATTenTion Transformer (T6) architecture, building upon the detailed exposition in Section 3.
-TPA QKV Factorization Recap.
-As defined in Section 3.1, for each token’s hidden-state vector $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, TPA projects the input into three tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\;\mathbb{R}^{T\times h\times d_{h}}$.
-The slices for the $t$-th token, $\mathbf{Q}_{t},\;\mathbf{K}_{t},\;\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$, are formed by sums of tensor products.
-The factor components $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})$ are produced by linear transformations on $\mathbf{x}_{t}$.
-For instance, for queries, with learnable weight matrices $\bm{W}^{a^{Q}}_{r}\in\mathbb{R}^{h\times d_{\text{model}}}$ and $\bm{W}^{b^{Q}}_{r}\in\mathbb{R}^{d_{h}\times d_{\text{model}}}$, these factors are:
-$\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}_{r}\,\mathbf{x}_{t},\quad\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}_{r}\,\mathbf{x}_{t}.$
-In practice, these are often computed by merging ranks into a single linear layer followed by reshaping, as detailed in Section 3.1.
-The factorization for keys ($\mathbf{K}_{t}$) and values ($\mathbf{V}_{t}$) follows the same pattern using their respective ranks $R_{K},R_{V}$ and projection matrices.
-Attention Step and Output Projection Recap.
-Once the factorized $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors are obtained for all tokens (with RoPE applied to the appropriate factors of $\mathbf{Q}$ and $\mathbf{K}$ as per Section 3.2), the attention output for each head $i\in\{1,\dots,h\}$ is computed using the scaled dot-product attention formula (Equation 3.2 from the main paper):
-$\textbf{head}_{i}=\operatorname{Softmax}\Bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,\mathbf{Q}_{i}\,(\mathbf{K}_{i})^{\top}\Bigr{)}\;\mathbf{V}_{i}.$
-Finally, the outputs from all $h$ heads are concatenated and projected back to the model dimension using an output weight matrix $\bm{W}^{O}$, as shown in Equation 3.3 from the main paper:
-$\operatorname{TPA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\bm{W}^{O}.$
-This constitutes the output of the TPA sub-layer in a T6 block.
-Appendix C Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
-In TPA, the query, key, and value tensors ($\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$) are constructed as sums of $R_{Q},R_{K},R_{V}$ rank-one tensors, respectively, as detailed in Equation 3.1.
-These sums are typically normalized by scaling factors $s_{Q}=1/R_{Q}$, $s_{K}=1/R_{K}$, and $s_{V}=1/R_{V}$.
-This section explores the computational advantages of performing attention calculations directly on these factorized representations, including their scaling factors, without explicitly materializing the full $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors.
-This approach can potentially reduce floating-point operations (FLOPs) and memory bandwidth.
-C.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$
-Consider a single head $i$. Each query vector $\mathbf{Q}^{(i)}_{t}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{Q}$ and scale $s_{Q}=1/R_{Q}$):
-$\mathbf{Q}^{(i)}_{t}~{}=~{}s_{Q}\sum_{r=1}^{R_{Q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
-and each key vector $\mathbf{K}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{K}$ and scale $s_{K}=1/R_{K}$):
-$\mathbf{K}^{(i)}_{\tau}~{}=~{}s_{K}\sum_{s=1}^{R_{K}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
-Their dot-product for tokens $t,\tau$ is
-$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}s_{Q}s_{K}\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (C.1)
-Similarly, the value vector $\mathbf{V}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{V}$ and scale $s_{V}=1/R_{V}$):
-$\mathbf{V}^{(i)}_{\tau}~{}=~{}s_{V}\sum_{u=1}^{R_{V}}a^{(u)}_{v,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
-C.2 Multi-Head Case
-For multi-head attention with $h$ heads, one repeats the factorization across all heads.
-The $\mathbf{b}_{q}^{(r)},\mathbf{b}_{k}^{(s)},\mathbf{b}_{v}^{(u)}$ vectors are shared across heads.
-The scaling factors $s_{Q},s_{K},s_{V}$ are applied per head as part of the definition.
-C.3 Complexity Analysis
-We compare the cost of standard multi-head attention versus TPA under two scenarios:
-  1. 1.
-    Naïve: Materialize $\mathbf{Q},\mathbf{K},\mathbf{V}$ from factors (including their scales), then perform the usual batched GEMM.
-  2. 2.
-    Specialized: Attempt to compute $\mathbf{Q}\,\mathbf{K}^{\top}$ and the final attention output directly from the rank-$(R_{Q},R_{K},R_{V})$ factors and their scales, without explicitly forming full $\mathbf{Q},\mathbf{K},\mathbf{V}$.
-Standard Multi-Head Attention. For batch size $B$ and sequence length $T$:
-  • Projection cost: $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}^{2}\bigr{)}$ or $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}\,d_{h}\bigr{)}$.
-  • Dot-product: $\mathbf{Q}\,(\mathbf{K})^{\top}\in\mathbb{R}^{(B\,h)\times T\times T}$ costs $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$.
-For large $T$, the $\mathcal{O}(B\,T^{2}\,d_{\text{model}})$ term dominates.
-TPA: Naïve Implementation.
-  • Constructing factors:   $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}\,\times(R_{Q}(h+d_{h})+R_{K}(h+d_{h})+R_{V}(h+d_{h}))\bigr{)}$.
-  • Materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$ (including scaling):   $\mathcal{O}\bigl{(}B\,T\,(R_{Q}\,h\,d_{h}+R_{K}\,h\,d_{h}+R_{V}\,h\,d_{h})\bigr{)}$.
-The scaling is a per-element multiplication.
-  • Dot-product $\mathbf{Q}\,(\mathbf{K})^{\top}$ and multiply by $\mathbf{V}$:   $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$ for scores, and $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$ for value aggregation.
-Typically $R_{Q},R_{K},R_{V}\ll h,d_{h}$, so the overhead of constructing factors and materializing is influenced by these ranks.
-The key benefit of TPA is in reducing KV cache size and potentially computation if specialized kernels are used.
-TPA: Specialized Implementation.
-If we bypass explicitly forming $\mathbf{Q},\mathbf{K},\mathbf{V}$, the computation involves the sums over ranks and the scaling factors.
-Below, we detail its complexity.
-Algorithm 2 Specialized TPA Computation (Without Materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$)
-1:Factorized Query components: $\mathbf{A}_{Q}(T,H,R_{Q})$, $\mathbf{B}_{Q}(T,R_{Q},d_{h})$
-2:Factorized Key components: $\mathbf{A}_{K}(T,H,R_{K})$, $\mathbf{B}_{K}(T,R_{K},d_{h})$
-3:Factorized Value components: $\mathbf{A}_{V}(T,H,R_{V})$, $\mathbf{B}_{V}(T,R_{V},d_{v})$ (Note: $d_{v}$ is value head dim, often $d_{v}=d_{h}$)
-4:Sequence length $T$, Number of heads $H$, Ranks $R_{Q},R_{K},R_{V}$, Head feature dimension $d_{h}$
-5:Scaling factors: $s_{Q}=1/R_{Q},s_{K}=1/R_{K},s_{V}=1/R_{V}$ (or other pre-defined scales)
-6:Output tensor $\mathbf{O}(T,H,d_{v})$
-7:// Phase 1: Attention Score Computation
-8$\triangleright$ Step 1.1: Compute shared $\mathbf{B}$-factor dot products $P$
-9:$P\leftarrow\text{einsum}(\text{"tqrd, tkrd}\rightarrow\text{tqtkrqrk"},\mathbf{B}_{Q},\mathbf{B}_{K})$
-10$\triangleright$ Indices: $t_{q}$=query token, $t_{k}$=key token, $r_{q}$=query rank, $r_{k}$=key rank, $d$=feature dim
-11$\triangleright$ Step 1.2: Combine $P$ with $\mathbf{A}_{Q},\mathbf{A}_{K}$ and apply $s_{Q},s_{K}$ to form scaled logits $\mathcal{L}^{\prime}$
-12:$\mathcal{L}_{\text{unscaled}}\leftarrow\text{einsum}(\text{"tqhrq, tkhrk, tqtkrqrk}\rightarrow\text{tqtkh"},\mathbf{A}_{Q},\mathbf{A}_{K},P)$
-13:$\mathcal{L}^{\prime}\leftarrow s_{Q}\cdot s_{K}\cdot\mathcal{L}_{\text{unscaled}}$
-14$\triangleright$ Indices: $h$=head
-15$\triangleright$ Step 1.3: Apply attention scaling ($1/\sqrt{d_{h}}$) and Softmax to get probabilities $\bm{\alpha}$
-16:$\mathcal{L}^{\prime\prime}_{\text{scaled}}\leftarrow\mathcal{L}^{\prime}/\sqrt{d_{h}}$
-17:$\bm{\alpha}\leftarrow\operatorname{Softmax}_{T_{k}}(\mathcal{L}^{\prime\prime}_{\text{scaled}})$triangleright Softmax over key tokens $T_{k}$ for each query $T_{q}$ and head $H$
-18$\triangleright$ Output $\bm{\alpha}(t_{q},t_{k},h)$
-19:
-20:// Phase 2: Value Aggregation
-21$\triangleright$ Step 2.1: Conceptually weight $\mathbf{A}_{V}$ by $\bm{\alpha}$ to form $W_{\mathbf{A}_{V}}$
-22:$W_{\mathbf{A}_{V}}\leftarrow\text{einsum}(\text{"tqtkh, tkhrv}\rightarrow\text{tqtkhrv"},\bm{\alpha},\mathbf{A}_{V})$
-23$\triangleright$ Indices: $r_{v}$=value rank
-24$\triangleright$ Step 2.2: Aggregate $W_{\mathbf{A}_{V}}$ with $\mathbf{B}_{V}$ and apply $s_{V}$ to get final output $\mathbf{O}$
-25:$\mathbf{O}_{\text{unscaled}}\leftarrow\text{einsum}(\text{"tqtkhrv, tkrvd}\rightarrow\text{tqhd"},W_{\mathbf{A}_{V}},\mathbf{B}_{V})$triangleright $d$ is value feature dim $d_{v}$
-26:$\mathbf{O}\leftarrow s_{V}\cdot\mathbf{O}_{\text{unscaled}}$
-27:return $\mathbf{O}$
-C.4 Complexity Analysis for the Specialized Implementation
-Complexity for a Single Query Token (Autoregressive Decoding).
-The dot product $\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}$ from Eq.
-(C.1) includes the $s_{Q}s_{K}$ scaling.
-For each pair $(r,s)$, we pay:
-  1. 1.
-    $\mathcal{O}(1)$ for multiplying $a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})$,
-  2. 2.
-    $\mathcal{O}(d_{h})$ for the dot product $\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\rangle$.
-The $s_{Q}s_{K}$ scaling is a single multiplication after the sums.
-Since $(r,s)$ runs over $R_{Q}\times R_{K}$, each token-pair $(t,\tau)$ costs roughly $\mathcal{O}\Bigl{(}R_{Q}\,R_{K}\,\bigl{(}1+d_{h}\bigr{)}\Bigr{)}\approx\mathcal{O}(R_{Q}\,R_{K}\,d_{h})$.
-For a single query token $t$ and $M$ cached key tokens $\tau=1\dots M$, this is $\mathcal{O}(M\,R_{Q}\,R_{K}\,d_{h})$ for a single head to compute all unscaled dot products.
-The scaling by $s_{Q}s_{K}$ is $\mathcal{O}(M)$ for the score vector.
-Multi-Head and Batches (Full Sequence Processing - reusing $\mathbf{b}$-Dot Products as in Algorithm 2). The $\mathbf{b}$-vectors can be shared across heads. Let $T$ be the sequence length.
-  1. 1.
-    $\mathbf{b}$-Dot-Product Stage ($P$):
-    Compute $P(t_{q},\tau_{k},r_{q},r_{k})=\langle\mathbf{b}_{q}^{(r_{q})}(\mathbf{x}_{t_{q}}),\mathbf{b}_{k}^{(r_{k})}(\mathbf{x}_{\tau_{k}})\rangle$.
-Cost: $\mathcal{O}\bigl{(}T^{2}\,R_{Q}R_{K}\,d_{h}\bigr{)}$.
-This is shared across heads.
-  2. 2.
-    Per-Head Score Computation:
-    For each head $h$, compute $\mathcal{L}_{t_{q},\tau_{k},h}=s_{Q}s_{K}\sum_{r_{q},r_{k}}a_{q,h}^{(r_{q})}(\mathbf{x}_{t_{q}})\,a_{k,h}^{(r_{k})}(\mathbf{x}_{\tau_{k}})\,P(t_{q},\tau_{k},r_{q},r_{k})$.
-    Cost: $\mathcal{O}\bigl{(}H\,T^{2}\,R_{Q}R_{K}\bigr{)}$.
-Putting these together for batch size $B$, the total cost for scores is
-$\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}R_{K}\,d_{h}\bigr{)}+\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}R_{K}\bigr{)}=\mathcal{O}\Bigl{(}B\,T^{2}\,R_{Q}R_{K}\bigl{(}d_{h}+H\bigr{)}\Bigr{)}$.
-This is for computing the scaled logits before softmax.
-The standard attention dot-product step is $\mathcal{O}\bigl{(}B\,T^{2}\,H\,d_{h}\bigr{)}$.
-For the specialized TPA to reduce flops in score computation: $R_{Q}R_{K}\,(d_{h}+H)<H\,d_{h}$.
-This implies $R_{Q}R_{K}<H/(1+H/d_{h})$.
-If $H\approx d_{h}$, then $R_{Q}R_{K}\approx H/2$.
-C.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
-We extend the idea to also avoid explicitly forming $\mathbf{V}$.
-After obtaining scaled logits $\mathcal{L}^{\prime}_{t_{q},\tau_{k}}=s_{Q}s_{K}(\mathbf{Q}_{t_{q}}\mathbf{K}_{\tau_{k}}^{\top})$ (per head), we apply $\alpha_{t_{q},\tau_{k}}=\operatorname{softmax}_{\tau_{k}}\bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,\mathcal{L}^{\prime}_{t_{q},\tau_{k}}\bigr{)}$.
-The final attention output at token $t_{q}$ (single head) using $\mathbf{V}_{\tau_{k}}=s_{V}\sum_{u=1}^{R_{V}}a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})$ is:
-$\operatorname{head}(t_{q})~{}=~{}\sum_{\tau_{k}=1}^{T}\alpha_{t_{q},\tau_{k}}~{}\mathbf{V}_{\tau_{k}}~{}=~{}s_{V}\sum_{\tau_{k}=1}^{T}\alpha_{t_{q},\tau_{k}}\sum_{u=1}^{R_{V}}a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}}).$
-Rearranging sums:
-$\operatorname{head}(t_{q})~{}=~{}s_{V}\sum_{u=1}^{R_{V}}\left[\sum_{\tau_{k}=1}^{T}\bigl{(}\alpha_{t_{q},\tau_{k}}\,a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\right].$
-The computation of $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})$ from $\mathbf{x}_{\tau_{k}}$ is $\mathcal{O}(T\,R_{V}\,d_{\text{model}}d_{v})$ if contextual, or loaded if fixed (where $d_{v}$ is the dimension of $\mathbf{b}_{v}$).
-Assuming $\mathbf{b}_{v}^{(u)}$ are available, the weighted summation for each $(t_{q},u)$ pair (inner sum over $\tau_{k}$) costs $\mathcal{O}(T\,d_{v})$.
-Summed over $t_{q}=1\ldots T$ and $u=1\ldots R_{V}$, this stage is $\mathcal{O}(T^{2}\,R_{V}\,d_{v})$.
-The final scaling by $s_{V}$ is $\mathcal{O}(Td_{v})$.
-C.6 Overall Complexity for Single-Head (Specialized)
-The dominant costs for a single head, assuming factors $\mathbf{A},\mathbf{B}$ are already computed from $\mathbf{X}$:
-  1. (i)
-    QK $\mathbf{B}$-Dot Product Stage ($P$): $\mathcal{O}(T^{2}\,R_{Q}\,R_{K}\,d_{h})$ (shared calculation if multi-head).
-  2. (ii)
-    QK $\mathbf{A}$-Factor Combination and Scaling Stage ($\mathcal{L}^{\prime}$): $\mathcal{O}(T^{2}\,R_{Q}\,R_{K})$. (Scaling by $s_{Q}s_{K}$ is part of this per-element operation on scores).
-  3. (iii)
-    Value Aggregation Stage (including $s_{V}$ scaling): $\mathcal{O}(T^{2}\,R_{V}\,d_{v})$.
-Total for single head, assuming $P$ is computed once: $\mathcal{O}\Bigl{(}T^{2}\,R_{Q}\,R_{K}\,d_{h}+T^{2}\,R_{Q}\,R_{K}+T^{2}\,R_{V}\,d_{v}\Bigr{)}$.
-C.7 Multi-Head and Batch Extensions (Reuse of $\mathbf{B}$-Dot Products)
-QK $\mathbf{B}$-Dot Products ($P$): Shared across heads. Cost: $\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}\,R_{K}\,d_{h}\bigr{)}$.
-Per-Head QK $\mathbf{A}$-Factor Combination and Scaling ($\mathcal{L}^{\prime}$): Each head combines its $\mathbf{A}_{Q},\mathbf{A}_{K}$ with $P$ and applies $s_{Q}s_{K}$.
-Cost: $\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}\,R_{K}\bigr{)}$.
-Value Aggregation (including $s_{V}$): Assuming $\mathbf{B}_{V}$ factors are shared and $\mathbf{A}_{V}$ are per-head.
-Cost: $\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{V}\,d_{h}\bigr{)}$.
-Assuming $d_{v}=d_{h}$.
-Total FLOPs for specialized multi-head TPA (full sequence processing) with batch size $B$:
-$\mathcal{F}_{\text{TPA-specialized}}=\underbrace{\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}\,R_{K}\,d_{h}\bigr{)}}_{\text{QK $\mathbf{B}$-dot products }P}+\underbrace{\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}\,R_{K}\bigr{)}}_{\text{Per-head QK }\mathbf{A}\text{-comb.
-+ }s_{Q}s_{K}\text{ scale}}+\underbrace{\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{V}\,d_{h}\bigr{)}}_{\text{Value Aggregation+ }s_{V}\text{ scale}}.$
-Discussion.
-By contrast, standard multi-head attention typically requires $\mathcal{F}_{\text{MHA}}=\mathcal{O}\bigl{(}2B\,T^{2}\,H\,d_{h}\bigr{)}$ FLOPs (for $\mathbf{Q}\mathbf{K}^{\top}$ and attention-weighted $\mathbf{V}$).
-The specialized TPA can be more efficient if $R_{Q}R_{K}d_{h}+HR_{Q}R_{K}+HR_{V}d_{h}<2Hd_{h}$.
-Dividing by $Hd_{h}$, this condition becomes $(R_{Q}R_{K}/H)+(R_{Q}R_{K}/d_{h})+R_{V}<2$.
-For example, if $R_{Q}=R_{K}=R_{V}=1$, the condition is $21/H+1/d_{h}+1<$ 2, or $11/H+1/d_{h}<$ 1, which is generally true for typical $H,d_{h}\geq 2$.
-Thus, with small ranks, TPA can offer computational savings.
-Actual speedups also depend critically on memory access patterns and kernel implementations, as demonstrated by FlashAttention-style approaches.
-By retaining $\mathbf{Q},\mathbf{K},$ and $\mathbf{V}$ in factorized form, one can bypass the explicit materialization of these large tensors:
-$\mathbf{x}_{t}\rightarrow\mathbf{Q}_{t},~{}\mathbf{K}_{\tau},\mathbf{V}_{\tau}\quad(\text{materialized})\rightarrow(\mathbf{Q}\,\mathbf{K}^{\top})\rightarrow\operatorname{softmax}(\mathbf{Q}\,\mathbf{K}^{\top})\,\mathbf{V}\rightarrow\text{final output}.$
-Instead, the large $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors (each of size $B\times T\times H\times d_{h}$) are not explicitly formed.
-The computation is restructured into rank-based $\mathbf{B}$-dot-product computations, per-head $\mathbf{A}$-factor combinations, and appropriate scaling.
-The challenge lies in choosing ranks $(R_{Q},R_{K},R_{V})$ small enough to ensure computational benefits while maintaining model quality, and in implementing the multi-stage kernels efficiently.
-When ranks are sufficiently small, this specialized approach can lead to gains in both computation and memory footprint.
 […]
-Figure 6: Data flow diagram for specialized TPA computation, avoiding materialization of full $\mathbf{Q},\mathbf{K},\mathbf{V}$ tensors.
-Phase 1 (Top): Attention Score Computation.
-Factorized $\mathbf{B}_{Q},\mathbf{B}_{K}$ produce shared dot-products $P$ (summing over feature dimension $d$).
-$P$ is combined with factorized $\mathbf{A}_{Q},\mathbf{A}_{K}$ and scaled by $s_{Q}s_{K}$ (summing over ranks $r_{q},r_{k}$) to form scaled logits $\mathcal{L}^{\prime}$.
-These are then scaled by $1/\sqrt{d_{h}}$ and passed through Softmax to yield attention probabilities $\bm{\alpha}$.
-Phase 2 (Bottom): Value Aggregation.
-$\bm{\alpha}$ weights the factorized $\mathbf{A}_{V}$ (conceptually forming $W_{\mathbf{A}_{V}}$ via an einsum-like combination), which then aggregates with $\mathbf{B}_{V}$ and is scaled by $s_{V}$ (summing over key tokens $t_{k}$ and value ranks $r_{v}$) to produce the final output $\mathbf{O}$.
-(Diagram simplifies batch dimensions $B$; $T$ denotes sequence length; $H$ is heads; $R_{X}$ are ranks $R_{Q},R_{K},R_{V}$ corresponding to indices $r_{q},r_{k},r_{v}$; $d_{h}$ is the feature dimension $d$ for $\mathbf{B}$ factors.
-Operations $\sum_{X}$ denote contractions over index $X$, and $\odot$ denotes an einsum-like combination.
-Scaling factors $s_{Q},s_{K},s_{V}$ are incorporated as described.).
-C.8 Inference Time Decoding of Different Attention Mechanisms
-In autoregressive decoding, we generate the output for the current token $\mathbf{x}_{M}$ (where $M$ is the current sequence length, previously denoted $T$ in this section for a generic single token, but $M$ is used in Alg.
-1) given cached keys and values from $M-1$ previous tokens.
-We analyze the FLOPs for computing the attention output for this single query token.
-For all mechanisms, we consider the FLOPs for computing scaled logits and aggregating values, excluding the initial projection of the current token’s hidden state $\mathbf{x}_{M}$ into Q, K, V representations.
-MHA, MQA, and GQA. For Multi-Head Attention (MHA), with $H$ query heads and $H$ distinct Key/Value heads:
-  • Logits: Each of the $H$ query vectors (dim $d_{h}$) interacts with its corresponding K-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
-  • Value Aggregation: Each of the $H$ attention patterns interacts with its V-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
-  • Total MHA: $2MHd_{h}$.
-Multi-Query Attention (MQA) uses $H$ query heads but shares a single Key/Value head ($H_{kv}=1$).
-  • Logits: Each of the $H$ query vectors (dim $d_{h}$) interacts with the single shared K-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
-  • Value Aggregation: Each of the $H$ attention patterns interacts with the single shared V-cache (size $M\times d_{h}$). Cost: $H\cdot(M\cdot d_{h})=MHd_{h}$.
-  • Total MQA: $2MHd_{h}$.
-Grouped-Query Attention (GQA) uses $H$ query heads and $N_{g}$ Key/Value head groups ($H_{kv}=N_{g}$).
-The $H$ query heads are partitioned into $N_{g}$ groups, each containing $H/N_{g}$ query heads that share one K/V head pair within that group.
-  • Logits: For each of the $N_{g}$ groups, $H/N_{g}$ query vectors interact with the group’s K-cache.
-Cost per group: $(H/N_{g})\cdot(M\cdot d_{h})$.
-Total for $N_{g}$ groups: $N_{g}\cdot(H/N_{g})\cdot Md_{h}=MHd_{h}$.
-  • Value Aggregation: Similarly, $MHd_{h}$.
-  • Total GQA: $2MHd_{h}$.
-MQA and GQA significantly reduce the KV cache size and memory bandwidth requirements compared to MHA.
-While the arithmetic FLOP count for the core attention computation (dot products and weighted sums) is $2MHd_{h}$ for all three if they have the same number of query heads $H$ and head dimension $d_{h}$, the practical speedups for MQA/GQA often arise from better memory access patterns due to smaller K/V caches.
-MLA.
-Multi-Head Latent Attention (MLA), as described in Appendix H.3, uses $H$ heads.
-For each head, the key $\mathbf{K}_{i}$ has dimension $d_{k}^{\prime}=d_{c}+d_{h}^{R}$ (compressed part + RoPE part), and the value $\mathbf{V}_{i}^{C}$ has dimension $d_{v}^{\prime}=d_{c}$ (only compressed part).
-  • Logits: $H$ query heads (dim $d_{k}^{\prime}$) interact with K-caches (dim $d_{k}^{\prime}$). Cost: $MHd_{k}^{\prime}$.
-  • Value Aggregation: $H$ attention patterns interact with V-caches (dim $d_{v}^{\prime}$). Cost: $MHd_{v}^{\prime}$.
-  • Total MLA: $MH(d_{k}^{\prime}+d_{v}^{\prime})=MH(d_{c}+d_{h}^{R}+d_{c})=MH(2d_{c}+d_{h}^{R})$.
+[…]
+4:scales $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$; mask $\mathsf{mask}\in\{0,-\infty\}^{B\times T_{q}\times T_{k}}$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+11:return $\mathrm{transpose}(\mathbf{O},~\text{"bhqe"}\to\text{"bqhe"})$
+A.3 Complexity of the specialized path
+Combining the terms above gives complexity $\mathcal{F}_{\text{TPA-spec}}=\Theta(BT^{2}R_{Q}R_{K}D)+\Theta(BHT^{2}R_{Q}R_{K})+\Theta(BHT^{2}R_{V}E)$, with the speed condition Eq. (A.4).
+For a single query ($T_{q}{=}1$) against a cache of length $M$, the specialized FLOPs are
+$\Theta(B\,M\,R_{Q}R_{K}\,D)\;+\;\Theta(B\,H\,M\,R_{Q}R_{K})\;+\;\Theta(B\,H\,M\,R_{V}\,E),$
+while MHA uses $2\,\Theta(BHMD)$. This matches the asymptotics embodied in FlashTPA (Section 5) and explains the regimes where $R_{Q}{\ll}D$ and $R_{K}{=}R_{V}\in\{1,2\}$ yield the largest gains.
+We apply the causal mask before softmax and use an online log-sum-exp update for numerical stability (as in FlashAttention).
+The intermediate $P\in\mathbb{R}^{B\times T_{q}\times T_{k}\times R_{Q}\times R_{K}}$ is evaluated blockwise in $T_{k}$ to keep peak memory linear in the block size; the same blocking naturally fuses with the masked softmax and the value aggregation step.
+The constants $s_{Q},s_{K},s_{V}$ can be absorbed into either $\mathbf{A}_{(\cdot)}$ or $\mathbf{B}_{(\cdot)}$ at training time.
+We expose them explicitly only to make Eq.
+(A.4) transparent; The choice has no effect on softmax invariance or gradients.
+The Triton kernel in Section 5 implements the blocked computation of $P$, the masked online softmax over $k$, and the fused value aggregation, mirroring Algorithm˜1.
+This avoids creating any $\mathbf{Q},\mathbf{K},\mathbf{V}$ or full $T_{q}{\times}T_{k}$ temporaries beyond working tiles.
+Compared with $2\,\Theta(BHT^{2}D)$ for MHA, the specialized path improves with small $(R_{Q},R_{K},R_{V})$ and benefits further from pre-rotating $\mathbf{B}_{K}$ for RoPE (cf.
+Section 3.2), which removes per-step rotations in decoding.
+Practical speed also depends on tiling, memory bandwidth, and kernel fusion; our measured gains in Section 6.2 align with the regime predicted by Eq.
+(A.4).
+A.4 Inference-time decoding cost across mechanisms
+In autoregressive decoding, we generate the output for the current token $\mathbf{x}_{T}$ given cached keys and values from T1T{-}1 previous tokens.
+We analyze the FLOPs for computing the attention output for this single query token and use $M$ for the current cache length.
+For all mechanisms, we analyze the total Floating Point Operations (FLOPs) and the number of parameters in the attention layer, including the cost of projecting the current token’s hidden state $\mathbf{x}_{T}$ into its respective Query, Key, and Value representations.
+The parameter count formulas are taken from Table 1.
+For Multi-Head Attention (MHA), with $H$ query heads and $H$ distinct Key/Value heads, the complexity is determined by the dot-product attention and value aggregation steps.
+  • Projection: Projecting $\mathbf{x}_{T}$ to get a query, key, and value vector for each of the $H$ heads costs $\Theta(d_{\text{model}}Hd_{h})$.
+  • Attention: The query vectors interact with a K/V cache of size $T$, costing $\Theta(MHd_{h})$.
+  • Total MHA: The complexity is $\Theta(d_{\text{model}}Hd_{h}+MHd_{h})$.
+Multi-Query Attention (MQA) uses $H$ query heads but shares a single Key/Value head ($H_{kv}=1$). The arithmetic complexity remains the same as MHA for the same number of query heads.
+  • Projection: Projecting for $H$ query heads and 1 shared K/V head costs $\Theta(d_{\text{model}}(Hd_{h}+2d_{h}))$.
+  • Attention: The interaction with the cache costs $\Theta(2MHd_{h})$.
+  • Total MQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2)+MHd_{h})$.
+Grouped-Query Attention (GQA) uses $H$ query heads and $N_{g}$ Key/Value head groups ($H_{kv}=N_{g}$). The arithmetic complexity is also identical to MHA.
+  • Projection: Projecting for $H$ query heads and $N_{g}$ K/V head groups costs $\Theta(d_{\text{model}}(Hd_{h}+2N_{g}d_{h}))$.
+  • Attention: The interaction with the cache costs $\Theta(2MHd_{h})$.
+  • Total GQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2N_{g})+MHd_{h})$.
+MQA and GQA significantly reduce the KV cache size and memory bandwidth compared to MHA.
+While the arithmetic FLOP count for the core attention computation (dot products and weighted sums) is $2MHd_{h}$ for all three (for fixed $H,d_{h}$), practical speedups for MQA/GQA arise from improved memory locality due to smaller K/V caches.
+Multi-Head Latent Attention (MLA), as described in Appendix F.3, uses $H$ heads.
+For each head, the key $\mathbf{K}_{i}$ has dimension $d_{k}^{\prime}=d_{c}+d_{h}^{R}$, and the value $\mathbf{V}_{i}^{C}$ has dimension $d_{v}^{\prime}=d_{c}$.
+  • Projection: Projecting $\mathbf{x}_{M}$ to get compressed and RoPE components for Q, K, and V costs $\Theta(d_{\text{model}}((d_{c}+d_{h}^{R})\cdot h+d_{c}+d_{h}^{R}))$.
+  • Attention: The query vectors interact with a K/V cache of size $M$, costing $\Theta(MH(2d_{c}+d_{h}^{R}))$.
+  • Total MLA: The complexity is $\Theta(d_{\text{model}}(d^{\prime}_{c}+2d_{c}+2d_{h}^{R})+MH(2d_{c}+d_{h}^{R}))$.
 TPA.
-We use the FlashTPA Decoding algorithm (Algorithm 1) for FLOPs analysis, with $N=1$ query token, $M$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$).
+We use the FlashTPA Decoding algorithm (Algorithm 2) for FLOPs analysis, with $N=1$ query token, $T$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$).
 For ranks $(R_{Q},R_{K},R_{V})$:
-  • Logits computation (Steps 1-3 of Alg. 1): $\mathcal{O}(M(R_{Q}R_{K}D+HR_{Q}R_{K}+HR_{K}))$.
-  • Value aggregation (Steps 5-6 of Alg. 1): $\mathcal{O}(M(HR_{V}+HR_{V}E))$.
-  • Total for TPA decoding: $\mathcal{O}(M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
-Example Comparison.
-Let’s use the ranks from FlashTPA experiments (Section 5.2): $(R_{Q},R_{K},R_{V})=(16,1,1)$, and typical dimensions $D=E=d_{h}=64$, $H=32$ (e.g., for a 2048 $d_{\text{model}}$).
-MHA Logits: $\quad M\cdot H\cdot d_{h}=M\cdot 32\cdot 64=2048M$
-MHA Value Agg.: $\quad M\cdot H\cdot d_{h}=M\cdot 32\cdot 64=2048M$
-MHA Total: $\quad 4096M$
-TPA $\text{TPA Logits}(R_{K}=1,R_{Q}=16):$ $\quad M(R_{Q}d_{h}+HR_{Q}+H)=M(16\cdot 64+32\cdot 16+32)$
-$=M(1024+512+32)=1568M$
+  • Projection: Projecting $\mathbf{x}_{M}$ to all Q, K, V factors costs $\Theta(d_{\text{model}}((R_{Q}+R_{K}+R_{V})(H+d_{h})))$.
+  • Attention: The decoding algorithm’s interaction with the cache costs $\Theta(M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
+  • Total for TPA decoding: $\Theta(d_{\text{model}}(R_{Q}+R_{K}+R_{V})(H+d_{h})+M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
+Example Comparison I.
+We compare the total Floating Point Operations (FLOPs) required to process a single token during autoregressive inference.
+This analysis separates the initial, constant projection cost from the attention cost, which scales linearly with the cache length $M$.
+The following parameters are used for the comparison:
+  • Model Dimension: $d_{\text{model}}=2048$
+  • Heads: $H=32$
+  • Head Dimension: $d_{h}=64$ (so $D=E=d_{h}$)
+  • GQA Groups: $N_{g}=4$
+  • MLA Dimensions: $d_{c}=256$, $d_{h}^{R}=32$, and $d^{\prime}_{c}=768$
+MHA (16.8M parameters):
+$\text{Parameters}=4d_{\text{model}}Hd_{h}=4\cdot 2048\cdot(32\cdot 64)\approx 16.8\times 10^{6}$
+$\text{Projection}=3\cdot d_{\text{model}}\cdot H\cdot d_{h}=3\cdot 2048\cdot 32\cdot 64\approx 12.6\times 10^{6}$
+$\text{Attention}=2\cdot M\cdot H\cdot d_{h}=4096M$
+GQA ($N_{g}=4$, 9.4M parameters):
+$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=2048\cdot 64\cdot(2\cdot 32+2\cdot 4)\approx 9.4\times 10^{6}$
+$\text{Projection}=d_{\text{model}}(H+2N_{g})d_{h}=2048\cdot(32+8)\cdot 64\approx 5.2\times 10^{6}$
+$\text{Attention}=2\cdot M\cdot H\cdot d_{h}=4096M$
+MLA (9.8M parameters):
+$\text{Parameters}=768(2048+2048+1024)+2048(32+2048)+256(2048+4096)\approx 9.8\times 10^{6}$
+$\text{Projection}=d_{\text{model}}((d_{c}+d_{h}^{R})\cdot h+d_{c}+d_{h}^{R})=2048\cdot((256+32)\cdot 32+256+32)\approx 19.5\times 10^{6}$
+$\text{Attention}=M\cdot H\cdot(2d_{c}+d_{h}^{R})=M\cdot 32\cdot(512+32)=17408M$
+TPA ($R_{Q}=16,R_{K}=1,R_{V}=1$, 7.7M parameters):
+$\text{Parameters}=d_{\text{model}}(16+1+1)(H+d_{h})+d_{\text{model}}Hd_{h}=2048(18)(96)+2048^{2}\approx 7.7\times 10^{6}$
+$\text{Projection}=d_{\text{model}}(16+1+1)(H+d_{h})=2048\cdot(18)\cdot(96)\approx 3.5\times 10^{6}$
+$\text{Attention}=M\cdot[1(1568)+1(2080)]=3648M$
+TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 8.1M parameters):
+$\text{Parameters}=d_{\text{model}}(16+2+2)(H+d_{h})+d_{\text{model}}Hd_{h}=2048(20)(96)+2048^{2}\approx 8.1\times 10^{6}$
+$\text{Projection}=d_{\text{model}}(16+2+2)(H+d_{h})=2048\cdot(20)\cdot(96)\approx 3.9\times 10^{6}$
+$\text{Attention}=M\cdot[2(1568)+2(2080)]=7296M$
+TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 6.2M parameters):
+$\text{Parameters}=d_{\text{model}}(8+1+1)(H+d_{h})+d_{\text{model}}Hd_{h}=2048(10)(96)+2048^{2}\approx 6.2\times 10^{6}$
+$\text{Projection}=d_{\text{model}}(8+1+1)(H+d_{h})=2048\cdot(10)\cdot(96)\approx 2.0\times 10^{6}$
+$\text{Attention}=M\cdot[1(800)+1(2080)]=2880M$
+TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 6.6M parameters):
+$\text{Parameters}=d_{\text{model}}(8+2+2)(H+d_{h})+d_{\text{model}}Hd_{h}=2048(12)(96)+2048^{2}\approx 6.6\times 10^{6}$
+$\text{Projection}=d_{\text{model}}(8+2+2)(H+d_{h})=2048\cdot(12)\cdot(96)\approx 2.4\times 10^{6}$
+$\text{Attention}=M\cdot[2(800)+2(2080)]=5760M$
+The analysis shows that TPA with low ranks offers a favorable trade-off.
+Reducing the query rank ($R_{Q}$) from 16 to 8 further decreases both the projection and attention costs, making the TPA ($R_{Q}{=}8,R_{K}{=}1,R_{V}{=}1$) configuration the most computationally efficient in this comparison.
+Increasing key/value ranks (e.g., to $R_{K}{=}2,R_{V}{=}2$) raises the attention cost linearly, remaining competitive with MHA for sufficiently long contexts where kernel fusion and blocking amortize memory traffic.
+Example Comparison II.
+We now repeat the analysis for a larger model configuration to observe how these trade-offs scale.
+The following parameters for a larger model are used for this comparison:
+  • Model Dimension: $d_{\text{model}}=4096$
+  • Heads: $H=32$
+  • Head Dimension: $d_{h}=128$ (so $D=E=d_{h}$)
+  • GQA Groups: $N_{g}=4$
+  • MLA Dimensions: $d_{c}=512$, $d_{h}^{R}=64$, and $d^{\prime}_{c}=1536$
+MHA (67.1M parameters):
+$\text{Parameters}=4d_{\text{model}}Hd_{h}=4\cdot 4096^{2}\approx 67.1\times 10^{6}$
+$\text{Projection}=3\cdot 4096\cdot 32\cdot 128\approx 50.3\times 10^{6}$
+$\text{Attention}=2\cdot M\cdot 32\cdot 128=8192M$
+GQA ($N_{g}=4$, 37.7M parameters):
+$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=4096\cdot 128\cdot(2\cdot 32+2\cdot 4)\approx 37.7\times 10^{6}$
+$\text{Projection}=4096\cdot(32+8)\cdot 128\approx 21.0\times 10^{6}$
+$\text{Attention}=2\cdot M\cdot 32\cdot 128=8192M$
+MLA (39.1M parameters):
+$\text{Parameters}=1536(4096+4096+2048)+4096(64+4096)+512(4096+8192)\approx 39.1\times 10^{6}$
+$\text{Projection}=4096\cdot((512+64)\cdot 32+512+64)\approx 77.9\times 10^{6}$
+$\text{Attention}=M\cdot 32\cdot(1024+64)=34816M$
+TPA ($R_{Q}=16,R_{K}=1,R_{V}=1$, 28.6M parameters):
+$\text{Parameters}=4096(18)(160)+4096^{2}\approx 28.6\times 10^{6}$
+$\text{Projection}=4096\cdot(16+1+1)\cdot(32+128)\approx 11.8\times 10^{6}$
+$\text{Attention}=M\cdot[1(2592)+1(4128)]=6720M$
+TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 29.9M parameters):
+$\text{Parameters}=4096(20)(160)+4096^{2}\approx 29.9\times 10^{6}$
+$\text{Projection}=4096\cdot(16+2+2)\cdot(32+128)\approx 13.1\times 10^{6}$
+$\text{Attention}=M\cdot[2(2592)+2(4128)]=13440M$
+TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 23.3M parameters):
+$\text{Parameters}=4096(10)(160)+4096^{2}\approx 23.3\times 10^{6}$
+$\text{Projection}=4096\cdot(8+1+1)\cdot(32+128)\approx 6.6\times 10^{6}$
+$\text{Attention}=M\cdot[1(1312)+1(4128)]=5440M$
+TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 24.6M parameters):
+$\text{Parameters}=4096(12)(160)+4096^{2}\approx 24.6\times 10^{6}$
+$\text{Projection}=4096\cdot(8+2+2)\cdot(32+128)\approx 7.9\times 10^{6}$
+$\text{Attention}=M\cdot[2(1312)+2(4128)]=10880M$
+For this larger configuration, TPA ($R_{Q}{=}8,R_{K}{=}1,R_{V}{=}1$) remains the clear leader in computational efficiency, with the lowest projection and attention costs.
+This highlights the value of tuning TPA ranks to balance expressiveness against compute.
+Example Comparison III.
 […]
-TPA Total: $\quad 1568M+2080M=3648M$
-For MQA and GQA, assuming $H=32$ query heads and $d_{h}=64$:
-MQA/GQA Total: $\quad 2MHd_{h}=2\cdot M\cdot 32\cdot 64=4096M$
-For MLA, with $H=32,d_{c}=256,d_{h}^{R}=32$:
-MLA Logits: $\quad MH(d_{c}+d_{h}^{R})=M\cdot 32\cdot(256+32)=M\cdot 32\cdot 288=9216M$
-MLA Value Agg.: $\quad MHd_{c}=M\cdot 32\cdot 256=8192M$
-MLA Total: $\quad M\cdot 32\cdot(2\cdot 256+32)=M\cdot 32\cdot 544=17408M$
-In this configuration, TPA requires fewer FLOPs ($3648M$) than MHA ($4096M$).
-The FLOPs for scaled logits in TPA ($1568M$) are less than MHA’s ($2048M$).
-The value aggregation FLOPs for TPA ($2080M$) are comparable to MHA’s ($2048M$).
-This reduction in FLOPs, particularly when $R_{K}$ and $R_{V}$ are small (e.g., 1), combined with optimized memory access patterns in algorithms like FlashTPA Decoding, contributes to TPA’s competitive decoding speed, especially for long sequences.
-The actual end-to-end wall-clock speedup also depends on kernel fusion, hardware specifics, and the efficiency of einsum implementations, but the factorized formulation offers a clear path to reduced computational load.
-Appendix D More on FlashTPA Decoding Algorithm
-This section provides further details on the FlashTPA Decoding algorithm (see Algorithm 1), including the setup of factorized components.
-Factorized Component Setup for FlashTPA Decoding.
-Let $B$ be the batch size, $N$ the number of query tokens (which is 1 for decoding, $N=1$), $M$ the current length of the KV cache, $H$ the number of attention heads, and $R_{Q}$ the rank of the query factorization.
-The feature dimensions for the factorized components are $D$ (for query $\mathbf{B}_{Q}$ and key $\bm{b}^{K}$) and $E$ (for value $\bm{b}^{V}$).
-Typically, $D$ and $E$ correspond to the head dimension $d_{h}$.
-The query for the current token $\mathbf{x}_{t}$ (where $t$ is the current time step, $N=1$) is factorized as $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$, where $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{H\times R_{Q}}$ (derived from the input tensor ‘Aq’ of shape $(B,1,H,R_{Q})$ for the current token) and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times D}$ (similarly from ‘Bq’ of shape $(B,1,R_{Q},D)$).
-The cached keys $\mathbf{K}_{s}$ and values $\mathbf{V}_{s}$ for past tokens $s\in[1,M]$ are stored in their factorized form.
-For clarity in the data flow diagram (Figure 2 in the main paper) and the pseudo-code in Algorithm 1, we present the case where cached keys and values use ranks $R_{K}=1$ and $R_{V}=1$.
-The general TPA formulation allows for $R_{K}>1$ and $R_{V}>1$ for cached items (by storing, for example, $\mathbf{A}_{K}(\mathbf{x}_{s})\in\mathbb{R}^{R_{K}\times H}$ and $\mathbf{B}_{K}(\mathbf{x}_{s})\in\mathbb{R}^{R_{K}\times D}$).
-The einsum operations detailed in Algorithm 1 can be extended to accommodate higher ranks for keys and values by including summations over these additional rank dimensions.
-For $R_{K}=1$ and $R_{V}=1$, the cached $s$-th key and value are: $\mathbf{K}_{s}=\mathbf{a}^{K}(\mathbf{x}_{s})\otimes\bm{b}^{K}(\mathbf{x}_{s})\quad\in\mathbb{R}^{H\times D}$, $\mathbf{V}_{s}=\mathbf{a}^{V}(\mathbf{x}_{s})\otimes\bm{b}^{V}(\mathbf{x}_{s})\quad\in\mathbb{R}^{H\times E}$.
-Here, $\mathbf{a}^{K}(\mathbf{x}_{s})\in\mathbb{R}^{H}$ (from the cached tensor $\mathbf{a}^{K}_{\text{cache}}$ of shape $(B,M,H)$), $\bm{b}^{K}(\mathbf{x}_{s})\in\mathbb{R}^{D}$ (from $\bm{b}^{K}_{\text{cache}}$ of shape $(B,M,D)$), $\mathbf{a}^{V}(\mathbf{x}_{s})\in\mathbb{R}^{H}$ (from $\mathbf{a}^{V}_{\text{cache}}$ of shape $(B,M,H)$), and $\bm{b}^{V}(\mathbf{x}_{s})\in\mathbb{R}^{E}$ (from $\bm{b}^{V}_{\text{cache}}$ of shape $(B,M,E)$).
-D.1 Detailed Computation Steps of FlashTPA Decoding Algorithm
-The algorithm proceeds through a series of einsum operations, which correspond to specific tensor contractions:
-  1. 1.
-    Compute $\mathbf{B}_{Q}$-$\bm{b}^{K}$ inner products ($S^{(1)}$): For each query (batch $b$, query index $n=0$), each query rank component $r$, and each cached key position $m$, we compute the dot product of the $D$-dimensional feature vector of the $r$-th query factor $(\mathbf{B}_{Q})_{b,n,r,:}$ with the $D$-dimensional feature vector of the $m$-th cached key $(\bm{b}^{K}_{\text{cache}})_{b,m,:}$.
-    $S^{(1)}_{b,n,m,r}=\sum_{d=1}^{D}(\mathbf{B}_{Q})_{b,n,r,d}\cdot(\bm{b}^{K}_{\text{cache}})_{b,m,d}.$
-    This step captures the similarity between the feature parts of the query factors and the cached key factors.
-  2. 2.
-    Combine with query factor $\mathbf{A}_{Q}$ ($S^{(2)}$): The result $S^{(1)}$ is then combined with the head-specific query factors $\mathbf{A}_{Q}$.
-For each head $h$, the contributions from all query ranks $r$ are summed up.
-    $S^{(2)}_{b,n,m,h}=\sum_{r=1}^{R_{Q}}(\mathbf{A}_{Q})_{b,n,h,r}\cdot S^{(1)}_{b,n,m,r}.$
-    This yields an intermediate score for each query, cached key, and head.
-  3. 3.
-    Incorporate cached key factor $\mathbf{a}^{K}_{\text{cache}}$ for full logits ($\mathcal{L}$): The head-specific factor of the cached keys, $\mathbf{a}^{K}_{\text{cache}}$, is multiplied with $S^{(2)}$ to produce the final unscaled attention logits.
-    $\mathcal{L}_{b,h,n,m}=S^{(2)}_{b,n,m,h}\cdot(\mathbf{a}^{K}_{\text{cache}})_{b,m,h}.$
-    This logit $\mathcal{L}_{b,h,n,m}$ represents $s_{Q}^{-1}\cdot(\mathbf{Q}_{t}\mathbf{K}_{m}^{\top})_{h}$, i.e., the dot product between the query $\mathbf{Q}_{t}$ and the $m$-th cached key $\mathbf{K}_{m}$ for head $h$, scaled by $R_{Q}$ (since $s_{Q}=1/R_{Q}$).
-The ‘einsum‘ operation also rearranges dimensions to $(B,H,N,M)$ for the subsequent softmax operation.
-  4. 4.
-    Apply Scaling and Softmax ($\bm{\alpha}$): The logits $\mathcal{L}$ are scaled by the product of $s_{Q}$ (e.g., $1/R_{Q}$), $s_{K}$ (e.g., $1/R_{K}$, which is 1 if $R_{K}=1$), and $s_{\text{total}}$ (e.g., $1/\sqrt{D}$, where $D$ is the dimension used for dot products, typically $d_{h}$).
-Softmax is then applied across the $M$ cached key positions for each head independently.
-    $\alpha_{b,h,n,m}=\operatorname{Softmax}_{m}\left(\mathcal{L}_{b,h,n,m}\cdot s_{Q}\cdot s_{K}\cdot s_{total}\right).$
-    These are the attention probabilities.
-  5. 5.
-    Compute weighted sum with cached value factor $\mathbf{a}^{V}_{\text{cache}}$ ($\mathbf{O}^{(A)}$): The attention probabilities $\bm{\alpha}$ are used to weight the head-specific factors of the cached values, $\mathbf{a}^{V}_{\text{cache}}$.
-    $\mathbf{O}^{(A)}_{b,n,m,h}=\alpha_{b,h,n,m}\cdot(\mathbf{a}^{V}_{\text{cache}})_{b,m,h}.$
-    This step prepares the value components for the final aggregation. The ‘einsum‘ operation also reorders dimensions to $(B,N,M,H)$.
-  6. 6.
-    Incorporate cached value factor $\bm{b}^{V}_{\text{cache}}$ and apply final scaling ($\mathbf{O}$): Finally, the intermediate weighted value factors $\mathbf{O}^{(A)}$ are combined with the feature-specific factors of the cached values, $\bm{b}^{V}_{\text{cache}}$, by summing over the $M$ cached positions.
-The result is then scaled by $s_{V}$ (e.g., $1/R_{V}$, which is 1 if $R_{V}=1$).
-    $\mathbf{O}_{b,n,h,e}=\left(\sum_{m=1}^{M}\mathbf{O}^{(A)}_{b,n,m,h}\cdot(\bm{b}^{V}_{\text{cache}})_{b,m,e}\right)\cdot s_{V}.$
-    This produces the final output tensor of shape $(B,1,H,E)$.
-D.2 Triton FlashTPA Decoding Kernel
-We implement the experiments using Triton language [56], and the detailed implementation pseudo code is displayed in Algorithm 3.
-The Triton FlashTPA decoding kernel supports parallelism across the number of heads, rank, and head dimensions.
-And in 3, we will show the experiment results with this kernel.
+The following parameters are used for this comparison:
+  • Model Dimension: $d_{\text{model}}=7168$
+  • Heads: $H=64$
+  • Head Dimension: $d_{h}=128$ (so $D=E=d_{h}$)
+  • GQA Groups: $N_{g}=8$
+  • MLA Dimensions: $d_{c}=512$, $d_{h}^{R}=64$, and $d^{\prime}_{c}=1536$
+MHA (235M parameters):
+$\text{Parameters}=4d_{\text{model}}Hd_{h}=4\cdot 7168\cdot 8192\approx 235\times 10^{6}$
+$\text{Projection}=3\cdot 7168\cdot 64\cdot 128\approx 176.2\times 10^{6}$
+$\text{Attention}=2\cdot M\cdot 64\cdot 128=16384M$
+GQA ($N_{g}=8$, 132M parameters):
+$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=7168\cdot 128\cdot(2\cdot 64+2\cdot 8)\approx 132\times 10^{6}$
+$\text{Projection}=7168\cdot(64+16)\cdot 128\approx 73.4\times 10^{6}$
+$\text{Attention}=2\cdot M\cdot 64\cdot 128=16384M$
+MLA (101M parameters):
+$\text{Parameters}=1536(7168+8192+4096)+7168(64+8192)+512(7168+16384)\approx 101\times 10^{6}$
+$\text{Projection}=7168\cdot((512+64)\cdot 64+512+64)=268.4\times 10^{6}$
+$\text{Attention}=M\cdot 64\cdot(2\cdot 512+64)=69632M$
+TPA ($R_{Q}=16,R_{K}=1,R_{V}=1$, 83M parameters):
+$\text{Parameters}=7168(18)(192)+7168\cdot 8192\approx 83.5\times 10^{6}$
+$\text{Projection}=7168\cdot(16+1+1)\cdot(64+128)\approx 24.8\times 10^{6}$
+$\text{Attention}=M\cdot[1(3136)+1(8256)]=11392M$
+TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 86.2M parameters):
+$\text{Parameters}=7168(20)(192)+7168\cdot 8192\approx 86.2\times 10^{6}$
+$\text{Projection}=7168\cdot(16+2+2)\cdot(64+128)\approx 27.5\times 10^{6}$
+$\text{Attention}=M\cdot[2(3136)+2(8256)]=22784M$
+TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 72.5M parameters):
+$\text{Parameters}=7168(10)(192)+7168\cdot 8192\approx 72.5\times 10^{6}$
+$\text{Projection}=7168\cdot(8+1+1)\cdot(64+128)\approx 13.8\times 10^{6}$
+$\text{Attention}=M\cdot[1(1600)+1(8256)]=9856M$
+TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 75.2M parameters):
+$\text{Parameters}=7168(12)(192)+7168\cdot 8192\approx 75.2\times 10^{6}$
+$\text{Projection}=7168\cdot(8+2+2)\cdot(64+128)\approx 16.5\times 10^{6}$
+$\text{Attention}=M\cdot[2(1600)+2(8256)]=19712M$
+At this very large scale, the cost of MHA projections becomes prohibitive.
+While MLA’s projection cost can be competitive, its attention cost scales with $(2d_{c}{+}d_{h}^{R})$ and exceeds MHA for long sequences.
+TPA with low ranks ($R_{Q}{=}8,R_{K}{=}1,R_{V}{=}1$) yields the lowest attention cost and a substantially smaller projection cost, strengthening its advantage as model size increases.
+Appendix B More on FlashTPA Decoding Algorithm
+In this section, we present FlashTPA for decoding in a hardware–friendly, numerically stable form and extend it to general key/value ranks $R_{K},R_{V}\geq 1$.
+The algorithm computes attention without materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$ or the full $N{\times}M$ attention matrix, by (i) forming head‑shared feature–space dot products, (ii) mixing them with head‑specific factors to obtain logits as in Eq.
+(A.2), and (iii) aggregating values as in Eq.
+(A.3) in a single online softmax pass.
+Notation and shapes.
+We allow $N$ query positions but decoding uses N=1N{=}1. Let $B$ be batch, $M$ the cache length, $H$ heads, $R_{Q},R_{K},R_{V}$ ranks, and $D,E$ feature sizes (typically $D{=}E{=}d_{h}$). Inputs:
+$\mathbf{A}_{Q}\!\in\!\mathbb{R}^{B\times N\times H\times R_{Q}},\;\mathbf{B}_{Q}\!\in\!\mathbb{R}^{B\times N\times R_{Q}\times D},\;\mathbf{A}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times H\times R_{K}},\;\mathbf{B}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{K}\times D},$
+$\mathbf{A}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times H\times R_{V}},\;\mathbf{B}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{V}\times E}.$
+We use scalings $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$, and $s_{\text{total}}{=}1/\sqrt{D}$.
+Let $\mathsf{mask}\!\in\!\{0,-\infty\}^{B\times N\times M}$ encode causality/padding.
+If RoPE pre‑rotation is used (Section 3.2), $\mathbf{B}_{K}^{\text{cache}}$ already includes positional phases; otherwise apply $\operatorname{RoPE}$ to $\mathbf{B}_{K}$ on load (no other change).
+Algorithm 2 FlashTPA Decoding (general $R_{K},R_{V}$, masked, online‑LSE)
+[…]
+[…]
+3:Initialize $\mathbf{y}\leftarrow 0^{B\times H\times N\times E}$, $\mathbf{lse}\leftarrow 0^{B\times H\times N}$, $\mathbf{m}\leftarrow(-\infty)^{B\times H\times N}$
+4:for each cache block $m{:}m{+}\Delta m\!\leq\!M$ do
+5:  Load $\mathbf{B}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{K}\times D}$, $\mathbf{A}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times H\times R_{K}}$
+6:  Load $\mathbf{A}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times H\times R_{V}}$, $\mathbf{B}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{V}\times E}$, $\mathsf{mask}_{\text{blk}}\!\in\!\mathbb{R}^{B\times N\times\Delta m}$
+7:  (1) Head‑shared feature dots:   $\mathbf{P}\leftarrow\mathrm{einsum}(\text{``bnrd,bmsd$\to$bnmrs''},~\mathbf{B}_{Q},~\mathbf{B}_{K,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times N\times\Delta m\times R_{Q}\times R_{K}}$
+8:  (2) Per‑head rank mixing to logits:
+9:  $\mathcal{L}_{\text{blk}}\leftarrow(s_{\text{total}}s_{Q}s_{K})\cdot\mathrm{einsum}(\text{``bnhr,bmhs,bnmrs$\to$bhnm''},~\mathbf{A}_{Q},~\mathbf{A}_{K,\text{blk}},~\mathbf{P})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times\Delta m}$
+10:  $\mathcal{L}_{\text{blk}}\leftarrow\mathcal{L}_{\text{blk}}+\mathrm{broadcast}(\mathsf{mask}_{\text{blk}})$
+11:  (3) Online softmax update (no $\bm{\alpha}$ materialization):
+12:  $\mathbf{m}_{\text{blk}}\leftarrow\max_{m}(\mathcal{L}_{\text{blk}})$; $\mathbf{p}_{\text{blk}}\leftarrow\exp(\mathcal{L}_{\text{blk}}-\mathbf{m}_{\text{blk}})$; $\mathbf{s}_{\text{blk}}\leftarrow\sum_{m}\mathbf{p}_{\text{blk}}$
+13:  (4) Block value mixing:   $\mathbf{V}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bmhu,bmue$\to$bhme''},~\mathbf{A}_{V,\text{blk}},~\mathbf{B}_{V,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times\Delta m\times E}$
+14:  $\mathbf{y}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bhnm,bhme$\to$bhne''},~\mathbf{p}_{\text{blk}},~\mathbf{V}_{\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times E}$
+15:  (5) Fuse blocks with log‑sum‑exp:
+16:  $\mathbf{m}_{\text{new}}\leftarrow\max(\mathbf{m},\mathbf{m}_{\text{blk}})$;   $\mathbf{y}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})[...,None]\odot\mathbf{y}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})[...,None]\odot\mathbf{y}_{\text{blk}}$
+17:  $\mathbf{lse}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})\odot\mathbf{lse}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})\odot\mathbf{s}_{\text{blk}}$;  $\mathbf{m}\leftarrow\mathbf{m}_{\text{new}}$
+18:end for
+19:return $\mathbf{O}\leftarrow s_{V}\cdot\frac{\mathbf{y}}{\mathbf{lse}[...,None]}$ permuted to $(B,N,H,E)$
+Step (1)–(2) implements Eq.
+(A.2); step (4)–(5) implements Eq.
+(A.3) while fusing the masked softmax with value aggregation via online log‑sum‑exp (as in FlashAttention), thereby avoiding any $\bm{\alpha}$ materialization.
+When $R_{K}{=}R_{V}{=}1$ the contractions reduce to the simpler einsums in Figure 2.
+Complexity and working set.
+Per block of $\Delta m$ cache items, the dominant FLOPs are
+$\Theta(B\,N\,\Delta m\,R_{Q}R_{K}D)\;+\;\Theta(B\,H\,N\,\Delta m\,R_{Q}R_{K})\;+\;\Theta(B\,H\,\Delta m\,R_{V}E),$
+matching the specialized analysis in Appendix A.2 and the decoding bounds in Appendix A.4.
+Peak memory scales with tiles of $\mathbf{B}_{K},\mathbf{A}_{K},\mathbf{A}_{V},\mathbf{B}_{V}$ and the small temporaries $\mathbf{P}$ and $\mathbf{V}_{\text{blk}}$; neither $\mathbf{Q},\mathbf{K},\mathbf{V}$ nor the full $N{\times}M$ attention matrix is formed.
+RoPE and masking.
+If keys are pre‑rotated (Eq.
+(3.5)), $\mathbf{B}_{K}^{\text{cache}}$ needs no decoding‑time rotation.
+Otherwise apply $\operatorname{RoPE}$ to $\mathbf{B}_{K,\text{blk}}$ row‑wise before step (1).
+The mask $\mathsf{mask}$ (zeros or -$\infty$) is added to logits in step (2) and supports both causal and padding masks.
+B.1 Triton FlashTPA Decoding Kernel
+We implement the experiments using Triton [57]; Algorithm 3 sketches the kernel corresponding to Algorithm 2.
+The provided kernel outline specializes to the frequently used case $R_{K}{=}R_{V}{=}1$; general ranks follow by tiling over $R_{K},R_{V}$ and replacing the rank‑1 vector–matrix products with the corresponding small GEMMs in steps $S1/S2$ and the value mixing path.
 Algorithm 3 Triton FlashTPA Decoding Kernel
@@
 6:
-7:$b\leftarrow p_{id_{B}}$; $h_{\text{start}}\leftarrow p_{id_{H}}\cdot B_{H}$
-8:$m_{\text{block}\_\text{start}}\leftarrow p_{id_{M}}\cdot M_{\text{block}}$; $m_{\text{block}\_\text{end}}\leftarrow\min((p_{id_{M}}+1)\cdot M_{\text{block}},M)$
-9$\triangleright$ $B_{H},B_{R},B_{D},B_{E}$ are tile sizes for dimensions H, R, D, E respectively.
+[…]
+[…]
+[…]
 10:
-11$\triangleright$ Initialize accumulators for the head block
-12:$\mathbf{o}_{\text{accum}}\leftarrow\mathbf{0}^{(E\times B_{H})}$; $\mathbf{m}_{\text{max}}\leftarrow-\infty^{(B_{H})}$; $\mathbf{s}_{\text{exp\_sum}}\leftarrow\mathbf{0}^{(B_{H})}$; $c_{\text{scale}}\leftarrow s_{\text{total}}\cdot s_{Q}\cdot s_{K}$
+[…]
+[…]
 13:
-14$\triangleright$ Load query factors (fixed for this program as N=1)
+[…]
 15:Load $\mathbf{A}_{Q,\text{local}}^{(R_{Q}\times B_{H})}$ from $\mathbf{A}_{Q}[b,0,h_{\text{start}}\dots,:]$
-16:Load $\mathbf{B}_{Q,\text{local}}^{(D\times R_{Q})}$ from $\mathbf{B}_{Q}[b,0,:,:]$triangleright Dimensions may be transposed after loading for matmul
+16:Load $\mathbf{B}_{Q,\text{local}}^{(D\times R_{Q})}$ from $\mathbf{B}_{Q}[b,0,:,:]$ $\triangleright$ Dimensions may be transposed after loading for matmul
 17:
-18$\triangleright$ Iterate over $M_{\text{chunk}}$-sized chunks within the K/V block
+[…]
 19:for $m_{\text{chunk\_start}}$ from $m_{\text{block}\_\text{start}}$ to $m_{\text{block}\_\text{end}}-1$ step $M_{\text{chunk}}$ do
@@
 22:  $\triangleright$ Load K/V factors for the current chunk
-23:    Load $\mathbf{a}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\mathbf{a}^{V}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\bm{b}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},D)$; $\bm{b}^{V}_{\text{chunk}}(E,M_{\text{curr}\_\text{chunk}})$triangleright Layouts optimized for memory access and matmuls
+23:  Load $\mathbf{a}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\mathbf{a}^{V}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\bm{b}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},D)$; $\bm{b}^{V}_{\text{chunk}}(E,M_{\text{curr}\_\text{chunk}})$ $\triangleright$ Layouts optimized for memory access and matmuls
 24:  $\bm{b}^{V}_{\text{chunk}}\leftarrow\bm{b}^{V}_{\text{chunk}}\cdot s_{V}$
 25:  $\triangleright$ Core TPA Score Calculation for the chunk
-26:    $S1_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{K}_{\text{chunk}},\mathbf{B}_{Q,\text{local}})$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},R_{Q})$
-27:    $S2_{\text{chunk}}\leftarrow\text{MatMul}(S1_{\text{chunk}},\mathbf{A}_{Q,\text{local}})$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
-28:    $S3_{\text{chunk}}\leftarrow S2_{\text{chunk}}\odot\mathbf{a}^{K}_{\text{chunk}}\cdot c_{\text{scale}}$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
+26:  $S1_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{K}_{\text{chunk}},\mathbf{B}_{Q,\text{local}})$ $\triangleright$ Shape: $(M_{\text{curr}\_\text{chunk}},R_{Q})$
+27:  $S2_{\text{chunk}}\leftarrow\text{MatMul}(S1_{\text{chunk}},\mathbf{A}_{Q,\text{local}})$ $\triangleright$ Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
+28:  $S3_{\text{chunk}}\leftarrow S2_{\text{chunk}}\odot\mathbf{a}^{K}_{\text{chunk}}\cdot c_{\text{scale}}$ $\triangleright$ Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
 29:  $\triangleright$ Online Softmax Update for the chunk
-30:    $\mathbf{m}_{\text{max\_local}}\leftarrow\max_{axis=0}(S3_{\text{chunk}})$triangleright Shape: $(B_{H})$
+30:  $\mathbf{m}_{\text{max\_local}}\leftarrow\max_{axis=0}(S3_{\text{chunk}})$ $\triangleright$ Shape: $(B_{H})$
 31:  $\mathbf{m}_{\text{max\_new}}\leftarrow\max(\mathbf{m}_{\text{max}},\mathbf{m}_{\text{max\_local}})$
@@
 34:  $\mathbf{p}_{\text{weighted\_av}}\leftarrow(\mathbf{p}_{\text{num}}/\mathbf{s}_{\text{exp\_sum\_local}}[\text{None},:])\odot\mathbf{a}^{V}_{\text{chunk}}$
-35:    $\mathbf{o}_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{V}_{\text{chunk}},\mathbf{p}_{\text{weighted\_av}})$triangleright Shape: $(E,B_{H})$
+35:  $\mathbf{o}_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{V}_{\text{chunk}},\mathbf{p}_{\text{weighted\_av}})$ $\triangleright$ Shape: $(E,B_{H})$
 36:  $\triangleright$ Update global (M-block level) accumulators
@@
 38:  $\mathbf{s}_{\text{exp\_sum}}\leftarrow\mathbf{s}_{\text{exp\_sum\_prev\_rescaled}}+\mathbf{s}_{\text{exp\_sum\_local}}$
-39:    $\text{ratio}\leftarrow\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum}}$triangleright This is $\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum\_new}}$
+39:  $\text{ratio}\leftarrow\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum}}$ $\triangleright$ This is $\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum\_new}}$
 40:  $\mathbf{o}_{\text{accum}}\leftarrow(1-\text{ratio})\cdot\mathbf{o}_{\text{accum}}+\text{ratio}\cdot\mathbf{o}_{\text{chunk}}$
@@
 43:
-44$\triangleright$ Store partial results for this program’s (batch, head_block, M_block)
+[…]
 45:Store $\mathbf{o}_{\text{accum}}$ into $\mathbf{O}_{\text{partial}}[b,p_{id_{M}},0,h_{\text{start}}\dots,:]$
-46:$\mathbf{LSE}_{\text{val}}\leftarrow\log(\mathbf{s}_{\text{exp\_sum}})+\mathbf{m}_{\text{max}}$
+[…]
 47:Store $\mathbf{LSE}_{\text{val}}$ into $\mathbf{LSE}_{\text{partial}}[b,p_{id_{M}},h_{\text{start}}\dots]$
-D.3 Additional Experimental Results
+B.2 Additional Experimental Results
 The following figures present additional speed comparisons for different embedding dimensions, with $d_{h}=64$ maintained.
 The y-axis represents $\log_{2}(\text{time})$ in seconds (lower is faster), and the x-axis represents $\log_{2}(\text{sequence length})$.
-Detailed Analysis of Figure 5 (Main Text; Embedding Dimension 2048): Figure 5 in the main paper depicts speed comparisons for an embedding dimension of 2048.
+Detailed Analysis of Figure 5 (Embedding Dimension 2048): Figure 5 in the main paper depicts speed comparisons for an embedding dimension of 2048.
 The results indicate that FlashTPA (blue line) is highly competitive.
@@
 Figure 8: Decoding time comparison of different attention mechanisms with an embedding dimension of 1024 and $d_{h}=64$.
-Appendix E Higher-Order Tensor Product Attention
+Appendix C Higher-Order Tensor Product Attention
 All prior discussions have focused on TPA where the query, key, and value matrices (e.g., $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$) are formed as a sum of $R_{Q}$ components.
@@
 One could also explore tying or sharing $\mathbf{c}^{Q}_{r}$ across queries, keys, and values to manage parameter overhead.
-From a memory perspective during inference, higher-order TPA maintains the benefit of factorized KV caching.
+From a memory perspective, during inference, higher-order TPA maintains the benefit of factorized KV caching.
 Only the constituent factors $\mathbf{a}_{K}(\mathbf{x}_{t}),\mathbf{b}_{K}(\mathbf{x}_{t}),\mathbf{c}_{K}(\mathbf{x}_{t})$ (and similarly for values) for each past token need to be stored.
@@
 Higher-order tensor decompositions can provide additional flexibility and potentially increased capacity.
-E.1 RoPE Compatibility in Higher-Order TPA
+C.1 RoPE Compatibility in Higher-Order TPA
 Rotary positional embeddings (RoPE) remain compatible with higher-order factorizations.
@@
 This is achieved by a specific RoPE transformation matrix $\mathbf{T}_{t}$ acting on the full $d_{h}$-dimensional vector, as stated in the following theorem.
-Theorem E.1 (RoPE Compatibility in Higher-Order TPA).
+Theorem C.1 (RoPE Compatibility in Higher-Order TPA).
 Consider the higher-order (3-order) Tensor Product Attention (TPA) query factorization
@@
 The transformation $\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top}$ operates on the $d_{h}$-dimensional vectorized features by post-multiplication.
-This structure of $\mathbf{T}_{t}$ ensures that the rotation effectively applied to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component (which is a column vector) corresponds to a pre-multiplication by $\mathbf{R}_{t}$, as detailed in the proof (Appendix F.2).
+This structure of $\mathbf{T}_{t}$ ensures that the rotation effectively applied to the $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$ component (which is a column vector) corresponds to a pre-multiplication by $\mathbf{R}_{t}$, as detailed in the proof (Appendix D.2).
 This preserves the structure induced by $\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})$ while rotating $\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
 Under these conditions, the RoPE-transformed query $\operatorname{RoPE}_{t}\bigl(\mathbf{Q}_{t}\bigr)$ admits a higher-order TPA factorization of the same rank $R_{Q}$:
-$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl{(}\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr{)}=\operatorname{RoPE}_{t}\bigl{(}\mathbf{Q}_{t}\bigr{)},$ (E.1)
+$\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\Bigl(\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\Bigr)=\operatorname{RoPE}_{t}\bigl(\mathbf{Q}_{t}\bigr),$ (C.1)
 where $\widetilde{\mathbf{b}}^{Q}_{r}(\mathbf{x}_{t})=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})$.
-Please see Appendix F.2 for the proof. For fourth-order or higher, this result still holds.
+Please see Appendix D.2 for the proof. For fourth-order or higher, this result still holds.
 To assess its empirical performance, we implemented third-order TPA.
@@
 Few-shot ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
-0-shot 49.24 24.91 57.06 34.01 31.80 63.33 50.59 23.23 66.90 44.56
-2-shot 53.37 25.34 48.78 34.00 29.20 62.79 52.33 26.41 75.30 45.28
-Table 5: The evaluation results of medium models with third-order TPA pre-trained using FineWeb-Edu 100B dataset with lm-evaluation-harness. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
-Few-shot ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
-0-shot 56.19 29.44 58.36 43.74 33.80 67.95 52.80 24.07 77.20 49.28
-2-shot 62.84 30.89 55.69 43.16 33.00 68.01 51.78 26.57 86.90 50.98
-Appendix F Proofs of Theorems
-F.1 Proof of Theorem 3.1
+0-shot 49.24 24.91 57.06 34.01 31.80 63.33 50.59 23.23 66.9 44.56
+2-shot 53.37 25.34 48.78 34.00 29.20 62.79 52.33 26.41 75.3 45.28
+Appendix D Proofs of Theorems
+D.1 Proof of Theorem 3.1
 Proof.
@@
 ∎
-F.2 Proof of Theorem E.1
-Theorem E.1 addresses the compatibility of RoPE with higher-order (specifically, 3rd-order) Tensor Product Attention. The theorem considers the query factorization:
+D.2 Proof of Theorem C.1
+Theorem C.1 addresses the compatibility of RoPE with higher-order (specifically, 3rd-order) Tensor Product Attention. The theorem considers the query factorization:
 $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}_{r}^{Q}(\mathbf{x}_{t})\otimes\operatorname{vec}\bigl(\mathbf{b}_{r}^{Q}(\mathbf{x}_{t})\otimes\mathbf{c}_{r}^{Q}(\mathbf{x}_{t})\bigr)\in\mathbb{R}^{h\times d_{h}},$
@@
 $\mathbf{T}_{t}=\mathbf{I}_{d_{c}}\otimes(\mathbf{R}_{t})^{\top},$
-where $\mathbf{I}_{d_{c}}$ is the $d_{c}\times d_{c}$ identity matrix, and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ is the standard RoPE block-diagonal matrix that pre-multiplies $d_{b}$-dimensional column vectors (as defined explicitly in the theorem statement in Section E).
+where $\mathbf{I}_{d_{c}}$ is the $d_{c}\times d_{c}$ identity matrix, and $\mathbf{R}_{t}\in\mathbb{R}^{d_{b}\times d_{b}}$ is the standard RoPE block-diagonal matrix that pre-multiplies $d_{b}$-dimensional column vectors (as defined explicitly in the theorem statement in Section C).
 The theorem claims that, under these conditions, $\widetilde{\mathbf{Q}}_{t}$ admits a higher-order TPA factorization:
@@
 ∎
-Appendix G More Related Work
+Appendix E More Related Work
 Transformers and Attention.
@@
 TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.
-Appendix H More on Attention Mechanisms
-H.1 Multi-Query Attention (MQA)
+Appendix F More on Attention Mechanisms
+F.1 Multi-Query Attention (MQA)
 Multi-Query Attention (MQA) [46] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection.
@@
 However, this comes at the cost of reduced model expressivity, as all heads must utilize the same key and value representations.
-H.2 Grouped Query Attention (GQA)
+F.2 Grouped Query Attention (GQA)
 Grouped Query Attention (GQA) [2] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups.
@@
 By varying $G$ from $1$ (equivalent to MQA) to $h$ (equivalent to MHA), GQA offers a trade-off between memory efficiency and model capacity.
-H.3 Multi-head Latent Attention (MLA)
+F.3 Multi-head Latent Attention (MLA)
 Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [32] and DeepSeek-V3 [33], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
@@
 In detail, if RoPE were ignored for the compressed components, the inner product $\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ (where $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}$) of the $i$-th head between $t$-th token query and $s$-th token key could be calculated using the current hidden state $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ and the cached latent state $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ for the $s$-th token:
-$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (H.1)
-$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (H.2)
+$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (F.1)
+$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (F.2)
 where $\bm{W}_{i}^{(\cdot)}$ denotes the $i$-th head’s portion of the respective weight matrix.
-The term $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding.
+The term $\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding.
 However, as noted by [51], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations.
-RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section H.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation H.4).
+RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section F.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation F.4).
 If RoPE were applied to the up-projected $Q^{C}$ and $K^{C}$:
-[…]
-Unlike Equation (H.2), acceleration by pre-computing the term $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}\mathbf{T}_{t-s}(\bm{W}^{UK}_{i})^{\top}]$ is not possible because it depends on the relative position $(t-s)$ and thus varies for different $(t,s)$ pairs.
+$\begin{aligned} \mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}&=[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t}}^{\top}(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{s}}^{\top}(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]\\ &=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t-s}}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s}.\end{aligned}$ (F.3)
+Unlike Equation (F.2), acceleration by pre-computing the term $\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}\mathbf{T}_{t-s}(\bm{W}^{UK}_{i})^{\top}]$ is not possible because it depends on the relative position (ts)(t-s) and thus varies for different $(t,s)$ pairs.
 To maintain RoPE compatibility while benefiting from compression, MLA introduces an additional, smaller key component $\mathbf{K}^{R}$ (and similarly $\mathbf{Q}^{R}$) to which RoPE is applied, while the main compressed components $\mathbf{K}^{C}$ and $\mathbf{V}^{C}$ (derived from $\mathbf{C}^{KV}$) remain RoPE-free.
 As we will demonstrate in Section 3.2 of the main paper, TPA offers a different approach to integrate RoPE efficiently with factorized attention through its tensor product formulation.
-H.4 Multi-matrix Factorization Attention (MFA)
+F.4 Multi-matrix Factorization Attention (MFA)
 [21] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
@@
 $\bm{W}^{DQ}\in\mathbb{R}^{d_{\text{model}}\times d_{c}},\quad\bm{W}^{UQ}_{i}\in\mathbb{R}^{d_{c}\times d_{c}},\quad\bm{W}^{K}_{\text{shared}},\bm{W}^{V}_{\text{shared}}\;\in\;\mathbb{R}^{\,d_{\text{model}}\times d_{c}}.$
-H.5 Rotary Position Embedding (RoPE)
-Many recent LLMs use rotary position embedding (RoPE; 51) to encode positional information in the query/key vectors.
+F.5 Rotary Position Embedding (RoPE)
+Many recent LLMs use rotary position embedding (RoPE; 52) to encode positional information in the query/key vectors.
 Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head).
@@
 This relies on the property:
-$\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (H.4)
-which embeds relative positional information $(t-s)$ into the attention scores.
-Appendix I More on TPA
+$\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (F.4)
+which embeds relative positional information (ts)(t-s) into the attention scores.
+Appendix G More on TPA
 Parameter Initialization for TPA Factors.
@@
 TPA with Non-contextual B.
-In Section A.1, we have introduced TPA with non-contextual A, where head-dimension factors $\mathbf{a}_{r}^{Q},\mathbf{a}_{r}^{K},\mathbf{a}_{r}^{V}\in\mathbb{R}^{h}$ are fixed.
+In Section 4.1, we have introduced TPA with non-contextual A, where head-dimension factors $\mathbf{a}_{r}^{Q},\mathbf{a}_{r}^{K},\mathbf{a}_{r}^{V}\in\mathbb{R}^{h}$ are fixed.
 Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to the input token $\mathbf{x}_{t}$.
@@
 This configuration might be effective if the fundamental token-level features (captured by $\mathbf{b}_{r}$) are relatively stable, while their combination across heads (captured by $\mathbf{a}_{r}(\mathbf{x}_{t})$) needs to adapt to the context.
-Performance comparisons for TPA with non-contextual A factors versus non-contextual B factors on small and medium-sized models are presented in Tables 6, 7, 8, and 9.
+Performance comparisons for TPA with non-contextual A factors versus non-contextual B factors on small and medium-sized models are presented in Tables 5, 6, 7, and 8.
 Table 5: Evaluation results of small models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
@@
 TPA KV with Shared $\mathbf{B}$. Further parameter reduction can be achieved by sharing the token-dimension factors $\mathbf{b}_{r}$ between keys and values:
-[…]
+$\mathbf{b}^{K}_{r}(\mathbf{x}_{t})=\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\quad(\text{if contextual}),\quad\text{or}\quad\mathbf{b}^{K}_{r}=\mathbf{b}^{V}_{r}\quad(\text{if non-contextual}).$
 This sharing reduces both parameter count and the KV cache footprint.
@@
 By carefully choosing which factor components (head-dimension or token-dimension) are contextual versus non-contextual, and by adjusting the ranks $(R_{Q},R_{K},R_{V})$, TPA can not only unify existing mechanisms like MHA, MQA, and GQA but also significantly reduce KV cache size—potentially by an order of magnitude—during autoregressive inference.
-Appendix J More on Experiments
-J.1 Experimental Settings
+Appendix H More on Experiments
+H.1 Experimental Settings
 We list the main architecture hyper-parameters and training devices in Table 9.
@@
 In particular, we use the AdamW [35] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$.
-We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [35] with $2{,}000$ warmup steps and a (total) global batch size of $480$.
+We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [36] with 2,0002{,}000 warmup steps and a (total) global batch size of $480$.
 For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
@@
 Model Size Parameters Devices Micro BS. GAS. #Layers $d_{\text{model}}$
-Small 124M 4$\times$ A100 GPUs 24 5 12 768
-Medium 353M 8$\times$ A100 GPUs 20 3 24 1024
-Large 772M 8$\times$ A100 GPUs 15 4 36 1280
-XL 1.55B 8$\times$ A100 GPUs 6 10 48 1600
+[…]
+[…]
+[…]
+[…]
 Table 10: The architecture hyper-parameters for different models.
@@
 $d_{c}^{\prime}$ (MLA) 512 1024 1024 1024
-J.2 Additional Experimental Results
-J.2.1 Perplexity Curves
+H.2 Additional Experimental Results
+H.2.1 Perplexity Curves
 We display the perplexity curves for medium, large, and XL size models in Figure 9.
@@
 Figure 9: The validation perplexity of medium-size (353M) models, large-size (773M), and XL-size (1.5B) models with different attention mechanisms on the FineWeb-Edu 100B dataset.
-J.2.2 Ablation Study on Different Ranks
+H.2.2 Ablation Study on Different Ranks
 Figure 10 illustrates the training loss, validation loss, and validation perplexity for XL-sized (1.5B parameters) TPA models with varying key/value ranks ($R_{K}=R_{V}=R$, as indicated in the figure legend), trained on the FineWeb-Edu 100B dataset.
@@
 Figure 10: The training loss, validation loss and validation perplexity curves of XL-size (1.5B) TPA models with different key/value ranks ($R_{K}=R_{V}=R$) on the FineWeb-Edu 100B dataset.
-J.2.3 0-shot Evaluation with lm-evaluation-harness
-We present 0-shot evaluation results using the lm-evaluation-harness for small (124M parameters) and XL (1.5B parameters) models in Tables 12 and 13, respectively.
+H.2.3 0-shot Evaluation with lm-evaluation-harness
+We present 0-shot evaluation results using the lm-evaluation-harness for small (124M parameters) and XL (1.5B parameters) models in Tables 11 and 12, respectively.
 Table 11: Evaluation results of small models (124M) with different attention mechanisms, pre-trained on FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness).
@@
 TPA 66.71 36.52 61.38 54.03 40.40 72.52 56.83 24.49 82.20 55.01
-J.2.4 2-shot Evaluation with lm-evaluation-harness
-Similarly, 2-shot evaluation results are provided in Tables 14 (Small), 15 (Medium), 16 (Large), and 17 (XL).
+H.2.4 2-shot Evaluation with lm-evaluation-harness
+Similarly, 2-shot evaluation results are provided in Tables 13 (Small), 14 (Medium), 15 (Large), and 16 (XL).
 Table 13: Evaluation results of small models (124M) with different attention mechanisms, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness).
@@
 TPA 71.76 39.16 61.25 53.74 37.80 72.80 55.49 23.86 90.70 56.28
-J.3 Ablation Studies on Learning Rates
+H.3 Ablation Studies on Learning Rates
 To assess sensitivity to learning rates, we conducted parallel experiments on medium-sized models using a learning rate of $3\times 10^{-4}$ (compared to the default $6\times 10^{-4}$ used for other medium model results).
 The training loss, validation loss, and validation perplexity curves are shown in Figure 11.
-Performance on standard benchmarks for these models trained with the $3\times 10^{-4}$ learning rate are reported in Tables 18 (0-shot) and 19 (2-shot).
+Performance on standard benchmarks for these models trained with the $3\times 10^{-4}$ learning rate are reported in Tables 17 (0-shot) and 18 (2-shot).
 The results demonstrate that TPA and TPA-KVonly maintain their performance advantages over other attention mechanisms even with this alternative learning rate.
@@
 TPA 66.54 34.47 58.96 45.35 33.00 69.21 53.99 24.51 91.30 53.04
+Appendix I Broader Impacts and Limitations
+This work allows for the processing of much longer sequences of information with limited hardware resources by reducing the KV cache size.
+This could make advanced AI capabilities accessible to entities with limited computational budgets, potentially fostering improvement on downstream tasks, including in-depth document analysis, complicated-context reasoning, and code generation, promoting innovation across various sectors in fields of scientific research, education, and software development.
+Although our work proposes a KV-cache efficient architecture for large language models, it may contain certain limitations.
+For instance, generalization to other modalities deserves more extensive investigation.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We describe all the contributions and scope in the abstract and introduction parts.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We discussed the limitations in Appendix I.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We list all the assumptions and proofs in Appendix D.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We listed all the experiment details in Section 6 for reproduction of our work.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: The code and data required to reproduce the main experimental results are provided at https://anonymous.4open.science/r/T6-anonymous-2025.
+The supplemental material will contain instructions for their use.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We just list all the training and test details in Section 6.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: The error bars are not reported because it would be too computationally expensive for repeated experiments on LLMs.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We just list all the computer resources in Section 6.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44.
+Justification: Our research only explores a novel framework for large language models with better KV-Cache efficiency.
+Therefore, the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We discussed the potential positive societal impacts and negative societal impacts in Appendix I.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: Our work proposes a novel framework of large language models. To our knowledge, this work has no direct path to any negative applications.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59.
+Justification: We add the citation to all the codes (nanoGPT and lm-evaluation-harness: MIT License) and datasets (FineWeb-Edu-100B: odc-by) that we used in this work.
+No other models are included in our work.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64.
+Justification: The code implementing our proposed TPA model and experimental setup is released at https://anonymous.4open.science/r/T6-anonymous-2025.
+This code will be documented to facilitate understanding and use by other researchers.
+No new datasets or pre-trained models are introduced beyond the code for the methods.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper only use open-source codes and datasets which do not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper only use open-source codes and datasets which do not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: This work aims at exploring more efficient architecture for large language models. Therefore, LLM architectures are well described in the main part of this paper.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v6

@@
 1 Introduction
-Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (brown2020language, ; chowdhery2022palm, ; touvron2023llama, ; bubeck2023sparks, ).
+Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (brown2020language; chowdhery2022palm; touvron2023llama; bubeck2023sparks).
 As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completion.
-However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (zhang2023h2o, ; liu2024kivi, ).
+However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (zhang2023h2o; liu2024kivi).
 Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
@@
 A variety of solutions have been explored to address this memory bottleneck.
-Some approaches compress or selectively prune cached states through sparse attention patterns (child2019generating, ) or token eviction strategies (zhang2023h2o, ; xiao2023efficient, ; ribar2023sparq, ), though such methods risk discarding tokens that may later prove important.
-Other work proposes off-chip storage of key-value states (he2024fastdecode, ), at the expense of increased I/O latency.
-Attention variants like Multi-Query Attention (MQA) (shazeer2019fast, ) and Grouped-Query Attention (GQA) (ainslie2023gqa, ) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications.
-Meanwhile, low-rank weight factorization methods such as LoRA (hu2021lora, ) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime.
-The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (liu2024deepseek, ) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (su2024roformer, ) integration, necessitating additional position-encoded parameters per head.
+Some approaches compress or selectively prune cached states through sparse attention patterns (child2019generating) or token eviction strategies (zhang2023h2o; xiao2023efficient; ribar2023sparq), though such methods risk discarding tokens that may later prove important.
+Other work proposes off-chip storage of key-value states (he2024fastdecode), at the expense of increased I/O latency.
+Attention variants like Multi-Query Attention (MQA) (shazeer2019fast) and Grouped-Query Attention (GQA) (ainslie2023gqa) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications.
+Meanwhile, low-rank weight factorization methods such as LoRA (hu2021lora) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime.
+The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (liu2024deepseek) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (su2024roformer) integration, necessitating additional position-encoded parameters per head.
 To overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), illustrated in Figure 1.
@@
 In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance.
-A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer, ), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama, ), Qwen (bai2023qwen, ), and Gemma (team2024gemma, ).
+A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer) and any possible position encodings, enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama), Qwen (bai2023qwen), and Gemma (team2024gemma).
 Our main contributions are summarized as follows:
@@
     We propose Tensor Product Attention (TPA), a mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor decompositions.
-This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (vaswani2017attention, ), MHA, MQA, GQA, and MLA, while also improving performance.
+This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (vaswani2017attention), MHA, MQA, GQA, and MLA, while also improving performance.
 In addition, we analyze existing attention mechanisms and reveal that MHA, MQA, and GQA can be expressed as non-contextual variants of TPA.
@@
   3. 3.
-    We demonstrate that TPA integrates seamlessly with RoPE (su2024roformer, ), facilitating its easy adoption in popular foundation model architectures like LLaMA and Gemma.
+    We demonstrate that TPA integrates seamlessly with RoPE (su2024roformer) and any possible position encodings as well as output gate and KV shifting, facilitating its easy adoption in popular foundation model architectures like LLaMA, Gemma, and Qwen.
   4. 4.
@@
 2 Background
-In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention, ), and introduce key notations.
-Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast, ), Grouped Query Attention (GQA) (ainslie2023gqa, ), Multi-head Latent Attention (MLA) (liu2024deepseek, ; liu2024deepseekv3, ), and Rotary Position Embedding (RoPE) (su2024roformer, ) are further discussed in the Appendix F.
+In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention), and introduce key notations.
+Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast), Grouped Query Attention (GQA) (ainslie2023gqa), Multi-head Latent Attention (MLA) (liu2024deepseek; liu2024deepseekv3), and Rotary Position Embedding (RoPE) (su2024roformer) are further discussed in the Appendix F.
 Notations.
@@
 2.1 Scaled Dot-Product Attention
-Scaled dot-product attention (vaswani2017attention, ) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$).
+Scaled dot-product attention (vaswani2017attention) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$).
 It produces a weighted combination of the values ($\mathbf{V}$).
@@
 2.2 Multi-Head Attention (MHA)
-Multi-Head Attention (MHA) (vaswani2017attention, ) extends scaled dot-product attention by dividing the model’s internal representation into several heads.
+Multi-Head Attention (MHA) (vaswani2017attention) extends scaled dot-product attention by dividing the model’s internal representation into several heads.
 Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information from different representational subspaces.
@@
 $\operatorname{TPA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl(\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr)\bm{W}^{O}.$ (3.3)
-Parameter Initialization. We use Xavier initialization (glorot2010understanding, ) for the factor weight matrices; details are in the Appendix G.
+Parameter Initialization. We use Xavier initialization (glorot2010understanding) for the factor weight matrices; details are in the Appendix G.
 3.2 RoPE Compatibility and Acceleration
@@
 We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention).
-Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (touvron2023llama, ) while changing the self-attention block to our TPA-based version.
-The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu, ; touvron2023llama, ).
+Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (touvron2023llama) while changing the self-attention block to our TPA-based version.
+The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu; touvron2023llama).
 Rotary Positional Embedding (RoPE).
-As discussed in Section 3.2, RoPE (su2024roformer, ) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
+As discussed in Section 3.2, RoPE (su2024roformer) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
 Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see Equation (3.5) and Theorem 3.1.
@@
 6.1 Language Modeling Tasks
-All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy2022, ), and we pretrain our models using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu, ).
+All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy2022), and we pretrain our models using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu).
 The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation.
-We compare T6 against the baseline Llama architecture (touvron2023llama, ) with SwiGLU activation (shazeer2020glu, ) and RoPE embeddings (su2024roformer, ), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention ) with Multi-Query Attention (MQA; shazeer2019fast ), Grouped Query Attention (GQA; ainslie2023gqa ), or Multi-head Latent Attention (MLA; liu2024deepseek ).
+We compare T6 against the baseline Llama architecture (touvron2023llama) with SwiGLU activation (shazeer2020glu) and RoPE embeddings (su2024roformer), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention) with Multi-Query Attention (MQA; shazeer2019fast), Grouped Query Attention (GQA; ainslie2023gqa), or Multi-head Latent Attention (MLA; liu2024deepseek).
 In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer.
@@
 For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages.
-Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek, ) (blue curves) generally trains more slowly and yields higher validation losses.
+Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek) (blue curves) generally trains more slowly and yields higher validation losses.
 Validation Perplexity.
@@
 Downstream Evaluation.
-We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC, ), BoolQ (BoolQ, ), HellaSwag (hellaswag, ), OBQA (OpenBookQA, ), PIQA (PIQA, ), WinoGrande (WinoGrande, ), and MMLU (MMLU, ), using the lm-evaluation-harness codebase (eval-harness, ).
+We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC), BoolQ (BoolQ), HellaSwag (hellaswag), OBQA (OpenBookQA), PIQA (PIQA), WinoGrande (WinoGrande), and MMLU (MMLU), using the lm-evaluation-harness codebase (eval-harness).
 For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy.
@@
 This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms.
-We benchmark FlashTPA against FlashMHA (shah2024flashattention, ), FlashGQA, FlashMQA, and FlashMLA (flashmla2025, ).
-It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton, ).
+We benchmark FlashTPA against FlashMHA (shah2024flashattention), FlashGQA, FlashMQA, and FlashMLA (flashmla2025).
+It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton).
 While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential.
@@
 We thank the anonymous reviewers and area chair for their helpful comments. We acknowledge the compute credits provided by Fetch.ai.
+[…]
 Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
@@
 B.1 Triton FlashTPA Decoding Kernel
-We implement the experiments using Triton [57]; Algorithm 3 sketches the kernel corresponding to Algorithm 2.
+We implement the experiments using Triton [tillet2019triton]; Algorithm 3 sketches the kernel corresponding to Algorithm 2.
 The provided kernel outline specializes to the frequently used case $R_{K}{=}R_{V}{=}1$; general ranks follow by tiling over $R_{K},R_{V}$ and replacing the rank‑1 vector–matrix products with the corresponding small GEMMs in steps $S1/S2$ and the value mixing path.
@@
 Transformers and Attention.
-As a sequence-to-sequence architecture, Transformer [60] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies.
-Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [10, 49, 16, 30, 27, 31], kernel-based projections [11], and linearized transformers [59, 25, 44, 69, 54, 67].
-To decrease memory usage and circumvent the limitation of memory bandwidth in training, [46] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head.
-To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [2] divides queries into several groups, and each group of queries shares a single key head and value head.
-Recently, DeepSeek-V2 [32] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value.
-Concurrently, [21] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q.
+As a sequence-to-sequence architecture, Transformer [vaswani2017attention] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies.
+Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [child2019generating, DBLP:conf/iclr/Shi0LR0LJ23, DBLP:conf/iclr/HanJKMWZ24, liang2024conv, li2024tighter, liang2024beyond], kernel-based projections [choromanski2020rethinking], and linearized transformers [tsai2019transformer, katharopoulos2020transformers, schlag2021linear, zhang2023trained, sun2023retentive, DBLP:conf/iclr/ZhangBKR24].
+To decrease memory usage and circumvent the limitation of memory bandwidth in training, [shazeer2019fast] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head.
+To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [ainslie2023gqa] divides queries into several groups, and each group of queries shares a single key head and value head.
+Recently, DeepSeek-V2 [liu2024deepseek] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value.
+Concurrently, [hu2024multi] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q.
 Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
 KV Cache Optimization.
-During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [40].
-This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [1].
+During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [ott2019fairseq].
+This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [adnan2024keyformer].
 Consequently, various studies have explored methods to mitigate these issues.
-These include KV cache eviction strategies that discard less significant tokens [70, 62, 8, 1], dynamic sparse attention mechanisms focusing on selected keys and values [42, 55, 50], offloading the KV cache to CPU memory [17, 26, 53], and quantizing the KV cache [61, 34, 19].
+These include KV cache eviction strategies that discard less significant tokens [zhang2023h2o, xiao2023efficient, cai2024pyramidkv, adnan2024keyformer], dynamic sparse attention mechanisms focusing on selected keys and values [ribar2023sparq, tang2024quest, singhania2024loki], offloading the KV cache to CPU memory [he2024fastdecode, lee2024infinigen, sun2024shadowkv], and quantizing the KV cache [xiao2023smoothquant, liu2024kivi, hooper2024kvquant].
 In contrast to these approaches, TPA focuses on reducing the intrinsic size of the KV cache by employing tensor-decomposed key and value representations.
@@
 Low-rank approximations are widely used to compress model parameters and reduce computational complexity.
-Notable examples include LoRA [20], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [28], MoRA [22]), long-context training (LongLoRA [9], SinkLoRA [66]), and continual training (InfLoRA [29], GS-LoRA [71], I-LoRA [41]).
+Notable examples include LoRA [hu2021lora], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [lialin2023relora], MoRA [jiang2024mora]), long-context training (LongLoRA [chen2023longlora], SinkLoRA [zhang2024sinklora]), and continual training (InfLoRA [liang2024inflora], GS-LoRA [zhao2024continual], I-LoRA [ren2024analyzing]).
 These methods generally produce static low-rank expansions that are independent of the input context.
-Theoretical justifications for the expressiveness of low-rank approximations have been provided by [38, 65].
-Initialization strategies for these factorization matrices have also been explored: OLoRA [7] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [48] employs LDU-decomposition to accelerate LoRA training.
-Furthermore, AdaLoRA [68] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks.
+Theoretical justifications for the expressiveness of low-rank approximations have been provided by [malladi2023kernel, zeng2023expressive].
+Initialization strategies for these factorization matrices have also been explored: OLoRA [buyukakyuz2024olora] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [shi2024loldu] employs LDU-decomposition to accelerate LoRA training.
+Furthermore, AdaLoRA [zhang2023adalora] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks.
 TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.
@@
 F.1 Multi-Query Attention (MQA)
-Multi-Query Attention (MQA) [46] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection.
+Multi-Query Attention (MQA) [shazeer2019fast] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection.
 Given a sequence of input embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$, the query, shared key, and shared value tensors are computed as:
@@
 F.2 Grouped Query Attention (GQA)
-Grouped Query Attention (GQA) [2] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups.
+Grouped Query Attention (GQA) [ainslie2023gqa] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups.
 Within each group, heads share a common key and value projection, while each head maintains its own unique query projection.
@@
 F.3 Multi-head Latent Attention (MLA)
-Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [32] and DeepSeek-V3 [33], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
+Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [liu2024deepseek] and DeepSeek-V3 [liu2024deepseekv3], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
 $\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},$
@@
 The RoPE transformation is applied to a separate key embedding $\mathbf{K}^{R}\in\mathbb{R}^{T\times d_{h}^{R}}$.
-Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached, reducing KV memory usage while largely preserving performance compared to standard MHA [60].
+Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached, reducing KV memory usage while largely preserving performance compared to standard MHA [vaswani2017attention].
 MLA also compresses the queries, lowering their training-time memory footprint:
@@
 The term $\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding.
-However, as noted by [51], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations.
+However, as noted by [kexuefm-10091], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations.
 RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section F.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation F.4).
@@
 F.4 Multi-matrix Factorization Attention (MFA)
-[21] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
+[hu2024multi] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
 $\mathbf{Q}_{i}=\mathbf{X}\bm{W}^{DQ}\bm{W}^{UQ}_{i},\quad\mathbf{K}_{\text{shared}}=\mathbf{X}\bm{W}^{K}_{\text{shared}},\quad\mathbf{V}_{\text{shared}}=\mathbf{X}\bm{W}^{V}_{\text{shared}},$
@@
 F.5 Rotary Position Embedding (RoPE)
-Many recent LLMs use rotary position embedding (RoPE; 52) to encode positional information in the query/key vectors.
+Many recent LLMs use rotary position embedding (RoPE; su2024roformer) to encode positional information in the query/key vectors.
 Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head).
@@
 Parameter Initialization for TPA Factors.
-We initialize the weight matrices for TPA factors, such as $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, and $\bm{W}_{r}^{b^{V}}$ (or their combined forms $\bm{W}^{a^{Q}}$, $\bm{W}^{b^{Q}}$, etc.), using Xavier initialization [15].
+We initialize the weight matrices for TPA factors, such as $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, and $\bm{W}_{r}^{b^{V}}$ (or their combined forms $\bm{W}^{a^{Q}}$, $\bm{W}^{b^{Q}}$, etc.), using Xavier initialization [glorot2010understanding].
 Specifically, each entry of a weight matrix is drawn from a uniform distribution $\mathcal{U}(-bound,bound)$, where $bound=\sqrt{6/(n_{\text{in}}+n_{\text{out}})}$.
@@
 Training Setup Details.
-We follow the nanoGPT training configuration [24].
-In particular, we use the AdamW [35] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$.
-We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [36] with 2,0002{,}000 warmup steps and a (total) global batch size of $480$.
+We follow the nanoGPT training configuration [Karpathy2022].
+In particular, we use the AdamW [loshchilov2017decoupled] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$.
+We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [loshchilov2016sgdr] with 2,0002{,}000 warmup steps and a (total) global batch size of $480$.
 For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).

### v7

@@
 In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, substantially shrinking the KV cache size at inference time.
-By factorizing these representations into contextual low-rank components and seamlessly integrating with Rotary Position Embedding (RoPE), TPA achieves improved model quality alongside memory efficiency.
+By factorizing these representations into contextual low-rank components and seamlessly integrating with RoPE and any possible position encoding mechanisms, TPA achieves improved model quality alongside memory efficiency.
 Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling.
@@
 1 Introduction
-Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (brown2020language; chowdhery2022palm; touvron2023llama; bubeck2023sparks).
+Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Bubeck et al., 2023).
 As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completion.
-However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (zhang2023h2o; liu2024kivi).
+However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (Zhang et al., 2023c; Liu et al., 2024c).
 Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
@@
 A variety of solutions have been explored to address this memory bottleneck.
-Some approaches compress or selectively prune cached states through sparse attention patterns (child2019generating) or token eviction strategies (zhang2023h2o; xiao2023efficient; ribar2023sparq), though such methods risk discarding tokens that may later prove important.
-Other work proposes off-chip storage of key-value states (he2024fastdecode), at the expense of increased I/O latency.
-Attention variants like Multi-Query Attention (MQA) (shazeer2019fast) and Grouped-Query Attention (GQA) (ainslie2023gqa) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications.
-Meanwhile, low-rank weight factorization methods such as LoRA (hu2021lora) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime.
-The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (liu2024deepseek) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (su2024roformer) integration, necessitating additional position-encoded parameters per head.
+Some approaches compress or selectively prune cached states through sparse attention patterns (Child et al., 2019) or token eviction strategies (Zhang et al., 2023c; Xiao et al., 2024; Ribar et al., 2024), though such methods risk discarding tokens that may later prove important.
+Other work proposes off-chip storage of key-value states (He and Zhai, 2024), at the expense of increased I/O latency.
+Attention variants like Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped-Query Attention (GQA) (Ainslie et al., 2023) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications.
+Meanwhile, low-rank weight factorization methods such as LoRA (Hu et al., 2022) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime.
+The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (Su et al., 2024) integration, necessitating additional position-encoded parameters per head.
 To overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), illustrated in Figure 1.
@@
 In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance.
-A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer) and any possible position encodings, enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama), Qwen (bai2023qwen), and Gemma (team2024gemma).
+A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024) and any possible position encodings, enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and Gemma (Team et al., 2024).
 Our main contributions are summarized as follows:
@@
     We propose Tensor Product Attention (TPA), a mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor decompositions.
-This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (vaswani2017attention), MHA, MQA, GQA, and MLA, while also improving performance.
+This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (Vaswani et al., 2017), MHA, MQA, GQA, and MLA, while also improving performance.
 In addition, we analyze existing attention mechanisms and reveal that MHA, MQA, and GQA can be expressed as non-contextual variants of TPA.
@@
   3. 3.
-    We demonstrate that TPA integrates seamlessly with RoPE (su2024roformer) and any possible position encodings as well as output gate and KV shifting, facilitating its easy adoption in popular foundation model architectures like LLaMA, Gemma, and Qwen.
+    We demonstrate that TPA integrates seamlessly with RoPE (Su et al., 2024) and any possible position encodings as well as output gate and KV shifting, facilitating its easy adoption in popular foundation model architectures like LLaMA, Gemma, and Qwen.
   4. 4.
@@
 2 Background
-In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention), and introduce key notations.
-Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast), Grouped Query Attention (GQA) (ainslie2023gqa), Multi-head Latent Attention (MLA) (liu2024deepseek; liu2024deepseekv3), and Rotary Position Embedding (RoPE) (su2024roformer) are further discussed in the Appendix F.
+In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (Vaswani et al., 2017), and introduce key notations.
+Other attention mechanisms like Multi-Query Attention (MQA) (Shazeer, 2019), Grouped Query Attention (GQA) (Ainslie et al., 2023), Multi-head Latent Attention (MLA) (Liu et al., 2024a, b), and Rotary Position Embedding (RoPE) (Su et al., 2024) are further discussed in the Appendix F.
 Notations.
@@
 We use $\top$ to denote the transpose of a vector or a matrix.
-Let $d_{\text{model}}$ be the embedding dimension, $h$ the number of attention heads, $d_{h}$ the dimension per head, $\mathbf{x}_{t}\in\mathbb{R}^{d}$ the input for the $t$-th token at a given attention layer, $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ denotes the input embeddings for $T$ tokens, and $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ denote the queries, keys, and values of $h$ heads for $T$ tokens.
+Let $d_{\text{model}}$ be the embedding dimension, $h$ the number of attention heads, $d_{h}$ the dimension per head, $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ the input for the $t$-th token at a given attention layer, $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ denotes the input embeddings for $T$ tokens, and $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ denote the queries, keys, and values of $h$ heads for $T$ tokens.
 With a little abuse of notation, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, $\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ denote the $i$-th head of queries, keys, and values, and $\mathbf{Q}_{t}$, $\mathbf{K}_{t}$, $\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the heads of the query, key, and value for $t$-th token.
 Throughout the paper, $\bm{W}^{Q},\bm{W}^{K},\bm{W}^{V}$ denote projection matrices for queries, keys, and values, respectively.
-In multi-head attention, each head is associated with its own set of $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}$, and each has dimension $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$, where $d_{k}$ is typically set to $d_{h}$, the dimension of each head.
+In multi-head attention, each head is associated with its own set of $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}$, and each has dimension $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{h}}$.
 [^5] Often, $h\times d_{h}=d_{\text{model}}$, so each head has query/key/value dimension $d_{h}$.
 Similarly, we have an output projection matrix $\bm{W}^{O}\in\mathbb{R}^{(h\cdot d_{h})\times d_{\text{model}}}$.
-For methods like MQA and GQA, some of these projection matrices are shared or partially shared across heads, but their shapes remain consistent.
 We define the tensor product of two vectors as follows: for vectors $\mathbf{a}\in\mathbb{R}^{m},\mathbf{b}\in\mathbb{R}^{n}$, the tensor product of $\mathbf{a}$ and $\mathbf{b}$ is: $\mathbf{a}\otimes\mathbf{b}=\mathbf{C}\in\mathbb{R}^{m\times n},\text{with}~C_{ij}=a_{i}b_{j}$, where $a_{i}$ is the $i$-th element of $\mathbf{a}$, $b_{j}$ is the $j$-th element of $\mathbf{b}$, and $C_{ij}$ is the $(i,j)$-th entry of $\mathbf{C}$.
@@
 2.1 Scaled Dot-Product Attention
-Scaled dot-product attention (vaswani2017attention) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$).
+Scaled dot-product attention (Vaswani et al., 2017) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$).
 It produces a weighted combination of the values ($\mathbf{V}$).
 Formally, the attention output is:
-$\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Softmax}\Bigl(\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\Bigr)\,\mathbf{V},$
-where each of $\mathbf{Q},\mathbf{K},\mathbf{V}$ is an $(n\times d_{k})$ matrix for $n$ tokens and key dimension $d_{k}$.
+$\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Softmax}\Bigl(\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{h}}}\Bigr)\,\mathbf{V},$
+where $\mathbf{Q}\in\mathbb{R}^{n\times d_{h}}$, $\mathbf{K}\in\mathbb{R}^{n\times d_{h}}$, and $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$ for $n$ tokens.
+The softmax is applied row-wise over the $n$ keys for each query.
 2.2 Multi-Head Attention (MHA)
-Multi-Head Attention (MHA) (vaswani2017attention) extends scaled dot-product attention by dividing the model’s internal representation into several heads.
+Multi-Head Attention (MHA) (Vaswani et al., 2017) extends scaled dot-product attention by dividing the model’s internal representation into several heads.
 Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information from different representational subspaces.
@@
 3.1 Tensor Factorization of Queries, Keys, and Values
-Let $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ for $t=1,\dots,T$ be the hidden-state vector corresponding to the $t$-th token in a sequence of length $T$.
-A typical multi-head attention block has $h$ heads, each of dimension $d_{h}$, satisfying $d_{\text{model}}=h\times d_{h}$.
+Let $d_{\text{attn}}:=h\,d_{h}$ denote the total attention projection dimension.
+Typically one sets $d_{\text{attn}}=d_{\text{model}}$, but this is not required: when $d_{\text{attn}}\neq d_{\text{model}}$, the projection matrices $\bm{W}^{Q},\bm{W}^{K},\bm{W}^{V}$ map from $\mathbb{R}^{d_{\text{model}}}$ into $\mathbb{R}^{d_{\text{attn}}}$ and $\bm{W}^{O}$ maps $\mathbb{R}^{d_{\text{attn}}}$ back to $\mathbb{R}^{d_{\text{model}}}$.
 Standard attention projects the entire sequence into three tensors, $\mathbf{Q},\;\mathbf{K},\;\mathbf{V}\;\in\mathbb{R}^{T\times h\times d_{h}}$, where $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the slices for the $t$-th token.
 Contextual Factorization.
-Instead of forming each head’s query, key, or value via a single linear map, TPA factorizes each $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$ into a sum of (contextual) tensor products whose ranks are $R_{q}$, $R_{k}$, and $R_{v}$, respectively and may differ.
+Instead of forming each head’s query, key, or value via a single linear map, TPA factorizes each $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$ into a sum of (contextual) tensor products whose ranks are $R_{Q}$, $R_{K}$, and $R_{V}$, respectively, and may differ.
 Specifically, for each token $t$, with a small abuse of notation, we define:
@@
 where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$.
-Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ (an outer product) contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$.
+Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$.
 Analogous definitions apply to the key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
 $\operatorname{TPA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl(\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr)\bm{W}^{O}.$ (3.3)
-Parameter Initialization. We use Xavier initialization (glorot2010understanding) for the factor weight matrices; details are in the Appendix G.
+Parameter Initialization. We use Xavier initialization (Glorot and Bengio, 2010) for the factor weight matrices; details are in the Appendix G.
 3.2 RoPE Compatibility and Acceleration
@@
 Direct Integration. A useful optimization is to integrate RoPE directly into the TPA factorization. For example, one can pre-rotate the token-dimension factors:
-$\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\;\longleftarrow\;\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{K}(\mathbf{x}_{t})\bigr),$ (3.5)
+$\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\;:=\;\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{K}(\mathbf{x}_{t})\bigr)=\mathbf{B}_{K}(\mathbf{x}_{t})\mathbf{T}_{t},$ (3.5)
 yielding a pre-rotated key representation:
-$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl(\mathbf{b}^{K}_{(r)}(\mathbf{x}_{t})\bigr)=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{K}(\mathbf{x}_{t})\bigr).$
-Here, $\operatorname{RoPE}_{t}$ is applied to each row of $\mathbf{B}_{K}(\mathbf{x}_{t})$ (i.e., to each $\mathbf{b}^{K}_{(r)}(\mathbf{x}_{t})$ vector).
-Thus, each $\mathbf{K}_{t}$ is effectively rotated before caching.
-This removes the need for explicit rotation at decoding time, accelerating autoregressive inference.
+$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl(\mathbf{b}^{K}_{r}(\mathbf{x}_{t})\bigr)=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t}).$
+Here, $\operatorname{RoPE}_{t}$ is applied to each row of $\mathbf{B}_{K}(\mathbf{x}_{t})$ (i.e., to each $\mathbf{b}^{K}_{r}(\mathbf{x}_{t})$ vector).
+Thus, each cached key factor corresponds to a RoPE-rotated key slice.
+This removes the need to rotate cached keys at decoding time; the current-step query (which is not cached) can still be rotated on the fly at negligible cost.
 Depending on hardware and performance requirements, different RoPE integration strategies can be adopted for training and inference.
@@
 where $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times h}$ and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times d_{h}}$. Then we have:
-$\operatorname{RoPE}({\mathbf{Q}}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),$ (3.6)
-where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr)$ (RoPE applied row-wise to $\mathbf{B}_{Q}(\mathbf{x}_{t})$).
-Furthermore, let $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ be factorized by TPA.
-Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$ be their RoPE-transformed versions.
-The relative positional encoding property of RoPE is preserved:
-$\operatorname{RoPE}_{t-s}(\mathbf{Q}_{t})\mathbf{K}_{s}^{\top}=\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top},$
-where $\operatorname{RoPE}_{t-s}$ denotes applying RoPE with relative position tst-s.
-Focusing on individual heads $i$, if $\mathbf{q}_{t,i}$ and $\mathbf{k}_{s,i}$ are the $i$-th head vectors from $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ respectively (as column vectors of dimension $d_{h}$), and $\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i},t)$ and $\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i},s)$, then the dot product for attention scores satisfies:
-$\widetilde{\mathbf{q}}_{t,i}^{\top}\,\widetilde{\mathbf{k}}_{s,i}=\mathbf{q}_{t,i}^{\top}\operatorname{RoPE}(\cdot,t-s)\mathbf{k}_{s,i}.$
+$\operatorname{RoPE}_{t}(\mathbf{Q}_{t})=\mathbf{Q}_{t}\mathbf{T}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),$ (3.6)
+where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}):=\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{T}_{t}=\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr)$ (RoPE applied row-wise to $\mathbf{B}_{Q}(\mathbf{x}_{t})$).
+Furthermore, let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})=\mathbf{Q}_{t}\mathbf{T}_{t}$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})=\mathbf{K}_{s}\mathbf{T}_{s}$ be the RoPE-transformed query/key slices.
+Then RoPE’s standard relative-position identity is preserved:
+$\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top}\;=\;\mathbf{Q}_{t}\mathbf{T}_{t-s}\mathbf{K}_{s}^{\top},\qquad\text{equivalently}\qquad\operatorname{RoPE}_{t-s}(\mathbf{Q}_{t})\,\mathbf{K}_{s}^{\top}\;=\;\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top},$
+where $\mathbf{T}_{t-s}:=\mathbf{T}_{t}\mathbf{T}_{s}^{\top}$.
+In particular, for any head $i$ (the $i$-th row), if $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{1\times d_{h}}$ and $\widetilde{\mathbf{q}}_{t,i}=\mathbf{q}_{t,i}\mathbf{T}_{t}$, $\widetilde{\mathbf{k}}_{s,i}=\mathbf{k}_{s,i}\mathbf{T}_{s}$, then $\widetilde{\mathbf{q}}_{t,i}\,\widetilde{\mathbf{k}}_{s,i}^{\top}\;=\;\mathbf{q}_{t,i}\mathbf{T}_{t-s}\mathbf{k}_{s,i}^{\top}$.
 Theorem 3.1 indicates that TPA does not break RoPE’s relative translational property. We prove it in the Appendix D.1.
@@
 For MLA, $d_{h}^{R}$ and $d_{h}$ are the dimensions for RoPE and non-RoPE parts; $d_{c}^{\prime}$ and $d_{c}$ are the dimensions of compressed vectors for query and key-value, respectively.
+The MLA parameter count includes the output projection.
 Method KV Cache # Parameters # Query Heads # KV Heads
-MHA $2hd_{h}$ $4d_{\text{model}}^{2}$ $h$ $h$
-MQA $2d_{h}$ $(2+2/h)d_{\text{model}}^{2}$ $h$ 11
-GQA $2gd_{h}$ $(2+2g/h)d_{\text{model}}^{2}$ $h$ $g$
-MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $d_{\text{model}}d_{h}^{R}+d_{c}(d_{\text{model}}+2hd_{h})$ $h$ $h$
+MHA $2hd_{h}$ $4d_{\text{model}}\,h\,d_{h}$ $h$ $h$
+MQA $2d_{h}$ $2d_{\text{model}}\,d_{h}\,(h+1)$ $h$ 11
+GQA $2Gd_{h}$ $2d_{\text{model}}\,d_{h}\,(h+G)$ $h$ $G$
+MLA $d_{c}+d_{h}^{R}$ $d_{c}^{\prime}(d_{\text{model}}+hd_{h}+hd_{h}^{R})$ $d_{c}(d_{\text{model}}+2hd_{h}) +$$d_{\text{model}}(hd_{h}+d_{h}^{R})$ $h$ $h$
 TPA $(R_{K}+R_{V})(h+d_{h})$ $d_{\text{model}}(R_{Q}+R_{K}+R_{V})(h+d_{h})+d_{\text{model}}\,hd_{h}$ $h$ $h$
@@
 We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention).
-Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (touvron2023llama) while changing the self-attention block to our TPA-based version.
-The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu; touvron2023llama).
+Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (Touvron et al., 2023) while changing the self-attention block to our TPA-based version.
+The feed-forward network (FFN) adopts a SwiGLU layer, as in (Shazeer, 2020; Touvron et al., 2023).
 Rotary Positional Embedding (RoPE).
-As discussed in Section 3.2, RoPE (su2024roformer) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
+As discussed in Section 3.2, RoPE (Su et al., 2024) is applied to the $\mathbf{Q}$ and $\mathbf{K}$.
 Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see Equation (3.5) and Theorem 3.1.
 SwiGLU Feed-Forward Network.
-Following shazeer2020glu; touvron2023llama, our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl[\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr]\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters.
+Following Shazeer (2020); Touvron et al.
+(2023), our T6 uses a SwiGLU-based Feed-Forward Network (FFN): $\operatorname{FFN}(\mathbf{x})=\bigl[\sigma(\mathbf{x}\,\bm{W}_{1})\,\odot\,(\mathbf{x}\,\bm{W}_{2})\bigr]\,\bm{W}_{3}$, where $\sigma$ is the SiLU (a.k.a., swish) nonlinearity, $\odot$ is element-wise product, and $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are learnable parameters.
 Note that other activation functions can also be used.
@@
 $\mathbf{a}^{V}_{\text{cache}}$
-$(M,H)$
+$(H,M)$
 $\odot$
 $\mathbf{O}^{(A)}$
-$(M,H)$
+$(H,M)$
 $\bm{b}^{V}_{\text{cache}}$
@@
 Rectangles represent tensors (blue for inputs, yellow for intermediates, red for final output), circles with $\sum$ or $\odot$ denote Einstein summation contractions or element-wise products respectively, and the green rounded rectangle is the softmax operation.
-Shapes are shown for a single query ($N=1$) interacting with $M$ cached items.
+Shapes are shown for a single query ($N=1$) interacting with $M$ cached items in the common rank-1 setting $R_{K}=R_{V}=1$.
+We use a head-first layout $(H,M)$ for logits and attention weights; the cached head factors $\mathbf{a}^{K}_{\text{cache}}$ and $\mathbf{a}^{V}_{\text{cache}}$ are shown transposed relative to their natural token-major layout for readability.
 $H$ is the number of heads, $R_{Q}$ is the query rank, and $D,E$ are respective feature dimensions for the $\mathbf{B}_{Q}/\bm{b}^{K}_{\text{cache}}$ and $\bm{b}^{V}_{\text{cache}}$ factors.
-Scaling factors in softmax are omitted for visual clarity.
+Scaling factors are omitted for visual clarity.
 This sequence of factorized operations allows FlashTPA Decoding to compute the attention output efficiently.
@@
 6.1 Language Modeling Tasks
-All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy2022), and we pretrain our models using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu).
+All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy, 2022), and we pretrain our models using the FineWeb-Edu 100B dataset (Lozhkov et al., 2024).
 The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation.
-We compare T6 against the baseline Llama architecture (touvron2023llama) with SwiGLU activation (shazeer2020glu) and RoPE embeddings (su2024roformer), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention) with Multi-Query Attention (MQA; shazeer2019fast), Grouped Query Attention (GQA; ainslie2023gqa), or Multi-head Latent Attention (MLA; liu2024deepseek).
+We compare T6 against the baseline Llama architecture (Touvron et al., 2023) with SwiGLU activation (Shazeer, 2020) and RoPE embeddings (Su et al., 2024), as well as Llama variants that replace Multi-Head Attention (MHA; Vaswani et al.
+(2017)) with Multi-Query Attention (MQA; Shazeer (2019)), Grouped Query Attention (GQA; Ainslie et al.
+(2023)), or Multi-head Latent Attention (MLA; Liu et al.
+(2024a)).
 In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer.
 We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B).
-We pretrain all models with 50B tokens using 1 epoch.
+We pretrain all models for 50B tokens (roughly half an epoch over FineWeb-Edu-100B).
 Details on architecture hyperparameters and training hardware are shown in Appendix H.1.
@@
 For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages.
-Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek) (blue curves) generally trains more slowly and yields higher validation losses.
+Meanwhile, Multi-Head Latent Attention (MLA) (Liu et al., 2024a) (blue curves) generally trains more slowly and yields higher validation losses.
 Validation Perplexity.
@@
 Downstream Evaluation.
-We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC), BoolQ (BoolQ), HellaSwag (hellaswag), OBQA (OpenBookQA), PIQA (PIQA), WinoGrande (WinoGrande), and MMLU (MMLU), using the lm-evaluation-harness codebase (eval-harness).
+We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020), and MMLU (Hendrycks et al., 2021), using the lm-evaluation-harness codebase (Gao et al., 2024).
 For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy.
@@
 This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms.
-We benchmark FlashTPA against FlashMHA (shah2024flashattention), FlashGQA, FlashMQA, and FlashMLA (flashmla2025).
-It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton).
+We benchmark FlashTPA against FlashMHA (Shah et al., 2024), FlashGQA, FlashMQA, and FlashMLA (Jiashi Li, 2025).
+It is important to note that our current FlashTPA implementation utilizes Triton (Tillet et al., 2019).
 While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential.
@@
 Acknowledgements
-We thank the anonymous reviewers and area chair for their helpful comments. We acknowledge the compute credits provided by Fetch.ai.
-[…]
+We thank the anonymous reviewers and area chairs for their helpful comments. We acknowledge the compute credits provided by Fetch.ai.
 Appendix A Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$
@@
 We make the batch/time/head/rank/value dimensions explicit and introduce the shorthands $D:=d_{h}$ and $E:=d_{v}$ (typically E=DE{=}D):
-$\mathbf{A}_{Q}\in\mathbb{R}^{B\times T_{q}\times h\times R_{Q}},\quad\mathbf{B}_{Q}\in\mathbb{R}^{B\times T_{q}\times R_{Q}\times D},\qquad\mathbf{A}_{K}\in\mathbb{R}^{B\times T_{k}\times h\times R_{K}},\quad\mathbf{B}_{K}\in\mathbb{R}^{B\times T_{k}\times R_{K}\times D},$
-$\mathbf{A}_{V}\in\mathbb{R}^{B\times T_{k}\times h\times R_{V}},\qquad\mathbf{B}_{V}\in\mathbb{R}^{B\times T_{k}\times R_{V}\times E}.$
+$\mathbf{A}_{Q}\in\mathbb{R}^{B\times T_{q}\times R_{Q}\times H},\quad\mathbf{B}_{Q}\in\mathbb{R}^{B\times T_{q}\times R_{Q}\times D},\qquad\mathbf{A}_{K}\in\mathbb{R}^{B\times T_{k}\times R_{K}\times H},\quad\mathbf{B}_{K}\in\mathbb{R}^{B\times T_{k}\times R_{K}\times D},$
+$\mathbf{A}_{V}\in\mathbb{R}^{B\times T_{k}\times R_{V}\times H},\qquad\mathbf{B}_{V}\in\mathbb{R}^{B\times T_{k}\times R_{V}\times E}.$
 Indices $b,q,k,h,r,s,u,d,e$ denote batch, query position, key position, head, query-rank, key-rank, value-rank, feature ($D$), and value feature ($E$).
 We write $T:=T_{q}\!=\!T_{k}$ for full-sequence attention; in decoding, $T_{q}\!=\!1$ and we denote the cache length by $M\!=\!T_{k}$.
+Convention.
+For a single token, the main text defines $\mathbf{A}_{*}(\mathbf{x}_{t})\in\mathbb{R}^{R_{*}\times H}$ and $\mathbf{B}_{*}(\mathbf{x}_{t})\in\mathbb{R}^{R_{*}\times D}$, with $\mathbf{Q}_{t}=\tfrac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$.
+Accordingly, throughout this appendix we index $\mathbf{A}_{Q}$ as $\mathbf{A}_{Q}[b,q,r,h]$ (rank-major).
+Some implementations may store $\mathbf{A}_{*}$ transposed as $(H\times R_{*})$ for memory layout, this is equivalent, since all uses contract over the rank index.
 High-level idea.
@@
 Single head. For a fixed head $h\!\in\![H]$ and token indices $(q,k)$, using $s_{Q}{=}1/R_{Q}$ and $s_{K}{=}1/R_{K}$ we have
-$\bigl[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}\bigr]_{q,k}\;=\;\frac{1}{R_{Q}R_{K}}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\,a^{(r)}_{q,h}(\mathbf{x}_{q})\,a^{(s)}_{k,h}(\mathbf{x}_{k})\;\big\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{q}),\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{k})\big\rangle,$ (A.1)
-and for values (with $s_{V}{=}1/R_{V}$), $\mathbf{V}^{(h)}_{k}=\frac{1}{R_{V}}\sum_{u=1}^{R_{V}}a^{(u)}_{k,h}(\mathbf{x}_{k})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{k})$.
+$\bigl[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}\bigr]_{q,k}\;=\;\frac{1}{R_{Q}R_{K}}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\,a^{Q,(r)}_{q,h}(\mathbf{x}_{q})\,a^{K,(s)}_{k,h}(\mathbf{x}_{k})\;\big\langle\mathbf{b}^{Q,(r)}(\mathbf{x}_{q}),\,\mathbf{b}^{K,(s)}(\mathbf{x}_{k})\big\rangle,$ (A.1)
+and for values (with $s_{V}{=}1/R_{V}$), $\mathbf{V}^{(h)}_{k}=\frac{1}{R_{V}}\sum_{u=1}^{R_{V}}a^{V,(u)}_{k,h}(\mathbf{x}_{k})\,\mathbf{b}^{V,(u)}(\mathbf{x}_{k})$.
 The per-head attention output at query position $q$ is then $\sum\nolimits_{k}\operatorname{softmax}\!\big(\frac{1}{\sqrt{D}}[\mathbf{Q}^{(h)}(\mathbf{K}^{(h)})^{\top}]_{q,:}\big)_{k}\,\mathbf{V}^{(h)}_{k}$.
@@
 With $\mathcal{S}\!=\!\frac{s_{Q}s_{K}}{\sqrt{D}}$, we compute
-$\mathcal{L}[b,h,q,k]$ $=\mathcal{S}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\mathbf{A}_{Q}[b,q,h,r]\;\mathbf{A}_{K}[b,k,h,s]\;P[b,q,k,r,s],$ (A.2)
-$\alpha[b,h,q,k]$ $=\operatorname{Softmax}_{k}\!\left(\mathcal{L}[b,h,q,k]+\log\mathsf{mask}[q,k]\right),$
-$\mathbf{O}[b,h,q,e]$ $=s_{V}\sum_{k=1}^{T_{k}}\sum_{u=1}^{R_{V}}\alpha[b,h,q,k]\;\mathbf{A}_{V}[b,k,h,u]\;\mathbf{B}_{V}[b,k,u,e].$ (A.3)
-Here $\mathsf{mask}[q,k]\!\in\!\{0,-\infty\}$ enforces causality and padding.
+$\mathcal{L}[b,h,q,k]$ $=\mathcal{S}\!\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}\mathbf{A}_{Q}[b,q,r,h]\;\mathbf{A}_{K}[b,k,s,h]\;P[b,q,k,r,s],$ (A.2)
+$\alpha[b,h,q,k]$ $=\operatorname{Softmax}_{k}\!\left(\mathcal{L}[b,h,q,k]+\mathsf{mask}[b,q,k]\right),$
+$\mathbf{O}[b,h,q,e]$ $=s_{V}\sum_{k=1}^{T_{k}}\sum_{u=1}^{R_{V}}\alpha[b,h,q,k]\;\mathbf{A}_{V}[b,k,u,h]\;\mathbf{B}_{V}[b,k,u,e].$ (A.3)
+Here $\mathsf{mask}[b,q,k]\!\in\!\{0,-\infty\}$ is an additive mask in logit space that enforces causality and padding.
 Eqs.
-(A.2)–(A.3) make explicit that (i) feature-space dot products $P$ are head-shared, and (ii) all $1/R_{(\cdot)}$ scalings can be folded into a single multiplicative factor per phase without affecting numerical stability (softmax shift invariance).
+(A.2)–(A.3) make explicit that (i) feature-space dot products $P$ are head-shared, and (ii) the rank normalizations $s_{Q},s_{K},s_{V}$ can be absorbed into the corresponding factor tensors (or into the scalar prefactors) without changing the computed attention output.
 A.2 Complexity: materialized vs. specialized computation
@@
   • Projection: Projecting $\mathbf{x}_{T}$ to get a query, key, and value vector for each of the $H$ heads costs $\Theta(d_{\text{model}}Hd_{h})$.
-  • Attention: The query vectors interact with a K/V cache of size $T$, costing $\Theta(MHd_{h})$.
-  • Total MHA: The complexity is $\Theta(d_{\text{model}}Hd_{h}+MHd_{h})$.
+  • Attention: Dot products and value aggregation over a cache of length $M$ cost $\Theta(2MHd_{h})$ (ignoring softmax constants).
+  • Total MHA: The complexity is $\Theta(d_{\text{model}}Hd_{h}+2MHd_{h})$.
 Multi-Query Attention (MQA) uses $H$ query heads but shares a single Key/Value head ($H_{kv}=1$). The arithmetic complexity remains the same as MHA for the same number of query heads.
@@
   • Attention: The interaction with the cache costs $\Theta(2MHd_{h})$.
-  • Total MQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2)+MHd_{h})$.
-Grouped-Query Attention (GQA) uses $H$ query heads and $N_{g}$ Key/Value head groups ($H_{kv}=N_{g}$). The arithmetic complexity is also identical to MHA.
-  • Projection: Projecting for $H$ query heads and $N_{g}$ K/V head groups costs $\Theta(d_{\text{model}}(Hd_{h}+2N_{g}d_{h}))$.
+  • Total MQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2)+2MHd_{h})$.
+Grouped-Query Attention (GQA) uses $H$ query heads and $G$ Key/Value head groups ($H_{kv}=G$). The arithmetic complexity is also identical to MHA.
+  • Projection: Projecting for $H$ query heads and $G$ K/V head groups costs $\Theta(d_{\text{model}}(Hd_{h}+2Gd_{h}))$.
   • Attention: The interaction with the cache costs $\Theta(2MHd_{h})$.
-  • Total GQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2N_{g})+MHd_{h})$.
+  • Total GQA: The complexity is $\Theta(d_{\text{model}}d_{h}(H+2G)+2MHd_{h})$.
 MQA and GQA significantly reduce the KV cache size and memory bandwidth compared to MHA.
@@
 Multi-Head Latent Attention (MLA), as described in Appendix F.3, uses $H$ heads.
-For each head, the key $\mathbf{K}_{i}$ has dimension $d_{k}^{\prime}=d_{c}+d_{h}^{R}$, and the value $\mathbf{V}_{i}^{C}$ has dimension $d_{v}^{\prime}=d_{c}$.
-  • Projection: Projecting $\mathbf{x}_{M}$ to get compressed and RoPE components for Q, K, and V costs $\Theta(d_{\text{model}}((d_{c}+d_{h}^{R})\cdot h+d_{c}+d_{h}^{R}))$.
-  • Attention: The query vectors interact with a K/V cache of size $M$, costing $\Theta(MH(2d_{c}+d_{h}^{R}))$.
-  • Total MLA: The complexity is $\Theta(d_{\text{model}}(d^{\prime}_{c}+2d_{c}+2d_{h}^{R})+MH(2d_{c}+d_{h}^{R}))$.
+Each head’s (up-projected) query/key vectors have dimension $d_{h}+d_{h}^{R}$.
+During decoding, however, the score computation against the cache can be decomposed into (i) a dot product in the cached latent space $\mathbb{R}^{d_{c}}$ for the content part and (ii) an additional RoPE dot product in $\mathbb{R}^{d_{h}^{R}}$ for the positional part.
+Concretely, MLA caches $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ per past token $s$, aggregates values in $\mathbb{R}^{d_{c}}$, and then up-projects once per step.
+  • Cached state: MLA caches the compressed KV latent $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ and the shared RoPE key component $\mathbf{k}_{s}^{R}\in\mathbb{R}^{d_{h}^{R}}$ per past token $s$.
+  • Projection (current token): Computing the query latents and the new cache entry (up to constant factors) costs
+    $\Theta\bigl(d_{\text{model}}d_{c}^{\prime}\;+\;d_{c}^{\prime}H(d_{h}+d_{h}^{R})\;+\;d_{\text{model}}(d_{c}+d_{h}^{R})\bigr),$
+    corresponding to forming $\mathbf{c}^{Q}$, $\mathbf{Q}^{C}$, $\mathbf{Q}^{R}$, and computing/storing $\mathbf{c}^{KV}$ and $\mathbf{k}^{R}$ for the current token.
+  • Attention (cache interaction): Using the identity $\mathbf{q}_{t,i}^{C\top}\mathbf{k}_{s,i}^{C}=(\bm{W}^{UK}_{i}\mathbf{q}_{t,i}^{C})^{\top}\mathbf{c}_{s}^{KV}$, the score against each cached token can be computed via a dot product in $\mathbb{R}^{d_{c}}$ plus the RoPE dot product in $\mathbb{R}^{d_{h}^{R}}$.
+The latent value can be aggregated in $\mathbb{R}^{d_{c}}$ and then up-projected once.
+The dominant cache-dependent cost is
+    $\Theta\bigl(MH(2d_{c}+d_{h}^{R})\bigr),$
+    up to lower-order per-step terms such as $\Theta(Hd_{c}d_{h})$.
+  • Total MLA: $\Theta\bigl(d_{\text{model}}d_{c}^{\prime}+d_{c}^{\prime}H(d_{h}+d_{h}^{R})+d_{\text{model}}(d_{c}+d_{h}^{R})+MH(2d_{c}+d_{h}^{R})\bigr)$.
 TPA.
-We use the FlashTPA Decoding algorithm (Algorithm 2) for FLOPs analysis, with $N=1$ query token, $T$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$).
+We use the FlashTPA Decoding algorithm (Algorithm 2) for FLOPs analysis, with $N=1$ query token, $M$ cached items, $D$ as feature dimension for $\mathbf{B}_{Q}/\bm{b}^{K}$ (typically $d_{h}$), and $E$ for $\bm{b}^{V}$ (typically $d_{h}$).
 For ranks $(R_{Q},R_{K},R_{V})$:
-  • Projection: Projecting $\mathbf{x}_{M}$ to all Q, K, V factors costs $\Theta(d_{\text{model}}((R_{Q}+R_{K}+R_{V})(H+d_{h})))$.
-  • Attention: The decoding algorithm’s interaction with the cache costs $\Theta(M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
-  • Total for TPA decoding: $\Theta(d_{\text{model}}(R_{Q}+R_{K}+R_{V})(H+d_{h})+M[R_{K}(R_{Q}D+HR_{Q}+H)+R_{V}H(1+E)])$.
+  • Projection: Projecting the current token $\mathbf{x}_{T}$ to all Q/K/V factors costs $\Theta\bigl(d_{\text{model}}(R_{Q}{+}R_{K}{+}R_{V})(H{+}d_{h})\bigr)$.
+  • Attention (cache interaction): Using Algorithm 2 with cache length $M$, the dominant cache-dependent FLOPs are
+    $\Theta\bigl(M\,(R_{Q}R_{K}D\;+\;HR_{Q}R_{K}\;+\;HR_{V}E)\bigr),$
+    up to lower-order terms (masking/element-wise products and online-softmax bookkeeping).
+  • Total for TPA decoding: $\Theta\bigl(d_{\text{model}}(R_{Q}{+}R_{K}{+}R_{V})(H{+}d_{h})\;+\;M(R_{Q}R_{K}D+HR_{Q}R_{K}+HR_{V}E)\bigr)$.
 Example Comparison I.
@@
   • Head Dimension: $d_{h}=64$ (so $D=E=d_{h}$)
-  • GQA Groups: $N_{g}=4$
+  • GQA Groups: $G=4$
   • MLA Dimensions: $d_{c}=256$, $d_{h}^{R}=32$, and $d^{\prime}_{c}=768$
@@
 $\text{Attention}=2\cdot M\cdot H\cdot d_{h}=4096M$
-GQA ($N_{g}=4$, 9.4M parameters):
-$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=2048\cdot 64\cdot(2\cdot 32+2\cdot 4)\approx 9.4\times 10^{6}$
-$\text{Projection}=d_{\text{model}}(H+2N_{g})d_{h}=2048\cdot(32+8)\cdot 64\approx 5.2\times 10^{6}$
+GQA ($G=4$, 9.4M parameters):
+$\text{Parameters}=d_{\text{model}}d_{h}(2H+2G)=2048\cdot 64\cdot(2\cdot 32+2\cdot 4)\approx 9.4\times 10^{6}$
+$\text{Projection}=d_{\text{model}}(H+2G)d_{h}=2048\cdot(32+8)\cdot 64\approx 5.2\times 10^{6}$
 $\text{Attention}=2\cdot M\cdot H\cdot d_{h}=4096M$
@@
 $\text{Parameters}=768(2048+2048+1024)+2048(32+2048)+256(2048+4096)\approx 9.8\times 10^{6}$
-$\text{Projection}=d_{\text{model}}((d_{c}+d_{h}^{R})\cdot h+d_{c}+d_{h}^{R})=2048\cdot((256+32)\cdot 32+256+32)\approx 19.5\times 10^{6}$
+$\text{Projection}\approx d_{\text{model}}d_{c}^{\prime}+d_{c}^{\prime}H(d_{h}+d_{h}^{R})+d_{\text{model}}(d_{c}+d_{h}^{R})+Hd_{c}d_{h}$
+$\hskip 42.19432pt=2048\cdot 768+768\cdot 32\cdot(64+32)+2048\cdot(256+32)+32\cdot 256\cdot 64$
+$\hskip 42.19432pt\approx 5.0\times 10^{6}$
 $\text{Attention}=M\cdot H\cdot(2d_{c}+d_{h}^{R})=M\cdot 32\cdot(512+32)=17408M$
@@
 $\text{Projection}=d_{\text{model}}(16+1+1)(H+d_{h})=2048\cdot(18)\cdot(96)\approx 3.5\times 10^{6}$
-$\text{Attention}=M\cdot[1(1568)+1(2080)]=3648M$
+$\text{Attention}=M\cdot[1(1536)+1(2048)]=3584M$
 TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 8.1M parameters):
@@
 $\text{Projection}=d_{\text{model}}(16+2+2)(H+d_{h})=2048\cdot(20)\cdot(96)\approx 3.9\times 10^{6}$
-$\text{Attention}=M\cdot[2(1568)+2(2080)]=7296M$
+$\text{Attention}=M\cdot[2(1536)+2(2048)]=7168M$
 TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 6.2M parameters):
@@
 $\text{Projection}=d_{\text{model}}(8+1+1)(H+d_{h})=2048\cdot(10)\cdot(96)\approx 2.0\times 10^{6}$
-$\text{Attention}=M\cdot[1(800)+1(2080)]=2880M$
+$\text{Attention}=M\cdot[1(768)+1(2048)]=2816M$
 TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 6.6M parameters):
@@
 $\text{Projection}=d_{\text{model}}(8+2+2)(H+d_{h})=2048\cdot(12)\cdot(96)\approx 2.4\times 10^{6}$
-$\text{Attention}=M\cdot[2(800)+2(2080)]=5760M$
+$\text{Attention}=M\cdot[2(768)+2(2048)]=5632M$
 The analysis shows that TPA with low ranks offers a favorable trade-off.
@@
   • Head Dimension: $d_{h}=128$ (so $D=E=d_{h}$)
-  • GQA Groups: $N_{g}=4$
+  • GQA Groups: $G=4$
   • MLA Dimensions: $d_{c}=512$, $d_{h}^{R}=64$, and $d^{\prime}_{c}=1536$
@@
 $\text{Attention}=2\cdot M\cdot 32\cdot 128=8192M$
-GQA ($N_{g}=4$, 37.7M parameters):
-$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=4096\cdot 128\cdot(2\cdot 32+2\cdot 4)\approx 37.7\times 10^{6}$
+GQA ($G=4$, 37.7M parameters):
+$\text{Parameters}=d_{\text{model}}d_{h}(2H+2G)=4096\cdot 128\cdot(2\cdot 32+2\cdot 4)\approx 37.7\times 10^{6}$
 $\text{Projection}=4096\cdot(32+8)\cdot 128\approx 21.0\times 10^{6}$
@@
 $\text{Parameters}=1536(4096+4096+2048)+4096(64+4096)+512(4096+8192)\approx 39.1\times 10^{6}$
-$\text{Projection}=4096\cdot((512+64)\cdot 32+512+64)\approx 77.9\times 10^{6}$
+$\text{Projection}\approx d_{\text{model}}d_{c}^{\prime}+d_{c}^{\prime}H(d_{h}+d_{h}^{R})+d_{\text{model}}(d_{c}+d_{h}^{R})+Hd_{c}d_{h}$
+$\hskip 42.19432pt=4096\cdot 1536+1536\cdot 32\cdot(128+64)+4096\cdot(512+64)+32\cdot 512\cdot 128$
+$\hskip 42.19432pt\approx 20.2\times 10^{6}$
 $\text{Attention}=M\cdot 32\cdot(1024+64)=34816M$
@@
 $\text{Projection}=4096\cdot(16+1+1)\cdot(32+128)\approx 11.8\times 10^{6}$
-$\text{Attention}=M\cdot[1(2592)+1(4128)]=6720M$
+$\text{Attention}=M\cdot[1(2560)+1(4096)]=6656M$
 TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 29.9M parameters):
@@
 $\text{Projection}=4096\cdot(16+2+2)\cdot(32+128)\approx 13.1\times 10^{6}$
-$\text{Attention}=M\cdot[2(2592)+2(4128)]=13440M$
+$\text{Attention}=M\cdot[2(2560)+2(4096)]=13312M$
 TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 23.3M parameters):
@@
 $\text{Projection}=4096\cdot(8+1+1)\cdot(32+128)\approx 6.6\times 10^{6}$
-$\text{Attention}=M\cdot[1(1312)+1(4128)]=5440M$
+$\text{Attention}=M\cdot[1(1280)+1(4096)]=5376M$
 TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 24.6M parameters):
@@
 $\text{Projection}=4096\cdot(8+2+2)\cdot(32+128)\approx 7.9\times 10^{6}$
-$\text{Attention}=M\cdot[2(1312)+2(4128)]=10880M$
+$\text{Attention}=M\cdot[2(1280)+2(4096)]=10752M$
 For this larger configuration, TPA ($R_{Q}{=}8,R_{K}{=}1,R_{V}{=}1$) remains the clear leader in computational efficiency, with the lowest projection and attention costs.
@@
   • Head Dimension: $d_{h}=128$ (so $D=E=d_{h}$)
-  • GQA Groups: $N_{g}=8$
+  • GQA Groups: $G=8$
   • MLA Dimensions: $d_{c}=512$, $d_{h}^{R}=64$, and $d^{\prime}_{c}=1536$
@@
 $\text{Attention}=2\cdot M\cdot 64\cdot 128=16384M$
-GQA ($N_{g}=8$, 132M parameters):
-$\text{Parameters}=d_{\text{model}}d_{h}(2H+2N_{g})=7168\cdot 128\cdot(2\cdot 64+2\cdot 8)\approx 132\times 10^{6}$
+GQA ($G=8$, 132M parameters):
+$\text{Parameters}=d_{\text{model}}d_{h}(2H+2G)=7168\cdot 128\cdot(2\cdot 64+2\cdot 8)\approx 132\times 10^{6}$
 $\text{Projection}=7168\cdot(64+16)\cdot 128\approx 73.4\times 10^{6}$
@@
 $\text{Parameters}=1536(7168+8192+4096)+7168(64+8192)+512(7168+16384)\approx 101\times 10^{6}$
-$\text{Projection}=7168\cdot((512+64)\cdot 64+512+64)=268.4\times 10^{6}$
+$\text{Projection}\approx d_{\text{model}}d_{c}^{\prime}+d_{c}^{\prime}H(d_{h}+d_{h}^{R})+d_{\text{model}}(d_{c}+d_{h}^{R})+Hd_{c}d_{h}$
+$\hskip 42.19432pt=7168\cdot 1536+1536\cdot 64\cdot(128+64)+7168\cdot(512+64)+64\cdot 512\cdot 128$
+$\hskip 42.19432pt\approx 38.2\times 10^{6}$
 $\text{Attention}=M\cdot 64\cdot(2\cdot 512+64)=69632M$
@@
 $\text{Projection}=7168\cdot(16+1+1)\cdot(64+128)\approx 24.8\times 10^{6}$
-$\text{Attention}=M\cdot[1(3136)+1(8256)]=11392M$
+$\text{Attention}=M\cdot[1(3072)+1(8192)]=11264M$
 TPA ($R_{Q}=16,R_{K}=2,R_{V}=2$, 86.2M parameters):
@@
 $\text{Projection}=7168\cdot(16+2+2)\cdot(64+128)\approx 27.5\times 10^{6}$
-$\text{Attention}=M\cdot[2(3136)+2(8256)]=22784M$
+$\text{Attention}=M\cdot[2(3072)+2(8192)]=22528M$
 TPA ($R_{Q}=8,R_{K}=1,R_{V}=1$, 72.5M parameters):
@@
 $\text{Projection}=7168\cdot(8+1+1)\cdot(64+128)\approx 13.8\times 10^{6}$
-$\text{Attention}=M\cdot[1(1600)+1(8256)]=9856M$
+$\text{Attention}=M\cdot[1(1536)+1(8192)]=9728M$
 TPA ($R_{Q}=8,R_{K}=2,R_{V}=2$, 75.2M parameters):
@@
 $\text{Projection}=7168\cdot(8+2+2)\cdot(64+128)\approx 16.5\times 10^{6}$
-$\text{Attention}=M\cdot[2(1600)+2(8256)]=19712M$
+$\text{Attention}=M\cdot[2(1536)+2(8192)]=19456M$
 At this very large scale, the cost of MHA projections becomes prohibitive.
@@
 We allow $N$ query positions but decoding uses N=1N{=}1. Let $B$ be batch, $M$ the cache length, $H$ heads, $R_{Q},R_{K},R_{V}$ ranks, and $D,E$ feature sizes (typically $D{=}E{=}d_{h}$). Inputs:
-$\mathbf{A}_{Q}\!\in\!\mathbb{R}^{B\times N\times H\times R_{Q}},\;\mathbf{B}_{Q}\!\in\!\mathbb{R}^{B\times N\times R_{Q}\times D},\;\mathbf{A}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times H\times R_{K}},\;\mathbf{B}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{K}\times D},$
-$\mathbf{A}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times H\times R_{V}},\;\mathbf{B}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{V}\times E}.$
+$\mathbf{A}_{Q}\!\in\!\mathbb{R}^{B\times N\times R_{Q}\times H},\;\mathbf{B}_{Q}\!\in\!\mathbb{R}^{B\times N\times R_{Q}\times D},\;\mathbf{A}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{K}\times H},\;\mathbf{B}_{K}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{K}\times D},$
+$\mathbf{A}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{V}\times H},\;\mathbf{B}_{V}^{\text{cache}}\!\in\!\mathbb{R}^{B\times M\times R_{V}\times E}.$
 We use scalings $s_{Q}{=}1/R_{Q}$, $s_{K}{=}1/R_{K}$, $s_{V}{=}1/R_{V}$, and $s_{\text{total}}{=}1/\sqrt{D}$.
 Let $\mathsf{mask}\!\in\!\{0,-\infty\}^{B\times N\times M}$ encode causality/padding.
-If RoPE pre‑rotation is used (Section 3.2), $\mathbf{B}_{K}^{\text{cache}}$ already includes positional phases; otherwise apply $\operatorname{RoPE}$ to $\mathbf{B}_{K}$ on load (no other change).
+If RoPE pre‑rotation is used (Section 3.2), $\mathbf{B}_{K}^{\text{cache}}$ already includes positional phases; otherwise apply $\operatorname{RoPE}$ to $\mathbf{B}_{K}$ on load.
 Algorithm 2 FlashTPA Decoding (general $R_{K},R_{V}$, masked, online‑LSE)
@@
 […]
-3:Initialize $\mathbf{y}\leftarrow 0^{B\times H\times N\times E}$, $\mathbf{lse}\leftarrow 0^{B\times H\times N}$, $\mathbf{m}\leftarrow(-\infty)^{B\times H\times N}$
+3:Initialize $\mathbf{y}\leftarrow 0^{B\times H\times N\times E}$, $\mathbf{s}\leftarrow 0^{B\times H\times N}$, $\mathbf{m}\leftarrow(-\infty)^{B\times H\times N}$ $\triangleright$ $\mathbf{s}$ accumulates $\sum\exp(\cdot)$; log-sum-exp is $\log\mathbf{s}+\mathbf{m}$
 4:for each cache block $m{:}m{+}\Delta m\!\leq\!M$ do
-5:  Load $\mathbf{B}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{K}\times D}$, $\mathbf{A}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times H\times R_{K}}$
-6:  Load $\mathbf{A}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times H\times R_{V}}$, $\mathbf{B}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{V}\times E}$, $\mathsf{mask}_{\text{blk}}\!\in\!\mathbb{R}^{B\times N\times\Delta m}$
+5:  Load $\mathbf{B}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{K}\times D}$, $\mathbf{A}_{K,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{K}\times H}$
+6:  Load $\mathbf{A}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{V}\times H}$, $\mathbf{B}_{V,\text{blk}}\!\in\!\mathbb{R}^{B\times\Delta m\times R_{V}\times E}$, $\mathsf{mask}_{\text{blk}}\!\in\!\mathbb{R}^{B\times N\times\Delta m}$
 7:  (1) Head‑shared feature dots:   $\mathbf{P}\leftarrow\mathrm{einsum}(\text{``bnrd,bmsd$\to$bnmrs''},~\mathbf{B}_{Q},~\mathbf{B}_{K,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times N\times\Delta m\times R_{Q}\times R_{K}}$
 8:  (2) Per‑head rank mixing to logits:
-9:  $\mathcal{L}_{\text{blk}}\leftarrow(s_{\text{total}}s_{Q}s_{K})\cdot\mathrm{einsum}(\text{``bnhr,bmhs,bnmrs$\to$bhnm''},~\mathbf{A}_{Q},~\mathbf{A}_{K,\text{blk}},~\mathbf{P})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times\Delta m}$
+9:  $\mathcal{L}_{\text{blk}}\leftarrow(s_{\text{total}}s_{Q}s_{K})\cdot\mathrm{einsum}(\text{``bnrh,bmsh,bnmrs$\to$bhnm''},~\mathbf{A}_{Q},~\mathbf{A}_{K,\text{blk}},~\mathbf{P})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times\Delta m}$
 10:  $\mathcal{L}_{\text{blk}}\leftarrow\mathcal{L}_{\text{blk}}+\mathrm{broadcast}(\mathsf{mask}_{\text{blk}})$
@@
 12:  $\mathbf{m}_{\text{blk}}\leftarrow\max_{m}(\mathcal{L}_{\text{blk}})$; $\mathbf{p}_{\text{blk}}\leftarrow\exp(\mathcal{L}_{\text{blk}}-\mathbf{m}_{\text{blk}})$; $\mathbf{s}_{\text{blk}}\leftarrow\sum_{m}\mathbf{p}_{\text{blk}}$
-13:  (4) Block value mixing:   $\mathbf{V}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bmhu,bmue$\to$bhme''},~\mathbf{A}_{V,\text{blk}},~\mathbf{B}_{V,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times\Delta m\times E}$
-14:  $\mathbf{y}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bhnm,bhme$\to$bhne''},~\mathbf{p}_{\text{blk}},~\mathbf{V}_{\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times E}$
+13:  (4) Block value aggregation (fused over $m,u$):
+14:  $\mathbf{y}_{\text{blk}}\leftarrow\mathrm{einsum}(\text{``bhnm,bmuh,bmue$\to$bhne''},~\mathbf{p}_{\text{blk}},~\mathbf{A}_{V,\text{blk}},~\mathbf{B}_{V,\text{blk}})$ $\triangleright$ $\mathbb{R}^{B\times H\times N\times E}$
 15:  (5) Fuse blocks with log‑sum‑exp:
 16:  $\mathbf{m}_{\text{new}}\leftarrow\max(\mathbf{m},\mathbf{m}_{\text{blk}})$;   $\mathbf{y}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})[...,None]\odot\mathbf{y}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})[...,None]\odot\mathbf{y}_{\text{blk}}$
-17:  $\mathbf{lse}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})\odot\mathbf{lse}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})\odot\mathbf{s}_{\text{blk}}$;  $\mathbf{m}\leftarrow\mathbf{m}_{\text{new}}$
+17:  $\mathbf{s}\leftarrow\exp(\mathbf{m}-\mathbf{m}_{\text{new}})\odot\mathbf{s}~+~\exp(\mathbf{m}_{\text{blk}}-\mathbf{m}_{\text{new}})\odot\mathbf{s}_{\text{blk}}$;  $\mathbf{m}\leftarrow\mathbf{m}_{\text{new}}$
 18:end for
-19:return $\mathbf{O}\leftarrow s_{V}\cdot\frac{\mathbf{y}}{\mathbf{lse}[...,None]}$ permuted to $(B,N,H,E)$
+19:return $\mathbf{O}\leftarrow s_{V}\cdot\frac{\mathbf{y}}{\mathbf{s}[...,None]}$ permuted to $(B,N,H,E)$
 Step (1)–(2) implements Eq.
@@
 B.1 Triton FlashTPA Decoding Kernel
-We implement the experiments using Triton [tillet2019triton]; Algorithm 3 sketches the kernel corresponding to Algorithm 2.
+We implement the experiments using Triton [57]; Algorithm 3 sketches the kernel corresponding to Algorithm 2.
 The provided kernel outline specializes to the frequently used case $R_{K}{=}R_{V}{=}1$; general ranks follow by tiling over $R_{K},R_{V}$ and replacing the rank‑1 vector–matrix products with the corresponding small GEMMs in steps $S1/S2$ and the value mixing path.
 Algorithm 3 Triton FlashTPA Decoding Kernel
-1:Input Tensors: $\mathbf{A}_{Q}(B,N,H,R_{Q})$, $\mathbf{a}^{K}(B,M,H)$, $\mathbf{a}^{V}(B,M,H)$, $\mathbf{B}_{Q}(B,N,R_{Q},D)$, $\bm{b}^{K}(B,M,D)$, $\bm{b}^{V}(B,M,E)$
+1:Input Tensors: $\mathbf{A}_{Q}(B,N,R_{Q},H)$, $\mathbf{a}^{K}(B,M,H)$, $\mathbf{a}^{V}(B,M,H)$, $\mathbf{B}_{Q}(B,N,R_{Q},D)$, $\bm{b}^{K}(B,M,D)$, $\bm{b}^{V}(B,M,E)$
 2:Scaling factors: $s_{\text{total}},s_{Q},s_{K},s_{V}$; Dimensions: $B,N(=1),M,H,R_{Q},D,E$
@@
 […]
-15:Load $\mathbf{A}_{Q,\text{local}}^{(R_{Q}\times B_{H})}$ from $\mathbf{A}_{Q}[b,0,h_{\text{start}}\dots,:]$
+15:Load $\mathbf{A}_{Q,\text{local}}^{(R_{Q}\times B_{H})}$ from $\mathbf{A}_{Q}[b,0,:,h_{\text{start}}\dots]$
 16:Load $\mathbf{B}_{Q,\text{local}}^{(D\times R_{Q})}$ from $\mathbf{B}_{Q}[b,0,:,:]$ $\triangleright$ Dimensions may be transposed after loading for matmul
@@
   • MHA (orange line) is consistently the slowest mechanism, with its decoding time increasing most rapidly with sequence length.
-  • MQA (green line) and GQA (red line) offer significant speedups over MHA and perform very similarly to each other, often overlapping in the plots.
-  • MLA (purple line) demonstrates strong performance, generally being faster than MQA/GQA, particularly at longer sequence lengths.
+  • MQA (purple line) and GQA (green line) offer significant speedups over MHA and perform very similarly to each other, often overlapping in the plots.
+  • MLA (blue line) demonstrates strong performance, generally being faster than GQA, particularly at longer sequence lengths.
   • FlashTPA shows excellent scalability.
 While at very short sequence lengths (e.g., $2^{12}$ to $2^{13}$), its performance is comparable to MQA/GQA and MLA, its decoding time increases at a notably slower rate with sequence length.
-Consequently, FlashTPA becomes significantly faster than MQA/GQA for sequences longer than approximately $2^{14}$.
+Consequently, FlashTPA becomes significantly faster than GQA for sequences longer than approximately $2^{14}$.
   • Compared to MLA, FlashTPA is consistently among the top two performers.
@@
 Figure 7 (Embedding Dimension 3072): With a larger embedding dimension of 3072, the relative performance trends observed in Figure 5 largely persist.
-  • FlashTPA (blue line) remains one of the most efficient decoding methods.
-MHA (orange line) is consistently the slowest, while MQA (green line) and GQA (red line) offer considerable improvements over MHA.
-  • MLA (purple line) and FlashTPA are the top two performers.
+  • FlashTPA (red line) remains one of the most efficient decoding methods.
+MHA (orange line) is consistently the slowest, while MQA (purple line) and GQA (green line) offer considerable improvements over MHA.
+  • MLA (blue line) and FlashTPA are the top two performers.
 FlashTPA consistently matches or exceeds the speed of MLA, particularly at longer sequence lengths (e.g., beyond $2^{15}$ or $2^{16}$ depending on the batch size).
@@
 Figure 8 (Embedding Dimension 1024): For a smaller embedding dimension of 1024, similar trends are observed:
-  • FlashTPA (blue line) is highly competitive. MHA (orange line) remains the least performant. MQA (green line) and GQA (red line) are faster than MHA.
-  • At very short sequence lengths (around $2^{12}$ to $2^{13}$), MQA/GQA can be slightly faster than or comparable to TPA and MLA, especially for smaller batch sizes (e.g., Batch Size 1).
-  • However, as sequence length increases, both MLA (purple line) and FlashTPA demonstrate superior scalability.
+  • FlashTPA (red line) is highly competitive. MHA (orange line) remains the least performant. MQA (purple line) and GQA (green line) are faster than MHA.
+  • However, as sequence length increases, both MLA (blue line) and FlashTPA demonstrate superior scalability.
 FlashTPA generally matches or outperforms MLA, particularly for sequences longer than $2^{15}$.
@@
 thereby obtaining
-$\operatorname{RoPE}(\mathbf{Q}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}).$
+$\operatorname{RoPE}_{t}(\mathbf{Q}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}).$
 Similarly, for the key tensor $\mathbf{K}_{s}$, we have
@@
 and thus
-$\operatorname{RoPE}(\mathbf{K}_{s})=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{s}).$
+$\operatorname{RoPE}_{s}(\mathbf{K}_{s})=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{s})^{\top}\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{s}).$
 Now, consider the product of the rotated queries and keys:
@@
 Focusing on individual heads $i$, the above matrix equality implies:
-$\operatorname{RoPE}_{t-s}(\mathbf{q}_{t,i})^{\top}\mathbf{k}_{s,i}=\widetilde{\mathbf{q}}_{t,i}^{\top}\widetilde{\mathbf{k}}_{s,i},$
+$(\mathbf{q}_{t,i}\mathbf{T}_{t-s})\,\mathbf{k}_{s,i}^{\top}=(\mathbf{q}_{t,i}\mathbf{T}_{t})\,(\mathbf{k}_{s,i}\mathbf{T}_{s})^{\top},$
 where
-$\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}},\quad\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
+$\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}_{t}(\mathbf{q}_{t,i})=\mathbf{q}_{t,i}\mathbf{T}_{t}\in\mathbb{R}^{1\times d_{h}},\quad\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}_{s}(\mathbf{k}_{s,i})=\mathbf{k}_{s,i}\mathbf{T}_{s}\in\mathbb{R}^{1\times d_{h}}.$
 This equality confirms that the relative positional encoding between queries and keys is preserved under TPA’s factorization and RoPE’s rotation.
@@
 Transformers and Attention.
-As a sequence-to-sequence architecture, Transformer [vaswani2017attention] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies.
-Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [child2019generating, DBLP:conf/iclr/Shi0LR0LJ23, DBLP:conf/iclr/HanJKMWZ24, liang2024conv, li2024tighter, liang2024beyond], kernel-based projections [choromanski2020rethinking], and linearized transformers [tsai2019transformer, katharopoulos2020transformers, schlag2021linear, zhang2023trained, sun2023retentive, DBLP:conf/iclr/ZhangBKR24].
-To decrease memory usage and circumvent the limitation of memory bandwidth in training, [shazeer2019fast] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head.
-To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [ainslie2023gqa] divides queries into several groups, and each group of queries shares a single key head and value head.
-Recently, DeepSeek-V2 [liu2024deepseek] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value.
-Concurrently, [hu2024multi] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q.
+As a sequence-to-sequence architecture, Transformer [60] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies.
+Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [10, 49, 16, 30, 27, 31], kernel-based projections [11], and linearized transformers [59, 25, 44, 69, 54, 67].
+To decrease memory usage and circumvent the limitation of memory bandwidth in training, [46] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head.
+To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [2] divides queries into several groups, and each group of queries shares a single key head and value head.
+Recently, DeepSeek-V2 [32] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value.
+Concurrently, [21] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q.
 Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
 KV Cache Optimization.
-During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [ott2019fairseq].
-This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [adnan2024keyformer].
+During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [40].
+This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [1].
 Consequently, various studies have explored methods to mitigate these issues.
-These include KV cache eviction strategies that discard less significant tokens [zhang2023h2o, xiao2023efficient, cai2024pyramidkv, adnan2024keyformer], dynamic sparse attention mechanisms focusing on selected keys and values [ribar2023sparq, tang2024quest, singhania2024loki], offloading the KV cache to CPU memory [he2024fastdecode, lee2024infinigen, sun2024shadowkv], and quantizing the KV cache [xiao2023smoothquant, liu2024kivi, hooper2024kvquant].
+These include KV cache eviction strategies that discard less significant tokens [70, 62, 8, 1], dynamic sparse attention mechanisms focusing on selected keys and values [42, 55, 50], offloading the KV cache to CPU memory [17, 26, 53], and quantizing the KV cache [61, 34, 19].
 In contrast to these approaches, TPA focuses on reducing the intrinsic size of the KV cache by employing tensor-decomposed key and value representations.
@@
 Low-rank approximations are widely used to compress model parameters and reduce computational complexity.
-Notable examples include LoRA [hu2021lora], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [lialin2023relora], MoRA [jiang2024mora]), long-context training (LongLoRA [chen2023longlora], SinkLoRA [zhang2024sinklora]), and continual training (InfLoRA [liang2024inflora], GS-LoRA [zhao2024continual], I-LoRA [ren2024analyzing]).
+Notable examples include LoRA [20], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [28], MoRA [22]), long-context training (LongLoRA [9], SinkLoRA [66]), and continual training (InfLoRA [29], GS-LoRA [71], I-LoRA [41]).
 These methods generally produce static low-rank expansions that are independent of the input context.
-Theoretical justifications for the expressiveness of low-rank approximations have been provided by [malladi2023kernel, zeng2023expressive].
-Initialization strategies for these factorization matrices have also been explored: OLoRA [buyukakyuz2024olora] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [shi2024loldu] employs LDU-decomposition to accelerate LoRA training.
-Furthermore, AdaLoRA [zhang2023adalora] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks.
+Theoretical justifications for the expressiveness of low-rank approximations have been provided by [38, 65].
+Initialization strategies for these factorization matrices have also been explored: OLoRA [7] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [48] employs LDU-decomposition to accelerate LoRA training.
+Furthermore, AdaLoRA [68] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks.
 TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.
@@
 F.1 Multi-Query Attention (MQA)
-Multi-Query Attention (MQA) [shazeer2019fast] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection.
+Multi-Query Attention (MQA) [46] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection.
 Given a sequence of input embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$, the query, shared key, and shared value tensors are computed as:
@@
 F.2 Grouped Query Attention (GQA)
-Grouped Query Attention (GQA) [ainslie2023gqa] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups.
+Grouped Query Attention (GQA) [2] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups.
 Within each group, heads share a common key and value projection, while each head maintains its own unique query projection.
@@
 F.3 Multi-head Latent Attention (MLA)
-Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [liu2024deepseek] and DeepSeek-V3 [liu2024deepseekv3], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
+Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [32] and DeepSeek-V3 [33], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
 $\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},$
@@
 The RoPE transformation is applied to a separate key embedding $\mathbf{K}^{R}\in\mathbb{R}^{T\times d_{h}^{R}}$.
-Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached, reducing KV memory usage while largely preserving performance compared to standard MHA [vaswani2017attention].
+Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached, reducing KV memory usage while largely preserving performance compared to standard MHA [60].
 MLA also compresses the queries, lowering their training-time memory footprint:
@@
 The term $\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding.
-However, as noted by [kexuefm-10091], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations.
+However, as noted by [52], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations.
 RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section F.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation F.4).
@@
 F.4 Multi-matrix Factorization Attention (MFA)
-[hu2024multi] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
+[21] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
 $\mathbf{Q}_{i}=\mathbf{X}\bm{W}^{DQ}\bm{W}^{UQ}_{i},\quad\mathbf{K}_{\text{shared}}=\mathbf{X}\bm{W}^{K}_{\text{shared}},\quad\mathbf{V}_{\text{shared}}=\mathbf{X}\bm{W}^{V}_{\text{shared}},$
@@
 F.5 Rotary Position Embedding (RoPE)
-Many recent LLMs use rotary position embedding (RoPE; su2024roformer) to encode positional information in the query/key vectors.
+Many recent LLMs use rotary position embedding (RoPE; 51) to encode positional information in the query/key vectors.
 Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head).
@@
 Parameter Initialization for TPA Factors.
-We initialize the weight matrices for TPA factors, such as $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, and $\bm{W}_{r}^{b^{V}}$ (or their combined forms $\bm{W}^{a^{Q}}$, $\bm{W}^{b^{Q}}$, etc.), using Xavier initialization [glorot2010understanding].
+We initialize the weight matrices for TPA factors, such as $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, and $\bm{W}_{r}^{b^{V}}$ (or their combined forms $\bm{W}^{a^{Q}}$, $\bm{W}^{b^{Q}}$, etc.), using Xavier initialization [15].
 Specifically, each entry of a weight matrix is drawn from a uniform distribution $\mathcal{U}(-bound,bound)$, where $bound=\sqrt{6/(n_{\text{in}}+n_{\text{out}})}$.
@@
 Training Setup Details.
-We follow the nanoGPT training configuration [Karpathy2022].
-In particular, we use the AdamW [loshchilov2017decoupled] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$.
-We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [loshchilov2016sgdr] with 2,0002{,}000 warmup steps and a (total) global batch size of $480$.
+We follow the nanoGPT training configuration [24].
+In particular, we use the AdamW [35] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$.
+We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [36] with 2,0002{,}000 warmup steps and a (total) global batch size of $480$.
 For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
@@
 Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
-MHA 64.44 32.85 59.05 44.18 33.20 68.72 50.12 26.01 87.40 49.44
+MHA 64.44 32.85 59.05 44.18 33.20 68.72 50.12 26.01 87.40 51.77
 MQA 64.27 32.94 57.71 44.36 31.80 68.01 51.70 25.99 86.00 51.42